A Complete Bibliography of Publications in Fish Physiology and Biochemistry: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

07 April 2025 Version 1.08

Title word cross-reference

−150 [1230]. ^{+} [962, 656, 619, 180, 277, 578, 179, 197, 1252, 818, 890]. ^{−} [619, 277, 1164]. ^{1} [1324]. ^{2+} [656, 578, 804, 460]. _{1} [1199, 882, 643, 1249]. _{2} [1364, 470, 154, 108, 643, 18, 897]. _{3} [996, 1071, 129, 1048]. _{3}^{−} [1164]. _{4} [612, 151]. _{A} [1262, 1164, 645]. _{crit} [1335]. _{w} [154]. α [934, 962, 100, 748, 128, 1207, 451, 214, 753, 271, 1044, 526, 490, 1193, 749, 898, 1218, 886, 831, 1179, 138, 818, 958, 836, 631]. β [125, 1280, 748, 114, 90, 128, 987, 480, 20, 570, 328, 724, 725, 609, 861, 1193, 862, 611, 18, 1394, 1169, 1172, 941, 586]. Δ [918]. γ [62, 1262, 461]. N [586]. n − 3 [918]. ′ [611]. × [119, 1052, 142, 1084, 1210, 1150, 623, 83, 28, 1370, 1007]. U_{crit} [470]. y^{+} [547].

-1 [609]. -1b [1387]. -acetyl- [586]. -Alanine [724]. -amylase [271]. -ATPase [962, 180, 277, 179, 197, 1252, 818, 890, 619, 1164]. -binding [478]. -carotene [480]. -catenin [1169]. -contaminated [1199, 1249]. -cypermethrin [898]. -d-glucosaminidase [586]. -estradiol [987, 611].

-ethynylestradiol [214]. -glucan [941]. -glucuronidase [114, 90]. -glutamylcarboxylase [62]. -gustducin [749]. -hydroxylase [100]. -hydroxylase/17 [100]. -induced [612, 151, 897]. -K [818]. -like [1030]. -methyltestosterone [631]. -naphthoflavone [725]. -oxidation [862]. -RA [737]. -regulated [996]. -stimulated [318]. -subunit [451]. -tocopherol [753]. -Transducin [749, 526]. -tubulin [831].

[?]2ClChandrasekar:2014:ABW. /activin/nodal [328]. /HCO [1164]. /K [962, 180, 179, 197, 1252, 619].

1 [934, 913, 742, 1297, 613, 852, 490, 1372, 860, 719, 145, 1218, 1347, 886]. 1-methyl-4-phenyl-1 [707]. 10-phenanthroline [318]. 11-ketotestosterone [396]. 12 [584]. 1254 [91]. 133a [789]. 14643 [481]. 17- [18]. 1758 [445]. 1b [1387, 1086].

2 [677, 786, 1068, 330, 893, 1385, 1396, 525]. 2-like [818]. 2-oleoyl-glycerol [841]. 2-phenoxyethanol [362]. 20 [344]. 20-lyase [100]. 28a [1154]. 2a [1377]. 2AA [1196]. 2Cl [619, 277]. 2D [1025]. 2D-DIGE [1025]. 2s [1377].

3 [602, 382, 861, 461]. 3/1 [609]. 34a/Sirtuin [1372]. 3A [497].

4-D [728]. 4-hydroxylase [1381]. 4-hydroxynonenal [494]. 4-nonylphenol [1192, 663, 664, 1065]. 462/731 [1045]. 47 [487].

50 [1137].

6-bisphosphatase [1086, 767]. 6-glucan [609]. 6-phosphate [447]. 6-tetrahydropyridine [707]. 64 [52].

7 [1154, 1330]. 70 [445]. 731 [1045].

9 [893]. 9- [737].

a-histological [664]. ability [600, 1285, 1361, 772, 1109, 1058]. abnormalities [1085]. Abramis [181]. Abrupt [1053]. absorption [1379, 421, 1367]. absorptive [1248, 1109, 887, 1370]. abundance [619]. ac [103]. Acanthopagrus [74, 390, 1377]. acceleration [902, 946]. acceptability [827]. access [1342]. acclimated [512, 280, 471, 994, 809, 890]. Acclimation [619, 550, 9, 301, 904, 562, 416, 902, 529, 249, 618, 339, 1122]. acclimatised [331]. Accumulation [503, 1133, 1265, 977, 593, 1295, 1394, 990]. acetaminophen [482]. acetaminophen-fed [482]. acetate [753]. acetate-induced [753]. acetyl [1058, 586]. acetylated [831]. acetylcholinesterase [132, 185, 751]. acetylcysteine [1128]. Achieving [237]. acid [325, 1085, 280, 1173, 188, 177, 842, 141, 708, 1176, 473, 791, 724, 637, 464, 186, 48, 1024, 517, 1149, 918, 1133, 1329, 50, 736, 84, 1037, 777, 862, 744, 217, 199, 496, 1310, 895, 82, 54, 499, 582, 697, 274, 672, 965, 516, 1051, 41, 249, 702, 763, 1033, 1141, 772, 1080, 1349, 1183, 547, 505, 1050, 1367]. acid-binding [842]. acid-deficient [1183]. acid-induced [1176]. acid-supplemented [895]. acid-to-eicosapentaenoic [199]. acidic [937]. acidilactici [1203]. acids [348, 1063, 796, 950, 1074, 312, 1098, 637, 70, 130, 988, 421, 54, 780, 460, 442, 171, 974]. Acipenser [907, 949, 365, 951, 1355, 399, 999, 175, 1382, 877, 570, 929, 1084, 1066, 263, 1224, 585, 262, 367, 688, 1335, 1324, 1112, 4, 969, 538, 1102, 1048, 1236, 1321, 302, 1081, 1078, 1151, 890]. Acknowledgments [863]. acrylate [1312]. acrylate-induced [1312]. act [724]. actin [20]. action [482, 705]. actions [27]. activate [115]. activated [1050]. activation [1063, 756]. active [486, 555, 1212]. activin [328]. activities [272, 568, 951, 136, 651, 1299, 709, 788, 805, 609, 1250, 872, 142, 1149, 74, 256, 78, 871, 870, 1205, 825, 594, 781, 604, 868, 773, 381, 715, 475, 921, 1302, 281, 358, 603, 247, 1183, 826, 1370]. Activity [1067, 801, 1168, 278, 532, 3, 700, 1126, 1382, 851, 1291, 180, 132, 970, 1010, 521, 77, 342, 1293, 343, 447, 1038, 537, 492, 271, 1145, 1306, 163, 1008, 562, 695, 1037, 777, 454, 1070, 61, 610, 767, 434, 640, 1326, 474, 566, 1257, 1022, 1156, 1147, 1378, 1346, 1381, 899, 1343, 378, 185, 931, 783, 346, 940, 1249, 1167, 816, 751, 193, 1252, 1072, 379, 558, 1027, 890, 313, 173, 269]. aculeatus [1342, 884]. Acute [934, 725, 1182, 261, 1036, 606, 742, 1139, 1098, 361, 800, 1012, 1093, 1137, 910, 1210, 1272, 986, 694, 807, 1226, 174, 1335, 1190, 1051, 255, 1349, 1359, 1143, 941, 649, 96]. Acyl [862, 918, 1179]. Acyl-coenzyme [862]. adaptation [1152, 1086]. adaptations [35, 808]. adapted [696]. Adaptive [890]. adaptor [366]. addition [735]. additive [1032, 807, 527, 1343]. additives [51]. adenohypophyseal [630, 615]. adenosine [810]. adequate [1163]. adhesion [1237]. adipocytes [903, 1133, 659, 1097]. adipogenesis [1097]. adiponectin [1206]. adipose [1207, 1176, 413, 771, 208]. adjustments [320]. administration [1135, 642, 275, 765, 582, 310, 1343]. Adriatic [1234]. adspersus [75]. adult [423, 636, 914, 725, 190, 594, 732, 817, 702, 1116]. aeneus [1218]. aerial [1027]. Aerobic [899, 73, 1078]. Aeromonas [1390, 1076, 1103, 1056, 408]. aeruginosa [503, 298]. aff [1032]. affect [849, 1285, 695, 232, 887]. affected [1299, 270, 1044, 199, 981, 1254, 631]. affects [993, 494, 520, 950, 1288, 1071, 251, 319, 1104, 1301, 728, 448, 69, 892, 1370]. affinis [40, 1157]. Affinity [318, 1191, 539, 704]. aflatoxin [1199, 882, 148, 1249]. Africa [141]. African [1192, 568, 58, 1113, 541, 61, 54, 332, 921]. after [365, 1135, 642, 1207, 240, 275, 1247, 209, 805, 609, 170, 184, 583, 252, 910, 852, 1086, 377, 839, 560, 445, 1228, 1005, 1344, 481, 333, 1303, 836]. against [1390, 1064, 1113, 1250, 296, 1275, 1103, 1272, 36, 1075, 445, 569, 274, 644, 897, 408, 151, 1059]. agalactiae [891, 661]. Age [1257, 1067, 1003, 1146]. Age- [1257]. age-related [1003, 1146]. agent [1358, 732]. agents [817, 689]. ages [796]. aggregation [215]. aggression [291]. aggressive [1352, 181]. aggressiveness [1036]. Aging [1003]. agmatine [1375]. agonists [722]. AGP [697]. agricultural [332]. Air [414, 2, 1314, 199, 1282, 743, 194]. Air-breathing [414, 2, 1314, 743, 194]. akaara [323]. AKG [1163]. Akt [1301]. AKTs [1034]. AKTs/PKBs [1034]. Alanine [724]. alarm [993]. alba [288, 835, 1119, 1258, 597]. albinism [1096]. albino [566, 1096]. albonubes [20, 284, 18]. albuginea [635]. albula [42]. albumin [43]. Alburnus [1136, 1266]. albus [821, 353, 681, 959, 634, 182, 22, 171]. aldrin [1228]. alewives [249]. alexandri [1282]. alfonsino [899]. alginic [188]. Alismatis [1176]. alkaline [791, 77, 169, 1095, 505]. alkalinity [818]. all-plant-based [608]. Allenbatrachus [402]. alleviation [1049, 989]. alligator [784]. allodiploid [1118]. allogynogenetic [119, 857, 408]. allometric [853, 1001]. allostatic [583]. allotriploid [1118]. Aloe [805, 465]. along [761, 263, 537, 492, 75]. Alopias [351]. Alosa [1010, 1037, 249]. Aloysia [566, 1155]. alpha [697, 704, 481, 1163, 582]. alpha-1-acid [697]. alpha-ketoglutarate [1163]. alter [1209]. Alteration [852, 516, 357, 1312, 460]. Alterations [338, 709, 178, 1167, 663, 664, 502, 605, 64, 102, 1354, 438, 890]. altered [795, 1200, 1220]. altering [287]. alternative [1230, 1107, 1020]. alters [423, 1110, 708, 613, 550, 327, 291]. altiparanae [587, 937, 1060]. Aluminum [86, 937]. Aluminum-induced [86]. always [161]. AM [906]. Amatitlania [738]. Amazon [934, 1138]. Amazonian [180, 451]. amazonicus [1273, 291, 819, 1036, 349]. ambient [521, 343, 306]. amblycephala [1208, 608, 953, 1028, 1333, 556, 1045, 1278, 1330, 264, 765, 736, 1086, 1184, 1301, 1271, 1372, 983, 1051, 1056, 737, 887, 28, 718, 1050, 1296, 1251, 513]. ameliorates [1144]. Ameliorating [480]. Ameliorative [1329, 1334]. amenable [828]. American [1010, 1037, 1174]. amino [325, 950, 177, 724, 186, 48, 1024, 50, 130, 421, 895, 54, 763, 1080, 1349, 1183, 547, 1367]. aminopeptidase [95]. Ammonia [984, 1115, 119, 742, 1175, 1314, 254, 299, 718, 1027, 1254]. among [1290]. amperometric [419]. amphetamine [476, 1151, 1315]. amphetamine-regulated [1151, 1315]. amphioxus [381]. Amphiprion [33, 430]. AMPK [958]. Amur [262, 688, 1112]. Amursky [203]. amylase [271]. amylolytic [521]. amylopectin [657]. amylose [657]. amylose/amylopectin [657]. Anabas [244, 504, 9]. Anablepidae [837]. anadromous [557]. anaemia [133]. anaerobic [899]. Anaesthesia [236, 1138, 354, 1359]. anaesthetic [252]. analyser [1364]. analyses [734, 459, 14, 1125, 869, 1193, 1153, 479]. Analysis [535, 772, 1177, 3, 1274, 126, 17, 125, 52, 677, 908, 1280, 368, 786, 973, 432, 953, 183, 1309, 190, 1395, 187, 556, 674, 904, 290, 1171, 812, 1241, 318, 1330, 202, 1316, 1222, 404, 1013, 1393, 1239, 418, 515, 499, 329, 394, 704, 661, 840, 1346, 976, 820, 1119, 201, 266, 497, 680, 1056, 121, 511, 1368, 1340, 1095, 1206, 138, 737, 845, 1371, 1303, 28, 815, 355, 315, 591, 958, 957, 1089, 1377, 998, 269]. analytic [474]. analyzing [871, 870]. Anarhichas [563]. ancient [1197]. and/or [1034]. And4 [1316]. androgen [922, 413, 187, 420, 775, 441, 1099]. Androgen-induced [1099]. androgenesis [535, 395]. androgens [365, 181]. anemonefish [33]. Anesthesia [71, 470, 1188, 1189, 1258, 597]. Anesthetic [566, 362, 1119]. anesthetics [1119]. Angelica [1248]. Anguilla [8, 287, 904, 273, 469, 31]. anguillarum [377, 445]. anguillicaudatus [1118, 754, 1316, 506]. anguillid [1152]. anhydrase [1205, 818]. animal [225, 673, 1336]. animals [1008, 346]. anion [700]. anisitsi [414]. Annual [186, 367, 439, 1003, 1120]. Anonymous [719]. Anoplopoma [1092, 848]. anorexigenic [1139, 896]. anoxic [167]. antagonist [427]. antagonists [722]. Antarctic [1357, 529, 797, 809]. antennata [624]. anterior [590]. Anti [608, 966, 786, 967, 970, 752]. anti- [967]. anti-apoptotic [786]. Anti-oxidative [966, 752]. anti-proteolytic [970]. Anti-stress [608]. antiapoptotic [903]. antibacterial [1343]. antibiotic [625]. antibiotics [144]. anticonsolidation [920]. antidandruff [1360]. antifouler [1360]. antifreeze [1159, 1236]. antimicrobial [1358, 778]. Antioxidant [581, 1250, 19, 1344, 1168, 348, 568, 1114, 951, 1355, 758, 443, 1134, 258, 1382, 851, 682, 877, 882, 906, 625, 480, 1031, 440, 805, 609, 1084, 753, 6, 296, 793, 1144, 627, 1312, 555, 594, 595, 736, 1248, 179, 1223, 530, 569, 1257, 968, 350, 253, 1361, 886, 715, 654, 921, 1320, 69, 644, 751, 358, 1321, 151, 1011, 1261, 718, 1370, 313, 990, 1296]. antioxidant-related [793]. antioxidants [154, 735, 1200, 685]. Antioxidative [36, 338, 1150, 1068, 305, 778, 1058]. antipsychotics [810]. antitumor [379]. antixoidant [1192]. Anxiogenic [214]. AOF [1255]. aorta [240, 1012]. aorta-cannulated [1012]. apical [486]. Apium [482]. apolipoprotein [1376]. apoptosis [245, 849, 1133, 1329, 535, 804, 983, 669, 286, 1385, 1396, 1331]. apoptotic [786]. appear [985]. appetite [792, 588, 412, 1002, 1151, 1370]. appetite-regulating [588]. Application [500, 303, 387, 1372, 1096]. applications [238]. applied [1232]. approach [126, 1342, 196, 386, 221, 554, 889, 1273, 954]. approaches [791]. appropriate [734]. aquacultural [228]. aquaculture [620, 431]. Aquaculture(R) [732]. aquaporin [388]. aquarium [181]. aquatic [1240, 985]. aqueous [1326]. arachidonic [280, 199, 496, 1033, 1141]. Arapaima [451]. architectural [148]. area [621, 369, 342, 899]. areus [142]. argenteus [277, 110]. Argentina [279]. Arginine [1288, 1180, 466, 33, 421, 1042, 91, 85, 315]. Argulus [456]. Argulus-infested [456]. argus [1150, 1395, 46, 988, 123, 813]. Argyrosomus [1160, 331, 994, 12, 820, 567, 475, 508]. Aristichthys [1297, 19, 1380, 379]. arjuna [866]. armata [1061]. Aroclor [91]. aromatase [894, 851]. arrested [1178]. Arsenic [146]. Artemia [56, 815]. arterial [154, 108]. artery [252]. Arthrospira [1383]. artificial [1117, 1229, 287, 303]. artificially [1009, 370]. asafoetida [1343]. asbt [582]. ascorbate [74]. ascorbic [1085, 274]. Asejire [58]. Asian [353, 409, 1287, 82, 629]. asiaticus [1103, 1091]. Asparagopsis [1061]. aspects [1043]. assay [563, 504, 1030, 203]. assess [884, 800, 1105, 485]. Assessing [952, 158, 16].

Assessment [258, 369, 751, 107, 883, 459, 1231, 238, 691, 592, 106, 991]. assessments [528]. assisted [896]. Asso [475]. associated [328, 1008, 221, 1354, 756, 770, 961]. association [977]. assurance [239]. astaxanthin [1134]. Astragalus [906, 296]. Astronotus [934, 207, 1255]. Astyanax [1279, 587, 1032, 937, 1060, 1259]. asymmetry [1003]. ATGL [1207]. Atlantic [1063, 45, 1341, 1017, 1300, 467, 35, 1364, 115, 1067, 621, 493, 240, 267, 637, 342, 484, 785, 673, 160, 1024, 152, 517, 103, 583, 822, 1073, 154, 252, 205, 108, 241, 62, 87, 841, 254, 299, 554, 895, 304, 47, 495, 206, 133, 1021, 947, 705, 401, 121, 83, 901, 200, 71]. atlanticus [899]. ATPase [136, 962, 619, 180, 277, 179, 1164, 886, 197, 1252, 818, 890]. ATPases [426]. Atractosteus [278, 788, 874, 784, 622]. atresia [907, 949, 762]. atripinnis [179]. Atrogin [1347]. Atrogin-1 [1347]. atrophy [221]. atrophy-associated [221]. attaching [979]. attachments [1021]. attempt [109]. attenuates [15, 446, 612, 983]. audiometry [1373]. auditory [731, 1373]. augments [1068]. aurata [1214, 1014, 620, 1283, 602, 945, 1061, 883, 371, 89, 11, 186, 806, 787, 130, 1317, 295, 217, 1140, 860, 924, 134, 719, 1022, 449, 635, 536, 477, 1042, 527, 1166, 442, 1025, 940, 359, 168, 553, 1383]. auratus [119, 857, 1265, 1262, 977, 645, 305, 408, 286, 618, 836, 961, 1315, 1376, 1110, 677, 913, 1280, 1197, 510, 896, 259, 854, 1055, 936, 456, 465, 145, 756, 638, 699, 1277, 567, 476, 1349, 28, 1089, 750, 631, 294]. aureus [1052, 472, 1007]. aurita [143]. Australia [956]. autophagy [1132]. availability [602]. available [673]. avermectin [836]. avidin [324]. avoidance [316]. awoara [988, 191]. axis [742, 583, 924, 994, 801, 1313, 1035, 1258, 1209, 1079, 1370]. axis-related [1370]. azadirachtin [456]. azo [1167].

B [689, 663, 696, 1020, 1199, 882, 1366, 1249, 772, 740]. b-histological [663]. b-wave [696]. b5 [454, 378]. Bach [719]. Bacillus [1098, 872, 358]. background [412]. bacteria [87, 97]. bacterial [786, 333]. baerii [951, 1355, 399, 1224, 1048, 1151]. balance [1376, 209]. Balkan [1233]. bambooshark [404]. banded [988]. barb [1253, 974]. Barbonymus [1253, 974]. Barbus [817]. barfin [290, 203]. bark [866]. barostatic [927]. barramundi [532, 1038, 575]. barrier [1076]. base [750]. based [1305, 196, 608, 973, 1098, 1241, 1195, 257, 721, 1030, 305, 892, 247, 1268, 436]. basin [519, 279]. basis [161]. bass [3, 962, 846, 1187, 443, 748, 1071, 132, 317, 473, 616, 409, 550, 246, 471, 872, 376, 526, 98, 82, 733, 445, 768, 611, 350, 38, 1361, 565, 1141, 1327, 720, 573, 168, 452]. batch [880]. bathed [1061]. batrachus [2, 146, 1016, 194, 692]. Bax [1054]. Bax1 [939]. Bax2 [939]. Bay [203, 193]. BCL [786]. BCL-2-like [786]. BDE [487]. BDE-47 [487]. be [862, 961, 1315]. bean [1098]. bean- [1098]. beans [1295]. Beclin [839]. beet [1190]. before [1267, 583]. Behavior [803, 1352, 510, 1336, 476, 1155, 414]. Behavioral [1360, 218, 1100, 434, 383, 309]. behaviour [214, 181, 398]. Behavioural [226]. belcheri [381]. belina [1346]. Beluga [176, 184, 260, 80, 1203, 139, 549]. Beneficial [1263]. benefits [1208, 1296]. benthic [216]. benthivorous [808]. benzo [1125]. Berberine [983, 1251]. Beryx [899]. bet [461]. Beta [431, 125, 450, 284, 1324, 24]. beta-cyclodextrin [1324]. Beta-glucan [431]. betaine [1114, 307, 428, 1190]. Betanin [612]. Betta [93, 268, 1201]. between [340, 614, 189, 72, 10, 637, 342, 744, 181, 689, 1021]. biased [1395]. bidirectional [1151]. big [221]. big-snout [221]. bighead [340, 1297, 19, 665, 1373, 1380, 379]. bigmouth [340]. bile [1221, 582]. bimaculatus [1032, 360, 921]. binary [987, 699]. bind [600]. Binding [668, 701, 382, 677, 913, 842, 172, 152, 531, 46, 1387, 584, 318, 478, 704, 679, 525]. bioaccumulation [928, 1049, 989]. bioactivation [530]. Biochemical [435, 64, 628, 1164, 60, 110, 325, 512, 212, 422, 503, 8, 842, 219, 218, 1176, 1360, 1031, 791, 76, 1203, 739, 1293, 184, 139, 298, 1038, 675, 456, 762, 910, 1216, 776, 137, 560, 728, 370, 1069, 759, 1263, 310, 629, 1202, 965, 1051, 834, 1220, 660, 321, 1167, 357, 407, 408, 1348, 943, 1091, 79]. biochemistry [192, 597, 1359]. Bioenergetics [1369, 5, 457, 880]. biofloc [868]. Bioinformatic [1153]. bioinformatics [1366]. Biological [692, 1286, 298]. biologically [1212]. biology [42]. biomarker [1342, 973, 132, 75, 592, 1338, 16, 1273, 441, 650, 106]. biomarkers [1192, 928, 1322, 1003, 70, 884, 762, 910, 1389, 938, 856, 644]. biomass [675, 1170]. biopsy [952]. biorefinery [1327]. biosensor [572]. biosynthesis [312, 841, 123, 1179]. biotechnological [386]. Biotechnology [1232]. biotin [581, 324]. biphenyls [954]. Bisphenol [954, 1247, 26, 1397, 27]. bisphenol-A [1247]. bisphosphatase [1086, 767]. bitterling [902]. black [119, 74, 641, 733, 926, 1167, 390, 1206, 1183, 1377]. blackspot [111, 112]. bladder [1381, 567]. Bleeker [1253, 1245, 1244, 974]. blend [846]. blended [376]. blennie [678]. blenny [398]. blind [412]. blind-side [412]. Bloch [113, 325, 696, 1298, 244, 504, 47, 710, 704, 360, 629, 9, 60, 995, 479]. Blood [648, 147, 1392, 212, 1191, 126, 177, 621, 992, 184, 260, 1107, 1294, 849, 1012, 1093, 154, 250, 283, 1272, 57, 97, 856, 133, 178, 4, 1281, 1080, 1349, 79, 452]. blue [823, 1277, 472, 1011]. bluefin [1017, 1300, 421, 195, 300, 39, 540, 683, 942, 47]. bluegill [906]. blunt [1114, 1207, 608, 953, 1028, 1333, 1278, 264, 765, 736, 1086, 1184, 1301, 1271, 983, 1051, 887, 28, 718, 1050, 1130, 1296, 1251, 513]. blunt-snout [765]. blunts [1138]. BMP [893]. bocourti [944]. body [113, 462, 1160, 443, 246, 120, 828, 322, 1306, 283, 853, 1001, 1216, 437, 292, 54, 66, 1156, 1170, 247, 1348, 1143, 1027, 1089, 760]. bogaraveo [111, 112]. Bogue [88]. Boleophthalmus [1083, 1177]. bonariensis [396, 574]. bonasus [342]. Bone [1013, 972, 1275, 149, 1021]. bones [979]. bonito [47]. bonnethead [650]. bonytongue [54]. Boops [88]. borax [1227]. borne [209]. Bostrichthys [643]. botanical [178]. both [1131, 146, 555, 769]. bovine [662, 1328]. box [591]. box25 [1347]. Boxshall [850]. brachysoma [301, 44, 1245, 1244]. brackish [619, 471, 65, 890]. Brain [285, 348, 1227, 1304, 222, 1289, 1352, 100, 1247, 380, 922, 33, 852, 1164, 400, 1313, 91, 85, 60, 810, 751, 476, 1349, 966, 676]. brain-pituitary-gonadal [1313]. brama [181]. bran [1198]. branched [724]. branched-chain [724]. Branchial [165, 619, 15, 277, 593, 210]. Branchiostoma [381]. Brandt [951, 1224]. brasiliensis [615]. Bravo [850]. Brazil [369]. Brazilian [967, 1115, 1223]. breaks [750]. bream [1114, 620, 1219, 602, 1299, 883, 858, 270, 1207, 992, 1264, 15, 608, 89, 275, 406, 953, 186, 1028, 806, 1333, 1049, 989, 243, 1293, 1107, 1294, 1379, 74, 1278, 99, 1101, 130, 264, 765, 736, 1086, 1184, 1301, 1271, 983, 295, 1140, 860, 924, 1105, 449, 181, 527, 442, 448, 1025, 1051, 940, 1281, 946, 887, 28, 168, 718, 1050, 1130, 1296, 1251, 513]. breathing [2, 1314, 743, 194, 414]. breeding [369, 1047, 181, 1245, 1244]. Brevimyrus [1099]. brevirostrum [585, 367, 1335, 1078]. brewer [1383]. brill [411]. British [1387]. broadband [1373]. brominated [487]. bronchialis [834]. bronze [437, 1218]. brood [109]. brooding [85]. broodstock [1214, 980, 217, 1239, 1245, 1244, 820]. brook [656, 159]. brown [960, 417, 433, 413, 1073, 862, 1257, 1005, 83]. brown-marbled [960, 417]. Brunnich [112]. Brycon [129, 1273, 717, 291, 819, 1036, 349]. buffalo [340]. bulk [928]. bullhead [251, 441]. buoyancy [567]. burbot [537, 492]. Burchell [58, 1085, 61, 1228]. Burkitt [668, 701]. burst [1126, 1345, 888]. butter [360]. Butyl [1312]. butyrate [527, 1269]. by-product [1293, 332].

C [1230, 605, 56, 1334, 274, 671, 1087, 794, 751]. c-Fos [751]. c-myc [794]. C-terminal [1087]. Ca [804, 460, 656, 578]. Cadmium [423, 458, 448, 913, 405, 258, 209, 11, 1049, 989, 140, 539, 284, 306, 839, 1193, 1265, 731, 149, 137, 40, 552, 966]. cadmium-induced [40]. caerulescens [509, 453, 1158]. Caffeine [1289, 1325]. caffeine-supplemented [1325]. cage [629, 931]. caging [1342, 884]. calcarifer [532, 409, 1038, 1287, 82, 629, 575]. calcification [259]. calcitonin [548]. calcium [215, 600, 172, 306, 533]. calcium-binding [172]. California [505]. californicus [505]. Caligus [850]. Callichthyidae [1218]. CaM [807]. camelina [517, 517]. Camellia [1084]. cAMP [1018]. campechanus [48]. Can [862, 841, 1361, 1349, 483]. Canada [1387]. canadum [432, 1323]. canaliculatus [1179]. candidate [1203]. Cannabinoid [501]. cannulated [1012]. cannulation [240, 252]. capabilities [887]. capability [999, 1261, 718, 1296]. capacities [73, 1341, 791, 1248]. capacity [1168, 1114, 1283, 520, 1134, 621, 1031, 440, 793, 1144, 1272, 1335, 654, 1320, 809, 1081, 1367, 1370]. captive [907, 949, 126, 571, 1245, 1244, 485]. captivity [439, 331, 814, 1185]. carangid [334]. Carassius [1376, 1392, 1110, 119, 677, 913, 1280, 1197, 857, 696, 510, 896, 259, 854, 1055, 936, 456, 1265, 1262, 465, 145, 756, 977, 645, 638, 699, 1277, 305, 407, 476, 1349, 408, 286, 1268, 28, 618, 1089, 750, 836, 631, 294, 961, 1315]. carbamate [136]. carbofuran [1279, 1113]. carbohydrate [119, 1299, 858, 89, 1391, 963, 531, 1278, 265, 1086, 965, 1206]. carbohydrate-binding [531]. carbohydrate-rich [1299]. Carbon [419, 1279, 467, 612, 296, 1075, 151]. carbonate [818]. Carbonic [818, 1205]. carboxykinase [1323]. carboxylesterase [592]. carcass [436]. Carcharias [189]. Cardiac [133, 878, 544, 415, 1116]. cardiac-type [1171]. cardiorespiratory [725]. cardiosomatic [1116]. cardiovascular [878]. carnitine [1064, 811, 897]. carnosine [724]. carotene [480]. carotenoids [654]. carp [1168, 365, 340, 1392, 636, 119, 1229, 382, 542, 1064, 1288, 114, 599, 851, 1176, 1260, 625, 1391, 857, 380, 440, 1297, 696, 10, 86, 724, 14, 729, 612, 1358, 105, 343, 522, 447, 752, 811, 917, 140, 296, 1148, 1093, 1133, 1329, 1137, 1241, 164, 458, 539, 1195, 675, 116, 1041, 19, 327, 594, 804, 852, 1002, 1248, 36, 57, 659, 771, 1075, 374, 781, 1204, 665, 1393, 149, 551, 1365, 370, 868, 1129, 1373, 545, 879, 1200, 1356, 926, 55, 801, 1343, 1202, 965, 516, 685, 1196, 321, 1108, 1167, 1020, 1351, 680, 1046, 1033, 1295, 407, 461, 559, 939, 897, 865, 358, 408]. carp [1206, 1380, 997, 1097, 1058, 686, 1371, 1307, 1348, 1303, 1268, 28, 483, 547, 379, 151, 463, 955, 1261, 618, 1089, 1367, 1122, 313, 631, 95, 24, 294, 714, 961, 1269, 1315, 589]. carpet [167]. carpio [1260, 140, 1168, 636, 119, 1229, 382, 851, 1176, 440, 209, 102, 86, 510, 920, 523, 612, 105, 343, 522, 752, 296, 1148, 1137, 1241, 164, 458, 539, 1195, 675, 116, 327, 804, 852, 1248, 1204, 149, 551, 1164, 868, 569, 55, 1343, 1202, 685, 321, 1351, 680, 1307, 1348, 151, 1261, 714]. carps [973, 920, 117]. carrying [34, 621, 1272]. CART [476, 1151]. case [237]. caspase [430]. caspase-3 [430]. caspases [804]. Caspian [915, 671]. caspicus [671]. catabolism [724, 50]. catalase [102]. Catalasic [474]. catalytic [676]. catecholamines [577]. catenin [1169]. catfish [1192, 211, 439, 568, 113, 325, 58, 288, 2, 285, 795, 833, 996, 301, 944, 708, 464, 1113, 1015, 849, 187, 1144, 716, 336, 762, 933, 1298, 1314, 910, 84, 25, 1178, 824, 1016, 541, 1282, 61, 728, 764, 691, 332, 534, 566, 856, 66, 360, 44, 178, 743, 1313, 194, 692, 981, 1220, 823, 1174, 528, 1119, 1258, 1249, 866, 921, 597, 712, 403, 652, 603, 355, 315, 1172, 124, 347, 147, 1331, 1254, 1370, 649, 345, 1034, 1026]. Cathepsin [377, 201, 816, 740]. cationic [547]. Catla [1247, 1125, 879, 1200, 886, 357, 436]. catshark [425]. caudal [252, 1255]. caught [1384, 47]. caused [947, 727, 294]. causes [1314]. cavity [1226]. CB1 [501]. CB2 [501]. CB2-like [501]. CC [17, 368, 432]. CCl [612, 151]. Cd [64, 336]. CD231 [355]. CD36 [1046]. CD79a [1103]. CD79b [1103]. CDK [1385, 1396]. CDK-2 [1385, 1396]. cDNA [677, 80, 716, 336, 1125, 153, 976, 511]. ceca [293]. celery [482]. cell [1305, 2, 746, 1332, 375, 509, 453, 1132, 854, 1055, 535, 1255, 386, 131, 825, 323, 754, 1124, 774, 1233, 1262, 540, 1105, 1090, 689, 1174, 1077, 344, 755, 831, 122, 118, 865, 1096, 652, 208, 311, 68, 363, 513]. cell-specific [746]. cells [1392, 960, 126, 599, 180, 486, 630, 1158, 103, 668, 701, 670, 849, 277, 909, 526, 1000, 39, 749, 1006, 982, 392, 1090, 1211, 823, 1351, 897, 147, 890, 615]. cellular [277, 574]. cellularity [1104]. centers [933]. central [1139]. Centropomus [256, 745]. Centropristis [733]. cephalus [548, 330, 669]. cerebellar [790]. cerebral [1199, 1249]. cerevisiae [350, 1383]. cerium [36]. certain [329, 516, 1167, 892]. cervinus [858]. cGnRH [80]. cGnRH-II [80]. chain [455, 1074, 312, 724, 637, 806, 918, 1171, 49]. chains [112].

Chalcalburnus [535, 1205]. chalcone [871, 870]. challenge [786, 738, 486, 805, 1357, 555, 377, 694, 807, 210, 819]. challenged [568, 381]. chamber [584]. chamber-specific [584]. change [372, 430, 392, 1335, 216, 335, 21]. changed [282, 1336]. Changes [514, 907, 949, 1199, 1242, 651, 795, 176, 48, 1293, 154, 777, 604, 195, 297, 1302, 723, 193, 826, 1027, 435, 930, 2, 619, 1311, 285, 913, 1003, 616, 1031, 1010, 186, 120, 646, 598, 343, 557, 782, 1092, 221, 1060, 935, 924, 148, 1162, 498, 60, 981, 727, 875, 964, 475, 751, 279, 407, 1380, 1143, 1122]. changing [1362]. Channa [1150, 700, 222, 930, 215, 577, 258, 918, 46, 1150, 988, 710, 704, 840, 59, 60, 261, 813, 995]. Channel [355, 84, 66, 823, 315, 345]. characid [1259]. Characidae [1060, 814, 349]. characiform [519]. Characiformes [349, 717, 615]. Characterisation [1350, 88, 596]. characteristics [58, 365, 402, 184, 56, 63, 137, 976, 1281, 79]. Characterization [1014, 1159, 960, 996, 967, 970, 953, 1154, 1028, 1009, 323, 1193, 641, 683, 611, 592, 843, 982, 416, 684, 710, 661, 1035, 366, 680, 1082, 1083, 1371, 1307, 425, 558, 676, 962, 1186, 1311, 17, 382, 677, 542, 885, 855, 908, 368, 786, 748, 738, 175, 842, 561, 23, 987, 1391, 409, 380, 1297, 798, 1332, 1374, 1309, 622, 963, 334, 447, 187, 556, 674, 844, 1094, 1278, 290, 1055, 99, 143, 1287, 762, 1255, 131, 450, 754, 1103, 1086, 1184, 1386, 1271, 284, 293, 410, 802, 932, 839, 1124, 563, 469, 688, 134, 942, 1356, 364, 420, 1347, 507, 660, 916, 755, 816, 81, 127]. characterization [403, 122, 118, 191, 511, 679, 939, 900, 919, 1096, 1276, 652, 182, 573, 1206, 311, 547, 379, 355, 315, 1172, 463, 958, 30, 213, 1062, 836, 363, 24, 740, 269, 513, 1034, 1026, 349]. characterizations [1054]. Characterizing [833]. charr [159]. Chelon [1147]. chemical [993, 250, 853, 1001, 442]. chemicals [1240, 276]. chemistry [992, 1107, 283, 367, 648, 133, 1005]. Chemokine [885, 17, 52, 368, 432]. chemotypes [1119, 1258]. Chen [124, 347]. chicken [80]. Chilean [75]. chilling [267]. Chiloscyllium [167, 1039, 404]. Chinese [175, 1131, 1103, 1040, 946, 1082, 302, 68, 1091, 1081, 826, 1027, 714]. Chinook [104]. Chirostoma [312]. chitosan [915]. chloride [166, 480, 546, 197, 751]. chlorpyrifos [274]. chlorpyrifos-induced [274]. choice [920]. cholecystokinin [380]. cholesterol [690, 1052, 465, 572]. choline [307, 428]. cholinesterases [1014, 1350, 676]. Chondrichthyes [180]. chopped [571]. choriogenin [59]. chorzoxazone [327]. chromatin [1022]. chromatophores [577]. Chromis [666]. chromium [1216, 57]. Chronic [319, 629, 1285, 852, 254, 299, 623, 1088, 174, 261, 727, 1108, 720, 943, 649]. Chrysiptera [707]. chuatsi [1131, 1375, 919, 468]. chub [911, 7, 297]. chum [470, 782, 156]. chyme [163]. chymotrypsin [403]. Cichlasoma [596, 706, 150, 723, 769]. Cichlid [446, 738, 596, 706, 150, 654, 723, 1079, 769]. Cichlidae [640, 32, 384]. cilia [831]. cimetidine [522]. cinnamon [430]. circadian [434, 1185]. Circulating [782, 411, 181]. Cirrhinus [592, 1350, 516]. cirrosa [462]. cis [737]. Cited3 [1333]. Citral [1119, 1258]. Cl [1164]. Clarias [1192, 211, 568, 58, 166, 1085, 2, 1113, 716, 146, 1314, 1016, 1389, 541, 61, 1228, 332, 534, 1346, 194, 692, 1065, 921]. class [517, 202, 404, 138]. Claudin [686]. cleave [896]. clinical [1364]. clofibrate [729]. clofibric [516]. clone [1196]. Cloning [1186, 17, 100, 677, 643, 450, 284, 1013, 418, 515, 1356, 169, 18, 511, 679, 939, 135, 468, 547, 463, 591, 125, 885, 913, 1280, 368, 748, 738, 128, 501, 1207, 23, 380, 1374, 80, 963, 46, 674, 290, 336, 430, 747, 377, 1271, 153, 986, 404, 688, 420, 743, 976, 660, 266, 497, 1056, 1046, 276, 461, 655, 900, 1181, 923, 472, 138, 686, 713, 815, 525, 1377, 631, 24]. cloud [18]. cloudy [425]. Clove [1138, 184, 260, 354]. clown [33]. clownfish [430]. Clupea [1369, 1006]. cluster [1045]. cMyc [1094]. co [619, 672, 1364]. co-exposure [672]. co-transporter [619]. coast [141, 75]. coastal [800, 782]. cobia [432, 1323]. Cobitis [53]. cocaine [1151, 1315]. cocaine- [1151, 1315]. cod [1364, 115, 188, 493, 1024, 517, 103, 154, 252, 87, 841, 554, 133, 1021, 705, 979, 71]. coenzyme [862]. cognate [330, 511]. cognate-mediated [330]. Coho [1387, 1127, 562, 1328]. coioides [6, 444, 377, 293, 410, 986, 807, 400, 590, 16, 679]. cold [1159, 1292, 618]. coli [568]. colipase [346]. collagen [1381]. Color [1380, 1260, 412, 1089, 1260, 140]. Color-induced [1380]. coloration [1042]. colored [140]. Colossoma [927, 1336]. colostrum [662]. Columbia [1387]. column [785, 206]. combats [223]. combination [64, 168]. combinations [922, 96]. combined [55, 633, 1319]. Commercial [728, 1283, 241]. commercially [673]. Common [256, 55, 1168, 636, 1229, 382, 1064, 851, 351, 1260, 86, 612, 105, 343, 522, 140, 296, 1137, 1241, 164, 458, 539, 1195, 675, 116, 852, 1204, 149, 749, 551, 868, 1129, 181, 1202, 965, 685, 321, 1108, 1307, 151]. commonly [973]. compact [1121]. company [239]. Comparative [1260, 1395, 549, 571, 1330, 765, 1129, 329, 1198, 41, 820, 210, 1368, 28, 1089, 369, 192, 929, 1066, 1125, 130, 1248, 519]. Comparing [1074]. Comparison [72, 454, 47, 721, 141, 189, 800, 918, 1290, 1188, 1021, 946]. compensatory [1246]. complex [166, 202, 404, 138]. complexed [1336]. complexes [672]. component [465]. components [853, 1001, 1021]. composition [113, 514, 951, 422, 925, 1117, 1160, 443, 37, 708, 322, 517, 531, 628, 1306, 250, 130, 283, 736, 1216, 84, 1037, 292, 895, 1156, 1253, 249, 574, 1170, 247, 1348, 1091, 436, 1050, 110, 760]. Compositions [22, 796, 98, 533]. compound [1161]. concentrate [1208, 1294, 776, 824, 895]. concentrating [900, 1307]. concentration [215, 342, 343, 460, 898, 572, 357]. concentrations [435, 1229, 405, 176, 598, 146, 262, 1228, 174, 85, 362, 1209, 1349, 485]. Concerto [719]. condensate [995]. condition [564, 281, 487]. conditions [1289, 1364, 996, 903, 10, 186, 1044, 1363, 1317, 1200, 719, 635]. confinement [280, 1342]. Confirmation [538, 896]. Conjugated [708]. conjugating [855]. conjugation [494]. connective [1356, 979]. connexin43 [94]. consciousness [225]. consensus [237]. conservation [131, 1233, 1232]. consider [132]. consistent [945]. conspecific [993]. constant [1142, 544, 895, 434, 902, 946]. constituents [529]. constraints [884]. consumption [512, 119, 133, 1336, 9, 1380, 124]. containing [835, 298, 909, 84, 728, 321]. Contaminant [889]. Contaminant-induced [889]. contaminated [1199, 76, 1060, 1249]. content [308, 613, 53, 146, 1210, 217, 898, 947]. contents [285, 790, 933, 1027]. Continuous [730, 982]. contractile [1121]. contractility [635, 553]. contraction [351, 1138]. contributions [1015]. control [792, 611, 567, 666, 28]. conversion [479]. convict [738]. Coordinated [1004]. Coorong [956]. copepodids [850]. Coping [1383, 945]. copper [211, 928, 166, 1227, 34, 908, 102, 192, 11, 343, 1182, 458, 539, 426, 1125, 78, 284, 689, 533, 552, 1059, 589]. copper-induced [1227, 1059]. core [905]. Coregonus [734, 794, 42]. coriiceps [797]. Coriolus [408]. corn [173]. corncobs [827]. corneal [948]. cornstarch [874]. Correction [1397, 1245, 1396]. correlated [85]. correlation [930, 1266]. correlations [1042]. corticotrophin [1166]. corticotrophin-releasing [1166]. corticotropin [694, 655]. corticotropin-releasing [694, 655]. Cortisol [234, 361, 1384, 199, 616, 176, 828, 170, 262, 695, 496, 817, 174, 434, 66, 335, 819, 830]. Corydoras [1218]. costs [314, 31]. cosupplementation [1306]. cotransporter [833, 996, 277]. cottonseed [690, 313]. Cottus [251, 441]. course [210]. cownose [342]. cox1 [304]. cox2a [304]. cox2b [304]. CpG [115, 1266, 864, 770]. CpG-rich [864]. Cre [524]. creatine [1249]. CREB [756]. CRH [655]. cricket [921]. Critical [94, 946, 1143]. critically [1158]. croaker [368, 1309, 1094, 221, 861, 499, 420, 875, 201, 366, 49, 1368, 941, 1053, 990, 1029]. crocea [1309, 1094, 861, 1222, 499, 420, 1368, 941, 1053, 1087, 990, 1029]. crowding [734, 45, 606, 688]. crucian [1392, 119, 857, 14, 36, 370, 408, 28, 618, 1089, 631, 294]. crude [298, 891]. Cryopreservation [1158, 1230, 735, 1236]. cryopreserved [1324]. cryoprotectant [1324]. Ctenopharygodon [1133]. Ctenopharyngodon [542, 599, 1391, 380, 630, 729, 447, 811, 917, 1093, 1329, 1076, 594, 1002, 57, 659, 771, 374, 781, 1393, 1365, 1196, 1020, 1046, 1033, 1295, 461, 939, 897, 358, 997, 1097, 1058, 686, 1371, 1303, 483, 547, 68, 955, 1367, 1122, 313, 95, 24, 1269, 589]. ctgf [1356]. Cu [166, 336, 1172]. cuchia [984]. Cucurbita [824]. cues [993]. culter [987, 1266]. culture [617, 375, 281, 158]. cultured [973, 63, 1363, 295, 691, 538, 110, 998]. cultures [793, 1105]. curcas [116, 1129]. Curcuma [1032]. curcumin [882, 1306, 1068, 244, 504]. current [237]. curve [120]. cut [1075]. Cuvier [678, 991]. CXC [52]. CXC-64 [52]. CXCR3 [885]. cyanea [707]. cyanobacteria [503]. cyanobacterial [675, 321]. cyanurate [672]. cyanuric [672]. cycle [439, 730, 1142, 352, 1120, 76, 290, 50, 1047, 217, 418, 515, 297, 536, 477, 693, 1307, 553]. cycles [434, 940]. cyclic [1110]. cyclicity [485]. cyclodextrin [1324, 1336]. cyclooxygenase [1033]. cyclooxygenase-mediated [1033]. cyclooxygenases [304]. Cydonia [1065]. cyhalothrin [1273]. Cylindrospermopsin [1077]. Cymbopogon [1215]. Cynoglossus [642, 100, 450, 1104, 1013, 864, 1035, 755, 118, 1082, 1181, 138, 770, 363].

CYP [68]. cyp19a1a [905, 894]. CYP1A [1125]. CYP450 [562]. cypermethrin [192, 898, 1108, 357]. cyprinid [1305, 1158, 407]. cyprinidae [510, 630, 1334, 284, 1162]. cyprinids [853, 1001, 1100]. Cypriniformes [630]. Cyprinodon [1285, 881]. Cyprinodontiformes [837]. Cyprinus [1168, 636, 119, 1229, 382, 851, 1176, 1260, 440, 209, 102, 86, 510, 920, 523, 612, 105, 343, 522, 752, 140, 296, 1148, 1137, 1241, 164, 458, 539, 1195, 675, 116, 327, 804, 852, 1248, 1204, 149, 551, 1164, 868, 569, 55, 1343, 1202, 685, 321, 1351, 680, 1307, 1348, 151, 1261, 714]. cysteamine [987, 444]. cysteine [1058]. Cytoarchitectonic [222]. cytochemistry [44]. Cytochrome [317, 1196, 852, 454, 378, 497, 220]. cytogenotoxicity [727]. cytoglobin [1311]. cytokeratin [587]. cytokine [668, 701, 752]. cytokine-inducing [668, 701]. cytometric [248]. cytoprotection [1157].

D [728, 586, 996, 1071, 1048]. Dabry [826]. dabryanus [646, 1118, 135, 1321, 826, 1027]. dace [1290]. Daily [452, 186, 1066]. dam [369, 367]. dam-impeded [367]. damage [1325, 599, 1358, 658, 140, 296, 1076, 595, 1075, 61, 1065, 1277, 16, 888, 151, 750, 1059, 990, 294]. damaged [318]. dancena [825]. Danio [758, 636, 52, 687, 725, 120, 190, 356, 318, 306, 1212, 1194, 898, 968, 185, 1030, 266, 763, 954, 966, 1235, 1011]. dark [696, 434]. dark-adapted [696]. darkbarbel [708]. Dasyatis [621, 342]. data [257, 991]. daubed [341]. day [1180, 1140]. day/night [1140]. days [728]. de-adhesion [1237]. de-oiled [1198]. dead [991, 746]. deaminase [810, 81]. Dec2 [67]. decoction [714]. decrease [1392, 267, 1131, 1144, 1361, 359]. decreased [1133, 1367]. decreases [532]. Defatted [1327]. defence [595, 886]. defense [443, 258, 6, 968, 69]. defenses [555, 644, 1321]. deficiency [780, 324]. Deficient [849, 1183]. definitive [94]. deformities [228, 40]. degenerated [374]. degradation [1073]. dehydrogenase [447, 1278, 871, 870]. deiodinase [670]. deiodinases [1004]. delays [252]. delineatus [248]. delivery [1225]. Delta [981, 847]. deltamethrin [598]. Demand [230]. demethylation [68]. demonstrates [1222]. Demonstration [831, 558]. denatured [534]. denervation [927]. density [1242, 489, 741, 412, 262, 1088, 1090, 1112, 629, 678, 1082]. dental [196]. dentatus [894, 623]. Dentex [320]. deoxygenation [1392]. dependence [78]. dependent [833, 996, 1221, 1362, 62, 454, 1257, 378]. Deposition [518, 708, 588, 917, 918, 861, 1112, 1361, 955, 436]. deprivation [1267, 1160, 177, 706, 1145, 1211]. derivative [1161]. derivatives [871, 870]. derived [1055, 1255, 87, 825, 1018, 425]. desaturase [918]. desaturases [1179]. Description [126]. detailed [1119]. detect [772]. detecting [650, 830]. detection [1194, 645]. Determination [4, 248, 419, 25, 563, 390]. determined [1373]. determining [110]. Detoxification [882, 123]. detoxified [116]. developing [785, 48, 1030]. Development [3, 788, 375, 828, 1234, 131, 524, 150, 39, 610, 799, 1043, 1136, 1105, 572, 122, 830, 311, 663, 664, 462, 707, 399, 960, 314, 1208, 795, 100, 493, 331, 1120, 1154, 791, 1374, 1309, 488, 120, 670, 681, 556, 1132, 1318, 1266, 549, 535, 1330, 129, 410, 771, 1204, 694, 632, 1016, 626, 563, 604, 972, 469, 683, 601, 698, 881, 875, 169, 112, 789, 344, 487, 162, 559, 779, 1102, 1276, 552, 713, 1019, 826]. Developmental [1010, 829, 67, 380, 103, 1125, 61, 964, 281, 720]. devil [707]. devoid [892]. dexamethasone [933]. dextrin [89, 767]. di-2-ethylhexyl [829]. diacanthus [1381]. diagnosing [126]. diazinon [435, 598]. Dicentrarchus [827, 962, 846, 1187, 1071, 473, 616, 471, 872, 526, 1226, 445, 648, 350, 38, 1141, 720, 168, 452]. dicer [798]. diclofenac [516]. Diel [1110, 616]. diencephalon [1180]. diesel [373]. Diet [198, 1114, 951, 1199, 1325, 925, 371, 489, 608, 1032, 613, 523, 263, 822, 579, 675, 50, 217, 938, 601, 449, 1198, 1343, 1249, 1116, 892, 1368, 995, 1348, 1268, 1130, 110, 508]. diet-fed [1130]. Dietary [532, 325, 1085, 950, 231, 1264, 805, 343, 1107, 1133, 1306, 980, 116, 307, 428, 1301, 564, 1068, 465, 1343, 654, 702, 291, 997, 483, 436, 1370, 1296, 1390, 1168, 1114, 1227, 1063, 796, 758, 1300, 443, 1134, 37, 502, 858, 270, 690, 1074, 605, 1391, 709, 440, 588, 929, 176, 711, 637, 77, 1221, 609, 963, 646, 1084, 322, 1293, 105, 752, 849, 6, 1012, 1038, 74, 1278, 549, 891, 1334, 223, 647, 424, 283, 265, 98, 264, 594, 595, 736, 1002, 1150, 57, 657, 1246, 848, 244, 504, 292, 1213, 199, 496, 137, 551, 895, 148, 54, 915]. dietary [767, 310, 569, 780, 1190, 1378, 460, 1381, 671, 968, 1202, 947, 350, 324, 1025, 1051, 834, 1281, 1215, 255, 921, 1046, 487, 1080, 1206, 1109, 1172, 943, 955, 1050, 1331, 760, 1007, 171]. Diethylstilbestrol [1178]. diets [278, 543, 422, 119, 1208, 1135, 1299, 992, 690, 89, 473, 874, 1098, 1203, 729, 673, 250, 518, 264, 1086, 84, 983, 332, 856, 721, 965, 1196, 715, 1141, 1327, 1170, 247, 997, 1163, 1183, 1091, 1251]. differ [340, 452]. Differences [342, 744, 73, 614, 684, 689]. Different [1011, 990, 278, 455, 543, 462, 1355, 796, 1364, 119, 1135, 1160, 37, 1067, 501, 1382, 1074, 111, 464, 186, 537, 492, 1224, 911, 1306, 250, 146, 518, 736, 1002, 1184, 1037, 1388, 1363, 777, 217, 817, 924, 994, 768, 867, 1100, 329, 174, 635, 801, 590, 1147, 1211, 289, 498, 362, 964, 1196, 809, 946, 1102, 803, 815, 1269, 553]. Differential [903, 1039, 649, 126, 634, 330, 1356, 391, 713]. differentially [242, 684, 698, 1303]. differentiated [396]. differentiation [894, 196, 486, 488, 630, 1015, 556, 1148, 959, 25, 848, 1225, 813, 1217, 998]. difloxacin [68]. DIGE [1025]. digest [528]. digestibility [270, 1107, 1294, 971]. digestion [1283, 999, 1101, 356, 988, 107]. Digestive [278, 791, 256, 773, 931, 281, 1168, 3, 925, 651, 846, 1299, 596, 1291, 709, 788, 1010, 622, 609, 872, 673, 243, 322, 1293, 537, 492, 1144, 1298, 104, 78, 808, 264, 1248, 657, 1037, 150, 632, 117, 295, 604, 610, 935, 469, 683, 868, 640, 1200, 601, 1022, 360, 1147, 1378, 1343, 875, 964, 346, 940, 475, 745, 1302, 723, 193, 809, 1170, 603, 247, 1109, 887, 1072, 505, 347, 826, 1370, 1007]. Dim [911]. Dim-light [911]. dimensional [565, 1303]. dimerus [706, 769]. dimethoate [218]. Dimorphic [1102, 92, 413, 1178, 641, 743, 1042, 923, 1217]. diodes [1011]. dipeptide [478]. dipeptidyl [896]. diphenyl [551, 938, 954]. Diplodus [858, 275, 406, 99]. diploid [493, 1118, 754]. direct [488]. direction [920]. discussion [225]. disease [944, 759, 801, 1296]. diseases [126, 303]. diselenide [551, 938]. disjunctivus [403]. dismutase [700].

Dispersant [415]. dispersion [545]. displacement [577]. displaying [206]. Disrupting [27, 1240, 276]. disruption [423, 75, 1194, 1065, 954]. disruptive [80]. dissolved [272, 289, 1122]. Dissostichus [807]. Distinct [26, 1120, 185, 680, 1397]. Distribution [645, 587, 23, 180, 271, 1278, 1271, 774, 560, 662, 1090, 1351, 900, 832, 686, 505, 30]. diurnal [452]. Dmrt1 [641, 399, 1266]. dmrt1s [631]. DMSO [717]. DNA [1358, 140, 53, 318, 1104, 504, 1070, 61, 1030, 864, 16, 829, 135, 436, 750, 770, 349]. Dnah9 [843]. dnd [746]. Do [161, 654]. docosahexaenoic [637, 1133, 1329]. Does [1142, 920, 232, 1187, 724, 1285, 695]. Domain [531, 478]. domain-2 [478]. domains [478]. dominance [536, 477]. donors [337, 429]. dopamine [1352, 1131, 790, 427]. dopamine-related [1352]. dorsal [240, 1012]. dose [338, 838, 765, 990, 1029]. doses [11, 784]. dots [305]. dovii [640, 971]. downregulation [108]. downstream [369, 782, 157]. dried [1170]. drum [462, 128, 1221, 1123, 757, 816, 1163]. dual [407]. ducts [682]. due [629]. dummy [665, 1057]. Dunaliella [543]. duodenal [134]. duplicated [1028, 1333]. Duplication [1169]. duration [616, 838]. during [439, 907, 949, 734, 905, 399, 1267, 35, 1242, 960, 651, 1342, 100, 877, 903, 1120, 588, 788, 486, 1154, 1010, 1374, 1309, 120, 630, 484, 670, 681, 556, 1266, 1009, 256, 290, 782, 535, 430, 108, 1330, 1316, 1047, 410, 694, 604, 1310, 367, 61, 924, 418, 515, 1308, 982, 392, 297, 1136, 640, 157, 566, 698, 1042, 1121, 1021, 498, 881, 1245, 1244, 775, 875, 964, 169, 112, 1119, 789, 475, 1302, 644, 723, 281, 880, 193, 813, 919, 1276, 573, 506, 1236, 713, 1307, 468, 302, 1004, 213, 525, 826, 1027, 1087]. dwarf [222]. dye [1167]. Dynamic [905, 694, 705]. Dynamics [471, 1110, 332]. dysfunction [814, 1029].

E-free [518]. E. [1210]. Early [910, 907, 949, 663, 664, 423, 462, 912, 1291, 722, 788, 486, 1154, 1010, 556, 256, 812, 241, 1125, 129, 410, 1204, 694, 61, 400, 881, 789, 475, 1302, 49, 281, 506, 468, 1217]. early-stage [812]. east [141]. eastern [491]. Echinacea [310]. Echium [918]. eckloni [1072]. Ecological [398, 329]. ecophysiological [617]. ecosystem [985]. ecotoxicity [884]. ectonucleotidase [810]. ectothermic [773]. EDS [973]. EDS-based [973]. eel [73, 8, 821, 353, 488, 984, 681, 904, 387, 273, 959, 634, 469, 1165, 182, 22, 171]. eels [1152, 287, 31, 157]. Effect [211, 1390, 827, 1114, 912, 951, 580, 443, 742, 1134, 37, 599, 219, 240, 1291, 605, 406, 312, 929, 711, 521, 77, 142, 1049, 989, 322, 517, 105, 752, 917, 1132, 354, 822, 74, 1093, 579, 1357, 1127, 1195, 1101, 456, 98, 262, 594, 1150, 57, 735, 776, 437, 1375, 268, 292, 938, 915, 316, 856, 1147, 527, 898, 216, 38, 408, 603, 1058, 286, 1321, 124, 347, 313, 760, 51, 345, 974, 1168, 166, 1023, 1061, 1175, 258, 480, 380, 510, 80, 246, 523, 1113, 484, 673, 373, 522, 583, 1329, 716, 1334, 585, 356, 647, 257, 578, 264]. effect [978, 1353, 284, 1075, 149, 117, 244, 504, 148, 582, 601, 703, 495, 873, 1156, 635, 1051, 638, 864, 1281, 866, 921, 1320, 880, 971, 678, 407, 247, 220, 607, 1078, 508]. effective [995, 239]. effectiveness [566]. Effects [272, 1173, 1063, 1219, 707, 212, 1279, 45, 1352, 796, 636, 1180, 999, 846, 1160, 1197, 1139, 851, 690, 1052, 944, 351, 1176, 440, 687, 207, 806, 729, 609, 376, 646, 1084, 470, 1024, 152, 6, 1149, 1012, 259, 263, 793, 936, 627, 277, 891, 937, 241, 424, 562, 444, 736, 853, 1001, 1210, 1216, 254, 299, 306, 657, 1037, 1246, 781, 665, 807, 1213, 217, 496, 61, 137, 817, 1328, 1088, 868, 767, 332, 310, 569, 1022, 449, 427, 1190, 1005, 1378, 1112, 1121, 1381, 1202, 442, 350, 965, 253, 321, 783, 249, 533, 699, 1040, 1277, 1215, 255]. Effects [763, 705, 487, 809, 1116, 574, 901, 897, 888, 1048, 358, 1163, 68, 1143, 1081, 1261, 618, 941, 718, 1050, 1007, 173, 171, 1251, 1227, 653, 663, 664, 280, 422, 1191, 1300, 503, 8, 894, 371, 502, 177, 489, 621, 973, 1074, 882, 625, 473, 570, 1360, 1391, 616, 192, 361, 725, 896, 884, 872, 668, 701, 447, 296, 549, 847, 205, 116, 762, 933, 78, 871, 870, 264, 595, 1248, 986, 1265, 731, 848, 1204, 1226, 551, 895, 415, 935, 1161, 1228, 691, 1100, 1263, 174, 465, 31, 926, 902, 1090, 633, 593, 671, 1057, 1343, 261, 810, 969]. effects [1077, 1119, 1056, 597, 751, 1141, 16, 335, 726, 1273, 717, 476, 276, 1080, 1181, 1319, 1252, 158, 552, 472, 151, 1091, 1011, 1155, 1151, 339, 479, 1059, 649, 990, 1029, 714, 837]. Efficacy [260, 1188]. efficiency [109, 833, 588, 116, 107, 436, 479, 676]. efficient [1096]. effluent [1085, 973]. EGCG [69]. Egfl6 [779]. EGFP [1000, 1000]. Egg [54, 331, 812, 880, 1235]. eggs [620, 48, 531, 628, 571, 47, 1237]. egr [145]. egr-1 [145]. Egyptian [298]. eicosanoids [304]. eicosapentaenoic [199]. elaboration [1305]. elasmobranch [909]. elasmobranchs [216]. Electric [1230]. electrical [554, 383, 309]. electrolytes [338, 178]. electromagnetic [1137, 759]. electron [546]. electrophile [34]. electrophile-responsive [34]. electrophoresis [565, 1303]. Electrophoretic [3, 102]. electroretinogram [696, 1336]. Eleginops [678, 809]. element [34, 533]. elements [10]. Elevated [91, 251, 424, 85, 997]. elicit [1128]. elicits [160]. eliminates [927]. ELISA [203]. ELISAs [1016]. Embiotocidae [697]. embryo [1023]. embryogenesis [1009, 919, 506, 302]. Embryonic [314, 103, 825, 1317, 1276, 552, 713]. embryos [914, 722, 741, 556, 812, 318, 869, 1194]. EMG [156]. emitting [1011]. emphasis [587, 162]. encode [722]. encoded [698]. encoding [922, 416, 81]. end [746]. endangered [1158, 1234, 1324, 311]. endemic [1162]. Endocrine [1240, 434, 1090, 1065, 954, 423, 280, 1387, 75, 265, 924, 763, 276, 887]. endocrine- [763]. Endocrine-disrupting [1240, 276]. endogenous [144, 721, 880]. endoplasmic [1331]. endosulfan [307, 337, 428, 429, 60, 261]. endosulfan-induced [307, 428]. endothelial [831]. endothermic [351]. endothermy [773]. energetic [1199, 31]. Energy [209, 113, 729, 1384, 1272, 1372, 1310, 12, 898, 1046, 941]. Enhanced [1025, 419]. Enhancement [1214, 460, 423]. enhancer [732]. enhances [906]. enhancing [647]. enkephalin [1079]. enriched [56, 856]. Enrichment [371]. enrichments [815]. enrofloxacin [1393]. enterocytes [1288, 658, 793, 1059]. Enteroendocrine [526]. environment [1240, 550, 1123, 181, 65]. environmental [930, 314, 984, 277, 924, 1166, 194, 106, 1254]. Environmentally [1209, 1228]. environments [1060, 768]. Enzymatic [624, 278, 1038, 1323, 623, 899, 107, 1072]. Enzyme [1346, 203, 272, 3, 136, 651, 855, 1382, 561, 1291, 709, 788, 805, 1010, 609, 872, 1293, 343, 256, 204, 337, 429, 871, 870, 1037, 563, 604, 610, 868, 773, 1006, 1156, 44, 1378, 931, 715, 475, 921, 1302, 763, 281, 193, 979, 603, 247, 1109, 1183, 826, 1370, 1007]. Enzyme-linked [203, 563]. enzyme-related [763]. enzyme-treated [1109]. enzymes [1168, 568, 34, 846, 1299, 1067, 338, 176, 1369, 859, 243, 322, 537, 492, 50, 871, 870, 562, 594, 1184, 657, 988, 179, 150, 781, 1068, 117, 61, 935, 662, 469, 683, 1200, 1257, 1022, 1147, 1198, 1042, 1343, 4, 940, 1167, 1351, 723, 892, 313]. enzymological [516]. ep [643]. EPA [1041]. epaulette [167]. EpCAM [865]. epidermal [514]. epidermis [1350]. Epigallocatechin [69]. Epigallocatechin-3-gallate [69]. Epinephelus [960, 1135, 417, 23, 6, 323, 444, 1210, 377, 293, 410, 986, 988, 1124, 807, 400, 1338, 500, 590, 16, 335, 191, 679]. epinephrine [1121]. epithelia [590, 197]. epithelial [1211, 831]. epithelium [546]. epoxide [761]. EpRE [34]. equilibrium [165]. equol [549]. erectus [1047]. ergic [1164]. Ergosan [322]. erinacea [485]. Eriochrome [1167]. ERK [316]. Erratum [384, 949, 1049, 509, 701, 537, 428, 429, 871, 1001, 299, 536, 383]. Eryngii [944, 944]. Erythroculter [987]. Erythrocyte [294, 53]. erythrocytes [804, 504, 378, 1220, 200, 750]. ES-like [375]. escape [902]. Escherichia [568]. Esox [1142, 1237, 880]. essential [288, 835, 566, 1119, 1215, 1258, 597, 712, 559, 1183]. Establishment [1332, 453, 854, 1055, 1255, 754, 1124, 755, 118, 652, 363, 513, 1096, 509]. estimated [325]. estimation [54]. estor [312]. Estradiol [242, 987, 570, 611, 391, 18, 485]. estradiol-17 [570]. Estrogen [722, 26, 1397, 450, 862, 400, 704, 607]. estrogenic [465]. estrogens [704, 59, 699]. estuarine [308, 548, 956, 1362]. estuarine-dependent [1362]. Estuary [956, 438]. ethanolic [219]. ether [954]. ethoxyquin [1248]. ethylhexyl [829]. ethynylestradiol [214]. etidronate [259]. etomidate [1189]. Etroplus [619]. eugenol [288, 1188]. European [962, 1186, 8, 846, 1187, 1071, 251, 473, 616, 248, 488, 273, 257, 526, 1226, 31, 157, 648, 1189, 1141, 1327, 720, 441, 452]. Euryglossa [976]. euryhaline [1166, 197]. euthanasia [1359]. evacuation [1127]. evaluating [833]. Evaluation [1364, 673, 551, 840, 1253, 757, 975, 650, 928, 501, 625, 824, 5, 1198, 878, 939, 1083, 1348, 508]. everted [421]. Evidence [251, 1317, 961, 446, 488, 1238, 1148, 1151]. Evidences [1019]. Evoked [554, 1373]. evolution [1243, 1162]. evolutionary [842, 397, 1377]. ExcelMOS(R) [1214]. excess [849, 124]. exchange [868]. excision [1030]. excitatory [1131]. excretion [119]. excretions [464]. exercise [161, 246, 155, 1005, 38, 124, 1078, 339]. exercise-induced [161]. exhaustive [1005, 38, 124]. Existence [395]. exocrine [99]. exogenous [1198, 528, 247, 1270]. exon [864]. exons [1020]. experiment [651]. Experimental [691, 257, 133, 479]. experimentally [568]. experiments [120, 884]. Exploration [666]. exposed [211, 166, 1227, 1085, 340, 1325, 435, 136, 314, 795, 93, 405, 176, 11, 165, 464, 922, 784, 523, 598, 1358, 343, 936, 154, 1334, 1123, 426, 223, 337, 429, 146, 802, 179, 1223, 728, 938, 924, 434, 289, 516, 60, 38, 1167, 975, 712, 357, 803, 168, 65, 750]. Exposure [867, 720, 663, 664, 167, 913, 908, 742, 338, 1247, 761, 209, 102, 361, 609, 184, 917, 1182, 298, 140, 1285, 319, 336, 910, 852, 284, 839, 1193, 1115, 199, 1282, 560, 1228, 592, 672, 1335, 391, 91, 1121, 898, 442, 185, 516, 448, 123, 1108, 533, 1344, 1209, 1056, 487, 481, 1273, 441, 650, 829, 966, 1027, 649, 1029, 589, 414, 1383]. exposures [627, 254, 299, 727, 96]. expressed [501, 537, 492, 318, 684, 698, 1303]. expresses [1000]. Expression [399, 821, 914, 1316, 626, 1259, 112, 191, 29, 269, 1284, 734, 92, 1159, 700, 1274, 245, 422, 602, 796, 1017, 962, 960, 925, 1180, 1117, 794, 1186, 894, 1311, 17, 100, 67, 382, 677, 833, 913, 996, 1280, 368, 742, 738, 175, 1067, 614, 128, 1139, 501, 599, 1207, 851, 15, 706, 987, 1176, 746, 882, 1247, 1391, 380, 432, 1297, 761, 722, 953, 1154, 459, 328, 613, 183, 1374, 1309, 1366, 1039, 80, 1221, 14, 142, 963, 103, 670, 752, 849, 26, 287, 556, 1149, 1397, 918, 674, 1318, 1266, 793, 1094, 27, 290, 99, 277, 716, 336, 430, 1123, 356]. expression [1044, 424, 643, 1330, 562, 202, 450, 747, 959, 1104, 978, 377, 1184, 1386, 153, 490, 410, 802, 839, 869, 986, 1193, 781, 774, 694, 404, 641, 777, 1013, 1354, 744, 1243, 496, 634, 935, 860, 733, 418, 515, 582, 688, 767, 416, 297, 304, 330, 1356, 206, 873, 1022, 856, 661, 420, 1147, 840, 1162, 145, 893, 391, 743, 1112, 977, 633, 536, 477, 1042, 1381, 1166, 976, 324, 448, 1218, 1051, 169, 1196, 864, 886, 1040, 1258, 789, 201, 266, 497, 757, 816, 1209, 366, 680, 1320, 1056, 1225, 702, 763, 1046, 1033, 121, 481, 127, 476, 18, 511, 461, 679, 655, 813, 939, 888]. expression [919, 1102, 1179, 1181, 1276, 208, 923, 829, 966, 832, 1153, 1169, 1206, 506, 472, 138, 713, 737, 845, 1371, 1307, 468, 1217, 815, 818, 302, 315, 1172, 1004, 463, 639, 1261, 591, 958, 30, 525, 1377, 770, 836, 1370, 313, 631, 173, 24, 1251, 1034, 1026]. Expressions [1347, 608, 736, 1210, 306, 1378, 718, 1050]. external [1324]. extra [153]. extra-pituitary [153]. Extracellular [693, 215, 761, 316, 1021]. extract [365, 219, 1176, 1250, 787, 1365, 1161, 1326, 178, 321, 866, 151]. extracted [298, 19]. extraction [828, 66]. extracts [1126, 891, 1248]. extrahepatic [649]. extreme [211]. extremely [1137, 759]. eye [254, 299].

f [862, 1347]. F-box25 [1347]. F. [72]. F1 [820]. faba [1295, 1295]. Fabp10 [842]. facilitative [1184]. factor [382, 677, 913, 1207, 548, 1297, 103, 556, 1318, 782, 1387, 1044, 295, 634, 860, 1292, 1356, 303, 1019, 1371, 1348, 525]. factor-1 [1044, 634]. factor-binding [913]. factor-inhibiting [1297]. factor-mRNA [1292]. factors [23, 132, 1047, 935, 1030, 1276, 713]. Fads2 [918]. faecalis [1296]. fails [1128]. false [33]. families [614, 1165]. Family [1362, 393, 355]. fario [413, 862]. farmed [734, 230, 237, 160, 517, 549, 224, 233, 1246, 226, 1021, 229, 41, 162, 979, 236, 239]. farming [629, 931]. FAS [1149]. fasciatus [243, 1379, 1272, 1259, 1319]. fast [161, 1337, 1310, 902]. fast-start [902]. fasting [999, 651, 380, 1149, 716, 356, 582, 1090, 1040, 832, 1307, 339]. Fat [955, 1114, 613, 983, 1361, 1116, 1130, 1251, 1046]. FAT/CD36 [1046]. fate [101]. fathead [731, 829]. fatigue [1081]. Fatty [473, 70, 348, 1063, 796, 842, 141, 708, 1074, 312, 637, 517, 1149, 918, 736, 84, 1037, 374, 777, 862, 744, 217, 1310, 82, 780, 460, 442, 41, 249, 702, 1141, 1179, 1050, 171, 974]. Fe [279]. fear [225]. fecundity [519]. fed [1114, 278, 543, 1355, 1017, 503, 119, 1208, 1135, 443, 992, 608, 690, 709, 1031, 581, 1098, 729, 1084, 139, 1038, 56, 571, 50, 518, 84, 983, 749, 662, 721, 1346, 965, 482, 715, 1249, 1033, 1295, 1170, 1368, 997, 1163, 1183, 815, 1130, 1251, 1269, 1383]. Feed [177, 1219, 1300, 1160, 625, 787, 1145, 1386, 657, 807, 268, 1005, 527, 1343, 528, 1327, 1296]. Feeding [667, 887, 211, 1283, 230, 999, 542, 37, 111, 1224, 232, 1100, 434, 640, 1200, 926, 1090, 590, 898, 528, 880, 892, 476, 655, 347, 1254, 173]. feeds [1061, 1248]. feedstuffs [971]. female [126, 119, 189, 1084, 1395, 1158, 1132, 549, 933, 98, 1210, 623, 744, 817, 297, 1259, 498, 692, 1245, 1244, 1099, 344, 485, 135, 845, 1370, 171]. female-specific [135]. females [907, 949, 730, 70, 1118, 571, 341]. feminization [25]. fenitrothion [605]. fermented [1038, 1156, 1198]. Fernandez [899]. Ferritin [778]. ferrugineus [326]. Fertility [1118]. fertilizing [1304]. ferula [1343, 1343]. FGF6 [1371]. FGF9 [344]. FGF9/16/20 [344]. fiber [1171]. fibroblast [1332, 831, 1371]. fibroblast-like [1332]. field [681, 1157, 182, 171]. fields [1137, 759]. fighting [93, 268, 1201]. FIH [1297]. FIH-1 [1297]. fillet [532, 744]. fimbria [1092, 848]. fin [413, 1255, 754, 1316, 893, 513]. fin-derived [1255]. final [901]. fine [459]. finfish [234]. fingerling [325, 1137, 57, 1378, 1051, 436]. fingerlings [827, 1031, 207, 223, 307, 337, 428, 429, 647, 264, 736, 148, 915, 868, 310, 247]. finishing [588]. fins [972]. fipronil [523, 938]. First [1238, 1283, 783]. fischeri [568]. Fish [1322, 155, 227, 597, 995, 239, 109, 348, 532, 166, 1305, 1304, 1219, 230, 494, 136, 422, 237, 215, 577, 1286, 161, 1208, 1240, 945, 846, 600, 883, 388, 352, 1311, 443, 1256, 914, 93, 1, 231, 967, 225, 617, 706, 1052, 393, 172, 1003, 903, 235, 473, 1247, 1360, 451, 581, 1138, 1231, 446, 10, 521, 466, 1369, 828, 376, 334, 1084, 1293, 952, 787, 419, 517, 668, 701, 658, 1337, 263, 354, 793, 627, 277, 1290, 535, 204, 1111, 571, 224, 1171, 1362, 584, 426, 50, 78, 163, 808, 871, 870, 1205, 386, 131, 395, 265]. fish [555, 595, 1248, 84, 524, 233, 1178, 179, 232, 774, 889, 1375, 149, 268, 244, 226, 560, 938, 1164, 415, 1069, 238, 697, 228, 982, 144, 332, 1136, 47, 474, 1022, 519, 1259, 1162, 145, 893, 391, 766, 529, 1201, 968, 947, 693, 261, 229, 41, 1196, 482, 43, 1077, 123, 1119, 654, 572, 1225, 81, 401, 303, 809, 726, 407, 162, 1079, 831, 954, 122, 919, 1102, 1368, 158, 830, 832, 1109, 468, 769, 1359, 68, 1091, 1261, 236, 220, 110, 1059, 363, 414, 349, 676, 1232]. fish-culture [158]. fishes [956, 948, 797]. Fishmeal [1294, 1107, 283, 776, 721, 1346, 1327, 1268]. fishmeal-based [1268]. five [512, 1100, 315, 1331]. FK506 [584]. FK506-binding [584]. flagellum [1087]. flame [487]. flanking [611]. flat [548]. flavescens [170, 1044]. flesh [162]. flesus [81]. flexuosus [1215].

Florida [491]. flounder [894, 37, 798, 459, 1009, 290, 412, 283, 75, 978, 623, 932, 869, 774, 1115, 1223, 843, 416, 1080, 1096, 1340, 923, 713, 1019, 1004, 525, 957]. Flow [248]. flows [158]. fluid [1376, 682, 877]. fluids [381]. fluoride [529]. fluoxetine [1036]. flutamide [922]. fluviatilis [130, 578, 1189]. FMOs [562]. focus [1369]. folic [1051]. follicle [1132, 844, 418, 769, 29]. follicle-stimulating [844, 418, 769, 29]. follicles [352, 695]. follicular [670, 762]. folliculogenesis [905]. following [962, 908, 1175, 706, 627, 1193, 554, 1393, 541, 499, 672, 1335, 829, 200, 124, 96]. fontinalis [656, 454, 159]. Food [792, 827, 136, 1267, 1342, 706, 1131, 356, 978, 1375, 900, 1155, 479, 961, 1315]. Foreword [13, 385]. form [569]. formaldehyde [586]. formalin [975]. formation [1275, 1089, 1087]. forms [1304, 23, 53, 1016, 1269, 1315]. formulated [268]. formulation [728, 1025, 714]. formulations [867]. Fos [751]. fossilis [113, 325, 285, 795, 762, 933, 1298, 178, 743, 1313, 866]. four [853, 1001, 1351, 772, 686, 558]. fox [639]. Foxl2 [515, 276, 813]. fraction [373]. fragmentation [504]. fragments [349]. free [177, 240, 48, 419, 154, 130, 518, 898, 1349, 830]. free-feeding [898]. free-swimming [240, 154]. freeze [561, 940, 1170]. freeze-dried [1170]. freeze-resistant [561]. freeze/thaw [940]. frequency [1137, 759, 1254]. fresh [63, 204]. fresh-water [63]. Freshwater [528, 455, 308, 577, 402, 258, 180, 1360, 521, 1290, 1123, 426, 50, 1008, 1226, 938, 764, 329, 982, 416, 710, 178, 59, 261, 41, 482, 249, 866, 726, 1273, 349]. frisii [915]. froth [22]. fructooligosaccharide [718]. fructooligosaccharides [806]. fructose [1086, 767]. fructose-1 [1086, 767]. fry [1168, 1023, 523, 1334, 268, 66, 671]. FSH [748, 611, 29, 24]. fucoidan [188, 1281]. fuel [165]. Fugu [173]. full [1125]. full-length [1125]. Fullerene [348]. fulvidraco [833, 996, 1015, 849, 336, 1178, 1378, 603, 1172, 1331, 1254, 649, 1034, 1026]. function [725, 948, 1076, 1316, 735, 731, 697, 495, 567, 878, 717, 655, 347, 1177]. Functional [932, 475, 663, 664, 1305, 1274, 602, 125, 786, 738, 798, 1013, 1313, 1095]. functionality [632, 1170]. functionally [1120]. functions [1203, 290, 1287, 629, 1281, 966, 887, 22]. Fundulus [272, 314, 72, 780]. furcatus [823]. further [1305]. furunculosis [644]. furzeri [1146]. fuscoguttatus [960, 417, 1210, 1338]. fusions [206]. future [237, 224].

G [976, 191]. G6PD [447]. GABA [1262, 1164, 645, 836]. gachua [222]. Gadus [1364, 115, 493, 1024, 517, 103, 154, 252, 87, 841, 133, 1021, 705, 979, 71]. Gaimard [824]. Galeocerdo [991]. gallate [69]. Gambusia [491, 1360, 40, 1157]. gamete [257, 1245, 1244]. gametes [1240, 1118, 1204]. Gangetic [460]. ganglion [1305, 540]. Ganoderma [1075]. GAPDH [245]. gar [278, 788, 874, 1366, 784, 622]. gariepinus [1192, 211, 568, 58, 166, 1085, 1113, 1389, 541, 61, 1228, 332, 534, 1346, 1065, 921]. garlic [882, 579]. garlic-supplemented [579]. Gasterosteus [1342, 884]. gastric [1127, 468]. Gastrointestinal [1183, 308, 356, 424, 526, 87, 1200, 1106]. gastrointestinal-derived [87]. gastroprotective [1326]. GATA [461]. GATA-3 [461]. GCRV [461]. Gdnf [1243]. gel [565, 1303]. gelatinized [50]. gender [563]. Gene [1125, 1166, 734, 92, 1159, 700, 1274, 1352, 602, 796, 1017, 925, 17, 52, 542, 833, 855, 368, 786, 1054, 738, 1067, 614, 1207, 851, 393, 1176, 882, 20, 1297, 798, 953, 459, 1039, 80, 14, 670, 681, 557, 46, 674, 1148, 1266, 27, 1171, 1123, 1314, 1044, 736, 377, 153, 284, 490, 802, 932, 404, 641, 744, 496, 634, 935, 582, 688, 767, 364, 661, 420, 1147, 840, 1378, 1313, 1381, 976, 324, 1051, 864, 1258, 201, 497, 757, 1209, 1320, 1056, 81, 702, 1046, 121, 18, 276, 813, 1169, 472, 138, 845, 815, 818, 463, 1261, 30, 213, 770, 836, 1370]. gene/protein [1046]. general [1189]. generated [827, 1020]. Generation [1000, 576]. generations [1312]. genes [34, 734, 1159, 422, 1311, 742, 175, 1139, 1339, 1391, 722, 459, 1309, 922, 1028, 1333, 1015, 1395, 752, 1009, 424, 202, 747, 1193, 1372, 694, 1354, 1243, 242, 394, 416, 684, 1022, 840, 391, 743, 698, 1112, 448, 1347, 638, 886, 1035, 1225, 763, 1033, 481, 939, 919, 1102, 829, 333, 1153, 686, 1371, 468, 1217, 315, 1235, 639, 958, 558, 1331, 1251]. Genetic [364, 864, 1233]. genetically [1246]. genistein [549]. Genome [639, 248]. Genome-wide [639]. Genomic [1386, 202, 843, 366, 138]. genomics [227]. genotoxicity [1250]. Georgia [1387]. Germ [746, 386, 960, 535, 1233, 1174, 344]. Gfra1 [1243]. Gfra1/Gdnf [1243]. GH [742, 792, 706, 152, 1370, 821, 924, 994]. GH-IGF [152]. GH-transgenic [792]. GH/IGF [742]. Gh/Igf1 [924, 994]. GHR1 [742, 770]. GHR2 [742]. ghrelin [152, 961]. giant [23, 451, 292, 500]. gibbosus [1384]. gibel [407, 559, 1268, 961, 1315]. gibelio [119, 857, 696, 1265, 1262, 977, 645, 305, 407, 408, 286, 1268, 618, 836, 961, 1315]. GIFT [805, 777, 702]. GIFT-tilapia [805]. gigas [451]. gill [621, 180, 342, 598, 904, 871, 870, 1205, 1060, 144, 1105, 969, 197, 379, 890, 927]. gills [11, 217, 1167, 666, 279, 966, 1303]. gilt [442]. gilt-head [442]. Gilthead [719, 536, 477, 1214, 1014, 620, 1283, 602, 1299, 883, 371, 270, 89, 806, 787, 130, 295, 217, 1140, 860, 924, 134, 449, 635, 1042, 448, 1025, 940, 168, 553, 1383, 186]. ginger [579]. ginger- [579]. Ginsburg [799]. glabra [151]. gland [446, 860, 881, 615]. glands [346]. Global [190]. glucan [609, 431, 941]. glucocorticoid [422, 722]. glucokinase [767]. glucomannan [760]. gluconeogenesis [50, 194]. glucosaminidase [586]. Glucose [1278, 1135, 89, 275, 406, 176, 186, 447, 1093, 871, 870, 1002, 1086, 1210, 1184, 1301, 986, 781, 767, 174, 434, 1165, 335, 1368, 1130, 452]. Glucose-6-phosphate [1278, 871, 870]. glucuronidase [114, 90]. GLUT2 [1184]. glutamate [793]. glutamine [658, 752]. glutamylcarboxylase [62]. Glutathione [859, 532, 494, 1128, 882, 102, 1038, 1306, 871, 870, 534, 1058]. gluten [173]. Glycemic [1135, 275]. Glycerol [561, 1041, 841, 766]. glycerol-3-phosphatase [561]. glycerol-3-phosphate [766]. glycine [1058]. glycinin [1367]. glycogen [146, 1210]. glycol [717]. glycolysis [50, 974]. glycolytic [272]. glycometabolism [1184]. Glycoprotein [397, 128, 697, 977, 266]. glycoprotein-like [697]. glycosaminoglycans [785]. glycosidases [163]. Glycyrrhiza [151]. glyphosate [1195, 867]. glyphosate-based [1195]. Glyptosternum [79, 147]. Gnathopogon [509, 453, 1158]. GnRH [536, 477]. GnRH2 [733]. GnRH3 [733]. GnRHa [365, 748, 331]. GnRHa-induced [331]. gobio [251]. Gobiocypris [746, 153, 802, 839, 1193, 276, 607]. goby [908]. gold [145]. golden [1284, 1386, 1388, 632, 775, 964, 815]. goldfish [1376, 1110, 677, 913, 1280, 1197, 501, 922, 896, 1182, 259, 790, 854, 1055, 936, 456, 242, 465, 756, 638, 699, 1277, 476, 1349, 750]. gonad [439, 92, 960, 285, 100, 488, 859, 681, 1266, 634, 341, 332, 755, 1217]. gonad-expression [634]. Gonadal [848, 663, 664, 399, 493, 630, 1132, 1148, 891, 1354, 626, 814, 1313, 536, 477, 1042, 487, 813, 1102, 845, 1217, 639]. Gonadogenesis [998]. gonadotrophic [451]. gonadotropic [982]. gonadotropin [1304, 1186, 748, 80, 297]. gonadotropin-releasing [1304, 80]. gonadotropins [1117, 548, 1374, 814, 692]. gonads [396, 1395, 1259, 957]. gonial [1158]. gonionotus [1253]. Goodea [179, 530]. Gordonia [834]. GPx [882]. GR [695]. gracilis [530]. graded [581]. grandis [314, 72, 780]. graphene [1344]. grass [542, 1288, 599, 1391, 380, 729, 447, 811, 917, 1093, 1133, 1329, 1041, 594, 1002, 57, 659, 771, 374, 781, 1393, 1365, 1356, 1196, 1020, 1046, 1033, 1295, 461, 939, 897, 865, 358, 997, 1097, 1058, 686, 1371, 1303, 483, 547, 463, 955, 1367, 1122, 313, 95, 24, 1269, 589]. graveolens [482]. gray [566, 1089]. gray-to-red [1089]. grayling [1234, 257]. great [373, 56, 362, 65]. green [1197, 1084, 27, 936, 1007]. green-wavelength [1197, 936]. grey [167, 548, 330, 1147]. group [591]. grouper [960, 126, 1135, 417, 23, 6, 323, 444, 1210, 377, 293, 410, 986, 988, 1124, 807, 604, 400, 500, 590, 633, 16, 335, 191, 679]. groupers [21]. groups [1067]. grow [161, 588]. grow-out [588]. growing [556]. Growth [1175, 15, 493, 23, 89, 111, 556, 1038, 1224, 1092, 84, 1388, 1363, 834, 715, 1170, 769, 1269, 1390, 1168, 1322, 827, 1114, 113, 325, 423, 462, 912, 951, 45, 161, 1208, 950, 1061, 352, 382, 443, 677, 821, 913, 1288, 1134, 37, 548, 706, 219, 690, 1052, 944, 708, 1074, 1339, 906, 570, 207, 459, 183, 711, 120, 1221, 806, 523, 872, 376, 1084, 322, 517, 1107, 1294, 752, 849, 263, 1318, 793, 27, 549, 579, 782, 1387, 412, 891, 1127, 1306, 1101, 250, 116, 337, 429, 647, 283, 262, 444, 594, 736, 978, 1216, 1248, 1353, 1386, 1150, 57, 657, 1037]. growth [1178, 1246, 564, 776, 744, 292, 1213, 195, 634, 1365, 895, 860, 1328, 915, 868, 691, 759, 1263, 1212, 640, 310, 1356, 926, 364, 1156, 1378, 633, 527, 671, 1057, 1343, 1202, 965, 324, 864, 1281, 921, 1141, 487, 809, 162, 1048, 1080, 1276, 573, 603, 247, 997, 1058, 713, 887, 1019, 1163, 1371, 1348, 483, 1155, 436, 525, 1050, 1367, 479, 770, 1370, 313, 760, 1007, 51, 171, 1251]. growth-related [459]. GRP94 [698]. grunniens [402]. grunting [402]. Gryllus [921]. GSK [861]. GSK-3 [861]. GSK3 [1394]. GSTP1 [1039]. GTH [748, 1245, 1244]. guanosine [1107]. guar [406]. gudgeon [437]. guentheri [1003]. Gulf [314, 780, 485]. gum [406]. Gunther [817, 44, 301, 716, 1099]. guppies [214]. gustducin [749]. Gut [295, 879, 1299, 564, 345]. guttatus [610, 935, 479]. guttulatus [601]. Gymnocypris [1072]. gynogenetic [1009].

H [1324, 897]. H. [1373]. habitat [1100]. Haemato [1031, 148, 325, 512, 261]. Haemato-biochemical [1031, 325, 512]. haemato-immunological [261]. Haemato-immunology [148]. Haematococcus [253]. Haematological [79, 211, 1364, 8, 338, 139, 456, 1140, 137, 921, 357, 574]. Haematology [44, 827, 133]. Haematopoiesis [164]. haemoglobins [1317]. haemolytic [133]. Haemopoiesis [32, 384]. hagfish [196, 397]. hake [1186, 842]. hakonensis [1290]. half [450, 1104, 864, 755, 118, 1181, 138, 770, 363]. half-smooth [450, 755, 1181, 138, 770, 363]. halibut [202, 254, 299, 505, 71]. Halichoeres [1180, 392]. halomethanes [530]. Ham. [222, 375, 545, 311]. Hamilton [114, 90, 546, 647, 1314, 1198, 775, 247, 436]. handling [1187, 1139, 267, 741, 442, 498, 1359]. hapuku [812]. hardness [464]. harengus [1369, 1006]. harmful [1105]. harvest [233]. harveyi [213]. hasta [908, 1293, 533, 1183]. hatched [1252]. hatcheri [503]. hatching [880, 22]. hCG [748]. HCO [1164]. hdac4 [269]. head [32, 384, 115, 188, 548, 752, 164, 458, 539, 933, 87, 1161, 442, 574, 363]. healing [822, 893]. Health [229, 1214, 711, 1294, 1111, 1353, 254, 299, 1189, 997, 1367, 1269]. hearing [1373]. heart [544, 141, 342, 1055, 584, 975, 481, 118]. heart-derived [1055]. hearts [1173]. Heat [1354, 445, 669, 1284, 353, 599, 301, 1015, 223, 747, 623, 802, 1068, 1051, 834, 511, 888, 213, 558]. heat-killed [834]. Heat-shock [669, 1284, 301, 747, 558]. heat-tolerant [353]. heavily [850]. heavy [455, 165, 336, 1171, 1205, 49]. Heckel [248]. held [1140]. Hemato [543, 1343, 1202, 1220]. hemato-biochemical [1220]. Hemato-immunological [543, 1343, 1202]. Hematocrit [96]. hematologic [345]. Hematological [218, 739, 598, 727, 928, 973, 480, 207, 373, 322, 260, 579, 1195, 776, 292, 915, 310, 321, 866, 408]. Hematology [65, 550, 376, 367, 313]. hematopoiesis [94]. hematopoietic [458]. heme [1028]. Hemibagrus [439]. Hemiscyllium [167]. Hemoglobin [1392, 342]. Hemolysis [1069]. hemopoietic [1195]. Hepatic [1326, 951, 1300, 1187, 114, 599, 806, 1049, 989, 265, 327, 861, 1075, 324, 1394, 1183]. hepatocyte [1176, 296, 1329, 53, 286, 151]. hepatocytes [2, 374, 781, 983, 504, 669, 1077, 359, 589]. hepatopancreas [581, 811, 917, 714]. Hepatoprotective [482, 151, 714, 805, 296]. Hepatotoxic [1192]. Hepcidin [1275, 738]. herbal [714]. herbicide [1195]. heritable [684]. hermaphroditic [821, 959, 733, 1099, 21]. Heros [543]. herring [1369, 1006]. Herzenstein [1072]. Hesperozygis [597, 712]. heteroclitus [272, 72]. Heteropneustes [113, 325, 285, 795, 762, 933, 1298, 178, 743, 1313, 866]. Heterotis [54]. hibernated [370]. HIF [934, 1297, 490, 1218, 886]. HIF-1 [934, 1297, 490, 1218, 886]. High [613, 539, 223, 1114, 514, 1134, 489, 729, 984, 1144, 936, 1086, 983, 1088, 840, 629, 40, 1051, 1116, 591, 718, 1130, 990, 1029, 1251, 508]. high-carbohydrate [1086]. high-dose [990, 1029]. High-fat [613, 1114, 983, 1116, 1130, 1251]. high-mobility [591]. high-quality [489]. high-temperature [840]. higher [773]. highly [82]. hilarii [814]. Hippocampus [1047, 799, 1043, 601]. Hippoglossus [254, 299, 71]. HIRA [559]. histidine [440]. histo [148]. histo-architectural [148]. Histochemical [662, 1351, 1192, 762, 273, 982, 799]. histocompatibility [202, 404, 138]. Histological [623, 1106, 326, 928, 663, 664, 435, 930, 405, 76, 1389, 867, 799, 1346, 60, 875, 528]. histology [439, 322, 250, 933, 776, 915, 868, 1263, 834, 533, 1072, 1251]. Histomorphological [964]. histomorphology [846, 1299]. histones [144]. Histopathological [438, 279, 607, 1031, 598, 261, 751]. histopathology [1074, 1265, 203]. history [1187]. Hnf4 [1179]. hoki [104]. Hokkaido [156, 782]. holbrooki [491, 1360]. homeostasis [1199, 388, 424]. homogeneity [538]. homolog [960, 1385, 1396]. homologs [916, 876]. Homologue [582, 506, 591]. homology [1287]. honeysuckle [1365]. honmoroko [1158, 509, 453]. Hoplias [560, 1077]. Hoplostethus [899]. Horabagrus [301, 44]. hormonal [439, 1243, 525]. hormone [1304, 125, 821, 128, 15, 23, 451, 183, 142, 419, 1149, 674, 844, 1318, 27, 1127, 762, 444, 978, 1386, 694, 860, 418, 1328, 364, 856, 1194, 282, 1020, 1209, 702, 397, 476, 655, 900, 1019, 1307, 1348, 769, 29, 1004, 1270, 770]. hormone-sensitive [142, 1149, 1020, 702]. hormone/insulin [1318, 1348]. hormone/insulin-like [1318, 1348]. hormones [502, 588, 80, 922, 670, 356, 262, 817, 1212, 1090, 1166, 881, 1340]. hornorum [183]. hospital [1085]. hot [1159, 985]. housed [706]. housekeeping [734, 840]. HPG [1035]. HPI [583, 801, 1258]. HPI-axis [1258]. HSC70 [747, 511]. HSC70/HSP70 [747]. HSL [1149]. HSP [852]. hsp47 [985, 888]. HSP70 [15, 849, 747, 445, 330, 718, 984, 558, 34]. HSP70s [608]. HSP90 [284, 919, 718, 984, 985]. HSPs [688, 1051, 1068]. hubbsi [842]. Huelva [438]. HUFA [56, 659, 217]. HUFA-enriched [56]. human [883]. humic [464]. humoral [606, 381]. humphead [213]. hus [549]. Huso [80, 176, 1203, 1084, 373, 184, 260, 139, 56, 549, 579, 292, 362, 65]. hyaluronidases [624]. hybrid [1074, 1084, 518, 1210, 1150, 1262, 971, 1370, 1007]. hybrids [623, 83, 28]. hydrochemical [10]. hydrochloride [987]. hydroelectric [369]. hydrolysate [1101, 1261]. hydromineral [165, 739]. hydrophila [1390, 1076, 1103, 1056, 408]. hydrostatic [790]. hydroxyl [804]. hydroxylase [1381, 476]. hydroxylase/17 [100]. hydroxylation [220]. hydroxynonenal [494]. hydroxyproline [1381]. Hydroyeast [732]. Hyper [1314]. Hyper-ammonia [1314]. hyperglycemia [1210, 986]. hypermelanosis [412]. hyperoxia [154, 108]. hypo [1341]. hypo-osmoregulatory [1341]. hypochlorite [1237]. Hypomesus [847]. Hypophthalmichthys [665, 1373]. hypophthalmus [598, 691, 981, 1220]. hypothalamic [170, 583, 1209, 819]. hypothalamus [801]. hypotonic [200]. Hypoxia [544, 621, 1322, 1289, 1325, 934, 494, 1110, 602, 1311, 1045, 1285, 1044, 555, 1272, 490, 688, 174, 1162, 1121, 720, 483, 339, 414, 1383]. hypoxia-inducible [1044]. hypoxia-sensitive [555]. hypoxia-tolerance [483]. hypoxia-tolerant [934, 494, 555]. hypoxic [903, 1297, 1028, 1333, 555, 1222, 1317, 1303, 941]. Hz [1137].

i-STAT [1364]. I. [823]. ice [1308]. ichthyotoxicity [1105]. Ictalurus [84, 66, 823, 355, 315, 345]. ide [552]. ideal [431]. idella [1391, 630, 447, 811, 917, 1093, 1133, 1329, 1076, 594, 781, 1365, 1196, 1020, 1046, 1058, 1303, 1367, 24, 1269]. idellal [729]. idellus [542, 599, 380, 1002, 57, 659, 771, 374, 1393, 1033, 1295, 461, 939, 897, 358, 997, 1097, 686, 1371, 483, 547, 68, 955, 1122, 313, 95, 589]. Identification [734, 67, 542, 855, 786, 172, 987, 432, 328, 183, 1333, 1148, 825, 490, 1262, 478, 634, 733, 144, 1030, 916, 789, 816, 1206, 1303, 1217, 302, 1062, 1331, 95, 798, 630, 1266, 75, 1184, 500, 823, 1174, 813, 919, 832, 135, 355, 639]. identified [221, 1385, 1396]. idle [68]. idus [552]. IFN [461]. IFN- [461]. IGF [742, 706, 152, 860, 1259, 208, 1370]. IGF-1 [742, 860]. IGF-binding [152]. IGF-I [706, 1259]. IGF-II [1259]. IGF/IGFBP [208]. Igf1 [924, 994, 1104]. Igf1r [634]. IGFBP [382, 677, 913, 1387, 208, 525, 1377]. IGFBP-1 [913]. IGFBP-2 [677, 525]. IGFBP-2a [1377]. IGFBP-2s [1377]. IGFBP-3 [382]. IgM [633, 49]. II [34, 3, 80, 290, 1205, 694, 1328, 1259, 713]. IIA [202]. IIB [202, 404, 833, 996]. ilishaeformis [987]. illumination [911]. Imbrasia [1346]. immature [302]. immune [1159, 1289, 188, 8, 371, 606, 1280, 1054, 489, 1264, 1071, 1339, 1098, 1203, 376, 579, 1092, 1306, 1353, 1047, 1037, 1213, 895, 329, 759, 801, 671, 968, 886, 1281, 69, 49, 1033, 1319, 333, 168, 943, 876, 718, 51]. immune-related [1159, 1280, 1339, 1033]. immunity [1214, 1134, 944, 1031, 1107, 1294, 752, 936, 980, 633, 358, 1163, 1296]. Immuno [1355]. Immuno-physiological [1355]. Immunocytochemical [630]. Immunohistochemical [909, 982]. immunolocalisation [1259]. Immunolocalization [197]. immunological [1322, 827, 543, 1137, 250, 1263, 381, 629, 1343, 1202, 261, 981]. immunology [148]. Immunomodulatory [522, 1161, 495, 617]. immunoreactive [526, 749]. immunoreactivity [33]. immunosensor [419, 830]. immunosorbent [563, 203]. immunostimulant [1098, 431]. immunostimulation [1390]. Immunostimulatory [873]. Immunotoxic [1108]. immunotoxicity [1064]. Impact [1342, 250, 1189, 305, 1359, 925, 886, 702]. Impacts [1373, 228, 720]. impair [1199]. impaired [1367]. impairment [706, 316]. impairs [1110, 1341]. impeded [367]. implanted [665, 1057]. implants [365, 484]. implicated [1171]. implications [602, 812, 1245, 1244, 397, 1359, 236]. importance [130]. important [1120, 985]. impounded [70]. improve [109, 1361, 654, 162, 483]. improved [1272, 1246, 460]. improvement [1289, 474]. improves [805, 1203, 287, 1145, 980, 997, 1130]. inability [896]. Inactivation [586, 316]. inchonensis [1263]. including [89, 1170]. Inclusion [1203, 1012, 1144, 116, 1213, 834]. inclusions [1190]. incomplete [710]. incorporation [711, 787]. Increase [790, 400, 1128, 521, 445, 1349]. Increased [245, 1220, 918, 426, 768, 206, 282, 194, 783]. increases [532, 544, 303, 819, 1036]. increasing [1221, 895]. index [1387, 106]. Indian [113, 325, 114, 1358, 146, 1016, 117, 710, 59, 516, 692]. indicate [1310]. indicative [245]. indicator [648, 229]. indicators [226, 1069, 442]. indices [1176, 373, 675, 728, 310]. indicum [178]. indispensable [950, 54]. Individual [1319, 699, 485]. individually [706, 64]. individuals [206, 302]. induce [140]. induced [1227, 215, 161, 352, 1064, 502, 1207, 1176, 331, 605, 64, 446, 86, 361, 546, 214, 1250, 612, 753, 296, 1275, 1009, 1329, 307, 428, 1312, 804, 1075, 889, 1354, 1389, 551, 316, 545, 274, 133, 145, 40, 1065, 197, 897, 1380, 151, 1270, 1059, 1029, 714, 1099]. inducer [689]. inducers [68]. induces [1138, 984, 1182, 629, 71, 1053]. inducible [1314, 1044, 558]. inducing [419, 668, 701, 1144, 1133, 1189]. Induction [34, 381, 750, 365, 1023, 301, 548, 317, 1314, 75, 454, 499, 18]. inert [528]. infected [1338, 873]. infection [1390, 891, 1103, 879, 661, 1165, 1056, 678, 461, 939, 333, 213]. infectious [854, 303]. infested [850, 456]. inflammation [1329, 1022, 1261]. inflammation-related [1022, 1261]. inflammatory [1289, 1339]. Influence [2, 1117, 188, 301, 1098, 412, 204, 163, 265, 1146, 759, 194, 40, 940, 1079, 1159, 833, 369, 920, 787, 1306, 675, 129, 1165, 1225]. Influences [324, 1107, 1294, 97, 9, 589]. ingredient [721]. ingredient-based [721]. ingredients [892]. inhabiting [329]. inhibit [1142]. inhibited [215]. inhibiting [1297]. Inhibition [893, 352, 447, 1275, 1249, 1394]. inhibitor [1054, 148]. inhibitor-1 [1054]. inhibitors [653, 220]. Inhibitory [1131, 1205, 117, 316, 971]. inhibits [730, 1018, 1097]. iniae [805]. initial [723, 193, 239]. initiates [1256]. injected [882, 19]. injection [570, 170]. injections [476]. Injuries [228]. injury [1176, 612, 1144, 714]. Innate [606, 489, 1071, 1098, 329, 759, 886, 168, 1296]. innocuous [554]. inorganic [1306, 1202]. inosine [1294]. insecticide [136, 938]. Insights [876, 34, 794, 388, 488, 787, 1111, 1316, 1243, 1179]. Insulin [913, 860, 713, 525, 382, 677, 548, 23, 275, 782, 1387, 781, 634, 1020, 1276, 1019, 1034, 1026]. Insulin-like [913, 860, 713, 525, 382, 677, 23, 1318, 782, 1387, 634, 1276, 1019, 1348]. intake [792, 1131, 356, 978, 1375, 1005, 900, 961, 1315]. Integrated [106]. integrity [1288, 735, 1354, 1070, 717, 508]. integrity-protecting [1070]. intense [1334]. intensity [1117, 240, 1182, 241, 55]. intensively [1201]. interacting [1262]. Interaction [529, 421, 43]. Interactions [637]. Interactive [489, 1080, 1252]. interference [1069]. interferon [1256]. interleukin [953]. interleukin-6 [953]. intermediary [806]. intermediate [1160, 929, 678]. intermediately [828]. internal [757]. interrenal [583, 801]. interruption [352]. Interspecific [1233, 1238]. Intestinal [421, 1074, 709, 521, 322, 1144, 1076, 163, 808, 594, 1353, 1265, 662, 915, 1263, 304, 590, 1211, 527, 834, 1225, 603, 1095, 686, 887, 1261, 1367, 1370, 760, 508]. intestine [1299, 380, 581, 537, 492, 273, 217, 421, 1090, 931, 1218, 1351]. intoxified [149]. intracellular [154]. Intrafollicular [1194]. intraperitoneal [275, 570, 170]. intraperitoneally [19]. intraperitoneally-injected [19]. intravenous [642]. inulin [633]. invasive [554]. inventory [1235]. inversions [452]. Investigating [161]. Investigation [82, 810, 1167]. investigations [673]. involved [925, 687, 1154, 804, 1301, 1222, 1243, 448, 1225, 779, 1179, 1217, 1315]. Involvement [1376, 1045, 1318, 1258, 344, 620, 970, 790, 1148, 814, 1019]. Iodine [282]. Ion [308, 1339, 1226]. ion-regulation [1339]. ions [343, 447]. IPNV [614]. Irf3 [1256]. iridescent [598]. Iron [185, 1275, 424, 1056]. IRS1 [1026]. IRS2 [1026]. isoform [958]. isoforms [455, 962, 996, 143, 536, 477, 1020, 481]. isolate [1031, 1129]. isolated [2, 103, 298, 1164, 1308]. Isolation [353, 20, 451, 409, 459, 14, 187, 1094, 836, 88, 318, 1174]. isoleucine [1080, 1059]. isotocin [1180, 466, 635]. isozymes [1205]. Isurus [773]. iTRAQ [904]. IV [1159].

Japan [782, 156, 203]. Japanese [748, 376, 387, 1290, 623, 932, 869, 469, 843, 1080, 1096, 1340, 1004, 525]. japonica [287, 1365, 469, 1385, 1396, 591]. japonicas [748, 376]. japonicus [911, 7, 297, 12, 567]. Jatropha [116, 1129, 1156]. javelin [908]. Jenynsia [837]. Jian [1176, 440, 752, 804, 680, 1261, 714]. JNK1 [330]. JNK1/2 [330]. Johnius [221]. Juan [899]. junction [1288, 686]. jundiá [528]. juvenile [1300, 161, 580, 999, 1135, 443, 405, 570, 709, 440, 209, 510, 165, 464, 376, 785, 1293, 517, 139, 849, 6, 1038, 1224, 579, 782, 277, 319, 585, 1092, 1362, 595, 852, 1002, 1248, 1353, 254, 299, 1301, 1150, 657, 1037, 1246, 807, 777, 1223, 744, 292, 7, 195, 300, 1282, 1365, 895, 540, 662, 688, 465, 601, 703, 926, 1335, 12, 1381, 350, 324, 40, 1361, 969, 1281, 1033, 809, 1048, 1319, 291, 1036, 1380, 803, 1109, 887, 1163, 1183, 483, 1091, 1143, 1081, 1261, 65, 124, 347, 1050, 1254, 1367, 110, 339, 479, 890, 313, 649, 1007, 974]. juveniles [512, 846, 1299, 502, 177, 858, 270, 89, 784, 806, 622, 105, 670, 56, 1044, 802, 1178, 776, 1115, 295, 1070, 449, 38, 715, 1141, 971, 1170].

K-dependent [62]. Kamenskii [915]. Karenia [1105]. Kaup [177, 434, 1170]. kDa [584, 445]. Keap1 [1053]. kelp [1124]. kept [1060]. kernel [116, 1156]. keta [470]. ketoglutarate [1163]. ketotestosterone [396]. key [561, 798, 50, 478, 366, 892]. kidney [32, 384, 115, 188, 625, 11, 752, 854, 164, 458, 539, 933, 87, 1124, 1161, 530, 672, 574, 441, 68, 363]. kidneys [977]. KIF3A [1177]. KIF3B [1177]. KIFC1 [1087]. killed [1263, 834]. killifish [314, 780]. kinase [1391, 1041, 1323, 316, 669, 1249]. kinds [1211]. kinesin [1087]. Kinetic [114, 58]. kinetics [925, 90, 351, 811, 220, 586]. kingfish [636, 161, 838, 1276, 394]. Kiss [394]. kiss2 [916, 1035]. kisspeptin [1247, 1035]. kissr2 [916, 1035]. kisutch [562, 1328]. Kit [1243]. Kit/Kitlg [1243]. Kitlg [1243]. Kleine [55, 719]. koi [1343, 1351, 1351]. konjac [603, 760]. Korean [418, 515]. Kryptolebias [1304]. kutum [915, 915]. Kützing [298].

L [912, 1067, 1073, 1195, 1204, 1043, 181, 1005, 325, 502, 105, 897, 291, 1064, 377, 445, 816]. l-carnitine [897, 1064]. l-tryptophan [325, 502, 105, 291]. L. [1390, 1168, 1214, 514, 462, 1341, 1017, 1300, 467, 35, 1342, 1229, 371, 548, 42, 1071, 240, 486, 1332, 186, 884, 411, 1250, 612, 673, 1084, 433, 160, 787, 184, 260, 522, 583, 822, 53, 1073, 164, 458, 539, 675, 250, 108, 241, 62, 116, 273, 1353, 1016, 149, 217, 1310, 895, 924, 868, 304, 47, 31, 206, 1022, 635, 133, 1021, 1189, 947, 931, 692, 38, 685, 321, 1295, 281, 83, 901, 979, 552, 553, 452]. L10a [626]. label [419, 830]. label-free [419, 830]. Labeo [114, 90, 338, 502, 1031, 546, 375, 1358, 1334, 50, 223, 307, 337, 428, 429, 647, 478, 148, 1350, 545, 1156, 1198, 41, 1167, 247]. Laboratory [5, 1157, 298]. labrax [827, 962, 846, 1187, 1071, 473, 616, 471, 872, 526, 1226, 445, 648, 350, 38, 1141, 720, 168, 452]. labrosus [1147]. Lack [108, 1335, 896]. lacking [217]. lactate [174, 38, 335]. lactic [772]. Lactobacillus [992, 944, 97, 633]. lake [58, 535, 871, 870, 1205, 179, 1136]. lakes [491]. lalandi [636, 161, 838, 1276]. lambari [587]. lambda [1273]. lambda-cyhalothrin [1273]. Lampetra [1385, 1396, 591]. lamprey [855, 101, 1310, 1385, 1396]. lanceolatus [23, 1210, 500]. lard [1196]. large [141, 1309, 1094, 861, 499, 420, 49, 1368, 1053, 990, 1029]. largemouth [443, 550, 246, 437, 768, 1361]. Larimichthys [1309, 1094, 1222, 420, 1368, 1087]. Larmichthys [861]. larvae [278, 620, 1284, 3, 462, 912, 1017, 636, 520, 950, 1061, 791, 872, 48, 243, 433, 670, 99, 571, 1298, 841, 306, 1363, 150, 199, 496, 972, 610, 935, 924, 469, 683, 640, 47, 1022, 360, 590, 282, 898, 448, 528, 1209, 745, 1106, 1225, 880, 971, 720, 1252, 1340, 505]. larval [462, 960, 1, 72, 1374, 604, 1136, 552, 826]. larviculture [564]. larvicultured [874]. LAT1 [547]. late [94]. Lateolabrax [748, 376, 1149, 1095]. Lates [532, 409, 1038, 1287, 82, 629, 575]. latipes [914, 1345, 980, 1312, 660, 1082, 818]. latipinna [204, 893]. lavaretus [794]. LC [588, 1179]. LC-PUFA [588, 1179]. Lead [593, 192, 753, 36, 569]. leaf [1219, 1250, 1161, 1326, 178]. leaffish [460]. learning [316, 145, 756]. learning-induced [145]. leaves [1390, 1213, 482]. lecithin [307, 428]. LECT2 [463]. lectin [531, 46, 379]. lectins [668, 701]. LED [936, 638]. LEDs [911]. Leiocassis [1370]. length [1125, 1143]. lenses [121]. leopard [126, 604, 633]. Lepisosteus [1366]. Lepomis [906, 1384]. Leporinus [289]. Leptin [963, 401, 374, 781]. Leptoclinus [341]. Lesueur [991]. let [1330]. let-7 [1154, 1330]. lethal [435, 784, 952, 1344, 357]. Leucaspius [248]. leucine [1301, 1079, 1080]. leucine-enkephalin [1079]. Leuciscus [552]. leucocyte [44]. leucocytes [1063]. Leucoraja [485]. leukocytes [1126, 115, 188, 87, 576]. levan [523]. level [1300, 270, 342, 1093, 336, 265, 852, 817, 1378, 775, 276, 845, 770, 239]. levels [113, 543, 1355, 934, 1352, 308, 795, 177, 1391, 581, 11, 464, 186, 411, 849, 6, 1149, 140, 1278, 154, 444, 736, 1002, 1323, 1184, 84, 1047, 1150, 1223, 895, 817, 427, 59, 289, 1165, 346, 1209, 702, 335, 819, 830, 997, 1163, 955, 1331, 990]. levothyroxine [1336]. LH [125, 748, 446, 24]. LHRH [472]. LHRH-A [472]. life [486, 783]. ligand [266, 679]. ligand-1 [266]. ligand-binding [679]. ligands [1164]. light [1117, 1197, 546, 160, 911, 936, 1145, 1334, 1363, 1233, 7, 434, 1185, 55, 638, 112, 1277, 1011]. light-emitting [1011]. light/dark [434]. lighting [730]. like [1063, 382, 677, 913, 786, 501, 23, 1332, 1366, 375, 1318, 782, 1387, 634, 860, 697, 873, 1030, 1276, 713, 1019, 1348, 818, 525]. limitations [238]. limiting [295]. Limmaeus [1284]. Lin [284, 1154]. lin-28a [1154]. lin28 [1154]. lin28/let [1154]. lin28/let-7 [1154]. linalool [1119, 1258]. line [1332, 1055, 1255, 825, 323, 1124, 1000, 689, 1174, 755, 122, 118, 652, 311, 363, 513]. lineatus [593, 279, 717, 576]. lines [509, 453, 854, 131, 754, 774, 801, 831, 1096]. linked [884, 563, 203]. Linn. [456, 146]. Linnaeus [999, 883, 850, 523, 571, 1388, 632, 292, 732, 137, 498, 65]. linoleic [708]. Linseed [1361, 918, 250]. lionfish [624]. Liopsetta [203]. Lipase [907, 949, 1207, 1221, 142, 1149, 271, 134, 1020, 702]. lipases [104]. Lipid [1283, 1017, 1295, 1114, 1173, 1110, 1300, 258, 141, 708, 406, 588, 805, 613, 1221, 729, 963, 1084, 517, 917, 1133, 1287, 264, 1002, 1248, 1150, 861, 179, 374, 217, 61, 1365, 1146, 47, 449, 1190, 1112, 947, 1320, 702, 1033, 1141, 1394, 997, 1091, 955, 1050, 1130, 1331, 1254, 649, 1007, 1251, 1269, 589]. lipid-lowering [729]. lipid-related [1112, 1251]. lipids [189, 98, 341, 171]. lipogenesis [74, 1301]. lipogenic [533]. lipoic [965]. lipolysis [74, 659]. Lipolytic [346]. Lipopolysaccharide [381, 885, 738, 170, 933, 333]. lipopolysaccharides [609, 1319]. lipoprotein [729, 142, 389, 1149, 702, 1082]. lipped [1147]. Lippia [288, 835, 1119, 1258, 597]. Listonella [445]. little [485]. Live [267, 1017, 1061, 268]. lived [1146]. Liver [189, 1393, 499, 272, 532, 58, 92, 435, 494, 245, 587, 908, 1074, 625, 1039, 11, 612, 849, 140, 1287, 250, 19, 1104, 36, 1075, 179, 776, 1262, 341, 1088, 868, 767, 684, 530, 474, 977, 460, 289, 947, 350, 1361, 203, 975, 279, 892, 575, 1179, 1349, 1091, 1011, 220, 990, 1251, 974]. livers [1325]. loach [646, 53, 826, 1027]. loaches [1118]. load [544, 583]. loading [1086, 1027]. local [613]. localisation [1350]. Localization [1304, 697, 380, 909, 1006, 1313, 208]. locally [298]. location [153]. locomotion [790]. locomotive [1100]. locomotor [351, 434]. Long [570, 1308, 892, 212, 1160, 1197, 312, 510, 637, 918, 1357, 1123, 108, 1222, 672, 601, 516]. long-chain [312, 637, 918]. Long-term [570, 1308, 892, 212, 1160, 510, 1357, 1123, 108, 672, 516]. longa [1032]. longfin [1106]. longirostris [1370]. Lonicera [1365]. Lophiosilurus [1282]. loricariid [414]. loss [1160, 721]. Lota [537, 492]. Low [140, 1104, 606, 996, 405, 992, 550, 724, 76, 277, 319, 1137, 205, 1044, 807, 759, 529, 1245, 1244, 1030, 1082, 1163, 990, 1029]. low- [990]. low-density [1082]. low-dose [1029]. low-frequency [1137]. low-molecular-weight [1030]. low-oxygen [1044]. low-phosphate [996]. Lowe [858]. lowering [729, 465]. loxP [524]. LPL [1149]. LPS [1207, 490, 381]. LPS-induced [1207]. LR [298, 503, 1064, 294]. Lrp13 [1082]. lucidum [1075]. lucioperca [571, 250, 1057, 498]. lucius [1142, 1237, 880]. luteinizing [125, 674, 418, 769]. Lutjanus [609, 48, 610, 935, 213]. lyase [100]. lycopene [1113]. lymphocytes [86]. lymphoma [668, 701]. lyophilized [662]. lysine [594, 421, 436, 1208, 1144]. lysosomal [660]. lysozyme [976].

M [778]. maccoyii [942]. macdonaldi [791, 776]. mackerel [911, 7, 297, 1245, 1244]. maclovinus [678, 809]. macrocephalus [716, 289]. macrochirus [906]. macromolecules [898]. macronutrient [929]. macrophage [933, 495]. macrophages [1161]. macropomum [927, 1336]. macrorhynus [221]. Macruronus [104]. maculata [1150]. maculatofasciatus [3]. maculatum [79, 147]. maculatus [1149, 341, 1095]. magnesium [263]. magnitude [161]. magur [1314, 1314]. mahseer [775, 964, 311]. main [898]. Maine [485]. maintaining [1376]. maintenance [113, 12, 721]. Major [202, 114, 1358, 404, 117, 516, 138, 992, 1264, 1049, 989, 1107, 1294, 1101, 735, 1308, 1105, 1281]. mako [773]. malabaricus [560, 1077, 335]. malachite [27]. MALDI [500]. male [930, 1352, 396, 126, 119, 491, 1120, 214, 1084, 1395, 1158, 26, 1397, 450, 1210, 1047, 623, 744, 242, 217, 1239, 1228, 1259, 1201, 693, 498, 775, 699, 1215, 401, 1370]. males [730, 1142, 1118, 290, 181]. malic [1006]. malondialdehyde [898]. maltose [89]. mammalian [34, 80]. management [780]. Manchurian [507]. mandarin [1256, 1375, 919, 468]. manganese [700, 937]. manipulation [1128, 1246, 1239]. mannan [292]. mannanoligosaccharide [603]. mapping [455]. maraena [734, 734]. marble [1234, 783]. marbled [960, 417, 904]. marginally [1299]. marginata [1326]. marine [308, 1, 1049, 989, 198, 277, 571, 1362, 825, 75, 265, 98, 1105, 1022, 346, 1225, 892, 122, 1179]. marinus [1310]. marker [823, 135]. markers [1325, 343, 672, 975]. marmorata [904]. marmoratus [1304, 1234, 783]. Masoten(R) [726]. masou [63]. Mass [896, 342, 318, 565]. masu [63]. maternal [1214]. Matrilin [206]. Matrilin-1 [206]. matrinchã [349]. matrinxa [1273, 291, 819, 1036]. matrix [761, 1021, 693]. matter [1231]. maturation [945, 548, 970, 361, 419, 287, 387, 257, 332, 31, 157, 427, 401, 901, 573]. maturation-inducing [419]. maturity [891, 563]. mawsoni [807]. mawsoni-CaM [807]. maxima [72]. maximum [1379]. maximus [514, 912, 17, 885, 1291, 1120, 1098, 1332, 1374, 165, 963, 628, 674, 844, 1318, 1353, 767, 1292, 497, 281, 481, 122, 511, 958, 998, 740]. may [1087, 961, 1315]. Mayan [596, 150]. maze [920]. MCH [900, 1307]. MDP [478]. meager [994]. meagre [1160, 331, 820, 475, 508, 1185]. meal [1208, 690, 1052, 1293, 1038, 263, 1144, 116, 749, 332, 1022, 1156, 1346, 921, 1320, 995, 1268, 1261, 313, 173]. meal-based [1268]. measurement [203]. measuring [1364, 1069]. mechanical [544]. Mechanism [387, 794, 1152, 443, 123, 955, 1089, 750]. mechanisms [666, 21]. mechanistic [889]. medaka [914, 1345, 980, 825, 660, 1082, 818]. mediate [557]. mediated [1372, 1068, 330, 1249, 1033]. mediation [446]. mediator [572]. mediator-type [572]. Medicinal [1126]. Mediterranean [994]. mef2aa [1000]. Megalobrama [1208, 608, 953, 1028, 1333, 556, 1045, 1278, 1330, 264, 765, 736, 1086, 1184, 1301, 1271, 1372, 983, 1051, 1056, 737, 887, 28, 718, 1050, 1296, 1251, 513]. Mekong [981]. Melamine [672]. melanin [933, 900, 1307]. melanin-concentrating [900, 1307]. Melanins [762, 933]. melano [933]. melano-macrophage [933]. melanocortin [1271]. melanocortin-5 [1271]. melanogenesis [1018]. melanophore [215]. melanophores [545]. melanopus [430]. Melatonin [494, 1180, 1117, 484, 936, 1185, 879, 1200]. member [1318]. members [1165]. membranaceus [906]. membrane [656, 735, 717, 355]. mep50 [914]. Mercuric [546, 166, 480]. mercury [727, 123, 1209, 650, 560]. meridionalis [187, 25, 652, 124, 347]. Merluccius [1186, 842]. mesopotamicus [662, 726]. Metabolic [35, 858, 1100, 320, 797, 726, 835, 1267, 1300, 314, 520, 1067, 89, 176, 523, 673, 1024, 585, 116, 337, 429, 264, 853, 1001, 657, 1282, 61, 1156, 1324, 947, 216, 38, 783, 763, 892, 1078]. metabolism [1114, 73, 1173, 951, 1279, 1017, 1160, 851, 690, 1052, 406, 473, 929, 725, 806, 729, 1049, 989, 1084, 917, 1287, 1384, 265, 327, 695, 1002, 1272, 1301, 1150, 1372, 374, 781, 777, 149, 744, 1365, 82, 449, 527, 1381, 253, 533, 1344, 1033, 1141, 678, 809, 1368, 803, 997, 1081, 941, 1130, 1254, 649, 1007, 1269, 589]. metabolism-related [1017, 744]. Metabolite [1338]. metabolites [1160, 177, 1224, 1033, 1048]. metabolized [1310]. metabolomic [196, 1359]. metal [11, 447, 627, 533]. Metallothionein [11, 123, 336, 650]. metals [192, 336, 1205]. metamorphosing [1270]. metamorphosis [881, 1004, 525]. methanol [717]. methemoglobin [454, 378]. methemoglobinemia [1392]. methionine [1144, 449]. method [1105, 474]. methodologies [1]. methods [828, 537, 492]. methoxychlor [1228]. Methyl [337, 429, 717, 707, 43]. methyl-parathion [43]. methylation [905, 1266, 1104, 864, 770]. methylmercury [80, 176]. methylotropic [29]. methylphenidate [1352]. methyltestosterone [848, 631]. methyltransferase [829, 315]. metomidate [252]. Mg [578]. mGnRHs [80]. Mianserin [993]. microalgae [371, 1105, 1327, 1170]. microbial [523, 879]. microbiota [1299, 709, 163, 564]. microcystin [503, 1064, 298, 321, 294]. microcystin-containing [321]. microcystin-LR [298, 503, 294]. microcystin-LR-induced [1064]. microcystins [298, 19]. Microcystis [503, 298]. microelectrode [1305]. microencapsulated [1215]. microflora [808, 1008, 603, 760]. micronucleus [1085, 504]. microorganism [1265]. microparticulate [874]. Micropterus [443, 550, 768, 1361]. microRNA [1292, 1340, 333]. microRNA-seq [1340]. microRNA-transcriptional [1292]. microRNAs [1330, 1062]. microscopy [546]. microsomes [327, 220]. middle [686]. migration [35, 782, 156, 157, 865]. migratory [1341, 101, 157, 519]. Miichthys [368, 201, 366]. miiuy [368, 875, 201, 366, 368, 201, 366]. milieu [1077]. mill [973]. mineral [600]. mineralization [761]. minerals [721, 1225]. mini [400, 398]. mini-review [400, 398]. Minimal [1201]. minimally [554]. mining [1366]. minnow [746, 1285, 153, 839, 985, 881, 18, 276, 829, 607]. minnows [1193, 731]. Minor [947, 563]. mint [715]. miR [1045, 1372, 789]. miR-133a [789]. miR-34a [1372]. miR-34a/Sirtuin-1 [1372]. miR-462 [1045]. miR-462/731 [1045]. miRNA [1039, 1372]. Misgurnus [1118, 754, 1316, 506]. Mitchill [159]. mitigate [1025]. mitigates [307, 428]. mitigating [647]. mitochondria [486, 983]. mitochondria-rich [486]. Mitochondrial [330, 700, 796, 735, 437, 717, 1083, 28, 1029]. Mitochondrion [180, 277, 890]. Mitochondrion-rich [180, 277, 890]. mitogenic [379]. mixed [1348]. mixture [987]. mixtures [699]. MK [316]. MK-801-induced [316]. MMP [893]. MMP-2 [893]. MMP-9 [893]. MO [470]. moara [1124]. mobility [591]. mobilization [1384]. model [34, 1128, 5, 457, 954]. modeling [1287]. moderate [720]. Modifications [1023, 1022, 1383]. modified [828]. modifies [1203, 185]. modulate [1126, 1187, 722, 242]. modulated [1264, 588]. modulates [1289, 1325, 33, 1384, 564, 669]. modulating [920]. Modulation [93, 1247, 675, 50, 695, 1002, 968, 1254, 395, 398]. Modulatory [1064, 1113]. mojavensis [872]. Molecular [340, 1274, 962, 1311, 125, 382, 1280, 368, 1054, 748, 738, 175, 128, 1207, 380, 1297, 1374, 80, 46, 674, 844, 1266, 290, 1287, 430, 747, 1103, 1086, 1184, 1271, 410, 802, 839, 986, 404, 924, 994, 942, 420, 1162, 743, 976, 660, 266, 497, 1056, 1046, 127, 276, 461, 813, 900, 1276, 923, 472, 138, 815, 315, 958, 213, 1377, 631, 24, 21, 740, 1186, 388, 677, 913, 501, 23, 1391, 791, 1309, 556, 1278, 336, 202, 1323, 377, 1388, 1356, 1030, 1196, 575, 511, 655, 1181, 573, 713, 547, 1172, 876, 30, 525, 1089, 1034, 1026]. molecule [793, 366]. molecules [687]. molitrix [1373]. monitor [572, 239]. monoamine [93, 170, 819]. monocrotophos [750]. Monodactylus [277]. monomeric [1212]. monophosphate [1107, 1294]. Monopterus [821, 353, 984, 681, 959, 634, 182, 22, 171]. monoxide [467]. morhua [1364, 115, 493, 1024, 517, 103, 154, 252, 87, 841, 133, 1021, 705, 979, 71]. Moringa [1250, 1213]. morio [1135]. Mormyridae [1099]. Morone [132, 317, 573]. Morpho [933, 1313]. morpho-functional [1313]. Morpho-histology [933]. morphogenesis [615]. morphogenetic [1013]. Morphologic [587]. Morphological [273, 875, 32, 384, 739, 164, 728, 1043, 460, 442]. Morphology [845, 1242, 709, 969, 865, 1072, 1261, 760, 345]. morphometry [180]. Mortality [235]. mortem [991]. MOS [292]. moschata [824]. moseri [290]. mosquitofish [491, 40, 1157]. mossambicus [423, 280, 352, 446, 383, 309, 210, 1079]. motility [653, 656, 1242, 1382, 1231, 578, 735, 685, 654, 717]. motivation [667]. motor [1087]. mould [148]. mount [1194]. mountain [18]. moving [160]. Mozambique [383, 309, 210]. Mozart [719]. MPTP [707]. mrigala [592, 1350, 516]. mRNA [934, 1180, 1117, 913, 128, 501, 706, 608, 987, 317, 1339, 1374, 142, 1149, 356, 444, 1104, 306, 1292, 1042, 864, 1040, 816, 763, 276, 888, 966, 832, 770, 173]. mRNAs [1318]. mt2 [966]. Mu [620]. mucilaginosa [1286]. mucosa [521, 1076, 78, 163, 1164]. mucosal [1264, 1213, 671, 1343]. mucous [329, 1350, 1351]. mucus [514, 992, 592]. mud [984]. mudskipper [948]. MUFA [473]. Mugil [548, 330, 669]. mullet [548, 1147]. mullets [330]. mulloway [12, 567]. Multi [405, 1342, 711, 1353, 1273]. multi-biomarker [1342, 1273]. Multi-organ [405]. multi-strain [711, 1353]. multidentata [837]. Multiple [1309, 389, 573, 1235, 1186, 800, 765, 297]. multiple-dose [765]. multiple-spawning [1186]. multiwalled [1279]. muramyl [478]. MuRFs [1347]. Murray [956]. murrel [700, 930, 258, 710, 840, 60, 261]. Muscle [462, 1310, 272, 455, 1219, 912, 35, 1208, 1160, 402, 1067, 141, 351, 625, 581, 459, 1138, 613, 1332, 952, 811, 917, 918, 1171, 146, 221, 1037, 1000, 764, 1308, 1190, 38, 112, 305, 565, 678, 162, 1349, 1371, 1062, 95]. muscle-specific [1067]. muscles [196, 1337, 1073, 974]. muscular [724]. mushroom [944]. music [55, 719]. mussel [749]. myc [794]. Mycteroperca [126, 604, 633]. MyD88 [366]. mykiss [1227, 544, 1023, 212, 435, 245, 925, 1126, 1382, 219, 690, 218, 722, 929, 328, 613, 741, 711, 322, 1066, 1132, 1145, 1306, 424, 626, 454, 1188, 1239, 582, 1161, 1088, 759, 1263, 144, 1212, 684, 310, 97, 873, 721, 698, 1121, 378, 834, 253, 715, 255, 1320, 69, 644, 975, 772, 127, 888, 208, 845, 943]. Mylopharyngdon [119]. Mylopharyngodon [926, 1206]. myoadenylate [81]. myocardium [1121]. myoglobin [353, 942, 1162]. myoglobins [334]. myosin [455, 1171, 112]. myostatin [716, 613]. myostatin-1 [613]. myrcene [1188]. Myxocyprinus [1103, 1091].

N [166, 1128, 473, 312, 588, 1312, 659, 780, 1058, 68, 1063]. n-3 [312, 588, 659, 780, 1063]. n-6 [473]. N-acetyl [1058]. N-acetylcysteine [1128]. n-Butyl [1312]. N-demethylation [68]. naccarii [929, 1066]. Nachtmusik [55, 719]. NAD [1006]. NAD-preferring [1006]. NADH [454, 378]. NADH-dependent [454, 378]. NADPH [378]. Naltrexone [446]. Nandus [460]. Nanjing [106]. Nannochloropsis [1327]. Nano [1225]. Nano-delivery [1225]. nanoparticle [1389]. nanoparticles [928, 1264]. nanoparticulate [1306, 1202]. nanotube [419]. nanotubes [1279]. naphthoflavone [725]. native [1279]. Natural [132, 186, 59, 820]. nauplii [56, 815]. necrosis [854]. need [721]. negative [361]. nematode [1165]. nemurus [439]. Neoditrema [697]. Neotropical [560, 1259, 1077, 927, 726]. Nephroprotective [753]. nephrotoxicity [1389]. NER [1030]. Nerium [178]. Network [1292, 1199]. neu4 [660]. neuroendocrine [922]. neuropeptide [978, 680, 679, 769]. Neuropeptides [466]. Neuroprotective [1227]. Neurotoxic [751]. neurotoxicity [132]. neurotoxin [707]. neurotransmission [667]. neurotransmitter [1131, 337, 429]. neurotransmitters [93]. newly [1252]. NG [1239]. Nibea [1381]. niches [329]. nickel [751]. nicotinate [166]. niger [1099]. night [1180, 1117, 1140]. nigrofasciata [738]. Nile [211, 928, 1390, 109, 663, 664, 1159, 1289, 1325, 993, 1352, 1339, 480, 192, 486, 361, 77, 859, 1250, 753, 26, 1397, 198, 891, 1216, 732, 1326, 1090, 661, 931, 324, 667, 344, 702, 210, 1153, 1169, 639, 30, 558, 586]. niloticus [32, 384, 928, 1390, 663, 664, 1159, 1289, 1325, 1352, 1052, 882, 605, 1339, 480, 64, 192, 486, 805, 361, 77, 859, 1250, 142, 753, 26, 1397, 298, 627, 891, 426, 1216, 1246, 777, 732, 137, 54, 1326, 274, 661, 931, 324, 727, 1347, 667, 344, 702, 107, 210, 639, 30, 558, 586, 1007, 51]. nine [141]. Nitric [215, 577, 1314, 395]. Nitrite [502, 338, 1223, 454]. Nitrite-induced [502]. nitrogen [231]. Nitrogenous [464]. nitrophenol [220]. nitrosative [245]. Nme [393]. NMR [196, 1324]. NMR-based [196]. No [719]. nobilis [1297, 19, 665, 1373, 1380, 379]. nocturnal [452]. NOD2 [478]. nodal [328, 843]. noise [883, 719]. non [8, 1134, 1031, 729, 952, 752, 1092, 50, 1353, 1198, 633, 1163, 943]. non-fermented [1198]. non-gelatinized [50]. non-lethal [952]. non-protein [729]. non-specific [8, 1134, 1031, 752, 1092, 1353, 633, 1163, 943]. noncoding [1222]. nonylphenol [1192, 663, 664, 26, 1397, 1065]. norepinephrine [790]. normal [1317, 1239, 4]. normalization [459, 681, 757]. normoxia [108]. North [1174]. northeastern [782]. northern [1142, 1242, 47, 1237, 880, 813]. notch [1148]. note [504]. Nothobranchius [1003, 1146]. notochord [779]. Notothenia [1357, 529, 797]. notothenioid [1357, 809]. novaezelandiae [104]. Novel [1, 700, 1054, 1385, 1396, 591]. novo [734, 908, 841]. noxious [554]. Nramp [424]. nrF [697, 1068]. nrf-2-mediated [1068]. nrF-AGP [697]. Nrf2 [1053]. Nrf2-Keap1 [1053]. ntcp [582]. nuclear [1085, 722, 1087]. nuclei [53]. Nucleic [988, 1098]. nucleotide [478, 1030, 943, 508]. nucleotide-binding [478]. nucleotides [255]. nutrient [602, 37, 1052, 111]. nutrients [947]. Nutrition [1284, 1, 1111]. Nutritional [1187, 1323, 824, 240, 459, 120, 1129, 648, 1211].

O [154, 897, 549]. O. [925, 1052, 142, 210, 1007]. objective [1231]. oblonga [1065]. obscure [657, 1252]. obscurus [742, 786, 1054, 1134, 1175, 557, 657, 1252, 173]. Observation [686]. observations [405]. obstructs [893]. obtained [34]. ocellaris [33]. ocellatum [167]. ocellatus [934, 128, 207, 1221, 1123, 1255, 902, 757, 816]. oculatus [1366]. Odontesthes [503, 396, 574]. off [985]. officinalis [787]. oil [532, 288, 835, 422, 846, 1299, 443, 473, 1138, 165, 376, 1084, 373, 517, 354, 918, 84, 415, 566, 947, 1196, 1361, 1119, 1215, 1258, 712, 1141, 1368, 1091]. oiled [1198]. oils [422, 250, 518, 715, 597]. olaquindox [714]. olaquindox-induced [714]. oleic [1176, 1329]. oleifera [1250, 1213]. oleoyl [841]. oleracea [1390]. Olfactomedin [1366]. Olfactomedin-like [1366]. olfactory [687, 546, 1164, 772]. OLFML2A [1366]. OLFML2B [1366]. oligodeoxynucleotides [115]. oligofructose [139]. oligomerization [478]. oligosaccharide [292]. oligosaccharides [827]. olivaceus [37, 798, 1154, 1009, 412, 283, 978, 623, 932, 869, 843, 169, 916, 789, 778, 1080, 1096, 1340, 923, 1004, 1270, 876, 525, 957]. olive [1219, 1355, 37, 1009, 283, 978, 774, 923, 957]. omega [348, 1338]. omega-3 [348]. omega-6 [348]. Ompok [360]. oncogene [1094]. Oncorhynchus [1227, 544, 1023, 212, 435, 245, 1126, 1382, 219, 690, 218, 722, 929, 328, 613, 741, 711, 470, 322, 1066, 1132, 63, 1145, 1306, 104, 424, 562, 626, 454, 1188, 1239, 582, 1328, 1161, 1088, 759, 1263, 144, 1212, 684, 310, 97, 873, 721, 698, 1121, 378, 834, 253, 715, 255, 1320, 69, 644, 975, 772, 127, 888, 208, 845, 943]. One [1151]. ongoing [1111]. onset [730, 1142, 741, 528]. ontogenesis [1310]. Ontogenetic [771, 632, 670, 469, 683]. ontogenic [601]. Ontogeny [243, 811, 1298, 935, 360, 745, 505, 821, 788, 1010, 256, 640, 475, 1302, 723, 193, 468, 615]. oocyte [388, 548, 419, 387, 427, 573]. oocytes [722, 287]. Oogenesis [341]. OP [318]. open [629]. openings [486]. opinion [162]. opioid [620]. opioidergic [446]. opiomelanocortin [153]. Oplegnathus [243, 1379]. opposite [197]. opsin [1197]. optical [32, 384]. Optimal [159]. Optimisation [1229, 121]. Optimum [1339]. oral [212, 1135, 642, 765]. orange [444, 377, 293, 410, 807, 400, 590, 899, 16, 679]. orange-spotted [444, 377, 293, 410, 807, 400, 590, 16, 679]. orbignyanus [129, 1115, 1223, 717]. Oreochromis [32, 384, 928, 1390, 663, 664, 1159, 423, 280, 1289, 1325, 1352, 352, 1052, 1074, 882, 605, 1339, 480, 64, 192, 486, 805, 446, 183, 361, 77, 859, 1250, 142, 753, 26, 1397, 298, 627, 891, 426, 518, 1216, 1246, 777, 732, 137, 1326, 274, 661, 383, 309, 931, 324, 727, 1347, 667, 344, 702, 107, 210, 1079, 472, 639, 30, 558, 586, 1007, 51]. orexigenic [1139]. orexigenic/anorexigenic [1139]. orexin [476]. Organ [136, 1029, 405, 437]. Organ-specific [136, 1029, 437]. organic [1074, 1038, 1306, 569, 1202]. Organization [52, 366]. Organogenesis [99]. organophosphate [1014, 592]. organs [501, 560]. oriental [754]. orientalis [421, 195, 300, 39, 540, 683, 976]. origin [34]. ornamental [402, 204]. ortholog [302]. orthologous [34]. Oryzias [914, 1345, 980, 825, 1312, 660, 1082, 818]. Oscar [934, 207, 1255]. osmolality [335, 479, 96, 837]. osmoregulation [969, 666, 1157, 809]. Osmoregulatory [981, 280, 1341, 962, 1152, 619, 426, 1136]. osmotic [486, 105]. ossification [1225]. ost [582]. ost-alpha [582]. osteoblast [1225]. osteocalcin [600]. osteological [1043]. OTC [625]. other [637, 381]. otoliths [10]. Oujiang [1260, 140]. out-season [1239]. output [351]. ova [98]. Ovarian [553, 707, 905, 894, 352, 795, 851, 1309, 488, 361, 509, 453, 389, 287, 1318, 549, 762, 695, 959, 392, 635, 538, 652, 1019]. ovaries [664, 556]. ovary [100, 190, 1330, 460, 1079, 897, 1011]. ovata [1284]. ovatus [1386, 1388, 632, 815]. overexpression [1127]. overfeeding [1116]. overland [541]. overload [1275, 185]. overload-induced [1275]. overview [1111]. ovulation [387]. ovulatory [814]. oxidant [530]. oxidases [862]. oxidation [1248, 862, 974]. Oxidative [568, 658, 1223, 1249, 532, 1325, 796, 758, 35, 1187, 1064, 1175, 599, 929, 86, 1250, 646, 1066, 752, 1182, 140, 1144, 1285, 1092, 595, 983, 889, 1115, 938, 867, 274, 672, 856, 442, 1065, 685, 783, 1281, 1225, 644, 975, 897, 888, 1252, 966, 286, 1321, 1053, 1059, 990]. oxide [928, 215, 577, 1314, 395, 1344]. oxidized [443, 376, 1084, 760]. Oxygen [119, 886, 272, 512, 73, 1191, 621, 550, 205, 1044, 1272, 133, 289, 1336, 9, 946, 1380, 124, 1122, 576]. oxygen-carrying [621, 1272]. oxygenase [1028]. oxygenase-1 [1028]. oxygeneios [812]. oxyrinchus [773]. oxytetracycline [625, 765].

P [154, 977, 266]. P-glycoprotein [977]. P-selectin [266]. P. [1154, 623, 971]. P450 [852, 1196, 497, 220]. P4501A [317]. P450arom [851]. P450c17 [100, 290]. P450c17-I [100, 290]. P450c17-II [290]. PACAP [864]. Pacific [609, 319, 75, 421, 195, 300, 39, 540, 683]. pacu [494, 662, 726]. paddlefish [709, 271]. Pagellus [111, 112]. Pagrus [651, 992, 1264, 1049, 989, 1107, 1294, 1101, 735, 1308, 1105, 1281, 567]. pain [225]. Pallas [4]. palm [532, 376, 1368]. palmitoyltransferase [811]. Pampus [110]. pancreas [99]. pancreatic [854, 469]. Pangasianodon [691, 981, 1220]. Pangasius [944, 598, 482]. paper [973, 162]. Para [220]. Para-nitrophenol [220]. Parachromis [640, 971]. Paraguay [519]. Paraguay-Paraná [519]. Paralabrax [3]. Paralichthys [894, 37, 798, 1009, 412, 283, 75, 978, 623, 932, 869, 1115, 1223, 843, 169, 916, 789, 778, 1080, 1096, 1340, 923, 1004, 1270, 876, 505, 525, 957]. parallelus [745]. paralog [613]. paralogues [680]. parameter [345]. parameters [827, 325, 1242, 1364, 8, 619, 606, 338, 489, 219, 480, 207, 76, 1203, 376, 322, 1293, 1024, 139, 549, 579, 154, 1195, 250, 456, 1216, 57, 777, 149, 292, 1213, 1140, 728, 915, 329, 759, 1263, 856, 1253, 1343, 1202, 516, 1051, 834, 1220, 321, 1215, 357, 878, 1116, 574, 1080, 408, 1348, 676, 1383]. Paramisgurnus [646, 1118, 135, 826, 1027]. Paraná [519]. parathion [43]. parents [1238]. parr [1067, 1073]. parrotfish [326]. Part [75]. Partial [596, 622, 469, 134, 81, 409, 1293, 563, 332, 1346, 1261]. partially [527]. participate [1087]. participates [861]. Parvalbumin [402]. paste [137]. pastoris [29]. Patagonian [503, 678]. paternal [85]. pathological [646]. pathway [1288, 798, 1154, 1148, 1301, 1018, 1378, 1053]. pathways [328, 1243, 843, 366, 1331]. pattern [905, 1017, 382, 738, 175, 722, 1309, 694, 688, 169, 1196, 506, 1217, 955, 1377]. patterns [326, 1180, 987, 1391, 102, 935, 1259, 1169, 302, 315, 525, 631]. pax3a [1009]. pax3b [1009]. Pcc [631]. Pcc-dmrt1s [631]. PCR [882, 459, 681, 1372, 840, 757]. Pdp11 [564]. peacock [398]. pearl [619]. pectinirostris [1083, 1177]. pectoral [972]. Pediococcus [1203]. pejerrey [503, 574]. pellucida [175, 1153]. Pelteobagrus [833, 996, 708, 1015, 849, 336, 1178, 1378, 603, 1172, 1331, 1254, 1370, 649, 1034, 1026]. pelvic [972]. pendimethalin [1250]. pendimethalin-induced [1250]. Pengze [631]. pentachlorophenol [802, 607]. pepsinogen [468]. pepsinogens [182]. pepsins [182]. PepT1 [356]. peptidase [896]. peptidase-4 [896]. peptidases [808]. Peptide [455, 544, 542, 896, 1018, 1090, 763]. Perca [170, 1044, 130, 578, 1189]. perch [170, 1131, 1044, 130, 578, 1189, 1040]. Percidae [571]. Perciformes [697, 1377]. Performance [1300, 1168, 827, 1114, 951, 45, 422, 580, 1208, 1134, 1175, 219, 944, 708, 906, 89, 207, 711, 637, 806, 872, 1084, 470, 1345, 322, 787, 752, 849, 263, 1224, 579, 1145, 1092, 1306, 1101, 241, 116, 736, 1353, 84, 57, 657, 1037, 776, 268, 292, 1365, 415, 994, 915, 868, 691, 759, 703, 926, 902, 1156, 1378, 1121, 671, 1057, 1343, 834, 715, 255, 921, 1141, 1170, 247, 1380, 997, 1058, 1163, 483, 1155, 1050, 1367, 339, 1370, 760, 51, 171, 1269]. performances [376, 427, 460, 603]. period [1242, 706, 880]. periodic [254, 299]. periods [588]. peripheral [126, 1220, 750]. periprandial [913, 832]. Péron [991]. peroxidase [532, 882, 1038, 1306]. peroxidation [258, 179, 61, 1146, 1190, 1320, 1091, 1050]. peroxisomal [862]. peroxisome [1050]. Persian [263, 1324, 969]. persicus [263, 1324, 969]. personality [667]. perturbations [947]. peru [609]. pesticide [192, 178]. pesticides [1014]. Petenia [193]. Petromyzon [1310]. PFOS [803]. PG [1239]. pH [1229, 464, 319, 78, 163, 1008, 295]. pharmaceutical [729]. Pharmacokinetics [642, 857, 765]. pharmacological [1271]. pharyngeal [1226]. phase [34, 600, 49, 452, 562]. phases [553]. phenanthroline [318]. phenol [851]. phenoloxidase [381]. Phenotype [614]. phenotypes [1099]. Phenotypic [1123]. phenoxyethanol [362]. phenyl [707]. phenylhydrazine [133]. pheromone [101, 433]. phosphatase [561, 169, 1095]. phosphatases [766, 505]. phosphate [833, 996, 447, 1278, 871, 870, 766]. phosphoenolpyruvate [1323]. phospholipid [796, 736, 841, 1150]. phosphorus [833, 464, 263, 895, 1163]. phosphorylase [136, 1210]. phosphorylated [824]. phosphorylation [752]. phosphotransfer [1199]. phosvitin [1018]. phosvitin-derived [1018]. photoperiod [1142, 462, 795, 186, 1239, 1040, 303]. photoperiods [498]. photopic [300, 1336]. photoreceptor [911, 1379]. photoreceptors [638]. photoresponse [911, 1379]. phthalate [829]. phylloquinone [761]. phylogenetic [1125, 355, 315]. phylogenic [976]. phylogeny [743, 639]. Physicochemical [58, 976]. Physio [1267, 1389]. physio-histological [1389]. Physio-metabolic [1267]. Physiological [166, 1152, 8, 992, 784, 372, 56, 847, 1362, 1282, 780, 383, 309, 498, 362, 1268, 168, 326, 211, 1322, 1355, 835, 602, 161, 167, 267, 1358, 1024, 936, 154, 205, 1123, 1388, 665, 777, 1188, 528, 1119, 720, 650, 1181, 414]. Physiology [484, 1061, 251, 570, 246, 787, 31, 593, 1057, 1189, 597, 803, 1155]. phytase [87]. phytoestrogens [93, 549]. phytoplanktivorous [19]. PI3K [1301]. Piaractus [662, 726]. piceus [119, 926, 1206]. Pichia [29]. picolinate [1216, 57]. pieces [55]. pigment [577]. pigmentation [638]. pike [1142, 1242, 1237, 880]. pikeperch [730, 571, 250, 1057, 498]. Pimephales [731, 829]. pin [1021, 979]. pink [567]. pinnifasciata [203]. pipefishes [91, 85]. piracanjuba [129]. pirarucu [451]. Pisces [640, 221]. Piscirickettsia [873, 678]. piscivorous [78]. Pistia [117]. Pituitary [615, 365, 1304, 451, 446, 583, 153, 860, 297, 856, 801, 1313, 1209, 397, 210, 1079]. pituitary-ovary [1079]. PKBs [1034]. plagiosum [1039, 404]. plaice [203]. planktivorous [808]. plant [1126, 608, 673, 891, 518, 721, 892, 247, 1348, 508]. plant-based [892, 247]. plantarum [944]. Plasma [439, 170, 1047, 930, 1110, 1229, 1160, 177, 219, 172, 369, 616, 805, 828, 1224, 1210, 367, 1069, 434, 1005, 1253, 1025, 775, 685, 203, 830, 1359, 479, 96]. plasmatic [186, 1140, 529]. plasticity [1123]. Plateau [1162, 147]. platensis [1320, 1383]. Platichthys [416, 81, 713, 1019]. platorynchus [92]. play [985]. Pleurotus [944]. ploidy [1118, 407]. pluvialis [253]. Poecilia [214, 204, 893]. pollutants [784]. polluted [438]. pollution [883, 330, 797]. polybrominated [954]. polychlorinated [954]. polyculture [198]. polymorphic [455]. polymorphism [202]. Polymorphisms [770, 864]. Polyodon [709, 271]. polyploid [1009, 53, 28]. Polyprion [812]. polyribocytidylic [499]. polyriboinosinic [499]. polysaccharide [1075, 345]. polysaccharides [1098, 296, 408]. polyunsaturated [312, 637, 918, 460]. pomace [1355]. pomfret [110]. pompano [1284, 1386, 1388, 632, 815]. pond [931, 981]. pond-raised [981]. populations [257, 174]. porgy [651, 390, 1183, 1377]. portable [1364]. Portugal [820]. Portulaca [1390]. positive [361]. possesses [778]. possible [794, 790, 697, 401, 1083, 158, 333]. post [462, 246, 637, 1012, 199, 496, 924, 538, 717, 124, 1078, 991]. post-exercise [246, 124, 1078]. post-larvae [462, 199, 496, 924]. Post-mortem [991]. post-prandial [1012]. post-smolt [637]. post-thaw [717]. post-vitellogenic [538]. postlarvae [912]. Postprandial [925, 575, 657, 763]. postsmolt [1387]. postsmolts [583]. posttranscriptional [794]. potamodromous [814]. Potamotrygonidae [180]. potassium [1229]. potencies [59]. Potential [1015, 833, 1074, 458, 78, 386, 1006, 1373, 228, 1338, 766, 757, 1327, 441]. potentials [554]. potentiates [337, 429]. Pou2 [103, 302]. poultry [532, 1293]. powder [1390, 1219, 184, 1343]. power [369, 351, 1138]. PPAR [736, 481, 1172, 1050]. PPAR-alpha [481]. practical [264]. prandial [1012]. Praziquantel [838]. PRDM1 [869]. Pre [835, 252, 396, 1029]. Pre-anaesthetic [252]. pre-differentiated [396]. pre-exposure [1029]. Pre-sedation [835]. preadipocytes [1041]. prebiotic [603]. precision [1075]. precision-cut [1075]. precursors [841]. predicted [1287]. preferences [1100]. preferred [159]. preferring [1006]. Preliminary [787, 1140, 534, 687]. Prenant [747]. prenanti [580, 747, 655, 900, 1394, 832, 1062, 760]. Preparation [1212]. preparations [358]. prepared [528]. preproghrelin [1040]. Presence [396, 467, 1218, 346]. present [224]. preservation [370, 618]. preservatives [1074]. pressure [790]. pressure-related [790]. prevent [1392]. Preventive [1059]. prevents [1325]. previtellogenesis [907, 949]. prey [1017, 808]. preys [78]. primary [1063, 287, 793, 1290, 624, 831, 1059]. primed [433]. primitive [94, 1102]. priorities [237]. Prl [994]. pro [903, 153, 636, 530]. pro- [903]. pro-opiomelanocortin [153]. Pro-oxidant [530]. Pro-Tex [636]. Probiotic [97, 207, 1203, 711, 198, 1353, 564]. Probiotics [617, 709, 872, 51]. procedure [828, 66]. process [1286, 925]. processing [573]. Prochilodus [593, 279, 717, 576]. produce [433]. produced [967, 1118, 87]. producing [1090, 1201]. product [1293, 332, 573]. Production [101, 507, 700, 20, 411, 387, 224, 241, 1314, 496, 228, 1245, 1244, 941]. profile [439, 212, 950, 908, 992, 380, 805, 1366, 48, 1107, 526, 626, 54, 97, 648, 289, 1141, 407, 1080, 171]. Profiles [670, 519, 501, 141, 906, 531, 668, 701, 1012, 277, 869, 1243, 1310, 1185, 47, 41, 366, 979, 461, 919, 1181, 1276, 829, 876]. profiling [1197, 1013, 764, 1338, 1324, 816, 565]. progeny [1238, 98]. progesterone [485]. Progestin [1120, 775]. progression [157]. prohibitin [1083]. projections [1305]. prolactin [860, 416, 495, 873, 30]. proliferation [1126, 535, 865]. proliferator [1050]. proliferator-activated [1050]. proline [1381]. prolonged [1267, 320]. promelas [731, 829]. Promoter [1171, 905, 125, 52, 953, 46, 1266, 490, 932, 1013, 558, 269]. promotes [564, 997, 1109]. promoting [1070]. proomii [872]. properties [608, 624, 104, 1205, 1164, 370, 976, 692, 778]. propionate [968]. Propiscin [1189]. propofol [739]. propolis [219]. proportion [929]. proportions [250]. prostaglandin [643]. Prostanoids [545]. Prostanoids-induced [545]. protandrous [430, 390]. Protease [653, 791, 271]. proteases [596, 622, 971]. protected [527]. protecting [983, 1070]. Protection [685, 1390, 758, 408]. Protective [1075, 149, 504, 1236, 1175, 244, 274, 1065]. protects [1275]. Protein [1073, 264, 12, 113, 514, 1219, 245, 602, 2, 1208, 520, 950, 382, 677, 913, 786, 489, 858, 270, 842, 599, 301, 317, 1031, 920, 77, 729, 673, 1107, 1294, 1387, 584, 223, 283, 747, 1002, 802, 776, 824, 1013, 626, 895, 445, 697, 1129, 500, 669, 364, 1378, 666, 1046, 107, 1327, 191, 995, 1109, 1348, 355, 315, 436, 213, 525, 558, 508]. protein-1 [913]. protein-2 [677, 525]. protein-3 [382]. Protein-sparing [264]. protein/carbohydrate [858]. proteinases [163]. proteins [1284, 925, 1288, 172, 903, 722, 102, 1015, 152, 904, 318, 221, 807, 1262, 1165, 1051, 1030, 573, 1236, 1303, 425, 1235]. proteoglycans [785]. proteolysis [295]. Proteolytic [1008, 970, 77, 78, 640]. proteome [1023, 120, 1393, 1088, 1025]. Proteomic [812, 764, 565, 1270, 904, 221, 499]. proteomics [190]. protexin [207]. proto [1094]. proto-oncogene [1094]. protocols [800, 1234, 640]. protogynous [326, 821, 959, 392]. proton [468]. provides [1316]. proximate [250, 84]. PRR13 [127]. Prussian [696]. Prx1 [876]. Psammoperca [427]. pseudo [1099]. pseudo-hermaphroditic [1099]. pseudoharengus [249]. pseudomale [755]. Pseudosciaena [499, 941, 1053, 990, 1029]. Pt5 [1018]. PTEN [287]. Pterois [624]. Pterygoplichthys [403, 414]. PUFA [473, 588, 1179]. puffer [657, 1252, 1143]. pufferfish [742, 786, 1054, 1134, 1175, 557, 1272]. pulp [973]. pulse [1273]. pump [468]. pumpkin [824]. pumpkinseed [1384]. punch [952]. punctata [59]. punctatum [167]. punctatus [930, 215, 577, 258, 84, 66, 710, 704, 60, 261, 823, 355, 315, 345]. puntazzo [99]. Puntius [1068, 985, 1354]. Purification [447, 63, 143, 104, 871, 870, 293, 1016, 563, 403, 182, 75, 191]. purified [1205]. Purinergic [359, 1289]. purpurea [310]. purslane [1390]. putatively [554]. putitora [512, 775, 964]. putrefaciens [564]. pyloric [293, 134]. pyloric-duodenal [134]. pylzovi [1311]. pyrene [1125]. pyrithione [1360, 977]. Pyruvate [1391, 1323]. pyruvic [772]. PYYb [542].

qingbo [703, 339]. Quality [370, 489, 331, 411, 628, 812, 257, 744, 1239, 1253, 666, 162, 1235, 837]. Quantification [85, 844, 144]. quantitative [1372, 840, 645, 757, 1235]. quantum [305]. quelen [288, 835, 967, 464, 739, 910, 824, 867, 566, 856, 528, 1119, 1215, 597, 712]. Quoy [824].

RA [737]. Rachycentron [432, 1323]. radiation [689]. radical [804]. radical-induced [804]. Rafinesque [355, 315]. raggedtooth [189]. Rainbow [1188, 1227, 544, 1023, 212, 435, 245, 1267, 925, 1126, 1382, 219, 690, 561, 722, 588, 328, 613, 741, 711, 322, 1066, 1132, 1012, 1145, 1306, 424, 695, 626, 744, 1243, 1239, 582, 1161, 1088, 759, 1263, 144, 1212, 684, 310, 97, 1190, 721, 1211, 698, 1121, 834, 253, 715, 255, 1320, 69, 644, 975, 751, 772, 127, 892, 888, 208, 845, 943]. raised [981]. Raji [668, 701]. rakers [1060]. range [1008]. rank [33, 536, 477]. ransonnetii [697]. rapamycin [752]. rapeseed [690, 1268]. Rapid [1237, 216]. rare [746, 153, 839, 1193, 276, 607]. rarus [746, 153, 802, 839, 1193, 276, 607]. Rastrelliger [1245, 1244]. rate [782, 1127, 853, 1001, 437]. rates [585, 216, 946, 887, 1078]. ratio [657, 199, 436]. ration [113]. ratios [858, 963]. rays [342]. rbp4 [737]. re [926]. re-feeding [926]. Reaction [327, 993]. reactive [73, 576]. readiness [319]. real [459, 681, 840, 757]. real-time [459, 681, 840, 757]. reared [462, 273, 1388, 777, 601, 282, 553]. Rearing [520, 806, 994, 684]. receptor [620, 1063, 1274, 422, 885, 501, 722, 922, 963, 187, 674, 844, 27, 1287, 643, 450, 1271, 781, 1262, 478, 634, 860, 418, 394, 427, 704, 420, 743, 1313, 645, 679, 1082, 815, 30, 836]. receptor-related [422]. receptors [1186, 501, 706, 1120, 183, 389, 26, 1397, 444, 1386, 416, 873, 359, 1206, 1004, 1050]. recirculating [55, 719]. recirculation [282]. recombinant [1197, 807, 1328, 1212, 507, 29]. recordings [1305]. recovery [136, 599, 240, 1115, 38, 1119, 1344, 1273]. recruitment [352]. Red [1221, 1392, 1219, 651, 1061, 1135, 128, 992, 1264, 1074, 609, 1049, 989, 48, 1107, 1294, 1123, 1101, 518, 323, 1248, 735, 1308, 1105, 1190, 1281, 757, 816, 1163, 28, 1011, 1089]. red-spotted [323]. Redox [289, 555, 1077]. redox-active [555]. reduce [458]. reduced [1085, 102, 585, 666]. reduces [613, 1145, 980, 893]. reducing [605, 566]. reductase [871, 870, 454, 378]. reductases [761]. Reevaluation [457]. refeeding [951, 651, 380, 356, 320, 1321]. Reference [681, 882, 459, 120, 1372]. references [840, 757]. referring [1109]. refining [121]. reflect [159]. reflex [927]. Regan [268, 1201]. regarding [488]. regenerating [1039, 893]. regeneration [471, 1316]. regimes [37, 111, 590]. region [46, 611, 394, 28]. regional [208]. regionally [351]. regions [864]. regius [1160, 331, 994, 820, 475, 508]. regulate [1284]. regulated [996, 862, 316, 1166, 1151, 1315]. regulates [934, 860]. regulating [1288, 588, 356, 669, 333]. Regulation [659, 1365, 602, 388, 748, 128, 903, 1339, 1297, 1039, 557, 852, 1272, 1323, 1386, 524, 490, 861, 1243, 1006, 460, 1196, 1035, 1020, 49, 1046, 655, 900, 1179, 737, 1371, 769, 1172, 1004, 525, 1315, 1026]. regulator [191]. Regulatory [865, 200, 1392, 794, 1243, 935, 394, 359, 1179]. reidi [799]. related [620, 73, 1159, 1352, 422, 1017, 1280, 1003, 1339, 459, 342, 859, 1015, 790, 793, 1184, 306, 781, 149, 744, 1146, 1022, 1112, 1381, 775, 638, 1215, 763, 1033, 1102, 208, 1261, 1370, 1251]. relation [619, 795, 411, 959, 530, 536, 477, 229, 1235]. relationship [10, 782, 756]. Relationships [181, 842, 1379]. Relative [59, 619]. release [544]. releasing [1304, 80, 694, 1166, 476, 655]. relevant [1228, 1209]. remarkable [641]. remodeling [693]. remodelling [133]. renal [90]. renaturation [534]. renin [909]. renin-containing [909]. reovirus [461, 939]. repair [658, 1030]. reparative [1059]. repeat [703]. Repeatability [1345]. repeated [45]. replace [1327]. replaced [1208, 473]. Replacement [1091, 532, 846, 1293, 517, 1294, 776, 332, 1346, 1261]. replacer [995]. replacing [1052, 263, 313]. reproduction [930, 1110, 1197, 491, 1287, 980, 386, 524, 31, 692, 1232]. Reproductive [331, 109, 439, 730, 1142, 423, 251, 1120, 446, 76, 922, 628, 290, 268, 242, 217, 732, 817, 418, 515, 297, 635, 460, 693, 398, 498, 1215, 407, 485, 1307, 171, 553]. reproductive-related [1215]. requirement [113, 325, 436]. requirements [54, 12]. rerio [758, 636, 52, 687, 725, 120, 190, 356, 318, 306, 1212, 1194, 898, 968, 185, 1030, 266, 763, 954, 966, 1235, 1011]. RESC [375]. research [1111]. reshaping [1087]. resistance [1264, 944, 1107, 1294, 56, 891, 915, 759, 1281, 1296]. resistant [614, 561]. resources [1233]. respiration [999, 437]. Respiratory [1122, 1126, 1218, 888]. responding [786, 958]. Response [113, 102, 426, 283, 768, 1200, 1322, 280, 45, 136, 1267, 1300, 1135, 883, 1311, 542, 1054, 738, 489, 992, 706, 987, 480, 616, 267, 486, 1098, 459, 446, 741, 1221, 784, 850, 963, 1293, 160, 372, 184, 956, 26, 1045, 1397, 1093, 627, 412, 252, 205, 336, 1123, 555, 19, 1086, 1353, 1184, 1037, 1222, 665, 199, 496, 415, 582, 1292, 759, 684, 879, 55, 719, 801, 1190, 661, 178, 756, 968, 448, 1347, 834, 255, 666, 921, 1225, 49, 712, 481, 210, 106, 461, 939, 1319, 200, 333, 1206, 168, 955, 718, 1122, 51]. Responses [87, 179, 1211, 689, 1192, 211, 543, 340, 1289, 1355, 835, 188, 8, 885, 913, 858, 617, 218, 275, 546, 739, 1028, 1333, 523, 170, 1358, 579, 1137, 1092, 1362, 1306, 584, 264, 736, 657, 623, 1388, 1115, 1223, 1282, 560, 895, 1188, 530, 465, 797, 671, 383, 309, 965, 261, 362, 1108, 1258, 197, 644, 751, 1157, 575, 650, 1368, 333, 832, 1268, 943, 876, 1011, 1331, 990, 1034, 414]. responsible [361]. responsive [34]. responsiveness [1278, 1125]. resting [853, 1001, 216, 1078]. restriction [1342]. result [773]. Resveratrol [1130, 980]. retardant [487]. retention [637, 436]. reticulata [214]. reticulum [1331]. retina [1366]. Retinal [540, 1305, 1277]. retinoid [815]. reveal [1371]. revealed [641]. reveals [1395, 1338, 772]. reversal [326, 945, 1042, 845]. review [354, 131, 25, 431, 400, 398]. Revisiting [555]. Rhamdia [288, 835, 967, 464, 739, 910, 824, 867, 566, 856, 528, 1119, 1215, 597, 712]. rhamnose [531, 668, 701, 46]. rhamnose-binding [531, 668, 701, 46]. rhamnosus [992, 97]. Rhinoptera [342]. Rhizoma [1176]. Rhizoprionodon [800]. rhodanese [58]. Rhodeus [902]. rhodomyrtone [1161]. Rhodomyrtus [1161]. Rhodotorula [1286]. rhombus [411]. rhythm [1117]. rhythmicity [1185]. rhythms [434, 452]. ribeirae [1279]. riboflavin [6]. ribosomal [626]. ribs [259]. rice [1208, 681, 1198, 182, 171]. ricefield [821, 959, 22]. rich [1299, 180, 486, 277, 864, 1141, 890]. Richardson [119, 529]. ringens [597, 712]. Risk [238]. risks [1240]. Rita [764]. River [106, 439, 70, 557, 156, 519, 1148, 279, 1348]. rivoliana [972, 1106, 1302]. RNA [798, 1222, 1308, 1340, 436]. RNA-seq [1340]. RNA/DNA [436]. Roach [76, 1117, 671]. rock [243, 1379]. rockfish [641, 418, 515]. rogercresseyi [850]. rohita [114, 90, 338, 502, 1031, 546, 375, 1358, 1334, 50, 223, 307, 337, 428, 429, 647, 478, 148, 1350, 545, 1156, 1198, 41, 1167, 247]. rohu [375, 647, 41, 247]. Role [577, 808, 1033, 1286, 656, 388, 1064, 1280, 94, 1111, 395, 1103, 36, 374, 985, 1006, 274, 1021, 1065, 667, 401, 900, 1083, 1179, 158, 1236]. Roles [758, 843, 865, 333, 1385, 1396]. Romanza [55, 719]. Ror1 [1274]. ROS [689, 941, 990]. rosacea [126, 604, 633]. rose [610, 935, 902]. rosemary [787]. Rosmarinus [787]. rossii [1357, 529, 797]. rotenone [340]. roughy [899]. Roundup [1204]. route [1341]. routine [585, 216]. RT [882, 681, 757]. RT-PCR [882, 681, 757]. rubber [1052, 1031]. rubripes [1363, 1217, 1143]. run [985, 367]. run-off [985]. runoff [981]. rusty [326]. ruthenus [907, 949, 365, 999, 1382, 877, 1084, 1102, 1236]. rutilus [671, 1117, 76, 915, 671]. rutin [856]. RXR [815].

S [981, 201]. S. [83]. sabina [621, 342]. sablefish [1092, 848]. Saccharomyces [350, 1383]. safeguard [239]. Safeguarding [233]. safety [838]. sage [715]. sailfin [403]. sakei [633]. Salado [279]. salar [1063, 45, 1341, 467, 35, 1067, 240, 637, 850, 785, 673, 160, 583, 822, 1073, 108, 241, 62, 454, 895, 304, 495, 206, 1021, 947, 83, 901, 979, 71]. salina [543]. salinities [314, 619, 1037, 1388, 924, 994, 691, 1147, 809]. Salinity [417, 956, 557, 969, 1168, 962, 992, 1339, 312, 372, 904, 847, 277, 1092, 1362, 426, 562, 1104, 694, 1335, 529, 1166, 194, 981, 197, 666, 1157, 335, 210, 1319, 1252, 1143, 958, 1053, 479, 96]. salinity-induced [197]. Salminus [814, 615]. Salmo [1063, 45, 1341, 467, 35, 1067, 240, 637, 850, 785, 673, 433, 160, 413, 583, 822, 1073, 1234, 108, 241, 62, 862, 454, 895, 304, 495, 1257, 206, 1005, 1021, 947, 783, 83, 901, 979, 71]. salmoide [443]. salmoides [550, 768, 1361]. salmon [1173, 544, 1063, 45, 1341, 467, 35, 1067, 614, 240, 267, 637, 484, 785, 673, 470, 160, 152, 583, 822, 782, 1387, 63, 1073, 319, 1127, 108, 241, 62, 104, 562, 156, 895, 1328, 304, 495, 206, 873, 1021, 947, 401, 121, 83, 901, 979, 200, 71]. salmonid [1234]. salmonis [873, 678]. salt [1221]. salt-dependent [1221]. Salvelinus [656, 454, 159]. sampling [240, 952, 184, 260, 121, 1359]. sand [3]. Sander [571, 250, 5, 457, 1057, 498]. sanguineus [213]. Santa [279]. sapidissima [1010, 1037]. sapphire [707]. sarafloxacin [857]. sarba [265]. Sarda [47]. Sardinella [143]. sardinelle [143]. sargus [275, 406]. sativa [517]. Sauvage [944, 598]. saves [1109]. saxatilis [132, 317, 573]. sazans [1070]. sbGnRH [1245, 1244]. scale [471]. scales [259]. scaling [853, 1001]. Scaphirhynchus [92]. Scaridae [326]. Scarus [326]. scat [1395, 123]. Scatophagus [1395, 123]. SCD [777]. scFOS [806]. Schizopygopsis [1311]. schizothoracin [747]. schizothoracine [1311, 1162]. Schizothorax [580, 747, 655, 900, 1394, 832, 1062, 760]. Schlegel [243, 980, 283, 540]. schlegeli [641, 418, 515, 390]. schlegelii [74, 1377]. schrenckii [262, 688, 1112]. schwanenfeldii [974]. Sciaenidae [221]. Sciaenops [128, 1221, 1123, 757, 816]. Scomber [911, 7, 297]. scombroid [47]. scope [1078]. Scophthalmus [514, 912, 17, 885, 1291, 1120, 1098, 1332, 1374, 165, 411, 963, 628, 674, 844, 1318, 1353, 767, 1292, 497, 281, 481, 122, 511, 958, 998, 740]. scotopic [195]. sCP [544]. Scp3 [959]. Screening [349, 333, 1359]. Scyliorhinus [425]. Sea [203, 620, 1219, 602, 962, 846, 1187, 1299, 883, 748, 858, 270, 992, 1264, 15, 1071, 89, 275, 406, 473, 616, 409, 101, 471, 806, 872, 376, 1049, 989, 1293, 1107, 1294, 74, 99, 1101, 241, 130, 526, 295, 1310, 1140, 860, 924, 82, 733, 445, 611, 1105, 449, 527, 629, 442, 350, 448, 1025, 38, 940, 1281, 565, 1141, 1327, 720, 168, 452, 186]. seabass [1149, 1287, 1226, 648, 629]. seabream [1214, 1014, 1283, 371, 111, 787, 27, 735, 1213, 217, 1308, 134, 719, 635, 536, 477, 1042, 112, 359, 553, 1383]. seahorse [1047, 799, 1043, 601]. seamounts [899]. seaperch [427]. season [628, 1239, 498, 1245, 1244, 775, 574]. Seasonal [491, 221, 1042, 167, 369, 762, 933, 982, 743, 981, 452]. Seasonality [930]. seasons [211, 189]. Seawater [83, 280, 471, 319]. seawater- [471]. seaweed [1061]. Sebastes [641, 418, 515]. second [30]. secondary [583, 1258, 390]. secretion [992, 446]. secretions [329, 1350]. secretory [49]. section [106]. sedated [739]. sedation [835, 252]. seed [1052, 387, 824]. segregation [1305]. selected [1286]. selectin [266]. Selection [1372, 459, 681]. selective [220]. selenium [443, 1264, 646, 753, 849, 1038, 1111, 1306, 569, 282, 1202, 727]. selenoprotein [758, 987]. self [1304, 434]. self-feeding [434]. self-fertilizing [1304]. SEM [973]. SEM- [973]. Semen [42, 130]. semilaevis [642, 100, 450, 1104, 1013, 864, 1035, 755, 118, 1082, 1181, 138, 770, 363]. seminal [1229, 172, 682, 877, 1253, 685]. Senegal [438]. senegalensis [422, 520, 177, 489, 372, 199, 496, 438, 434, 282, 1170]. Senegalese [422, 520, 950, 405, 177, 489, 372, 564, 496, 434, 282, 1170]. sensibility [967]. sensitive [340, 142, 1149, 555, 1020, 702]. sensitiveness [1185]. sensitivity [1014, 696, 1379, 7, 300, 1164, 772]. sensors [1272]. seq [1340]. Sequence [201, 1095, 737, 34, 187, 506]. Sequences [869]. sequencing [734, 743]. Seriola [636, 161, 838, 972, 1106, 1302, 1276]. Serjania [1326]. serotonergic [1070, 667]. serotonin [1352, 510, 920, 1131]. serotonin-modulating [920]. SERPINH1 [888]. Serum [184, 943, 338, 581, 64, 192, 1293, 343, 139, 262, 817, 460, 1202, 4, 43]. set [239]. several [376, 447, 971]. severus [543]. sewage [797]. Sex [73, 326, 439, 930, 894, 945, 285, 502, 369, 411, 1395, 430, 1047, 25, 848, 641, 392, 332, 1257, 519, 427, 743, 1042, 574, 813, 1102, 390, 845, 1217, 998, 21, 349]. sex-biased [1395]. sex-dependent [1257]. sex-dimorphic [641]. Sex-related [73, 1102]. sex-specific [349]. sexes [146, 769]. Sexual [326, 945, 1015, 1178, 563, 31, 157, 401, 901]. Sexually [92, 413, 1042, 923]. shad [1010, 1037]. shanny [341]. share [600]. shark [167, 141, 351, 598, 800, 773]. sharks [1191, 189, 773, 991]. sharpeyi [817]. sharpsnout [99]. sheepshead [1285, 881]. shellfish [617]. Shewanella [564]. shi [462]. shift [1341]. shishamo [531]. shock [1284, 599, 301, 1015, 223, 747, 802, 445, 669, 1051, 511, 213, 558]. Short [1145, 338, 1139, 1074, 209, 806, 372, 152, 1146, 1228, 1043, 516, 1245, 1244]. short- [516]. short-chain [1074, 806]. short-lived [1146]. short-snouted [1043]. Short-term [1145, 338, 1139, 209, 372, 152, 1228]. shortfin [773]. shortnose [585, 367, 1335, 1078, 96]. shovelnose [92]. shrimp [198]. sialidase [660]. Siamese [268, 1201]. Siberian [951, 1355, 399, 1224, 1048, 1151]. side [412, 1119]. Siganus [1179, 479]. sign [119, 1084, 1210, 623, 1370]. signal [687, 1154, 27, 316, 669]. signal-regulated [316]. signal-regulating [669]. signaling [1289, 1288, 328, 1148, 793, 1301, 862, 400, 843, 1378, 893, 366, 191, 1109, 1053]. signalling [756]. significance [602, 753]. silencing [700]. silico [394, 704, 954]. silicon [305]. silicon-based [305]. Siluriformes [336, 79, 147]. Silurus [187, 25, 652, 124, 347]. silver [340, 73, 288, 15, 857, 464, 910, 36, 824, 728, 1373, 31, 157, 566, 856, 1253, 1119, 1258, 1249, 597, 712, 110, 1389]. silvering [1165]. silverside [312]. silvery [1183]. silvery-black [1183]. Silymarin [1097, 212, 997]. similarly [1341]. sinensis [175, 1084, 643, 1248, 703, 803, 302, 1081, 339, 890]. Single [633, 800, 765, 880, 1273]. single-pulse [1273]. Siniperca [1131, 1375, 919, 468]. Sirtuin-1 [1372]. sisorid [147]. Sisoridae [79, 147]. site [76, 770]. situ [90, 884, 1313]. six [735]. size [1283, 616, 248, 246, 627, 437]. sized [828]. sizes [1224]. skates [485]. skeletal [272, 196, 459, 613, 811, 1337, 1073, 40, 1225, 678, 95]. skeletogenesis [448]. skin [8, 1260, 323, 329, 592, 144, 638, 1096]. skull [972]. Slaughter [467, 1267]. SLC2 [1165]. Slc34a2 [996, 833]. SLC5 [1165]. slices [1075]. slow [1171]. slow/cardiac [1171]. slow/cardiac-type [1171]. Smads [328]. small [798]. SMAP [920]. smelt [561, 531, 847]. smolt [637, 1073, 241, 83]. smoltification [484]. smolts [1341, 1067]. smooth [450, 1104, 864, 755, 118, 1181, 138, 770, 363]. smtB [966]. SMX [168]. snakehead [222, 215, 577, 918, 46, 1150, 988, 59, 813, 995]. snapper [609, 48, 610, 935, 567, 213]. snook [256, 193]. snout [1114, 1207, 608, 953, 1028, 1333, 1278, 854, 264, 765, 736, 1086, 1184, 1301, 1271, 221, 983, 601, 1051, 887, 28, 718, 1050, 1130, 1296, 1251, 513]. snouted [1043]. soaking [247]. sobaity [1293]. Social [33, 819]. sockeye [319]. SOD [700]. sodium [1229, 833, 996, 454, 1237, 968, 1269]. sodium-dependent [833, 996]. sole [422, 520, 950, 642, 177, 489, 372, 450, 1104, 564, 1013, 496, 749, 438, 434, 282, 864, 755, 1170, 118, 1082, 1181, 138, 770, 363]. Solea [422, 520, 177, 489, 372, 199, 496, 749, 438, 434, 282, 1170]. soles [405]. solids [16]. soluble [373]. solution [184, 260]. somatic [196, 1178, 392, 878]. somatolactin [1193]. somatolactin- [1193]. somatostatin [987]. Some [1137, 211, 455, 1159, 422, 489, 1369, 373, 139, 579, 891, 1008, 871, 870, 1205, 292, 1213, 137, 4, 346, 1033, 345]. sonic [402, 221]. sophore [1068, 985, 1354]. sound [1373]. source [1300, 89, 588, 77, 1327]. sources [673, 1306, 283, 1202]. South [956, 141, 75]. southern [942, 124, 347, 187, 25, 652]. Sox10 [506]. Sox2 [932]. sox9 [399]. Sox9a [14]. soy [1098, 549, 776, 895, 1129, 1109]. soya [1294]. Soybean [1367, 1038, 263, 1144, 1022, 1156, 1091, 313]. sp [1074, 518, 1327]. sp. [441]. Spain [438]. Sparidae [1362]. Sparidentex [1293, 1183]. sparing [264]. Sparus [1214, 1014, 620, 1283, 602, 945, 1061, 883, 371, 89, 11, 186, 806, 787, 130, 265, 1317, 295, 217, 1140, 860, 924, 134, 719, 1022, 449, 635, 536, 477, 1042, 527, 1166, 442, 1025, 940, 359, 168, 553, 1383]. spathula [709, 271]. Spatial [145, 756]. spatiotemporal [1243, 468]. spatula [784]. spawning [35, 1242, 1186, 352, 331, 367, 817, 297, 780, 519, 427, 775, 820]. spawns [820]. special [504]. Species [584, 73, 1279, 1186, 967, 141, 1138, 724, 466, 800, 1337, 1234, 624, 909, 555, 853, 1001, 938, 982, 1174, 407, 1102, 1377, 576]. Species- [584]. speciesism [225]. specific [272, 136, 602, 8, 388, 285, 1134, 1067, 258, 746, 1031, 1066, 103, 752, 1092, 584, 1353, 437, 1162, 633, 823, 705, 644, 127, 979, 919, 135, 1163, 943, 836, 631, 1029, 349]. specification [1171]. specificity [967]. spectra [638, 1277]. Spectral [696, 1379, 7, 1117, 318, 300]. Spectrofluorimetric [43]. spectrometry [896, 565]. spectrometry-assisted [896]. speed [1143]. speeds [159, 946]. Sperm [365, 411, 433, 1239, 653, 1242, 1382, 1230, 877, 1231, 1234, 1241, 578, 735, 1228, 1006, 654, 717, 1236, 837]. Spermatogenesis [775, 369, 1178, 982, 1083, 1177]. Spermatogonial [823]. spermatozoa [620, 656, 1369, 1324, 1253, 685]. spermatozoon [970]. spermiation [365]. spermiogenesis [1087]. Spexin [1181, 832]. Sphyrna [650]. spills [415]. spinal [206]. spined [1342, 884, 53, 174]. Spinibarbus [703, 803, 339]. spirulina [1383, 1320]. spleen [752, 574]. splendens [93, 268, 899, 1201]. splendida [193]. spongy [1121]. spot [619, 723]. spotted [3, 930, 189, 1366, 1395, 1149, 323, 202, 444, 377, 293, 410, 807, 563, 610, 935, 400, 590, 59, 60, 261, 123, 16, 679]. spp [1105]. spring [985]. squid [1101]. SRF [789]. stability [532, 1308, 158]. stage [285, 722, 103, 812, 241, 1217]. stage-specific [285, 103]. stages [423, 1291, 486, 1125, 217, 61, 924, 964, 783, 281, 901, 1102]. staining [1194]. starch [270, 89, 50, 1184]. starry [416, 713, 1019]. start [902]. starvation [951, 8, 142, 152, 926, 902, 320, 1347, 875, 763, 1321, 1143]. starved [245]. starving [211]. STAT [1364]. status [348, 796, 1061, 1187, 1264, 1071, 240, 581, 459, 929, 711, 1084, 1066, 1038, 337, 429, 265, 594, 1248, 84, 1150, 776, 867, 157, 648, 282, 981, 997, 1269]. Steindachner [610]. stellate [570]. stellatus [570, 416, 4, 713, 1019]. stem [103, 825, 823]. steps [239]. Sterlet [907, 949, 1084, 999, 1382, 682, 877, 970, 1102, 1236]. steroid [439, 285, 411, 817, 427, 845]. steroidogenesis [548, 937]. steroidogenic [1042]. steroids [930, 502, 369, 1047, 814, 332, 519]. stickleback [1342, 884, 174]. stimulated [318]. stimulating [844, 418, 769, 29]. Stimulation [1041, 554, 333]. stimulations [876]. Stimulatory [1168, 978, 638]. stimuli [719]. stimulus [160, 756, 383, 309]. stinging [1298]. stingray [621]. stingrays [180, 342]. stock [109]. Stocking [678, 489, 262, 1088, 1112, 629]. stomach [988, 749, 182]. storage [1230, 877, 724, 131, 1308, 1021, 940, 121, 618]. strain [711, 298, 1353, 702]. strains [1341, 1098, 566, 1268]. Strait [1387]. strand [750]. strategies [1224, 241, 407]. strategy [121]. Streptococcus [805, 891, 661, 1296]. Stress [352, 850, 647, 541, 304, 712, 1168, 1322, 568, 620, 734, 280, 1325, 45, 245, 758, 1187, 883, 606, 1064, 1134, 1175, 1139, 489, 599, 992, 1264, 608, 1339, 906, 616, 1297, 1098, 446, 86, 741, 1028, 1333, 1250, 984, 646, 1024, 800, 105, 343, 956, 1107, 1294, 1182, 1012, 56, 140, 1093, 1285, 936, 412, 252, 1092, 223, 307, 428, 1314, 1210, 623, 986, 983, 889, 1068, 1354, 1115, 1223, 496, 938, 915, 691, 688, 274, 672, 566, 1190, 840, 629, 442, 498, 448, 1065, 1025, 1051, 685, 783, 886, 1277, 1281, 1249, 255, 1225, 644, 975, 1157]. stress [106, 897, 888, 1252, 200, 991, 286, 1303, 1321, 818, 168, 1359, 943, 71, 941, 958, 1053, 718, 1331, 1383]. stress- [1339]. Stress-induced [352, 446]. stressed [330, 669]. stresses [919]. stressor [800]. stressors [361]. striata [918, 733, 995]. striatus [700, 840]. striped [132, 317, 918, 981, 1220, 995, 573]. strongly [892]. Structural [663, 664, 842, 486, 334, 379, 462, 912, 364]. Structure [948, 794, 165, 716, 1076, 1125, 202, 1386, 731, 843, 1220, 138, 686, 347]. structures [8, 1290, 624, 1116]. Studies [271, 25, 326, 166, 1173, 156, 1140, 534, 59, 298]. Study [146, 1072, 32, 384, 462, 912, 222, 237, 587, 114, 369, 687, 192, 838, 929, 791, 76, 546, 1066, 260, 139, 549, 1334, 571, 164, 1076, 273, 130, 1248, 524, 117, 982, 799, 519, 1313, 1025, 43, 1106, 359, 1157, 210, 1096, 294]. studying [686]. sturgeon [907, 949, 653, 92, 951, 1355, 399, 999, 175, 570, 929, 1084, 373, 1066, 1238, 263, 1224, 56, 549, 205, 585, 262, 292, 367, 1070, 688, 1335, 1324, 1112, 362, 969, 538, 1048, 1321, 302, 1081, 65, 1078, 1151, 96]. styles [945]. sub [435, 784, 594, 727, 1344, 357, 809]. sub-adult [594]. Sub-Antarctic [809]. sub-chronic [727]. sub-lethal [435, 784, 1344, 357]. subchronic [560]. subfamily [344]. subjected [280, 758, 741, 1224, 498]. Subjective [1231].

Sublethal [973, 731, 338, 209, 11, 523, 598, 1223, 691, 898]. submitted [951, 1187]. suboptimal [947]. subsequent [241, 1115]. substituting [268]. Substitution [1156, 422, 1107, 283]. subtype [643]. subtypes [679]. subunit [125, 451, 1262, 958, 836]. subunits [748, 128, 297, 24]. sucker [1103, 1091]. sudden [372, 335]. suh [1148]. suitability [952]. suitable [459]. sulfamethoxazole [168]. sulfamonomethoxine [642]. sulfate [34]. sulphate [1125]. Sulphated [785]. summer [894]. sunbleak [248]. sunfish [906]. sunflower [250]. superoxide [700]. supplement [1208]. Supplementation [523, 1141, 1109, 212, 443, 1264, 690, 605, 1221, 1049, 989, 1084, 1107, 1294, 1144, 1334, 1216, 564, 776, 1068, 938, 895, 449, 1198, 282, 253, 255, 1048, 997, 1058, 1050, 1130, 1007, 1296, 508]. supplementations [1024]. supplemented [543, 1219, 1325, 992, 198, 579, 518, 137, 895, 715]. supplementing [263]. suppresses [1394]. suppression [446]. suratensis [619]. surface [621, 342, 948]. surgery [470]. surgically [665, 1057]. surrounding [979]. survey [1003]. Survival [392, 479, 1061, 98, 1037, 1363, 268, 292, 1070, 282, 1057, 324, 1252]. survival-promoting [1070]. surviving [985]. susceptibility [854, 323, 801, 185, 303, 1096]. susceptible [614]. suspended [16]. sustained [160, 974]. sutchi [482]. Svalbard [341]. SW [438]. swamp [353, 634]. Swim [567, 1381]. swimbladder [1165]. Swimming [31, 161, 580, 999, 240, 1345, 160, 154, 1145, 703, 926, 902, 1335, 159, 946, 158, 1380, 803, 1143, 1081, 339, 974]. switch [390]. syndrome [324]. Synechogobius [908, 533]. Syngnathus [91, 85]. synthase [1314, 395, 1210]. synthase/nitric [395]. synthesis [2, 1208, 561, 1194, 1058]. synthetase [1149, 702]. system [371, 258, 682, 877, 791, 375, 152, 752, 198, 1241, 808, 395, 1312, 1262, 421, 1070, 868, 569, 55, 519, 282, 350, 1245, 1244, 253, 886, 305, 208, 830, 505, 347, 239]. systematic [1003]. systemic [1264, 1213]. systems [625, 627, 1073, 224, 931, 158].

T [1167, 920, 129, 461]. T-bet [461]. T-maze [920]. taenia [53]. tags [1057]. Takifugu [742, 786, 1054, 1134, 1175, 557, 1272, 657, 1363, 1319, 1252, 1217, 1143]. Talking [991]. tambaqui [927]. tandem [565]. Tanichthys [20, 284, 18]. tannin [117]. target [413, 752, 560, 481]. targeted [789]. tarichi [535, 1205, 1136]. taurine [1168, 1049, 989, 776, 483]. taurus [189]. TCA [50]. tea [1084, 449, 1007]. technicians [1231]. technique [1231]. technology [524]. tectum [1305]. telemetry [156]. telencephalic [790]. telencephalon [222, 316, 145]. Teleost [600, 927, 1128, 577, 945, 388, 90, 369, 1247, 1049, 989, 146, 1354, 733, 893, 1166, 81, 1273, 1179]. teleost-specific [388]. Teleostei [693, 1099, 32, 384, 814, 1162, 1218, 349, 615]. teleosts [455, 308, 389, 329, 197]. Temminck [243, 980, 283, 540]. temperate [1180]. Temperature [1191, 270, 585, 514, 1284, 912, 1341, 1364, 1180, 580, 520, 894, 795, 606, 417, 1134, 301, 351, 251, 267, 521, 186, 806, 822, 1285, 936, 205, 1092, 584, 337, 429, 78, 808, 257, 327, 694, 807, 1140, 1146, 768, 684, 703, 1005, 840, 216, 40, 1051, 38, 1220, 783, 940, 1040, 705, 901, 1349, 1252, 1143, 1081, 124, 618, 1078, 96]. temperatures [512, 167, 1160, 1382, 777, 902, 9, 803, 1122]. Temporal [157, 407]. tench [364]. tensions [946]. tenuispinis [1180]. term [212, 1160, 338, 1139, 570, 209, 510, 372, 152, 1357, 1145, 1123, 108, 1308, 1228, 672, 516, 892]. terminal [1087]. Terminalia [866]. terraenovae [800]. terrestris [204]. test [1241, 9]. testes [663, 682, 1228, 693]. Testicular [1174, 1120, 509, 453, 535, 204, 937, 1065]. testis [930, 922, 190, 395, 1243]. Testosterone [287, 922]. testudineus [244, 504, 9]. tetrachloride [612, 296, 1075, 151]. tetrachloride-induced [296, 1075]. tetradecylthioacetic [1173]. tetrahydropyridine [707]. tetraploid [754]. tetraspanin [355]. Tex [636]. TGF [328]. TGF- [328]. TH [476]. thaw [940, 717]. their [996, 748, 922, 1028, 1333, 1379, 290, 854, 891, 78, 1386, 410, 54, 1350, 228, 416, 684, 530, 772, 461, 919, 333, 1153, 1371, 1331, 631]. Thermal [512, 72, 1357, 768, 9, 1025, 249]. thermoregulatory [510]. Thermostability [1337]. thermostable [1290]. these [1074]. thiamin [595]. thick [1147]. thick-lipped [1147]. Thiersant [826]. threat [883]. threatened [44, 261]. Three [1315, 1304, 1342, 167, 175, 1138, 466, 884, 26, 1397, 1238, 143, 777, 174, 719, 519, 9, 1051, 702, 723, 1122, 649, 1269]. three-river [519]. three-spined [1342, 884, 174]. three-spot [723]. thresher [351]. thresholds [1373]. throughout [628, 626]. Thunnus [1017, 1300, 421, 195, 300, 39, 540, 683, 942, 47]. Thymallus [1234, 257]. Thymalus [257]. thyme [715]. thymosin [1280]. thymus [1256, 445]. thynnus [1017, 1300, 47]. thyroid [502, 670, 262, 1194, 282, 881, 1209, 487, 1340, 1004, 1270]. thyroidal [242]. thyrotropin [128, 670, 1166, 507, 476]. thyrotropin- [1166]. Thyroxine [1336]. Tibet [79]. Tibetan [1311, 1162, 147]. tiburo [650]. tiger [991, 1143]. tight [1288, 686]. Tilapia [859, 753, 32, 384, 928, 1390, 109, 663, 664, 1159, 280, 1289, 1325, 993, 1352, 1052, 1074, 1339, 480, 192, 486, 805, 183, 361, 77, 1250, 142, 670, 198, 891, 518, 1216, 1246, 777, 732, 1326, 1090, 661, 383, 309, 931, 324, 727, 1347, 667, 344, 702, 107, 210, 1153, 1169, 472, 639, 30, 558, 586, 1007, 51, 606]. tilapias [26, 1397]. time [459, 77, 681, 232, 840, 940, 757, 880, 210]. time-course [210]. Tinca [364]. tinfoil [974]. Tissue [602, 1066, 774, 644, 481, 1227, 1376, 435, 846, 913, 258, 23, 708, 952, 413, 1278, 871, 870, 1271, 771, 582, 1356, 1162, 743, 976, 60, 1165, 249, 1141, 121, 127, 979, 655, 900, 208, 832, 686, 1172, 30, 836, 171, 1034, 1026]. Tissue-specific [602, 1066, 644, 258, 1162, 127, 836]. tissues [962, 1182, 426, 518, 410, 869, 1136, 185, 702, 302, 649]. TLR [366]. TMP [168]. TMP-SMX [168]. TNF [1207]. TNF- [1207]. toadfish [402]. tocopherol [753, 518]. tocotrienol [518]. TOF [500]. tolerance [1168, 512, 167, 417, 105, 1285, 986, 1068, 9, 83, 483]. tolerant [340, 934, 494, 353, 555]. tolerate [1285]. Toll [1063, 873]. Toll-like [1063, 873]. tomato [137]. tomentosa [1161]. tongue [642, 450, 1104, 1013, 864, 755, 118, 1082, 1181, 138, 770, 363]. tool [239]. tools [878, 1232]. topmouth [987]. topography [540]. TOR [1288, 752, 1301, 1378, 1109, 512, 775, 964, 311]. TOR/PI3K/Akt [1301]. torazame [425]. Total [422, 141, 332, 1320]. totoaba [791, 776, 791, 776]. toxic [503, 675]. Toxicity [928, 977, 1227, 480, 546, 523, 1113, 1049, 989, 753, 36, 569]. Toxicological [866, 415]. toxicology [1241]. Toyohira [156]. tra [691]. trace [10, 1225]. Trachinotus [1284, 1386, 1388, 632, 815]. tract [308, 788, 243, 1298, 78, 526, 601, 360, 875, 964, 745, 1106, 1072]. training [339, 974]. trait [667]. traits [864, 720, 770]. transaminase [1027]. Transcript [1165, 336, 939, 1151, 1315]. transcription [1352, 1256, 103, 1030]. transcription-based [1030]. transcriptional [1086, 611, 1292, 1020, 481, 1172, 876, 558, 1331, 1034, 1026]. Transcriptome [1222, 957, 908, 1260, 1395, 1089]. transcriptomic [908, 1268]. Transcripts [922, 1339, 1313, 355, 1235, 213]. Transducin [749, 526]. transduction [687, 27]. transfection [1096]. transfer [962, 1226]. Transferability [625]. transferase [534]. transferrin [661, 1056]. transgenic [34, 792, 20, 1000]. transition [287]. transmitters [665]. transmontanus [538]. transpacificus [847]. transplantation [386, 1233, 1174]. transport [835, 656, 566, 1165, 1225]. Transportation [288, 541]. transporter [619, 1184, 986, 763, 547]. transporter- [763]. transporters [306, 1226, 582, 197]. treated [1340, 1109]. treatment [1173, 707, 838, 848, 1393, 178, 1036, 247, 836, 51]. TRH [476]. Tribolodon [1290]. Tribulus [204]. tributary [369]. tributyltin [852]. trichlorfon [726, 286]. triclosan [1358]. trigger [1115]. triglyceride [1207]. triiodothyronine [129, 242]. triiodothyroxine [781]. trimaculatum [723]. trimaculatus [392]. trimethoprim [168]. triphylla [566, 1155]. triploid [493, 14, 1132, 754, 1005, 845]. triploidy [1023, 1291]. trophic [70, 560, 529]. Tropical [278, 788, 874, 70, 622, 1337, 904, 50, 41, 676]. tropicus [278, 788, 874, 622]. tropomyosins [1337]. trout [1227, 544, 1023, 212, 435, 245, 1267, 656, 925, 1126, 1382, 219, 690, 722, 588, 929, 328, 613, 741, 711, 322, 433, 1066, 413, 1132, 1012, 1073, 1234, 1145, 1306, 424, 695, 1233, 862, 626, 744, 1243, 1188, 1239, 582, 1161, 1088, 759, 1263, 144, 1212, 684, 310, 672, 97, 1257, 1190, 721, 1005, 1211, 698, 1121, 834, 253, 507, 783, 715, 255, 1320, 69, 644, 975, 751, 83, 772, 127, 892, 888, 208, 845, 943]. TRP14 [876]. trutta [413, 433, 1073, 862, 1257, 1005, 83]. Trx1 [876]. Trypsin [88, 99, 63, 1290, 143, 293, 173]. trypsinogens [410]. trypsins [1290]. tryptophan [325, 502, 105, 343, 1334, 647, 291, 1370]. tshawytscha [104]. Tsukamurella [1263]. TTA [1173]. TTCF [311]. tubercles [181]. tubulin [831]. tuna [1017, 1300, 421, 195, 300, 39, 540, 683, 942, 47]. tunica [635]. turbot [514, 912, 17, 885, 1291, 1120, 1332, 1374, 165, 963, 628, 674, 844, 1318, 1353, 767, 1292, 497, 281, 122, 511, 998, 740]. turbots [1098]. turbulent [158]. turning [920]. Two [688, 766, 1020, 919, 1096, 1341, 1135, 608, 973, 23, 1260, 491, 192, 183, 854, 1234, 624, 909, 1312, 747, 410, 1262, 1016, 61, 938, 982, 640, 47, 55, 902, 698, 91, 85, 216, 1174, 680, 565, 679, 946, 1303, 1268]. two-dimensional [565, 1303]. two-hybrid [1262]. type [833, 996, 670, 1171, 1328, 572, 1096, 208, 1159]. types [455, 1171, 1386, 1351]. tyrosine [476].

U [1335]. UBE2A [855]. ubiquitin [855]. ubiquitin-conjugating [855]. UCP1 [602]. UCP2 [602]. UK [237]. ultrafreezer [1230]. Ultramorphological [1060]. ultrasonic [665]. ultrastructural [32, 384, 462, 912, 630]. Ultrastructure [590]. ultraviolet [7, 61]. Umbrina [462]. uncoupling [602]. undecimalis [256]. undergoing [452]. underlying [1089, 750, 21]. unfed [47]. Unmethylated [115]. unsaturated [82]. up-regulates [934]. up-regulation [49]. Updated [1305]. upon [911, 163, 228, 416, 966]. upregulation [1068, 661]. upstream [156]. uptake [306]. urea [687, 1027]. urmiana [56]. urophthalmus [596, 150]. urophysis [466]. urotensin [694]. usage [1020]. use [1074, 1310, 415, 1324, 650]. used [1346, 527]. Using [874, 485, 928, 1214, 132, 791, 221, 1262, 421, 1373, 1105, 640, 55, 1237, 968, 1030, 991]. utilisation [111]. utility [229]. Utilization [1387, 1300, 950, 833, 1052, 89, 657, 12, 1296]. UV [318, 474, 689]. UV-B [689]. UV-damaged-DNA [318].

vaccination [583, 445, 644]. vachelli [708, 1370]. Val [729]. Valenciennes [439, 972]. Validation [66]. valuable [1233]. value [1129]. values [4, 96]. var [1176, 440, 752, 804, 1248, 1351, 680, 1261, 1089, 631, 714]. var. [1260, 140]. variability [485]. variable [1182]. variables [512, 930]. variants [602]. variation [1066, 762, 367, 1257, 629, 1270]. Variations [537, 492, 186, 933, 1140, 743]. variegatus [1285, 1285, 202, 881]. varieties [1260]. various [619, 1224, 283, 84]. varying [314, 154]. vasa [960, 472]. vasotocin [1180, 795, 466, 33, 743, 1313, 1042, 91, 85]. VDAC2 [774]. Vegetable [1299, 422, 846]. VEGF [934]. velocities [717]. velocity [653]. vendace [42]. vera [805, 465]. Verasper [290, 202]. vermiculated [403]. versicolor [408]. vertebral [785, 206]. vertebrate [1197, 1146, 1179]. vertebrates [1153]. very [722]. Vessel [883]. veterinary [1359]. Vg1 [1030]. VgC [710]. via [1133, 1018, 983, 316, 1053]. viability [599, 1077]. viable [1238]. Vibrio [568, 377, 213]. vibriosis [1338]. Vicia [1295]. victims [225]. videometric [359]. Vietnam [981]. vinclozolin [922]. Violin [719]. Virgibacillus [872]. viridis [666]. virus [854, 323, 1096]. viscera [88, 143, 1101, 403]. visceral [773]. visible [7, 474]. vision [195]. Visual [300, 39]. Vitamin [1071, 62, 443, 996, 1175, 502, 605, 761, 1334, 518, 84, 137, 274, 671, 1048, 1050]. vitamins [56, 772]. vitellogenic [352, 70, 538]. Vitellogenin [409, 75, 391, 907, 949, 967, 1309, 1016, 563, 332, 500, 710, 59, 1030, 203, 18, 425]. vitellogenins [573, 1235]. vitreus [5, 457]. vitrification [1158, 1234]. vitro [1014, 1173, 544, 1382, 548, 877, 761, 447, 296, 1241, 695, 659, 117, 421, 454, 635, 59, 971, 1097, 294, 589, 508]. viviparous [697]. vivo [1173, 748, 851, 761, 296, 659, 244, 504, 704, 59, 333, 1097, 286]. volitans [624]. volume [1392, 2, 1242, 1201, 359, 200]. voluntary [1145]. vs [110]. vtgs [1309]. vulpinus [351].

W [987]. waigiensis [427]. Waigieu [427]. Walbaum [97, 975]. walking [2, 716, 1016, 194, 692]. walleye [5, 457]. warfarin [761]. warm [529]. waste [1007]. Water [388, 247, 514, 1284, 288, 835, 619, 1160, 251, 209, 464, 471, 373, 936, 63, 204, 937, 1092, 257, 179, 768, 868, 55, 719, 1201, 194, 1220, 1349, 65, 890, 51]. water-acclimated [471]. water-borne [209]. water-soluble [373]. waterborne [908, 93, 405, 917, 839, 1193, 1265, 990]. waterless [370, 618]. waters [341]. wave [696]. wavelength [1197, 936]. wavelengths [911]. weaned [1300]. weaning [935]. weather [1159]. weatherfish [754]. Weber [403]. weight [1160, 1030]. Welfare [224, 230, 237, 231, 234, 235, 267, 233, 232, 226, 238, 228, 227, 442, 229, 236, 239]. western [485]. wheat [1098]. wheat-based [1098]. white [275, 406, 98, 764, 719, 449, 305, 538, 162, 18]. whitefish [734, 794]. Whitespotted [404]. whole [120, 828, 1012, 54, 66, 1194, 91, 85, 572, 247]. whole-body [828, 54, 66, 247]. whole-mount [1194]. wide [1008, 639]. widely [501]. wild [1286, 1135, 331, 10, 343, 571, 273, 257, 1384, 530, 47, 1324, 775, 41, 820, 1096, 110]. wild-caught [1384]. wild-type [1096]. winter [1025]. wirelessly [572]. withdrawal [672]. without [1198]. wnt4 [923]. Wolffian [682]. wolffish [563]. wortmannin [1034]. wound [822, 893]. wrasse [1180, 392]. WY [481]. WY-14643 [481].

X [1287, 815]. xingguonensis [1248]. Xizang [79]. XXX [845]. xylan [827].

Ya [832, 680]. Ya-fish [832]. Yangtze [106, 1321]. Year [132, 96]. yeast [1012, 1262, 350, 29, 1261, 345, 1383]. yellow [833, 996, 1309, 170, 1015, 849, 1094, 1144, 336, 1044, 861, 1178, 499, 44, 420, 49, 191, 1368, 603, 1172, 941, 1053, 1331, 1254, 649, 990, 1029, 1034, 1026, 1148, 1348]. yellowtail [636, 161, 838, 1106, 1276, 394]. Yingchen [714]. yolk [573, 425]. young [1116, 96, 132]. young-of-year [96, 132]. Yuriria [179]. YY [896]. YYb [542].

Zanzibar [183]. Zealand [161, 104]. zebra [858, 968, 954]. Zebrafish [761, 1018, 34, 905, 1274, 796, 758, 636, 67, 125, 52, 1139, 792, 1230, 725, 120, 190, 94, 1275, 356, 318, 524, 306, 490, 694, 1000, 316, 1212, 1194, 689, 185, 810, 1030, 1344, 266, 1209, 763, 487, 878, 1116, 779, 333, 966, 1235, 29, 1011, 1155, 269, 394]. zero [868]. zero-water [868]. zillii [606]. zinc [1360, 917, 78, 910, 149, 977, 649, 990, 1029]. Zn [64, 336, 1172, 990]. zona [175, 1153]. Zuiew [182, 22]. zymogram [271].

References

Conceicao:2010:NMM

[1] Luis E. C. Conceição, Cláudia Aragão, and Jorge Dias. Novel methodologies in marine fish larval nutrition. Fish Physiology and Biochemistry, 36(1):1–16, March 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9373-z.

Biswas:2010:ICV

[2] Kuheli Biswas, Lucy M. Jyrwa, and Nirmalendu Saha. Influence of cell volume changes on protein synthesis in isolated hepatocytes of airbreathing walking catfish (Clarias batrachus). Fish Physiology and Biochemistry, 36(1):17–27, March 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9275-5.

Alvarez-Gonzalez:2010:DDE

[3] C. A. Alvarez-González, F. J. Moyano-López, and S. Dumas. Development of digestive enzyme activity in larvae of spotted sand bass Paralabrax maculatofasciatus II: Electrophoretic analysis. Fish Physiology and Biochemistry, 36(1):29–37, March 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9276-4.

Shahsavani:2010:DNV

[4] D. Shahsavani, M. Mohri, and H. Gholipour Kanani. Determination of normal values of some blood serum enzymes in Acipenser stellatus Pallas. Fish Physiology and Biochemistry, 36(1):39–43, March 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-008-9277-3.

Madenjian:2010:LEW

[5] Charles P. Madenjian, Chunfang Wang, and Richard G. Stickel. Laboratory evaluation of a walleye (Sander vitreus) bioenergetics model. Fish Physiology and Biochemistry, 36(1):45–53, March 2010. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9278-2.

Huang:2010:EDR

[6] Junwa Huang, Lixia Tian, and Yongjian Liu. Effects of dietary riboflavin levels on antioxidant defense of the juvenile grouper Epinephelus coioides. Fish Physiology and Biochemistry, 36(1):55–62, March 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-008-9279-1.

Matsumoto:2010:SSJ

[7] Taro Matsumoto, Hiroshi Ihara, and Yasunori Ishibashi. Spectral sensitivity of juvenile chub mackerel (Scomber japonicus) in visible and ultraviolet light. Fish Physiology and Biochemistry, 36(1):63–70, March 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9289-z.

Caruso:2010:PRS

[8] G. Caruso, G. Maricchiolo, and M. G. Denaro. Physiological responses to starvation in the European eel (Anguilla anguilla): effects on haematological, biochemical, non-specific immune parameters and skin structures. Fish Physiology and Biochemistry, 36(1):71–83, March 2010. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9290-6.

Sarma:2010:AEA

[9] Kamal Sarma, A. K. Pal, and Kartik Baruah. Acclimation of Anabas testudineus (Bloch) to three test temperatures influences thermal tolerance and oxygen consumption. Fish Physiology and Biochemistry, 36 (1):85–90, March 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9293-3.

Gao:2010:RBT

[10] Yonghua Gao, Qingling Feng, and Shengrong Li. The relationship between trace elements in fish otoliths of wild carp and hydrochemical conditions. Fish Physiology and Biochemistry, 36(1):91–100, March 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9294-2.

Ghedira:2010:MML

[11] J. Ghedira, J. Jebali, and H. Boussetta. Metallothionein and metal levels in liver, gills and kidney of Sparus aurata exposed to sublethal doses of cadmium and copper. Fish Physiology and Biochemistry, 36 (1):101–107, March 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9295-1.

Pirozzi:2010:PEU

[12] Igor Pirozzi, Mark A. Booth, and Geoff L. Allan. Protein and energy utilization and the requirements for maintenance in juvenile mulloway (Argyrosomus japonicus). Fish Physiology and Biochemistry, 36 (1):109–121, March 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9296-0.

Haoran:2010:F

[13] Lin Haoran. Foreword. Fish Physiology and Biochemistry, 36(2):123– 124, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-008-9242-1.

Guo:2010:IEA

[14] Xinhong Guo, Jinpeng Yan, and Yun Liu. Isolation and expression analyses of the Sox9a gene in triploid crucian carp. Fish Physiology and Biochemistry, 36(2):125–133, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9209-2.

Deane:2010:GHA

[15] Eddie E. Deane and Norman Y. S. Woo. Growth hormone attenuates branchial HSP70 expression in silver sea bream. Fish Physiology and Biochemistry, 36(2):135–140, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9229-y.

Tse:2010:DDB

[16] C. Y. Tse, K. M. Chan, and C. K. Wong. DNA damage as a biomarker for assessing the effects of suspended solids on the orange-spotted grouper, Epinephelus coioides. Fish Physiology and Biochemistry, 36 (2):141–146, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9243-0.

Chen:2010:CCE

[17] S. L. Chen, Y. Liu, and L. Meng. Cloning, characterization, and expression analysis of a CC chemokine gene from turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 36(2):147–155, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9218-1.

Wang:2010:CEI

[18] Ruilong Wang, Yun Gao, and Guangzhi Ma. Cloning, expression, and induction by 17-β estradiol (E_{2}) of a vitellogenin gene in the white cloud mountain minnow Tanichthys albonubes. Fish Physiology and Biochemistry, 36(2):157–164, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9222-5.

Li:2010:ARL

[19] Li Li, Ping Xie, and Longgen Guo. Antioxidant response in liver of the phytoplanktivorous bighead carp (Aristichthys nobilis) intraperitoneallyinjected with extracted microcystins. Fish Physiology and Biochemistry, 36(2):165–172, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9228-z.

Er-meng:2010:IET

[20] Yu Er-meng, Ye Xing, and Jian Qing. Isolation of Tanichthys albonubes β actin gene and production of transgenic Tanichthys albonubes. Fish Physiology and Biochemistry, 36(2):173–180, June 2010. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9238-x.

Zhou:2010:MMU

[21] Li Zhou and Jian-Fang Gui. Molecular mechanisms underlying sex change in hermaphroditic groupers. Fish Physiology and Biochemistry, 36(2):181–193, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9219-0.

Yin:2010:CFH

[22] Shaowu Yin and Yun Liu. Compositions and functions of the hatching froth from ricefield eel (Monopterus albus Zuiew). Fish Physiology and Biochemistry, 36(2):195–200, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-008-9274-6.

Dong:2010:GHT

[23] Haiyan Dong, Lingxian Zeng, and Haoran Lin. Growth hormone and two forms of insulin-like growth factors I in the giant grouper (Epinephelus lanceolatus): molecular cloning and characterization of tissue distribution. Fish Physiology and Biochemistry, 36(2):201–212, June 2010. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9231-4.

Zhou:2010:MCC

[24] Yi Zhou, Yandong Niu, and Jianzhong Li. Molecular cloning, characterization and expression of FSH and LH beta subunits from grass carp (Ctenopharyngodon idella). Fish Physiology and Biochemistry, 36 (2):213–221, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9223-4.

Liu:2010:SFS

[25] Z. H. Liu, Y. G. Zhang, and D. S. Wang. Studies on feminization, sex determination, and differentiation of the Southern catfish, Silurus meridionalis— a review. Fish Physiology and Biochemistry, 36 (2):223–235, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9281-7.

Huang:2010:DET

[26] Weiren Huang, Yong Zhang, and Haoran Lin. Distinct expression of three estrogen receptors in response to bisphenol A and nonylphenol in male Nile tilapias (Oreochromis niloticus). Fish Physiology and Biochemistry, 36(2):237–249, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9280-8. See correction [1397].

Jiao:2010:DAB

[27] Baowei Jiao and Christopher H. K. Cheng. Disrupting actions of bisphenol a and malachite green on growth hormone receptor gene expression and signal transduction in seabream. Fish Physiology and Biochemistry, 36(2):251–261, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9227-0.

Yan:2010:CAM

[28] Jinpeng Yan, Liangguo Liu, and Yun Liu. Comparative analysis of mitochondrial control region in polyploid hybrids of red crucian carp (Carassius auratus) × blunt snout bream (Megalobrama amblycephala). Fish Physiology and Biochemistry, 36(2):263–272, June 2010. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9251-0.

Yu:2010:ERZ

[29] Xiaobin Yu, Sze-Wah Lin, and Wei Ge. Expression of recombinant zebrafish follicle-stimulating hormone (FSH) in methylotropic yeast Pichia pastoris. Fish Physiology and Biochemistry, 36(2):273–281, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9244-z.

Zhang:2010:SPR

[30] Yong Zhang, Zijie Long, and Haoran Lin. The second prolactin receptor in Nile tilapia (Oreochromis niloticus): molecular characterization, tissue distribution and gene expression. Fish Physiology and Biochemistry, 36 (2):283–295, June 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9355-1.

Palstra:2010:SPE

[31] Arjan P. Palstra and Guido E. E. J. M. van den Thillart. Swimming physiology of European silver eels (Anguilla anguilla L.): energetic costs and effects on sexual maturation and reproduction. Fish Physiology and Biochemistry, 36(3):297–322, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9397-4.

Abdel-Aziz:2010:HHK

[32] El-Saydah H. Abdel-Aziz, Suzan B. S. Abdu, and Huda F. Fouad. Haemopoiesis in the head kidney of tilapia, Oreochromis niloticus (Teleostei: Cichlidae): a morphological (optical and ultrastructural) study. Fish Physiology and Biochemistry, 36(3):323–336, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-008-9297-z.

Iwata:2010:SRM

[33] Eri Iwata, Yukiko Nagai, and Hideaki Sasaki. Social rank modulates brain arginine vasotocin immunoreactivity in false clown anemonefish (Amphiprion ocellaris). Fish Physiology and Biochemistry, 36(3):337– 345, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9298-y.

Almeida:2010:IPI

[34] Daniela Volcan Almeida, Bruna Félix da Silva Nornberg, and Luis Fernando Marins. Induction of phase II enzymes and hsp70 genes by copper sulfate through the electrophile-responsive element (EpRE): insights obtained from a transgenic zebrafish model carrying an orthologous EpRE sequence of mammalian origin. Fish Physiology and Biochemistry, 36(3): 347–353, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9299-x.

Bombardier:2010:MAO

[35] E. Bombardier, R. K. Booth, and R. S. McKinley. Metabolic adaptations of oxidative muscle during spawning migration in the Atlantic salmon Salmo salar L. Fish Physiology and Biochemistry, 36(3):355– 365, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-008-9300-8.

Ling:2010:ARC

[36] Qufei Ling and Fashui Hong. Antioxidative role of cerium against the toxicity of lead in the liver of silver crucian carp. Fish Physiology and Biochemistry, 36(3):367–376, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-008-9301-7.

Cho:2010:EDN

[37] Sung Hwoan Cho, Choong-Il Kim, and Kyung-Duck Kim. Effect of dietary nutrient composition on the growth of olive flounder (Paralichthys olivaceus) with different feeding regimes. Fish Physiology and Biochemistry, 36(3):377–385, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-009-9303-0.

Sfakianakis:2010:ETM

[38] D. G. Sfakianakis and M. Kentouri. Effect of temperature on muscle lactate metabolic recovery in sea bass (Dicentrarchus labrax, L.) juveniles exposed to exhaustive exercise. Fish Physiology and Biochemistry, 36(3): 387–390, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9305-y.

Matsuura:2010:DVC

[39] R. Matsuura, Y. Sawada, and Y. Ishibashi. Development of visual cells in the Pacific bluefin tuna Thunnus orientalis. Fish Physiology and Biochemistry, 36(3):391–402, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-009-9306-x.

Sassi:2010:IHT

[40] Asma Sassi, Ali Annabi, and Imed Messaoudi. Influence of high temperature on cadmium-induced skeletal deformities in juvenile mosquitofish (Gambusia affinis). Fish Physiology and Biochemistry, 36(3):403–409, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-009-9307-9.

Sharma:2010:CFA

[41] Prakash Sharma, Vikas Kumar, and Gudipati Venkateshwarlu. Comparative fatty acid profiles of wild and farmed tropical freshwater fish rohu (Labeo rohita). Fish Physiology and Biochemistry, 36(3):411– 417, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9309-7.

Dietrich:2010:SBV

[42] Grzegorz J. Dietrich, Mariola Dietrich, and Andrzej Ciereszko. Semen biology of vendace (Coregonus albula L.). Fish Physiology and Biochemistry, 36(3):419–425, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-009-9311-0.

Silva:2010:SSI

[43] Dilson Silva, Madelayne Cortez-Moreira, and Célia Martins Cortez. Spectrofluorimetric study of the interaction of methyl-parathion with fish serum albumin. Fish Physiology and Biochemistry, 36(3):427– 433, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9312-z.

Prasad:2010:HLE

[44] G. Prasad and Sonia Charles. Haematology and leucocyte enzyme cytochemistry of a threatened yellow catfish Horabagrus brachysoma (Gunther 1864). Fish Physiology and Biochemistry, 36(3):435–443, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-009-9313-y.

Basrur:2010:ERC

[45] T. V. Basrur, R. Longland, and R. J. Wilkinson. Effects of repeated crowding on the stress response and growth performance in Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 36(3):445– 450, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9314-x.

Jia:2010:MCR

[46] W. Z. Jia, N. Shang, and Q. L. Guo. Molecular cloning of rhamnosebinding lectin gene and its promoter region from snakehead Channa argus. Fish Physiology and Biochemistry, 36(3):451–459, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-009-9315-9.

Ortega:2010:CLP

[47] Aurelio Ortega and Gabriel Mourente. Comparison of the lipid profiles from wild caught eggs and unfed larvae of two scombroid fish: northern bluefin tuna (Thunnus thynnus L., 1758) and Atlantic bonito (Sarda sarda Bloch, 1793). Fish Physiology and Biochemistry, 36(3): 461–471, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9316-8.

Hastey:2010:CFA

[48] R. P. Hastey, R. P. Phelps, and K. A. Cummins. Changes in free amino acid profile of red snapper Lutjanus campechanus, eggs, and developing larvae. Fish Physiology and Biochemistry, 36(3):473–481, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-009-9317-7.

Tian:2010:RLY

[49] Chen Tian, Xinhua Chen, and Jingqun Ao. The up-regulation of large yellow croaker secretory IgM heavy chain at early phase of immune response. Fish Physiology and Biochemistry, 36(3):483–490, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-009-9318-6.

Kumar:2010:MKE

[50] Vikas Kumar, N. P. Sahu, and Kartik Baruah. Modulation of key enzymes of glycolysis, gluconeogenesis, amino acid catabolism, and TCA cycle of the tropical freshwater fish Labeo rohita fed gelatinized and non-gelatinized starch diet. Fish Physiology and Biochemistry, 36(3): 491–499, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9319-5.

Zhou:2010:ETP

[51] Xuxia Zhou, Ziqiang Tian, and Weifen Li. Effect of treatment with probiotics as water additives on tilapia (Oreochromis niloticus) growth performance and immune response. Fish Physiology and Biochemistry, 36 (3):501–509, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9320-z.

Chen:2010:OPA

[52] Li-Chen Chen, Jen-Leih Wu, and Jyh-Yih Chen. Organization and promoter analysis of the zebrafish (Danio rerio) chemokine gene (CXC-64) promoter. Fish Physiology and Biochemistry, 36(3):511–521, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-009-9321-y.

Juchno:2010:DCH

[53] Dorota Juchno, Bozena Lackowska, and Wincenty Kilarski. DNA content of hepatocyte and erythrocyte nuclei of the spined loach (Cobitis taenia L.) and its polyploid forms. Fish Physiology and Biochemistry, 36(3): 523–529, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9322-x.

Monentcham:2010:EWB

[54] Serge-Eric Monentcham, Bernard Whatelet, and Patrick Kestemont. Egg and whole-body amino acid profile of African bonytongue (Heterotis niloticus) with an estimation of their dietary indispensable amino acids requirements. Fish Physiology and Biochemistry, 36(3):531–538, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-009-9323-9.

Papoutsoglou:2010:CCE

[55] Sofronios E. Papoutsoglou, Nafsika Karakatsouli, and Georgios Vasilikos. Common carp (Cyprinus carpio) response to two pieces of music (“Eine Kleine Nachtmusik” and “Romanza”) combined with light intensity, using recirculating water system. Fish Physiology and Biochemistry, 36(3): 539–554, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9324-8.

Jalali:2010:PCS

[56] Mohammad Ali Jalali, Seyed Abbas Hosseini, and Mohammad Reza Imanpour. Physiological characteristics and stress resistance of great sturgeon (Huso huso) juveniles fed with vitamins C, E, and HUFAenriched Artemia urmiana nauplii. Fish Physiology and Biochemistry, 36 (3):555–564, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9326-6.

Liu:2010:EDC

[57] Tailiang Liu, Hua Wen, and Wei Liu. Effect of dietary chromium picolinate on growth performance and blood parameters in grass carp fingerling, Ctenopharyngodon idellus. Fish Physiology and Biochemistry, 36(3): 565–572, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9327-5.

Akinsiku:2010:PKC

[58] Omolara Titilayo Akinsiku, Femi Kayode Agboola, and Adeyinka Afolayan. Physicochemical and kinetic characteristics of rhodanese from the liver of African catfish Clarias gariepinus Burchell in Asejire lake. Fish Physiology and Biochemistry, 36(3):573–586, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-009-9328-4.

Rani:2010:RPN

[59] K. V. Rani, N. Sehgal, and Om Prakash. Relative potencies of natural estrogens on vitellogenin and choriogenin levels in the Indian freshwater spotted snakehead, Channa punctata: in vivo and in vitro studies. Fish Physiology and Biochemistry, 36(3):587–595, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-009-9332-8.

Sarma:2010:BHC

[60] Kamal Sarma, A. K. Pal, and Kartik Baruah. Biochemical and histological changes in the brain tissue of spotted murrel, Channa punctatus (Bloch), exposed to endosulfan. Fish Physiology and Biochemistry, 36(3): 597–603, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9333-7.

Mekkawy:2010:EUA

[61] Imam A. A. Mekkawy, Usama M. Mahmoud, and Alaa El-Din H. Sayed. Effects of ultraviolet a on the activity of two metabolic enzymes, DNA damage and lipid peroxidation during early developmental stages of the African catfish, Clarias gariepinus (Burchell, 1822). Fish Physiology and Biochemistry, 36(3):605–626, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-009-9334-6.

Krossoy:2010:VKD

[62] Christel Krossøy, Erik-Jan Lock, and Robin Ørnsrud. Vitamin Kdependent γ-glutamylcarboxylase in Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 36(3):627–635, September 2010. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-009-9335-5.

Kanno:2010:PCT

[63] Gaku Kanno, Takahito Yamaguchi, and Hiroki Saeki. Purification and characteristics of trypsin from masu salmon (Oncorhynchus masou) cultured in fresh-water. Fish Physiology and Biochemistry, 36(3): 637–645, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9336-4.

Firat:2010:BAI

[64] Özgür Firat and Ferit Kargın. Biochemical alterations induced by Zn and Cd individually or in combination in the serum of Oreochromis niloticus. Fish Physiology and Biochemistry, 36(3):647–653, September 2010. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-009-9337-3.

Zarejabad:2010:HGS

[65] Asad Mohammadi Zarejabad, Mohammad Ali Jalali, and Somayeh Pouralimotlagh. Hematology of great sturgeon (Huso huso Linnaeus, 1758) juvenile exposed to brackish water environment. Fish Physiology and Biochemistry, 36(3):655–659, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-009-9339-1.

Peterson:2010:VWB

[66] Brian C. Peterson and Natha J. Booth. Validation of a whole-body cortisol extraction procedure for channel catfish (Ictalurus punctatus) fry. Fish Physiology and Biochemistry, 36(3):661–665, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-009-9340-8.

Chen:2010:IDE

[67] Li Chen, Jingyao Zhou, and Jinglun Xue. Identification and developmental expression of dec2 in zebrafish. Fish Physiology and Biochemistry, 36 (3):667–675, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9341-7.

Yu:2010:EFC

[68] Ling Zhi Yu, Xian Le Yang, and Kun Hu. Effects of fish CYP inducers on difloxacin N-demethylation in kidney cell of Chinese idle (Ctenopharyngodon idellus). Fish Physiology and Biochemistry, 36(3): 677–686, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9342-6.

Thawonsuwan:2010:EGE

[69] J. Thawonsuwan, V. Kiron, and V. Verlhac. Epigallocatechin-3-gallate (EGCG) affects the antioxidant and immune defense of the rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 36(3): 687–697, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9344-4.

Gomes:2010:FAT

[70] Aline D. Gomes, Tiago G. Correia, and Renata G. Moreira. Fatty acids as trophic biomarkers in vitellogenic females in an impounded tropical river. Fish Physiology and Biochemistry, 36(3):699–718, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-009-9345-3.

Zahl:2010:AIS

[71] Inger Hilde Zahl, Anders Kiessling, and Rolf Erik Olsen. Anesthesia induces stress in Atlantic salmon (Salmo salar), Atlantic cod (Gadus morhua) and Atlantic halibut (Hippoglossus hippoglossus). Fish Physiology and Biochemistry, 36(3):719–730, September 2010. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-009-9346-2.

Galleher:2010:CLT

[72] Stacy N. Galleher, Matthew R. Gilg, and Kelly J. Smith. Comparison of larval thermal maxima between Fundulus heteroclitus and F. grandis. Fish Physiology and Biochemistry, 36(3):731–740, September 2010. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-009-9347-1.

Amerand:2010:SRD

[73] A. Amérand, A. Vettier, and P. Sébert. Sex-related differences in aerobic capacities and reactive oxygen species metabolism in the silver eel. Fish Physiology and Biochemistry, 36(3):741–747, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-009-9348-0.

Ji:2010:EDA

[74] H. Ji, A. D. Om, and S. Sakamoto. Effect of dietary ascorbate on lipogenesis and lipolysis activities in black sea bream, Acanthopagrus schlegelii. Fish Physiology and Biochemistry, 36(3):749–755, September 2010. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-009-9349-z.

Leonardi:2010:VCF

[75] Maritza Leonardi, Jeanett Vera, and Violeta Morín. Vitellogenin of the Chilean flounder Paralichthys adspersus as a biomarker of endocrine disruption along the marine coast of the South Pacific. Part I: induction, purification, and identification. Fish Physiology and Biochemistry, 36(3): 757–765, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9350-6.

Geraudie:2010:RER

[76] Perrine Geraudie, Marie Gerbron, and Christophe Minier. Roach (Rutilus rutilus) reproductive cycle: a study of biochemical and histological parameters in a low contaminated site. Fish Physiology and Biochemistry, 36(3):767–777, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-009-9351-5.

Gonzalez-Felix:2010:EDP

[77] Mayra L. González-Félix, Francisco J. Castillo-Yañez, and José Lozano-Taylor. Effect of dietary protein source and time on alkaline proteolytic activity of Nile tilapia (Oreochromis niloticus). Fish Physiology and Biochemistry, 36(3):779–785, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-009-9352-4.

Kuzmina:2010:DTP

[78] V. V. Kuz’mina and N. V. Ushakova. The dependence on temperature and pH of the effects of zinc and copper on proteolytic activities of the digestive tract mucosa in piscivorous fish and their potential preys. Fish Physiology and Biochemistry, 36(3):787–795, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-009-9353-3.

Zhang:2010:HBB

[79] Huijuan Zhang, Congxin Xie, and Peng Shang. Haematological and blood biochemical characteristics of Glyptosternum maculatum (Siluriformes: Sisoridae) in Xizang (Tibet). Fish Physiology and Biochemistry, 36(3): 797–801, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9354-2.

Gharaei:2010:MCC

[80] Ahmad Gharaei, Fereidoun Mahboudi, and Saeed Keyvanshokooh. Molecular cloning of cDNA of mammalian and chicken II gonadotropinreleasing hormones (mGnRHs and cGnRH-II) in the beluga (Huso huso) and the disruptive effect of methylmercury on gene expression. Fish Physiology and Biochemistry, 36(3):803–817, September 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-009-9356-0.

Thebault:2010:PCG

[81] M. T. Thébault, A. Tanguy, and J. P. Raffin. Partial characterization of the gene encoding myoadenylate deaminase from the teleost fish Platichthys flesus. Fish Physiology and Biochemistry, 36(4):819– 825, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9358-y.

Mohd-Yusof:2010:IHU

[82] N. Y. Mohd-Yusof, O. Monroig, and D. R. Tocher. Investigation of highly unsaturated fatty acid metabolism in the Asian sea bass, Lates calcarifer. Fish Physiology and Biochemistry, 36(4):827–843, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9409-4.

Urke:2010:STA

[83] H. A. Urke, J. Koksvik, and T. Kristensen. Seawater tolerance in Atlantic salmon, Salmo salar L., brown trout, Salmo trutta L., and S. salar × S. trutta hybrids smolt. Fish Physiology and Biochemistry, 36(4): 845–853, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9359-x.

Lim:2010:GPV

[84] Chhorn Lim, Mediha Yildirim-Aksoy, and Phillip H. Klesius. Growth performance, vitamin E status, and proximate and fatty acid composition of channel catfish, Ictalurus punctatus, fed diets containing various levels of fish oil and vitamin E. Fish Physiology and Biochemistry, 36 (4):855–866, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9360-4.

Ripley:2010:QWB

[85] Jennifer L. Ripley and Christy M. Foran. Quantification of whole brain arginine vasotocin for two Syngnathus pipefishes: elevated concentrations correlated with paternal brooding. Fish Physiology and Biochemistry, 36 (4):867–874, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9361-3.

Garcia-Medina:2010:AIO

[86] Sandra García-Medina, Amparo Celene Razo-Estrada, and Marcela Galar-Martínez. Aluminum-induced oxidative stress in lymphocytes of common carp (Cyprinus carpio). Fish Physiology and Biochemistry, 36 (4):875–882, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9363-1.

Lazado:2010:RAC

[87] Carlo C. Lazado, Christopher Marlowe A. Caipang, and Viswanath Kiron. Responses of Atlantic cod Gadus morhua head kidney leukocytes to phytase produced by gastrointestinal-derived bacteria. Fish Physiology and Biochemistry, 36(4):883–891, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-009-9364-0.

Barkia:2010:TVB

[88] Ahmed Barkia, Ali Bougatef, and Moncef Nasri. Trypsin from the viscera of bogue (Boops boops): isolation and characterisation. Fish Physiology and Biochemistry, 36(4):893–902, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-009-9365-z.

Enes:2010:GPM

[89] P. Enes, H. Peres, and A. Oliva-Teles. Growth performance and metabolic utilization of diets including starch, dextrin, maltose or glucose as carbohydrate source by gilthead sea bream (Sparus aurata) juveniles. Fish Physiology and Biochemistry, 36(4):903–910, December 2010. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-009-9366-y.

Chilke:2010:SKR

[90] Arun M. Chilke. In situ kinetics of renal β-glucuronidase in teleost, Labeo rohita (Hamilton). Fish Physiology and Biochemistry, 36(4): 911–915, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9367-x.

Ripley:2010:EWB

[91] Jennifer L. Ripley and Christy M. Foran. Elevated whole brain arginine vasotocin with Aroclor 1254 exposure in two Syngnathus pipefishes. Fish Physiology and Biochemistry, 36(4):917–921, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-009-9368-9.

Amberg:2010:SDG

[92] Jon J. Amberg, Reuben Goforth, and Maria S. Sepúlveda. Sexually dimorphic gene expression in the gonad and liver of shovelnose sturgeon (Scaphirhynchus platorynchus). Fish Physiology and Biochemistry, 36 (4):923–932, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9369-8.

Clotfelter:2010:MMN

[93] Ethan D. Clotfelter, Meredith M. McNitt, and Cliff H. Summers. Modulation of monoamine neurotransmitters in fighting fish Betta splendens exposed to waterborne phytoestrogens. Fish Physiology and Biochemistry, 36(4):933–943, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-009-9370-2.

Jiang:2010:CRC

[94] Qiu Jiang, Dong Liu, and Houyan Song. Critical role of connexin43 in zebrafish late primitive and definitive hematopoiesis. Fish Physiology and Biochemistry, 36(4):945–951, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-009-9371-1.

Zhou:2010:IAS

[95] Li-Gen Zhou, Bing-Xin Liu, and Min-Jie Cao. Identification of an aminopeptidase from the skeletal muscle of grass carp (Ctenopharyngodon idellus). Fish Physiology and Biochemistry, 36(4):953–962, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-009-9372-0.

Ziegeweid:2010:HPO

[96] Jeffrey R. Ziegeweid and Marsha C. Black. Hematocrit and plasma osmolality values of young-of-year shortnose sturgeon following acute exposures to combinations of salinity and temperature. Fish Physiology and Biochemistry, 36(4):963–968, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-009-9374-y.

Panigrahi:2010:PBE

[97] A. Panigrahi, V. Kiron, and T. Watanabe. Probiotic bacteria Lactobacillus rhamnosus influences the blood profile in rainbow trout Oncorhynchus mykiss (Walbaum). Fish Physiology and Biochemistry, 36 (4):969–977, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-009-9375-x.

Lewis:2010:EDM

[98] H. A. Lewis, J. T. Trushenski, and C. C. Kohler. Effect of dietary marine lipids on female white bass ova compositions and progeny survival. Fish Physiology and Biochemistry, 36(4):979–992, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-009-9376-9.

Kamaci:2010:OEP

[99] H. Okan Kamaci, Cüneyt Suzer, and Kürşat Firat. Organogenesis of exocrine pancreas in sharpsnout sea bream (Diplodus puntazzo) larvae: characterization of trypsin expression. Fish Physiology and Biochemistry, 36(4):993–1000, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-009-9377-8.

Chen:2010:CEP

[100] Cai F. Chen, Hai S. Wen, and Na Li. Cloning and expression of P450c17-I (17α-hydroxylase/17,20-lyase) in brain and ovary during gonad development in Cynoglossus semilaevis. Fish Physiology and Biochemistry, 36(4):1001–1012, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-009-9378-7.

Fine:2010:PFS

[101] J. M. Fine and P. W. Sorensen. Production and fate of the sea lamprey migratory pheromone. Fish Physiology and Biochemistry, 36(4): 1013–1020, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9379-6.

Firat:2010:REC

[102] Özgür Fırat and Ferit Kargın. Response of Cyprinus carpio to copper exposure: alterations in reduced glutathione, catalase and proteins electrophoretic patterns. Fish Physiology and Biochemistry, 36(4):1021– 1028, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9380-0.

Holen:2010:ESC

[103] E. Holen, A. Kausland, and K. Skjærven. Embryonic stem cells isolated from Atlantic cod (Gadus morhua) and the developmental expression of a stage-specific transcription factor ac–Pou2. Fish Physiology and Biochemistry, 36(4):1029–1039, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9381-z.

Kurtovic:2010:PPD

[104] Ivan Kurtovic, Susan N. Marshall, and Benjamin K. Simpson. Purification and properties of digestive lipases from Chinook salmon (Oncorhynchus tshawytscha) and New Zealand hoki (Macruronus novaezelandiae). Fish Physiology and Biochemistry, 36(4):1041–1060, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-010-9382-y.

Hoseini:2010:EDT

[105] Seyyed Morteza Hoseini and Seyed Abbas Hosseini. Effect of dietary l-tryptophan on osmotic stress tolerance in common carp, Cyprinus carpio, juveniles. Fish Physiology and Biochemistry, 36(4):1061– 1067, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9383-x.

Wang:2010:IBR

[106] C. Wang, G. H. Lu, and P. F. Wang. Integrated biomarker response index for the assessment of environmental stress of the Yangtze River (Nanjing section). Fish Physiology and Biochemistry, 36(4):1069– 1078, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9384-9.

Uscanga:2010:AEE

[107] A. Uscanga, F. J. Moyano, and C. A. Alvarez. Assessment of enzymatic efficiency on protein digestion in the tilapia Oreochromis niloticus. Fish Physiology and Biochemistry, 36(4):1079–1085, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-010-9385-8.

Kristensen:2010:LAP

[108] T. Kristensen, B. O. Rosseland, and J. C. Massabau. Lack of arterial PO_{2} downregulation in Atlantic salmon (Salmo salar L.) during longterm normoxia and hyperoxia. Fish Physiology and Biochemistry, 36(4): 1087–1095, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9386-7.

Abdelhamid:2010:AIR

[109] Abdelhamid M. Abdelhamid, Ahmed I. Mehrim, and Mohamed A. El-Sharawy. An attempt to improve the reproductive efficiency of Nile tilapia brood stock fish. Fish Physiology and Biochemistry, 36(4):1097– 1104, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9387-6.

Zhao:2010:BCJ

[110] Feng Zhao, Ping Zhuang, and Zhaohong Shi. Biochemical composition of juvenile cultured vs. wild silver pomfret, Pampus argenteus: determining the diet for cultured fish. Fish Physiology and Biochemistry, 36(4):

1105–1111, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9388-5.

Figueiredo-Silva:2010:GNU

[111] A. Cláudia Figueiredo-Silva, Geneviève Corraze, and Luísa M. P. Valente. Growth and nutrient utilisation of blackspot seabream (Pagellus bogaraveo) under different feeding regimes. Fish Physiology and Biochemistry, 36(4):1113–1124, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9389-4.

Silva:2010:EML

[112] P. Silva, D. M. Power, and E. Rocha. Expression of the myosin light chains 1, 2 and 3 in the muscle of blackspot seabream (Pagellus bogaraveo, Brunnich), during development. Fish Physiology and Biochemistry, 36(4):1125–1132, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9390-y.

Ahmed:2010:RRL

[113] Imtiaz Ahmed. Response to the ration levels on growth, body composition, energy, and protein maintenance requirement of the Indian catfish (Heteropneustes fossilis— Bloch 1974). Fish Physiology and Biochemistry, 36(4):1133–1143, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9391-x.

Chilke:2010:KSH

[114] Arun M. Chilke. Kinetic study of hepatic β-glucuronidase in the Indian major carp, Labeo rohita (Hamilton). Fish Physiology and Biochemistry, 36(4):1145–1149, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9392-9.

Caipang:2010:UCO

[115] Christopher Marlowe A. Caipang, Sanchala Gallage, and Viswanath Kiron. Unmethylated CpG oligodeoxynucleotides activate head kidney leukocytes of Atlantic cod, Gadus morhua. Fish Physiology and Biochemistry, 36(4):1151–1158, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9393-8.

Kumar:2010:DID

[116] V. Kumar, H. P. S. Makkar, and K. Becker. Dietary inclusion of detoxified Jatropha curcas kernel meal: effects on growth performance and metabolic efficiency in common carp, Cyprinus carpio L. Fish Physiology and Biochemistry, 36(4):1159–1170, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9394-7.

Mandal:2010:IEE

[117] Sudipta Mandal and Koushik Ghosh. Inhibitory effect of Pistia tannin on digestive enzymes of Indian major carps: an in vitro study. Fish Physiology and Biochemistry, 36(4):1171–1180, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-010-9395-6.

Wang:2010:ECN

[118] X. L. Wang, N. Wang, and S. L. Chen. Establishment, characterization of a new cell line from heart of half smooth tongue sole (Cynoglossus semilaevis). Fish Physiology and Biochemistry, 36(4):1181–1189, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-010-9396-5.

Cai:2010:OCA

[119] Chun Fang Cai, Yuan Tu Ye, and Yong Ling Wang. Oxygen consumption and ammonia excretion of black carp (Mylopharyngdon piceus Richardson) and allogynogenetic crucian carp (Carassius auratus gibelio [female sign] × Cyprinus carpio [male sign]) fed different carbohydrate diets. Fish Physiology and Biochemistry, 36(4):1191–1198, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-010-9398-3.

Gomez-Requeni:2010:RGC

[120] P. Gómez-Requeni, L. E. C. Conceição, and I. Rønnestad. A reference growth curve for nutritional experiments in zebrafish (Danio rerio) and changes in whole body proteome during development. Fish Physiology and Biochemistry, 36(4):1199–1215, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9400-0.

Trosse:2010:OGE

[121] Christiane Tröße, Rune Waagbø, and Pål A. Olsvik. Optimisation of gene expression analysis in Atlantic salmon lenses by refining sampling strategy and tissue storage. Fish Physiology and Biochemistry, 36(4): 1217–1225, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9401-z.

Wang:2010:DCN

[122] N. Wang, X. L. Wang, and S. L. Chen. Development and characterization of a new marine fish cell line from turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 36(4):1227–1234, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-010-9402-y.

Sinaie:2010:MBD

[123] Mahmood Sinaie, Kazem Darvish Bastami, and Sara Haghparast. Metallothionein biosynthesis as a detoxification mechanism in mercury exposure in fish, spotted scat (Scatophagus argus). Fish Physiology and Biochemistry, 36(4):1235–1242, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9403-x.

Zeng:2010:ETE

[124] Ling-Qing Zeng, Yao-Guang Zhang, and Shi-Jian Fu. Effect of temperature on excess post-exercise oxygen consumption in juvenile southern catfish (Silurus meridionalis chen) following exhaustive exercise. Fish Physiology and Biochemistry, 36(4):1243–1252, December 2010. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9404-9.

Chen:2010:MCF

[125] Jyh-Yih Chen and Ming-Jyun Chiou. Molecular cloning and functional analysis of the zebrafish luteinizing hormone beta subunit (LHβ) promoter. Fish Physiology and Biochemistry, 36(4):1253–1262, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-010-9405-8.

Burgos-Aceves:2010:DPB

[126] Mario A. Burgos-Aceves, Rafael Campos-Ramos, and Danitzia A. Guerrero-Tortolero. Description of peripheral blood cells and differential blood analysis of captive female and male leopard grouper Mycteroperca rosacea as an approach for diagnosing diseases. Fish Physiology and Biochemistry, 36(4):1263–1269, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9406-7.

Verleih:2010:MCP

[127] Marieke Verleih, Alexander Rebl, and Tom Goldammer. Molecular characterization of PRR13 and its tissue-specific expression in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 36(4): 1271–1276, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9407-6.

Cohn:2010:MCR

[128] William B. Cohn, Richard A. Jones, and Duncan S. MacKenzie. Molecular cloning and regulation of mRNA expression of the thyrotropin β and glycoprotein hormone α subunits in red drum, Sciaenops ocellatus. Fish Physiology and Biochemistry, 36(4):1277–1290, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-010-9408-5.

Landines:2010:ITE

[129] M. A. Landines, A. I. Sanabria, and E. C. Urbinati. The influence of triiodothyronine (T_{3}) on the early development of piracanjuba (Brycon orbignyanus). Fish Physiology and Biochemistry, 36(4):1291– 1296, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9410-y.

Lahnsteiner:2010:CSC

[130] Franz Lahnsteiner. A comparative study on the composition and importance of free amino acids in semen of gilthead sea bream, Sparus aurata, and perch, Perca fluviatilis. Fish Physiology and Biochemistry, 36(4): 1297–1305, December 2010. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9442-3.

Lakra:2011:DCC

[131] W. S. Lakra, T. Raja Swaminathan, and K. P. Joy. Development, characterization, conservation and storage of fish cell lines: a review. Fish Physiology and Biochemistry, 37(1):1–20, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9411-x.

Durieux:2011:NFC

[132] Eric D. H. Durieux, Thomas B. Farver, and David J. Ostrach. Natural factors to consider when using acetylcholinesterase activity as neurotoxicity biomarker in Young-of-year striped bass (Morone saxatilis). Fish Physiology and Biochemistry, 37(1):21–29, March 2011. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-010-9412-9.

Powell:2011:CRB

[133] Mark D. Powell, Melissa S. Burke, and Dalia Dahle. Cardiac remodelling, blood chemistry, haematology and oxygen consumption of Atlantic cod, Gadus morhua L., induced by experimental haemolytic anaemia with phenylhydrazine. Fish Physiology and Biochemistry, 37 (1):31–41, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9413-8.

Nolasco:2011:PCP

[134] Héctor Nolasco, Francisco Moyano-López, and Fernando Vega-Villasante. Partial characterization of pyloric-duodenal lipase of gilthead seabream (Sparus aurata). Fish Physiology and Biochemistry, 37(1):43–52, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-010-9414-7.

Xia:2011:CIF

[135] Xiaohua Xia, Jie Zhao, and Zhongjie Chang. Cloning and identification of a female-specific DNA marker in Paramisgurnus dabryanus. Fish Physiology and Biochemistry, 37(1):53–59, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9415-6.

Begum:2011:OSA

[136] Ghousia Begum. Organ-specific ATPase and phosphorylase enzyme activities in a food fish exposed to a carbamate insecticide and recovery response. Fish Physiology and Biochemistry, 37(1):61–69, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9417-4.

Mekkawy:2011:ECS

[137] Imam A. A. Mekkawy, Usama M. Mahmoud, and Mervat Naguib. Effects of cadmium on some haematological and biochemical characteristics of Oreochromis niloticus (Linnaeus, 1758) dietary supplemented with tomato paste and vitamin E. Fish Physiology and Biochemistry, 37(1):71–84, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9418-3.

Xu:2011:MCG

[138] Tian jun Xu and Song lin Chen. Molecular cloning, genomic structure and expression analysis of major histocompatibility complex class Iα gene of half-smooth tongue sole (Cynoglossus semilaevis). Fish Physiology and Biochemistry, 37(1):85–90, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9419-2.

Hoseinifar:2011:SSH

[139] Seyed Hossein Hoseinifar, Alireza Mirvaghefi, and Kazem Darvish Bastami. The study of some haematological and serum biochemical parameters of juvenile beluga (Huso huso) fed oligofructose. Fish Physiology and Biochemistry, 37(1):91–96, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9420-9.

Jia:2011:LLC

[140] Xiuying Jia, Hangjun Zhang, and Xiaoxu Liu. Low levels of cadmium exposure induce DNA damage and oxidative stress in the liver of Oujiang colored common carp Cyprinus carpio var. color. Fish Physiology and Biochemistry, 37(1):97–103, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9416-5.

Davidson:2011:CHM

[141] Bruce Davidson, Jonathan Sidell, and Geremy Cliff. A comparison of the heart and muscle total lipid and fatty acid profiles of nine large shark species from the east coast of South Africa. Fish Physiology and Biochemistry, 37(1):105–112, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9421-8.

Han:2011:ESA

[142] Chunyan Han, Xiaobo Wen, and Haobo Li. Effect of starvation on activities and mRNA expression of lipoprotein lipase and hormone-sensitive lipase in tilapia (Oreochromis niloticus × O. areus). Fish Physiology and Biochemistry, 37(1):113–122, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9423-6.

Khaled:2011:PCT

[143] Hayet Ben Khaled, Kemel Jellouli, and Moncef Nasri. Purification and characterization of three trypsin isoforms from viscera of sardinelle (Sardinella aurita). Fish Physiology and Biochemistry, 37(1): 123–133, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9424-5.

Noga:2011:IHE

[144] Edward J. Noga, Paul J. Borron, and Jung-Kil Seo. Identification of histones as endogenous antibiotics in fish and quantification in rainbow trout (Oncorhynchus mykiss) skin and gill. Fish Physiology and Biochemistry, 37(1):135–152, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9422-7.

Rajan:2011:SLI

[145] K. Emmanuvel Rajan, A. Ganesh, and K. Radhakrishnan. Spatial learning-induced egr-1 expression in telencephalon of gold fish Carassius auratus. Fish Physiology and Biochemistry, 37(1):153–159, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9425-4.

Kumari:2011:SMG

[146] Bibha Kumari and Jawaid Ahsan. Study of muscle glycogen content in both sexes of an Indian teleost Clarias batrachus (Linn.) exposed to different concentrations of arsenic. Fish Physiology and Biochemistry, 37 (1):161–167, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9427-2.

Zhang:2011:BCS

[147] Hui Juan Zhang, Cong Xin Xie, and Xue Feng Yang. Blood cells of a sisorid catfish Glyptosternum maculatum (Siluriformes: Sisoridae), in Tibetan Plateau. Fish Physiology and Biochemistry, 37(1): 169–176, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9429-0.

Mohapatra:2011:HIH

[148] S. Mohapatra, N. P. Sahu, and Shivendra Kumar. Haemato-immunology and histo-architectural changes in Labeo rohita fingerlings: effect of dietary aflatoxin and mould inhibitor. Fish Physiology and Biochemistry, 37(1):177–186, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9428-1.

Malekpouri:2011:PEZ

[149] Pedram Malekpouri, Ali Asghar Moshtaghie, and Mehdi Soltani. Protective effect of zinc on related parameters to bone metabolism in common carp fish (Cyprinus carpio L.) intoxified with cadmium. Fish Physiology and Biochemistry, 37(1):187–196, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9430-7.

Lopez-Ramirez:2011:DDE

[150] G. López-Ramírez, C. A. Cuenca-Soria, and F. J. Moyano. Development of digestive enzymes in larvae of Mayan cichlid Cichlasoma urophthalmus. Fish Physiology and Biochemistry, 37(1):197–208, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-010-9431-6.

Yin:2011:HAE

[151] Guojun Yin, Liping Cao, and Chengping Lu. Hepatoprotective and antioxidant effects of Glycyrrhiza glabra extract against carbon tetrachloride (CCl_{4})-induced hepatocyte damage in common carp (Cyprinus carpio). Fish Physiology and Biochemistry, 37(1):209–216, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9436-1.

Hevroy:2011:EST

[152] E. M. Hevrøy, C. Azpeleta, and P. A. Olsvik. Effects of short-term starvation on ghrelin, GH-IGF system, and IGF-binding proteins in Atlantic salmon. Fish Physiology and Biochemistry, 37(1):217–232, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9434-3.

Liu:2011:CCP

[153] Xiaohong Liu, Biwen Xie, and Zhijian Wang. cDNA cloning, pituitary location, and extra-pituitary expression of pro-opiomelanocortin gene in rare minnow (Gobiocypris rarus). Fish Physiology and Biochemistry, 37 (1):233–247, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9433-4.

Karlsson:2011:CAE

[154] Anders Karlsson, Lene Sørlie Heier, and Anders Kiessling. Changes in arterial P_{w}O_{2}, physiological blood parameters and intracellular antioxidants in free-swimming Atlantic cod (Gadus morhua) exposed to varying levels of hyperoxia. Fish Physiology and Biochemistry, 37(1): 249–258, March 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9438-z.

Palstra:2011:FUE

[155] Arjan P. Palstra and Josep V. Planas. Fish under exercise. Fish Physiology and Biochemistry, 37(2):259–272, June 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9505-0.

Makiguchi:2011:ETS

[156] Yuya Makiguchi, Yoshifumi Konno, and Hiroshi Ueda. EMG telemetry studies on upstream migration of chum salmon in the Toyohira river, Hokkaido, Japan. Fish Physiology and Biochemistry, 37(2):273– 284, June 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-011-9495-y.

Palstra:2011:TPM

[157] Arjan P. Palstra, Ma. Angeles Guerrero, and Guido E. E. J. M. van den Thillart. Temporal progression in migratory status and sexual maturation in European silver eels during downstream migration. Fish Physiology and Biochemistry, 37(2):285–296, June 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9496-x.

Webb:2011:APE

[158] Paul W. Webb and Aline J. Cotel. Assessing possible effects of fishculture systems on fish swimming: the role of stability in turbulent flows. Fish Physiology and Biochemistry, 37(2):297–305, June 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9497-9.

Tudorache:2011:OSS

[159] Christian Tudorache, Robyn A. O’Keefe, and Tillmann J. Benfey. Optimal swimming speeds reflect preferred swimming speeds of brook charr (Salvelinus fontinalis Mitchill, 1874). Fish Physiology and Biochemistry, 37(2):307–315, June 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9498-8.

Herbert:2011:MLS

[160] Neill A. Herbert, Sunil Kadri, and Felicity A. Huntingford. A moving light stimulus elicits a sustained swimming response in farmed Atlantic salmon, Salmo salar L. Fish Physiology and Biochemistry, 37 (2):317–325, June 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9499-7.

Brown:2011:DSF

[161] Elliot J. Brown, Michael Bruce, and Neill A. Herbert. Do swimming fish always grow fast? Investigating the magnitude and physiological basis of exercise-induced growth in juvenile New Zealand yellowtail kingfish, Seriola lalandi. Fish Physiology and Biochemistry, 37(2):327–336, June 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9500-5.

Videler:2011:OPE

[162] J. J. Videler. An opinion paper: emphasis on white muscle development and growth to improve farmed fish flesh quality. Fish Physiology and Biochemistry, 37(2):337–343, June 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9501-4.

Kuzmina:2011:IPU

[163] V. V. Kuz’mina, E. G. Skvortsova, and V. A. Sheptitskiy. Influence of pH upon the activity of glycosidases and proteinases of intestinal mucosa, chyme and microbiota in fish. Fish Physiology and Biochemistry, 37(3): 345–353, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9426-3.

Kondera:2011:HHK

[164] Elżbieta Kondera. Haematopoiesis in the head kidney of common carp (Cyprinus carpio L.): a morphological study. Fish Physiology and Biochemistry, 37(3):355–362, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9432-5.

Goanvec:2011:BSH

[165] Christelle Goanvec, Elisabeth Poirier, and Michaël Theron. Branchial structure and hydromineral equilibrium in juvenile turbot (Scophthalmus maximus) exposed to heavy fuel oil. Fish Physiology and Biochemistry, 37(3):363–371, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9435-2.

Al-Salahy:2011:PSE

[166] M. Bassam Al-Salahy. Physiological studies on the effect of copper nicotinate (cu–n complex) on the fish, Clarias gariepinus, exposed to mercuric chloride. Fish Physiology and Biochemistry, 37(3):373– 385, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9437-0.

Chapman:2011:PTG

[167] Clint A. Chapman, Blake K. Harahush, and Gillian M. C. Renshaw. The physiological tolerance of the grey carpet shark (Chiloscyllium punctatum) and the epaulette shark (Hemiscyllium ocellatum) to anoxic exposure at three seasonal temperatures. Fish Physiology and Biochemistry, 37(3):387–399, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9439-y.

Yildiz:2011:PSI

[168] Hijran Yavuzcan Yildiz and Sermin Altunay. Physiological stress and innate immune response in gilthead sea bream (Sparus aurata) and sea bass (Dicentrarchus labrax) exposed to combination of trimethoprim and sulfamethoxazole (TMP-SMX). Fish Physiology and Biochemistry, 37(3): 401–409, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9440-5.

Shi:2011:CEP

[169] Z. Y. Shi, X. W. Chen, and Y. F. Gu. Cloning and expression pattern of alkaline phosphatase during the development of Paralichthys olivaceus. Fish Physiology and Biochemistry, 37(3):411–424, September 2011. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9441-4.

Haukenes:2011:PCH

[170] Alf H. Haukenes, Bruce A. Barton, and Kenneth J. Renner. Plasma cortisol and hypothalamic monoamine responses in yellow perch Perca flavescens after intraperitoneal injection of lipopolysaccharide. Fish Physiology and Biochemistry, 37(3):425–432, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-010-9443-2.

Zhou:2011:EDL

[171] Qiu-Bai Zhou, Hua-Dong Wu, and Xing-Hong Yan. Effects of dietary lipids on tissue fatty acids profile, growth and reproductive performance of female rice field eel (Monopterus albus). Fish Physiology and Biochemistry, 37(3):433–445, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9444-1.

Dietrich:2011:ICB

[172] Mariola A. Dietrich, Joanna Nynca, and Andrzej Ciereszko. Identification of calcium-binding proteins in fish seminal plasma. Fish Physiology and Biochemistry, 37(3):447–452, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9445-0.

Zhong:2011:EFC

[173] Guofang Zhong, Xi Qian, and Hongqi Zhou. Effects of feeding with corn gluten meal on trypsin activity and mRNA expression in Fugu obscurus. Fish Physiology and Biochemistry, 37(3):453–460, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9446-z.

OConnor:2011:EAC

[174] E. A. O’Connor, T. G. Pottinger, and L. U. Sneddon. The effects of acute and chronic hypoxia on cortisol, glucose and lactate concentrations in different populations of three-spined stickleback. Fish Physiology and Biochemistry, 37(3):461–469, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9447-y.

Chuang-Ju:2011:MCE

[175] Li Chuang-Ju, Wei Qi-Wei, and Gui Jian-Fang. Molecular characterization and expression pattern of three zona pellucida 3 genes in the Chinese sturgeon, Acipenser sinensis. Fish Physiology and Biochemistry, 37(3): 471–484, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9448-x.

Gharaei:2011:CME

[176] Ahmad Gharaei, Mostafa Ghaffari, and Reza Akrami. Changes in metabolic enzymes, cortisol and glucose concentrations of beluga (Huso huso) exposed to dietary methylmercury. Fish Physiology and Biochemistry, 37(3):485–493, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9450-3.

Costas:2011:FDS

[177] Benjamín Costas, Cláudia Aragão, and Luís E. C. Conceição. Feed deprivation in Senegalese sole (Solea senegalensis Kaup, 1858) juveniles: effects on blood plasma metabolites and free amino acid levels. Fish Physiology and Biochemistry, 37(3):495–504, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-010-9451-2.

Prasad:2011:ABE

[178] ManiRam Prasad, Abhishek Kumar, and Ajai K. Srivastav. Alterations in blood electrolytes of a freshwater catfish Heteropneustes fossilis in response to treatment with a botanical pesticide, Nerium indicum leaf extract. Fish Physiology and Biochemistry, 37(3):505–510, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9452-1.

Lopez-Lopez:2011:RAE

[179] Eugenia López-López, Jacinto Elías Sedeño-Díaz, and Liliana Favari. Responses of antioxidant enzymes, lipid peroxidation, and Na^{+} /K^{+}- ATPase in liver of the fish Goodea atripinnis exposed to Lake Yuriria water. Fish Physiology and Biochemistry, 37(3):511–522, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9453-0.

Duncan:2011:MRC

[180] Wallice P. Duncan, Naara F. Silva, and Marisa N. Fernandes. Mitochondrion-rich cells distribution, Na^{+} /K^{+}-ATPase activity and gill morphometry of the Amazonian freshwater stingrays (Chondrichthyes: Potamotrygonidae). Fish Physiology and Biochemistry, 37(3):523–531, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-010-9454-z.

Poncin:2011:RBC

[181] P. Poncin, B. Nzau Matondo, and J. C. Philippart. Relationships between circulating androgens, aggressive behaviour and breeding tubercles in males of the common bream Abramis brama l. in an aquarium environment. Fish Physiology and Biochemistry, 37(3):533–542, September 2011.

CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9455-y.

Weng:2011:PCP

[182] Wu-Yin Weng, Tao Wu, and Min-Jie Cao. Purification and characterization of pepsinogens and pepsins from the stomach of rice field eel (Monopterus albus Zuiew). Fish Physiology and Biochemistry, 37(3): 543–552, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9456-x.

Gao:2011:IEA

[183] Feng-Ying Gao, Mai-Xin Lu, and Li-Ping Yang. Identification and expression analysis of two growth hormone receptors in Zanzibar tilapia (Oreochromis hornorum). Fish Physiology and Biochemistry, 37(3): 553–565, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9457-9.

Hoseini:2011:SBC

[184] Seyyed Morteza Hoseini, Seyed Abbas Hosseini, and Ali Jafar Nodeh. Serum biochemical characteristics of beluga, Huso huso (L.), in response to blood sampling after clove powder solution exposure. Fish Physiology and Biochemistry, 37(3):567–572, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-010-9458-8.

SantAnna:2011:IEM

[185] M. C. B. Sant’Anna, Vanessa de Matas Soares, and Mauricio Reis Bogo. Iron exposure modifies acetylcholinesterase activity in zebrafish (Danio rerio) tissues: distinct susceptibility of tissues to iron overload. Fish Physiology and Biochemistry, 37(3):573–581, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-010-9459-7.

Gomez-Milan:2011:AVP

[186] E. Gómez-Milán, C. de Haro, and M. J. Sánchez-Muros. Annual variations of the plasmatic levels of glucose and amino acid and daily changes under different natural conditions of temperature and photoperiod in Gilthead Sea bream (Sparus aurata, L.). Fish Physiology and Biochemistry, 37(3):583–592, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9460-1.

Huang:2011:ISA

[187] B. F. Huang, Y. L. Sun, and D. S. Wang. Isolation, sequence analysis, and characterization of androgen receptor in Southern catfish, Silurus meridionalis. Fish Physiology and Biochemistry, 37(3):593– 601, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9461-0.

Caipang:2011:IAA

[188] Christopher Marlowe A. Caipang, Carlo C. Lazado, and Viswanath Kiron. Influence of alginic acid and fucoidan on the immune responses of head kidney leukocytes in cod. Fish Physiology and Biochemistry, 37(3): 603–612, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9462-z.

Davidson:2011:LLF

[189] Bruce Davidson and Geremy Cliff. Liver lipids of female Carcharias taurus (spotted raggedtooth) sharks: a comparison between seasons. Fish Physiology and Biochemistry, 37(3):613–618, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-010-9463-y.

Groh:2011:GPA

[190] Ksenia J. Groh, Victor J. Nesatyy, and Marc J.-F. Suter. Global proteomics analysis of testis and ovary in adult zebrafish (Danio rerio). Fish Physiology and Biochemistry, 37(3):619–647, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-010-9464-x.

Wang:2011:EPC

[191] Li Wang and Xinzhong Wu. Expression, purification and characterization of yellow grouper Epinephelus awoara regulator of G protein signaling 16 protein. Fish Physiology and Biochemistry, 37(3):649– 656, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-010-9465-9.

Firat:2011:CSE

[192] Özgür Fırat, Hikmet Y. Cogun, and Yasemin Kötemen. A comparative study on the effects of a pesticide (cypermethrin) and two metals (copper, lead) to serum biochemistry of Nile tilapia, Oreochromis niloticus. Fish Physiology and Biochemistry, 37(3):657–666, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9466-3.

Uscanga-Martinez:2011:CDE

[193] A. Uscanga-Martínez, N. Perales-García, and J. R. Indy. Changes in digestive enzyme activity during initial ontogeny of bay snook Petenia splendida. Fish Physiology and Biochemistry, 37(3):667–680, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-011-9467-2.

Saha:2011:IIE

[194] Nirmalendu Saha, Lucy M. Jyrwa, and Kuheli Biswas. Influence of increased environmental water salinity on gluconeogenesis in the airbreathing walking catfish, Clarias batrachus. Fish Physiology and Biochemistry, 37(3):681–692, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9468-1.

Matsumoto:2011:CSV

[195] Taro Matsumoto, Tokihiko Okada, and Yasunori Ishibashi. Changes in the scotopic vision of juvenile Pacific bluefin tuna (Thunnus orientalis) with growth. Fish Physiology and Biochemistry, 37(3):693– 700, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9469-0.

Chiu:2011:NBM

[196] Kuo-Hsun Chiu, Shangwu Ding, and Hin-Kiu Mok. A NMR-based metabolomic approach for differentiation of hagfish dental and somatic skeletal muscles. Fish Physiology and Biochemistry, 37(3):701– 707, September 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9470-7.

Tang:2011:ICT

[197] Cheng-Hao Tang, Lie-Yueh Hwang, and Tsung-Han Lee. Immunolocalization of chloride transporters to gill epithelia of euryhaline teleosts with opposite salinity-induced Na^{+} /K^{+}-ATPase responses. Fish Physiology and Biochemistry, 37(4):709–724, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9471-6.

Jatoba:2011:DSP

[198] Adolfo Jatobá, Felipe do Nascimento Vieira, and Edemar Roberto Andreatta. Diet supplemented with probiotic for Nile tilapia in polyculture system with marine shrimp. Fish Physiology and Biochemistry, 37(4): 725–732, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9472-5.

Martins:2011:CRA

[199] Dulce Alves Martins, Sofia Engrola, and Luis E. C. Conceição. Cortisol response to air exposure in Solea senegalensis post-larvae is affected by dietary arachidonic acid-to-eicosapentaenoic acid ratio. Fish Physiology and Biochemistry, 37(4):733–743, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9473-4.

Wormser:2011:RVR

[200] Chloe Wormser, Louise Z. Mason, and Douglas B. Light. Regulatory volume response following hypotonic stress in Atlantic salmon erythrocytes. Fish Physiology and Biochemistry, 37(4):745–759, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9474-3.

Sun:2011:SEA

[201] Yue na Sun, Tian jun Xu, and Ri xin Wang. Sequence and expression analysis of cathepsin S gene in the miiuy croaker Miichthys miiuy. Fish Physiology and Biochemistry, 37(4):761–765, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9475-2.

Li:2011:MHC

[202] Hongjun Li, Lianxin Jiang, and Chongbo He. Major histocompatibility complex class IIA and IIB genes of the spotted halibut Verasper variegatus: genomic structure, molecular polymorphism, and expression analysis. Fish Physiology and Biochemistry, 37(4):767–780, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9476-1.

Shved:2011:ELI

[203] Nikita Shved, Vadim Kumeiko, and Iraida Syasina. Enzyme-linked immunosorbent assay (ELISA) measurement of vitellogenin in plasma and liver histopathology in barfin plaice Liopsetta pinnifasciata from Amursky Bay, Sea of Japan. Fish Physiology and Biochemistry, 37(4): 781–799, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9477-0.

Kavitha:2011:IET

[204] P. Kavitha and P. Subramanian. Influence of Tribulus terrestris on testicular enzyme in fresh water ornamental fish Poecilia latipinna. Fish Physiology and Biochemistry, 37(4):801–807, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9478-z.

Kieffer:2011:ETP

[205] James D. Kieffer, Daniel W. Baker, and Christos N. Papadopoulos. The effects of temperature on the physiological response to low oxygen in Atlantic sturgeon. Fish Physiology and Biochemistry, 37(4): 809–819, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9479-y.

Pedersen:2011:MEI

[206] Mona E. Pedersen, Harald Takle, and Kirsten O. Hannesson. Matrilin-1 expression is increased in the vertebral column of Atlantic salmon (Salmo salar L.) individuals displaying spinal fusions. Fish Physiology and Biochemistry, 37(4):821–831, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9480-5.

Firouzbakhsh:2011:EPP

[207] F. Firouzbakhsh, F. Noori, and K. Jani-Khalili. Effects of a probiotic, protexin, on the growth performance and hematological parameters in the Oscar (Astronotus ocellatus) fingerlings. Fish Physiology and Biochemistry, 37(4):833–842, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9481-4.

Weil:2011:IIS

[208] Claudine Weil, Véronique Lebret, and Jean-Charles Gabillard. The IGF/ IGFBP system in rainbow trout (Oncorhynchus mykiss) adipose tissue: expression related to regional localization and cell type. Fish Physiology and Biochemistry, 37(4):843–852, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9482-3.

Ferrari:2011:EBJ

[209] L. Ferrari, B. L. Eissa, and A. Salibián. Energy balance of juvenile Cyprinus carpio after a short-term exposure to sublethal water-borne cadmium. Fish Physiology and Biochemistry, 37(4):853–862, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9483-2.

Velan:2011:CTC

[210] Ariel Velan, Gideon Hulata, and Avner Cnaani. Comparative time-course study on pituitary and branchial response to salinity challenge in Mozambique tilapia (Oreochromis mossambicus) and Nile tilapia (O. niloticus). Fish Physiology and Biochemistry, 37(4):863–873, December 2011. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9484-1.

Abdel-Hameid:2011:ESF

[211] Nassr-Allah H. Abdel-Hameid. Effect of starving and feeding on some haematological and physiological responses of the Nile catfish, Clarias gariepinus exposed to copper at extreme seasons. Fish Physiology and Biochemistry, 37(4):875–884, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9485-0.

Banaee:2011:ELT

[212] Mahdi Banaee, Antoni Sureda, and Golam Reza Rafei. Effects of longterm silymarin oral supplementation on the blood biochemical profile of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 37(4):885–896, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9486-z.

Zhang:2011:MCH

[213] Xinzhong Zhang, Huanying Pang, and Jichang Jian. Molecular characterization of heat shock protein 70 gene transcripts during Vibrio harveyi infection of humphead snapper, Lutjanus sanguineus. Fish Physiology and Biochemistry, 37(4):897–910, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9487-y.

Hallgren:2011:ABI

[214] Stefan Hallgren, Kristina Volkova, and Inger Porsch Hällström. Anxiogenic behaviour induced by 17α-ethynylestradiol in male guppies (Poecilia reticulata). Fish Physiology and Biochemistry, 37(4):911–918, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-011-9488-x.

Biswas:2011:NOI

[215] Saikat P. Biswas, Nikhil V. Palande, and Arun G. Jadhao. Nitric oxide inhibited the melanophore aggregation induced by extracellular calcium concentration in snakehead fish, Channa punctatus. Fish Physiology and Biochemistry, 37(4):919–927, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9489-9.

Santo:2011:ERT

[216] Valentina Di Santo and Wayne A. Bennett. Effect of rapid temperature change on resting routine metabolic rates of two benthic elasmobranchs. Fish Physiology and Biochemistry, 37(4):929–934, December 2011. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9490-3.

Martin:2011:EDL

[217] M. V. Martín, E. Almansa, and A. Lorenzo. Effects of a diet lacking HUFA on lipid and fatty acid content of intestine and gills of male gilthead seabream (Sparus aurata L.) broodstock at different stages of the reproductive cycle. Fish Physiology and Biochemistry, 37(4): 935–949, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9491-2.

Dogan:2011:HBB

[218] Demet Dogan and Canan Can. Hematological, biochemical, and behavioral responses of Oncorhynchus mykiss to dimethoate. Fish Physiology and Biochemistry, 37(4):951–958, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9492-1.

Deng:2011:EEE

[219] Junming Deng, Qingcong An, and Xi Zhang. Effect of ethanolic extract of propolis on growth performance and plasma biochemical parameters of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 37(4):959–967, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9493-0.

Zamaratskaia:2011:NHF

[220] Galia Zamaratskaia and Vladimir Zlabek. Para-nitrophenol hydroxylation by fish liver microsomes: kinetics and effect of selective cytochrome P450 inhibitors. Fish Physiology and Biochemistry, 37(4): 969–976, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9494-z.

Lin:2011:SCA

[221] Yuan-Chih Lin, Kuo-Hsun Chiu, and Hin-Kiu Mok. Seasonal changes in atrophy-associated proteins of the sonic muscle in the big-snout croaker, Johnius macrorhynus (Pisces, Sciaenidae), identified by using a proteomic approach. Fish Physiology and Biochemistry, 37(4): 977–991, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9502-3.

Baile:2011:CSB

[222] Vidya V. Baile and Pratap J. Patle. Cytoarchitectonic study of the brain of a dwarf snakehead, Channa gachua (Ham.). I. The telencephalon. Fish Physiology and Biochemistry, 37(4):993–1004, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9503-2.

Kumar:2011:HDP

[223] Shivendra Kumar, N. P. Sahu, and Vikas Kumar. High dietary protein combats the stress of Labeo rohita fingerlings exposed to heat shock. Fish Physiology and Biochemistry, 37(4):1005–1019, December 2011. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9504-1.

Kiessling:2012:WFF

[224] Anders Kiessling, Hans van de Vis, and Simon Mackenzie. Welfare of farmed fish in present and future production systems. Fish Physiology and Biochemistry, 38(1):1–3, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9594-9.

Cottee:2012:FVL

[225] Stephanie Yue Cottee. Are fish the victims of ‘speciesism’? A discussion about fear, pain and animal consciousness. Fish Physiology and Biochemistry, 38(1):5–15, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-010-9449-9.

Martins:2012:BIW

[226] Catarina I. M. Martins, Leonor Galhardo, and Tore Kristiansen. Behavioural indicators of welfare in farmed fish. Fish Physiology and Biochemistry, 38(1):17–41, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9518-8.

Prunet:2012:FWG

[227] P. Prunet, Ø. Øverli, and D. Baron. Fish welfare and genomics. Fish Physiology and Biochemistry, 38(1):43–60, February 2012. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-011-9522-z.

Noble:2012:IDF

[228] Chris Noble, Hernán A. Cañon Jones, and Stephanie Yue Cottee. Injuries and deformities in fish: their potential impacts upon aquacultural production and welfare. Fish Physiology and Biochemistry, 38 (1):61–83, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9557-1.

Segner:2012:HFF

[229] Helmut Segner, Henrik Sundh, and Lloyd Vaughan. Health of farmed fish: its relation to fish welfare and its utility as welfare indicator. Fish Physiology and Biochemistry, 38(1):85–105, February 2012. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-011-9517-9.

Attia:2012:DFW

[230] Joël Attia, Sandie Millot, and Børge Damsgård. Demand feeding and welfare in farmed fish. Fish Physiology and Biochemistry, 38(1): 107–118, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9538-4.

Conceicao:2012:DNF

[231] Luis E. C. Conceição, Cláudia Aragão, and Lluis Tort. Dietary nitrogen and fish welfare. Fish Physiology and Biochemistry, 38(1): 119–141, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9592-y.

Lopez-Olmeda:2012:DFT

[232] J. F. López-Olmeda, C. Noble, and F. J. Sánchez-Vázquez. Does feeding time affect fish welfare? Fish Physiology and Biochemistry, 38(1): 143–152, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9523-y.

Lines:2012:SWF

[233] J. A. Lines and J. Spence. Safeguarding the welfare of farmed fish at harvest. Fish Physiology and Biochemistry, 38(1):153–162, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9561-5.

Ellis:2012:CFW

[234] Tim Ellis, Hijran Yavuzcan Yildiz, and Catarina I. M. Martins. Cortisol and finfish welfare. Fish Physiology and Biochemistry, 38(1): 163–188, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9568-y.

Ellis:2012:MFW

[235] Tim Ellis, Iain Berrill, and Toby G. Knowles. Mortality and fish welfare. Fish Physiology and Biochemistry, 38(1):189–199, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9547-3.

Zahl:2012:AFF

[236] Inger Hilde Zahl, Ole Samuelsen, and Anders Kiessling. Anaesthesia of farmed fish: implications for welfare. Fish Physiology and Biochemistry, 38(1):201–218, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9565-1.

Berrill:2012:ACC

[237] I. K. Berrill, T. Cooper, and J. F. Turnbull. Achieving consensus on current and future priorities for farmed fish welfare: a case study from the UK. Fish Physiology and Biochemistry, 38(1):219–229, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-010-9399-2.

Muller-Graf:2012:RAF

[238] Christine Müller-Graf, Franck Berthe, and Ana Afonso. Risk assessment in fish welfare, applications and limitations. Fish Physiology and Biochemistry, 38(1):231–241, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9520-1.

vandeVis:2012:FWA

[239] J. W. van de Vis, M. Poelman, and M. Pilarczyk. Fish welfare assurance system: initial steps to set up an effective tool to safeguard and monitor farmed fish welfare at a company level. Fish Physiology and Biochemistry, 38(1):243–257, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9596-7.

Djordjevic:2012:ENS

[240] B. Djordjevic, T. Kristensen, and A. Kiessling. Effect of nutritional status and sampling intensity on recovery after dorsal aorta cannulation in free-swimming Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 38(1):259–272, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-009-9362-2.

Kristensen:2012:EPI

[241] T. Kristensen, T. O. Haugen, and B. O. Rosseland. Effects of production intensity and production strategies in commercial Atlantic salmon smolt (Salmo salar L.) production on subsequent performance in the early sea stage. Fish Physiology and Biochemistry, 38(1):273–282, February 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9566-0.

Marlatt:2012:ETD

[242] V. L. Marlatt, E. Gerrie, and V. L. Trudeau. Estradiol and triiodothyronine differentially modulate reproductive and thyroidal genes in male goldfish. Fish Physiology and Biochemistry, 38(2):283–296, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9506-z.

He:2012:ODT

[243] Tao He, Zhizhong Xiao, and Jun Li. Ontogeny of the digestive tract and enzymes in rock bream Oplegnathus fasciatus (Temminck et Schlegel 1844) larvae. Fish Physiology and Biochemistry, 38(2): 297–308, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9507-y.

Manju:2012:VPE

[244] Maniyan Manju, Mohammad Abdulkader Akbarsha, and Oommen Vilaverthottathil Oommen. In vivo protective effect of dietary curcumin in fish Anabas testudineus (Bloch). Fish Physiology and Biochemistry, 38(2):309–318, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9508-x.

Baumgarner:2012:IEG

[245] Bradley L. Baumgarner, Catherine P. Riley, and Jiri Adamec. Increased expression of GAPDH protein is not indicative of nitrosative stress or apoptosis in liver of starved rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 38(2):319–327, April 2012. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-011-9509-9.

Gingerich:2012:EBS

[246] Andrew J. Gingerich and Cory D. Suski. The effect of body size on post-exercise physiology in largemouth bass. Fish Physiology and Biochemistry, 38(2):329–340, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9510-3.

Xavier:2012:WSE

[247] Biji Xavier, N. P. Sahu, and Kartik Baruah. Water soaking and exogenous enzyme treatment of plant-based diets: effect on growth performance, whole-body composition, and digestive enzyme activities of rohu, Labeo rohita (Hamilton), fingerlings. Fish Physiology and Biochemistry, 38(2):341–353, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9511-2.

Filipiak:2012:FCD

[248] Marta Filipiak, Grzegorz Tylko, and Wincenty Kilarski. Flow cytometric determination of genome size in European sunbleak Leucaspius delineatus (Heckel, 1843). Fish Physiology and Biochemistry, 38(2): 355–362, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9512-1.

Snyder:2012:ETA

[249] Randal J. Snyder, William D. Schregel, and Yuanhong Wei. Effects of thermal acclimation on tissue fatty acid composition of freshwater alewives (Alosa pseudoharengus). Fish Physiology and Biochemistry, 38 (2):363–373, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9513-0.

Kowalska:2012:IDD

[250] Agata Kowalska, Zdzisław Zakeś, and Krystyna Demska-Zakeś. Impact of diets with different proportions of linseed and sunflower oils on the growth, liver histology, immunological and chemical blood parameters, and proximate composition of pikeperch Sander lucioperca (L.). Fish Physiology and Biochemistry, 38(2):375–388, April 2012. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-011-9514-z.

Dorts:2012:EEW

[251] Jennifer Dorts, Gaël Grenouillet, and Patrick Kestemont. Evidence that elevated water temperature affects the reproductive physiology of the European bullhead Cottus gobio. Fish Physiology and Biochemistry, 38 (2):389–399, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9515-y.

Karlsson:2012:PAM

[252] Anders Karlsson, Bjørn Olav Rosseland, and Anders Kiessling. Preanaesthetic metomidate sedation delays the stress response after caudal artery cannulation in Atlantic cod (Gadus morhua). Fish Physiology and Biochemistry, 38(2):401–411, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9516-x.

Sheikhzadeh:2012:EEH

[253] Najmeh Sheikhzadeh, Hossein Tayefi-Nasrabadi, and Mohammad Hamed Najafi Enferadi. Effects of Haematococcus pluvialis supplementation on antioxidant system and metabolism in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 38(2):413–419, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-011-9519-7.

Liakonis:2012:ECP

[254] K. M. Liakonis, R. Waagbø, and A. K. Imsland. Effects of chronic and periodic exposures to ammonia on the eye health in juvenile Atlantic halibut (Hippoglossus hippoglossus). Fish Physiology and Biochemistry, 38(2):421–430, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9521-0.

Tahmasebi-Kohyani:2012:EDN

[255] Ahmad Tahmasebi-Kohyani, Saeed Keyvanshokooh, and Hossein Pasha-Zanoosi. Effects of dietary nucleotides supplementation on rainbow trout (Oncorhynchus mykiss) performance and acute stress response. Fish Physiology and Biochemistry, 38(2):431–440, April 2012. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-011-9524-x.

Jimenez-Martinez:2012:DEA

[256] L. D. Jimenez-Martinez, C. A. Alvarez-González, and I. G. Palomino-Albarrán. Digestive enzyme activities during early ontogeny in common snook (Centropomus undecimalis). Fish Physiology and Biochemistry, 38(2):441–454, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9525-9.

Lahnsteiner:2012:EWT

[257] Franz Lahnsteiner and Manfred Kletzl. The effect of water temperature on gamete maturation and gamete quality in the European grayling (Thymalus thymallus) based on experimental data and on data from wild populations. Fish Physiology and Biochemistry, 38(2):455–467, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9526-8.

Dabas:2012:ATS

[258] Anurag Dabas, N. S. Nagpure, and W. S. Lakra. Assessment of tissuespecific effect of cadmium on antioxidant defense system and lipid peroxidation in freshwater murrel, Channa punctatus. Fish Physiology and Biochemistry, 38(2):469–482, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9527-7.

Iguchi:2012:EEC

[259] Masaharu Iguchi. Effects of etidronate on calcification of scales and ribs in the goldfish, Carassius auratus. Fish Physiology and Biochemistry, 38(2):483–491, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9528-6.

Hoseini:2012:ECS

[260] Seyyed Morteza Hoseini and Melika Ghelichpour. Efficacy of clove solution on blood sampling and hematological study in beluga, Huso huso (L.). Fish Physiology and Biochemistry, 38(2):493–498, April 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9529-5.

Sarma:2012:ACE

[261] Kamal Sarma, A. K. Pal, and Kartik Baruah. Acute and chronic effects of endosulfan on the haemato-immunological and histopathological responses of a threatened freshwater fish, spotted murrel, Channa punctatus. Fish Physiology and Biochemistry, 38(2):499–509, April 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9530-z.

Li:2012:ESD

[262] Dapeng Li, Zidong Liu, and Congxin Xie. Effect of stocking density on growth and serum concentrations of thyroid hormones and cortisol in Amur sturgeon, Acipenser schrenckii. Fish Physiology and Biochemistry, 38(2):511–520, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9531-y.

Imanpoor:2012:ERF

[263] Mohamad Reza Imanpoor and Tahere Bagheri. Effects of replacing fish meal by soybean meal along with supplementing phosphorus and magnesium in diet on growth performance of Persian sturgeon, Acipenser persicus. Fish Physiology and Biochemistry, 38(2):521–528, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9532-x.

Li:2012:PSE

[264] Xiangfei Li, Yangyang Jiang, and Xianping Ge. Protein-sparing effect of dietary lipid in practical diets for blunt snout bream (Megalobrama amblycephala) fingerlings: effects on digestive and metabolic responses. Fish Physiology and Biochemistry, 38(2):529–541, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9533-9.

Leung:2012:IDC

[265] L. Y. Leung and Norman Y. S. Woo. Influence of dietary carbohydrate level on endocrine status and hepatic carbohydrate metabolism in the marine fish Sparus sarba. Fish Physiology and Biochemistry, 38 (2):543–554, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9534-8.

Sun:2012:MCE

[266] Guijin Sun, Jie Pan, and Ximin Wang. Molecular cloning and expression analysis of P-selectin glycoprotein ligand-1 from zebrafish (Danio rerio). Fish Physiology and Biochemistry, 38(2):555–564, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9535-7.

Foss:2012:LCA

[267] A. Foss, E. Grimsbø, and B. Roth. Live chilling of Atlantic salmon: physiological response to handling and temperature decrease on welfare. Fish Physiology and Biochemistry, 38(2):565–571, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9536-6.

Mandal:2012:ESL

[268] Sagar C. Mandal, Mahinder P. S. Kohli, and Kartik Baruah. Effect of substituting live feed with formulated feed on the reproductive performance and fry survival of Siamese fighting fish, Betta splendens (Regan, 1910). Fish Physiology and Biochemistry, 38(2):573–584, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9539-3.

Zhu:2012:ECP

[269] Kecheng Zhu, Huanling Wang, and Min Wang. Expression characterization and the promoter activity analysis of zebrafish hdac4. Fish Physiology and Biochemistry, 38(2):585–593, April 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9540-x.

Couto:2012:TDS

[270] A. Couto, P. Enes, and A. Oliva-Teles. Temperature and dietary starch level affected protein but not starch digestibility in gilthead sea bream juveniles. Fish Physiology and Biochemistry, 38(3):595–601, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9537-5.

Ji:2012:SAD

[271] H. Ji, H. T. Sun, and D. M. Xiong. Studies on activity, distribution, and zymogram of protease, α-amylase, and lipase in the paddlefish Polyodon spathula. Fish Physiology and Biochemistry, 38(3):603–613, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9541-9.

Abbaraju:2012:EDO

[272] Naga V. Abbaraju and Bernard B. Rees. Effects of dissolved oxygen on glycolytic enzyme specific activities in liver and skeletal muscle of Fundulus heteroclitus. Fish Physiology and Biochemistry, 38 (3):615–624, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9542-8.

Kuzir:2012:MHS

[273] S. Kužir, E. Gjurčević, and Z. Kozarić. Morphological and histochemical study of intestine in wild and reared European eel (Anguilla anguilla L.). Fish Physiology and Biochemistry, 38(3):625–633, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9543-7.

Ozkan:2012:PRA

[274] Ferbal Özkan, Suna Gül Gündüz, and Serap Yalın. The protective role of ascorbic acid (vitamin C) against chlorpyrifos-induced oxidative stress in Oreochromis niloticus. Fish Physiology and Biochemistry, 38 (3):635–643, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9544-6.

Enes:2012:GIR

[275] P. Enes, H. Peres, and A. Oliva-Teles. Glycemic and insulin responses in white sea bream Diplodus sargus, after intraperitoneal administration of glucose. Fish Physiology and Biochemistry, 38(3):645–652, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9546-4.

Wang:2012:MCE

[276] Houpeng Wang, Tingting Wu, and Zaizhao Wang. Molecular cloning of Foxl2 gene and the effects of endocrine-disrupting chemicals on its mRNA level in rare minnow, Gobiocypris rarus. Fish Physiology and Biochemistry, 38(3):653–664, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9548-2.

Kang:2012:ELE

[277] Chao-Kai Kang, Fu-Chen Liu, and Tsung-Han Lee. Effects of low environmental salinity on the cellular profiles and expression of Na^{+}, K^{+}- ATPase and Na^{+}, K^{+}, 2Cl^{−} cotransporter 1 of branchial mitochondrionrich cells in the juvenile marine fish Monodactylus argenteus. Fish Physiology and Biochemistry, 38(3):665–678, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9549-1.

Aguilera:2012:DEA

[278] Carlos Aguilera, Roberto Mendoza, and Gabriel Marquez. Digestive enzymatic activity on tropical gar (Atractosteus tropicus) larvae fed different diets. Fish Physiology and Biochemistry, 38(3):679–691, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9550-8.

Troncoso:2012:HCG

[279] Ileana C. Troncoso, Jimena Cazenave, and María de los Ángeles Bistoni. Histopathological changes in the gills and liver of Prochilodus lineatus from the Salado River basin (Santa Fe, Argentina). Fish Physiology and Biochemistry, 38(3):693–702, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9551-7.

Anholt:2012:EAA

[280] R. D. Van Anholt, F. A. T. Spanings, and W. M. Koven. The effects of arachidonic acid on the endocrine and osmoregulatory response of tilapia (Oreochromis mossambicus) acclimated to seawater and subjected to confinement stress. Fish Physiology and Biochemistry, 38 (3):703–713, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9552-6.

Tong:2012:DEA

[281] X. H. Tong, S. H. Xu, and D. Y. Ma. Digestive enzyme activities of turbot (Scophthalmus maximus L.) during early developmental stages under culture condition. Fish Physiology and Biochemistry, 38 (3):715–724, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9553-5.

Ribeiro:2012:ISS

[282] A. R. A. Ribeiro, L. Ribeiro, and M. Moren. Iodine and selenium supplementation increased survival and changed thyroid hormone status in Senegalese sole (Solea senegalensis) larvae reared in a recirculation system. Fish Physiology and Biochemistry, 38(3):725–734, June 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9554-4.

Lee:2012:RDS

[283] J. Lee, I. C. Choi, and J. Y. Yoo. Response of dietary substitution of fishmeal with various protein sources on growth, body composition and blood chemistry of olive flounder (Paralichthys olivaceus, Temminck & Schlegel, 1846). Fish Physiology and Biochemistry, 38 (3):735–744, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9555-3.

Liu:2012:CCH

[284] Haichao Liu, Huihui Chen, and Xufa Ma. Cloning and characterization of the HSP90 beta gene from Tanichthys albonubes Lin (Cyprinidae): effect of copper and cadmium exposure. Fish Physiology and Biochemistry, 38(3):745–756, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9556-2.

Chaube:2012:BSC

[285] R. Chaube and S. Mishra. Brain steroid contents in the catfish Heteropneustes fossilis: sex and gonad stage-specific changes. Fish Physiology and Biochemistry, 38(3):757–767, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9558-0.

Xu:2012:ETO

[286] WeiNa Xu, WenBin Liu, and GuiFeng Li. Effect of trichlorfon on oxidative stress and hepatocyte apoptosis of Carassius auratus gibelio in vivo. Fish Physiology and Biochemistry, 38(3):769–775, June 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9559-z.

Huang:2012:TIT

[287] Yung-Sen Huang, Ya-Mei Chen, and Ching-Fong Chang. Testosterone improves the transition of primary oocytes in artificial maturation eels (Anguilla japonica) by altering ovarian PTEN expression. Fish Physiology and Biochemistry, 38(3):777–787, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9560-6.

Becker:2012:TSC

[288] Alexssandro G. Becker, Thaylise V. Parodi, and Bernardo Baldisserotto. Transportation of silver catfish, Rhamdia quelen, in water with eugenol and the essential oil of Lippia alba. Fish Physiology and Biochemistry, 38(3):789–796, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9562-4.

Riffel:2012:RPL

[289] A. P. K. Riffel, L. O. Garcia, and M. A. Pavanato. Redox profile in liver of Leporinus macrocephalus exposed to different dissolved oxygen levels. Fish Physiology and Biochemistry, 38(3):797–805, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9563-3.

Jin:2012:MCC

[290] Guo X. Jin, Hai S. Wen, and Na Li. Molecular cloning, characterization expression of P450c17-I and P450c17-II and their functions analysis during the reproductive cycle in males of barfin flounder (Verasper moseri). Fish Physiology and Biochemistry, 38(3):807–817, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9564-2.

Wolkers:2012:DTA

[291] Carla Patrícia Bejo Wolkers, Mônica Serra, and Elisabeth Criscuolo Urbinati. Dietary l-tryptophan alters aggression in juvenile matrinxã Brycon amazonicus. Fish Physiology and Biochemistry, 38(3):819–827, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-011-9569-x.

Mansour:2012:EDM

[292] M. Razeghi Mansour, R. Akrami, and A. Gharaei. Effect of dietary mannan oligosaccharide (MOS) on growth performance, survival, body composition, and some hematological parameters in giant sturgeon juvenile (Huso huso Linnaeus, 1754). Fish Physiology and Biochemistry, 38(3):829–835, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9570-4.

Liu:2012:PCT

[293] Chun-Hung Liu, Ya-Li Shiu, and Jue-Liang Hsu. Purification and characterization of trypsin from the pyloric ceca of orange-spotted grouper, Epinephelus coioides. Fish Physiology and Biochemistry, 38 (3):837–848, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9571-3.

Zhou:2012:EDC

[294] Wenshan Zhou, Hualei Liang, and Xuezhen Zhang. Erythrocyte damage of crucian carp (Carassius auratus) caused by microcystin-LR: in vitro study. Fish Physiology and Biochemistry, 38(3):849–858, June 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9572-2.

Marquez:2012:GPL

[295] Lorenzo Márquez, Rocío Robles, and Francisco J. Moyano. Gut pH as a limiting factor for digestive proteolysis in cultured juveniles of the gilthead sea bream (Sparus aurata). Fish Physiology and Biochemistry, 38(3):859–869, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9573-1.

Jia:2012:VVH

[296] Rui Jia, Liping Cao, and Guojun Yin. In vitro and in vivo hepatoprotective and antioxidant effects of Astragalus polysaccharides against carbon tetrachloride-induced hepatocyte damage in common carp (Cyprinus carpio). Fish Physiology and Biochemistry, 38(3):871–881, June 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9575-z.

Nyuji:2012:CEP

[297] Mitsuo Nyuji, Sethu Selvaraj, and Michiya Matsuyama. Changes in the expression of pituitary gonadotropin subunits during reproductive cycle of multiple spawning female chub mackerel Scomber japonicus. Fish Physiology and Biochemistry, 38(3):883–897, June 2012. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-011-9576-y.

Ibrahem:2012:LEE

[298] Mai D. Ibrahem, Hanan M. Khairy, and Marwa A. Ibrahim. Laboratory exposure of Oreochromis niloticus to crude microcystins (containing microcystin-LR) extracted from Egyptian locally isolated strain (Microcystis aeruginosa Kützing): biological and biochemical studies. Fish Physiology and Biochemistry, 38(3):899–908, June 2012. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-011-9577-x.

Liakonis:2012:EEC

[299] K. M. Liakonis, R. Waagbø, and A. K. Imsland. Erratum to: Effects of chronic and periodic exposures to ammonia on the eye health in juvenile Atlantic halibut (Hippoglossus hippoglossus). Fish Physiology and Biochemistry, 38(3):909–910, June 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9545-5.

Matsumoto:2012:VSS

[300] Taro Matsumoto, Tokihiko Okada, and Yasunori Ishibashi. Visual spectral sensitivity of photopic juvenile Pacific bluefin tuna (Thunnus orientalis). Fish Physiology and Biochemistry, 38(4):911–917, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9574-0.

Dalvi:2012:IAT

[301] Rishikesh S. Dalvi, Asim K. Pal, and Kartik Baruah. Influence of acclimation temperature on the induction of heat-shock protein 70 in the catfish Horabagrus brachysoma (Günther). Fish Physiology and Biochemistry, 38(4):919–927, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9578-9.

Ye:2012:IEP

[302] Huan Ye, Hao Du, and Chuang-Ju Li. Identification of a pou2 ortholog in Chinese sturgeon, Acipenser sinensis and its expression patterns in tissues, immature individuals and during embryogenesis. Fish Physiology and Biochemistry, 38(4):929–942, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-011-9579-8.

Valenzuela:2012:AAP

[303] Ariel Valenzuela, Victor Campos, and Ciro Oyarzun. Application of artificial photoperiod in fish: a factor that increases susceptibility to infectious diseases? Fish Physiology and Biochemistry, 38(4): 943–950, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9580-2.

Olsen:2012:SEC

[304] R. E. Olsen, A. Svardal, and A. Wargelius. Stress and expression of cyclooxygenases (cox1, cox2a, cox2b) and intestinal eicosanoids, in Atlantic salmon, Salmo salar L. Fish Physiology and Biochemistry, 38 (4):951–962, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9581-1.

Stanca:2012:ISB

[305] Loredana Stanca, Sorina Nicoleta Petrache, and Anca Dinischiotu. Impact of silicon-based quantum dots on the antioxidative system in white muscle of Carassius auratus gibelio. Fish Physiology and Biochemistry, 38(4):963–975, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9582-0.

Liu:2012:EAC

[306] Chih-Tsen Liu, Ming-Yi Chou, and Su Mei Wu. Effects of ambient cadmium with calcium on mRNA expressions of calcium uptake related transporters in zebrafish (Danio rerio) larvae. Fish Physiology and Biochemistry, 38(4):977–988, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9583-z.

Kumar:2012:DCB

[307] Neeraj Kumar, S. B. Jadhao, and R. S. Rana. Dietary choline, betaine and lecithin mitigates endosulfan-induced stress in Labeo rohita fingerlings. Fish Physiology and Biochemistry, 38(4):989–1000, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9584-y.

Becker:2012:ILG

[308] Alexssandro G. Becker, Jamile F. Gonçalves, and Bernardo Baldisserotto. Ion levels in the gastrointestinal tract content of freshwater and marine–estuarine teleosts. Fish Physiology and Biochemistry, 38(4): 1001–1017, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9585-x.

Roques:2012:PBR

[309] Jonathan A. C. Roques, Wout Abbink, and Gert Flik. Physiological and behavioral responses to an electrical stimulus in Mozambique tilapia (Oreochromis mossambicus). Fish Physiology and Biochemistry, 38(4): 1019–1028, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9586-9.

Oskoii:2012:EDA

[310] Somayeh Bohlouli Oskoii, Ahmad Tahmasebi Kohyani, and Ehsan Sadeghi. Effects of dietary administration of Echinacea purpurea on growth indices and biochemical and hematological indices in rainbow trout (Oncorhynchus mykiss) fingerlings. Fish Physiology and Biochemistry, 38(4):1029–1034, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9587-8.

Yadav:2012:DCC

[311] K. Yadav, W. S. Lakra, and Akhilesh Singh. Development and characterization of a cell line TTCF from endangered mahseer Tor tor (Ham.). Fish Physiology and Biochemistry, 38(4):1035–1045, August 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9588-7.

Fonseca-Madrigal:2012:ESB

[312] J. Fonseca-Madrigal, D. Pineda-Delgado, and D. R. Tocher. Effect of salinity on the biosynthesis of n-3 long-chain polyunsaturated fatty acids in silverside Chirostoma estor. Fish Physiology and Biochemistry, 38(4): 1047–1057, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-011-9589-6.

Zheng:2012:ERS

[313] Qingmei Zheng, Xiaobo Wen, and Xiaohui Xie. Effect of replacing soybean meal with cottonseed meal on growth, hematology, antioxidant enzymes activity and expression for juvenile grass carp, Ctenopharyngodon idellus. Fish Physiology and Biochemistry, 38(4):1059–1069, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-011-9590-0.

Brown:2012:EDM

[314] Charles A. Brown, Fernando Galvez, and Christopher C. Green. Embryonic development and metabolic costs in Gulf killifish Fundulus grandis exposed to varying environmental salinities. Fish Physiology and Biochemistry, 38(4):1071–1082, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-011-9591-z.

Yeh:2012:MCP

[315] Hung-Yueh Yeh and Phillip H. Klesius. Molecular characterization, phylogenetic analysis and expression patterns of five protein arginine methyltransferase genes of channel catfish, Ictalurus punctatus (Rafinesque). Fish Physiology and Biochemistry, 38(4):1083–1098, August 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9593-x.

Ng:2012:EMI

[316] Ming-Chong Ng, Chun-Po Hsu, and Kwok-Tung Lu. Effect of MK-801-induced impairment of inhibitory avoidance learning in zebrafish via inactivation of extracellular signal-regulated kinase (ERK) in telencephalon. Fish Physiology and Biochemistry, 38(4):1099–1106, August 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-011-9595-8.

Durieux:2012:CPM

[317] Eric D. H. Durieux, Richard E. Connon, and David J. Ostrach. Cytochrome P4501A mRNA and protein induction in striped bass (Morone saxatilis). Fish Physiology and Biochemistry, 38(4):1107–1116, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-011-9597-6.

Lai:2012:AIM

[318] Yi-Show Lai, Feng-Ju Hsieh, and Todd Hsu. Affinity isolation and mass spectral analysis of 1,10-phenanthroline (OP)-stimulated UV-damaged-DNA binding proteins expressed in zebrafish (Danio rerio) embryos. Fish Physiology and Biochemistry, 38(4):1117–1129, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9598-5.

Kennedy:2012:CLP

[319] Christopher J. Kennedy and Chris Picard. Chronic low pH exposure affects the seawater readiness of juvenile Pacific sockeye salmon. Fish Physiology and Biochemistry, 38(4):1131–1143, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9599-4.

Perez-Jimenez:2012:MAE

[320] A. Pérez-Jiménez, G. Cardenete, and A. E. Morales. Metabolic adjustments of Dentex dentex to prolonged starvation and refeeding. Fish Physiology and Biochemistry, 38(4):1145–1157, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9600-2.

Sieroslawska:2012:EMC

[321] Anna Sieroslawska, Anna Rymuszka, and Tadeusz Skowroński. Effects of microcystin-containing cyanobacterial extract on hematological and biochemical parameters of common carp (Cyprinus carpio L.). Fish Physiology and Biochemistry, 38(4):1159–1167, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-011-9601-1.

Heidarieh:2012:EDE

[322] Marzieh Heidarieh, Ali Reza Mirvaghefi, and Mehdi Behgar. Effect of dietary ergosan on growth performance, digestive enzymes, intestinal histology, hematological parameters and body composition of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 38(4): 1169–1174, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9602-8.

Lei:2012:CVS

[323] Xiao-Ying Lei, Zhong-Yuan Chen, and Qi-Ya Zhang. Characterization and virus susceptibility of a skin cell line from red-spotted grouper (Epinephelus akaara). Fish Physiology and Biochemistry, 38(4):1175– 1182, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-012-9603-7.

Sarker:2012:IDB

[324] Pallab Kumer Sarker, Rodrigue Yossa, and Grant W. Vandenberg. Influences of dietary biotin and avidin on growth, survival, deficiency syndrome and hepatic gene expression of juvenile Nile tilapia Oreochromis niloticus. Fish Physiology and Biochemistry, 38(4):1183–1193, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9604-6.

Ahmed:2012:DAA

[325] Imtiaz Ahmed. Dietary amino acid l-tryptophan requirement of fingerling Indian catfish, Heteropneustes fossilis (Bloch), estimated by growth and haemato-biochemical parameters. Fish Physiology and Biochemistry, 38 (4):1195–1209, August 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9609-1.

Abdel-Aziz:2012:SPP

[326] El-Sayedah H. Abdel-Aziz, Fayzah A. Bawazeer, and Mashael Al-Otaibi. Sexual patterns and protogynous sex reversal in the rusty parrotfish, Scarus ferrugineus (Scaridae): histological and physiological studies. Fish Physiology and Biochemistry, 38(4):1211–1224, August 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9610-8.

Li:2012:RTA

[327] Zaijian Li, Weiyuan Zhang, and Jiyue Cao. Reaction temperature alters chorzoxazone metabolism in carp (Cyprinus carpio) hepatic microsomes. Fish Physiology and Biochemistry, 38(5):1225–1231, October 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9605-5.

Gahr:2012:IES

[328] Scott A. Gahr, Gregory M. Weber, and Caird E. Rexroad III. Identification and expression of Smads associated with TGF-β/activin/nodal signaling pathways in the rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 38(5):1233–1244, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9611-7.

Nigam:2012:CAI

[329] Ashwini Kumar Nigam, Usha Kumari, and Ajay Kumar Mittal. Comparative analysis of innate immune parameters of the skin mucous secretions from certain freshwater teleosts, inhabiting different ecological niches. Fish Physiology and Biochemistry, 38(5):1245–1256, October 2012. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9613-5.

Padmini:2012:MHC

[330] E. Padmini and B. Vijaya Geetha. Mitochondrial HSP70 cognatemediated differential expression of JNK1/2 in the pollution stressed grey mullets, Mugil cephalus. Fish Physiology and Biochemistry, 38(5): 1257–1271, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9614-4.

Duncan:2012:RDG

[331] Neil Duncan, Alicia Estévez, and Constantinos C. Mylonas. Reproductive development, GnRHa-induced spawning and egg quality of wild meagre (Argyrosomus regius) acclimatised to captivity. Fish Physiology and Biochemistry, 38(5):1273–1286, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9615-3.

Nyina-wamwiza:2012:EPT

[332] L. Nyina-wamwiza, P. S. Defreyne, and P. Kestemont. Effects of partial or total fish meal replacement by agricultural by-product diets on gonad maturation, sex steroids and vitellogenin dynamics of African catfish (Clarias gariepinus). Fish Physiology and Biochemistry, 38(5): 1287–1298, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9616-2.

Wu:2012:VSZ

[333] Tsung-Han Wu, Chieh-Yu Pan, and Jyh-Yih Chen. In vivo screening of zebrafish microRNA responses to bacterial infection and their possible roles in regulating immune response genes after lipopolysaccharide stimulation. Fish Physiology and Biochemistry, 38(5):1299– 1310, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9617-1.

Hasan:2012:SCC

[334] Muhammad Mehedi Hasan, Shugo Watabe, and Yoshihiro Ochiai. Structural characterization of carangid fish myoglobins. Fish Physiology and Biochemistry, 38(5):1311–1322, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9619-z.

Tsui:2012:ESS

[335] Wen-Ching Tsui, Jiann-Chu Chen, and Sha-Yen Cheng. The effects of a sudden salinity change on cortisol, glucose, lactate, and osmolality levels in grouper Epinephelus malabaricus. Fish Physiology and Biochemistry, 38(5):1323–1329, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9620-6.

Kim:2012:YCE

[336] Jin-Hyoung Kim, Jae-Sung Rhee, and Jae-Seong Lee. The yellow catfish, Pelteobagrus fulvidraco (Siluriformes) metallothionein cDNA: molecular cloning and transcript expression level in response to exposure to the heavy metals Cd, Cu, and Zn. Fish Physiology and Biochemistry, 38(5): 1331–1342, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9621-5.

Kumar:2012:MDP

[337] Neeraj Kumar, S. B. Jadhao, and R. S. Rana. Methyl donors potentiates growth, metabolic status and neurotransmitter enzyme in Labeo rohita fingerlings exposed to endosulfan and temperature. Fish Physiology and Biochemistry, 38(5):1343–1353, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9622-4.

Ciji:2012:ASE

[338] Alexander Ciji, N. P. Sahu, and M. S. Akhtar. Alterations in serum electrolytes, antioxidative enzymes and haematological parameters of Labeo rohita on short-term exposure to sublethal dose of nitrite. Fish Physiology and Biochemistry, 38(5):1355–1365, October 2012. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-012-9623-3.

Zhao:2012:EHA

[339] Wen-Wen Zhao, Xu Pang, and Shi-Jian Fu. The effects of hypoxia acclimation, exercise training and fasting on swimming performance in juvenile qingbo (Spinibarbus sinensis). Fish Physiology and Biochemistry, 38 (5):1367–1377, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9624-2.

Amberg:2012:MRD

[340] Jon J. Amberg, Theresa M. Schreier, and Mark P. Gaikowski. Molecular responses differ between sensitive silver carp and tolerant bighead carp and bigmouth buffalo exposed to rotenone. Fish Physiology and Biochemistry, 38(5):1379–1391, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9625-1.

Murzina:2012:OLG

[341] Svetlana A. Murzina, Camilla A. Meyer Ottesen, and Oksana G. Poluektova. Oogenesis and lipids in gonad and liver of daubed shanny (Leptoclinus maculatus) females from Svalbard waters. Fish Physiology and Biochemistry, 38(5):1393–1407, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9627-z.

Grim:2012:DAL

[342] Jeffrey M. Grim, A. Adam Ding, and Wayne A. Bennett. Differences in activity level between cownose rays (Rhinoptera bonasus) and Atlantic stingrays (Dasyatis sabina) are related to differences in heart mass, hemoglobin concentration, and gill surface area. Fish Physiology and Biochemistry, 38(5):1409–1417, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9628-y.

Hoseini:2012:DTC

[343] Seyyed Morteza Hoseini, Seyed Abbas Hosseini, and Mohammad Soudagar. Dietary tryptophan changes serum stress markers, enzyme activity, and ions concentration of wild common carp Cyprinus carpio exposed to ambient copper. Fish Physiology and Biochemistry, 38(5): 1419–1426, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9629-x.

Sun:2012:IEF

[344] Yun-Lv Sun, Sheng Zeng, and De-Shou Wang. Involvement of FGF9/ 16/20 subfamily in female germ cell development of the Nile tilapia, Oreochromis niloticus. Fish Physiology and Biochemistry, 38(5):1427– 1439, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9630-4.

Zhu:2012:EYP

[345] Huiling Zhu, Huiyu Liu, and Lihe Liu. Effect of yeast polysaccharide on some hematologic parameter and gut morphology in channel catfish (Ictalurus punctatus). Fish Physiology and Biochemistry, 38(5): 1441–1447, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9631-3.

Smichi:2012:LAL

[346] Nabil Smichi, Ahmed Fendri, and Nabil Miled. Lipolytic activity levels and colipase presence in digestive glands of some marine animals. Fish Physiology and Biochemistry, 38(5):1449–1458, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9633-1.

Zeng:2012:EFF

[347] Ling-Qing Zeng, Feng-Jie Li, and Yao-Guang Zhang. Effect of feeding on the function and structure of the digestive system in juvenile southern catfish (Silurus meridionalis chen). Fish Physiology and Biochemistry, 38 (5):1459–1475, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9634-0.

Acosta:2012:FOO

[348] Daiane da Silva Acosta, Flávia Conde Kneip, and Laura Alicia Geracitano. Fullerene and omega-3 and omega-6 fatty acids on fish brain antioxidant status. Fish Physiology and Biochemistry, 38(5):1477–1485, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9635-z.

daSilva:2012:SCS

[349] Eder Marques da Silva, Marina Sek Lien Wong, and Adriane Pinto Wasko. Screening and characterization of sex-specific DNA fragments in the freshwater fish matrinchã, Brycon amazonicus (Teleostei: Characiformes: Characidae). Fish Physiology and Biochemistry, 38(5):1487– 1496, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9638-9.

Santacroce:2012:EDY

[350] Maria Pia Santacroce, Elisabetta Merra, and Elisabetta Casalino. Effects of dietary yeast Saccharomyces cerevisiae on the antioxidant system in the liver of juvenile sea bass Dicentrarchus labrax. Fish Physiology and Biochemistry, 38(5):1497–1505, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9640-2.

Donley:2012:ETP

[351] Jeanine M. Donley, Chugey A. Sepulveda, and Diego Bernal. Effects of temperature on power output and contraction kinetics in the locomotor muscle of the regionally endothermic common thresher shark (Alopias vulpinus). Fish Physiology and Biochemistry, 38(5):1507–1519, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9641-1.

Chabbi:2012:SII

[352] Ambarisha Chabbi and C. B. Ganesh. Stress-induced inhibition of recruitment of ovarian follicles for vitellogenic growth and interruption of spawning cycle in the fish Oreochromis mossambicus. Fish Physiology and Biochemistry, 38(5):1521–1532, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9643-z.

Chotichayapong:2012:IHT

[353] Chatrachatchaya Chotichayapong, Kittipong Wiengsamut, and Takahide Tsuchiya. Isolation of heat-tolerant myoglobin from Asian swamp eel Monopterus albus. Fish Physiology and Biochemistry, 38(5):1533–1543, October 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9644-y.

Javahery:2012:EAC

[354] Susan Javahery, Hamed Nekoubin, and Abdolmajid Haji Moradlu. Effect of anaesthesia with clove oil in fish (review). Fish Physiology and Biochemistry, 38(6):1545–1552, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9682-5.

Yeh:2012:CCE

[355] Hung-Yueh Yeh and Phillip H. Klesius. Channel catfish, Ictalurus punctatus (Rafinesque), tetraspanin membrane protein family: identification, characterization and phylogenetic analysis of tetraspanin 3 and tetraspanin 7 (CD231) transcripts. Fish Physiology and Biochemistry, 38(6):1553–1563, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-012-9645-x.

Koven:2012:EFR

[356] William Koven and Patricia Schulte. The effect of fasting and refeeding on mRNA expression of PepT1 and gastrointestinal hormones regulating digestion and food intake in zebrafish (Danio rerio). Fish Physiology and Biochemistry, 38(6):1565–1575, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9649-6.

Vani:2012:AHB

[357] T. Vani, N. Saharan, and Rajesh Kumar. Alteration in haematological and biochemical parameters of Catla catla exposed to sub-lethal concentration of cypermethrin. Fish Physiology and Biochemistry, 38(6): 1577–1584, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9650-0.

Weifen:2012:EEB

[358] Li Weifen, Zhang Xiaoping, and Yu Dongyou. Effects of Bacillus preparations on immunity and antioxidant activities in grass carp (Ctenopharyngodon idellus). Fish Physiology and Biochemistry, 38(6):1585–1592, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9652-y.

Torre:2012:PRR

[359] Agata Torre, Francesca Trischitta, and Caterina Faggio. Purinergic receptors and regulatory volume decrease in seabream (Sparus aurata) hepatocytes: a videometric study. Fish Physiology and Biochemistry, 38(6): 1593–1600, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9653-x.

Pradhan:2012:ODT

[360] P. K. Pradhan, J. K. Jena, and E. Gisbert. Ontogeny of the digestive tract in butter catfish Ompok bimaculatus (Bloch) larvae. Fish Physiology and Biochemistry, 38(6):1601–1617, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9655-8.

Gennotte:2012:CRP

[361] Vincent Gennotte, Philippe Sawadogo, and Carole Rougeot. Cortisol is responsible for positive and negative effects in the ovarian maturation induced by the exposure to acute stressors in Nile tilapia, Oreochromis niloticus. Fish Physiology and Biochemistry, 38(6):1619– 1626, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9656-7.

Shaluei:2012:PRG

[362] Fardin Shaluei, Aliakbar Hedayati, and Maryam Baghfalaki. Physiological responses of great sturgeon (Huso huso) to different concentrations of 2-phenoxyethanol as an anesthetic. Fish Physiology and Biochemistry, 38(6):1627–1634, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9659-4.

Zheng:2012:ECN

[363] Yuan Zheng, Na Wang, and Song-Lin Chen. Establishment and characterization of a new fish cell line from head kidney of half-smooth tongue sole (Cynoglossus semilaevis). Fish Physiology and Biochemistry, 38(6): 1635–1643, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9660-y.

Panicz:2012:GSC

[364] R. Panicz, J. Sadowski, and R. Drozd. Genetic and structural characterization of the growth hormone gene and protein from tench, Tinca tinca. Fish Physiology and Biochemistry, 38(6):1645–1653, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9661-x.

Alavi:2012:SCA

[365] Sayyed Mohammad Hadi Alavi, Azadeh Hatef, and Otomar Linhart. Sperm characteristics and androgens in Acipenser ruthenus after induction of spermiation by carp pituitary extract or GnRHa implants. Fish Physiology and Biochemistry, 38(6):1655–1666, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9662-9.

Tang:2012:CGO

[366] Da Tang, Yunhang Gao, and Tianjun Xu. Characterization, genomic organization, and expression profiles of MyD88, a key adaptor molecule in the TLR signaling pathways in miiuy croaker (Miichthys miiuy). Fish Physiology and Biochemistry, 38(6):1667–1677, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9663-8.

Matsche:2012:AVH

[367] Mark A. Matsche and Jarrett Gibbons. Annual variation of hematology and plasma chemistry in shortnose sturgeon, Acipenser brevirostrum, during a dam-impeded spawning run. Fish Physiology and Biochemistry, 38(6):1679–1696, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9664-7.

Cheng:2012:MCC

[368] Yuanzhi Cheng, Yuena Sun, and Tianjun Xu. Molecular cloning, characterization and expression analysis of a CC chemokine gene from miiuy croaker (Miichthys miiuy). Fish Physiology and Biochemistry, 38(6): 1697–1708, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9665-6.

Domingos:2012:ASP

[369] Fabricio F. T. Domingos, Ralph G. Thomé, and Elizete Rizzo. Assessment of spermatogenesis and plasma sex steroids in a seasonal breeding teleost: a comparative study in an area of influence of a tributary, downstream from a hydroelectric power dam, Brazil. Fish Physiology and Biochemistry, 38(6):1709–1719, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9668-3.

Mi:2012:QBP

[370] Hongbo Mi, Chunlu Qian, and Linchun Mao. Quality and biochemical properties of artificially hibernated crucian carp for waterless preservation. Fish Physiology and Biochemistry, 38(6):1721–1728, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9669-2.

Cerezuela:2012:EGS

[371] Rebeca Cerezuela, Francisco Antonio Guardiola, and M. Ángeles Esteban. Enrichment of gilthead seabream (Sparus aurata L.) diet with microalgae: effects on the immune system. Fish Physiology and Biochemistry, 38(6):1729–1739, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9670-9.

Herrera:2012:PST

[372] Marcelino Herrera, Cláudia Aragão, and Luis E. C. Conceição. Physiological short-term response to sudden salinity change in the Senegalese sole (Solea senegalensis). Fish Physiology and Biochemistry, 38(6): 1741–1751, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9671-8.

Hedayati:2012:EWS

[373] Aliakbar Hedayati and Abdolreza Jahanbakhshi. The effect of watersoluble fraction of diesel oil on some hematological indices in the great sturgeon Huso huso. Fish Physiology and Biochemistry, 38(6):1753– 1758, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9672-7.

Lu:2012:RLL

[374] Rong-Hua Lu, Xu-Fang Liang, and Yan He. The role of leptin in lipid metabolism in fatty degenerated hepatocytes of the grass carp Ctenopharyngodon idellus. Fish Physiology and Biochemistry, 38(6): 1759–1774, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9673-6.

Goswami:2012:DLC

[375] M. Goswami, W. S. Lakra, and J. K. Jena. Development of an ES-like cell culture system (RESC) from rohu, Labeo rohita (Ham.). Fish Physiology and Biochemistry, 38(6):1775–1783, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9674-5.

Han:2012:EPO

[376] Yu-Zhe Han, Tong-Jun Ren, and Connie-Fay Komilus. Effects of palm oil blended with oxidized fish oil on growth performances, hematology, and several immune parameters in juvenile Japanese sea bass, Lateolabrax japonicas. Fish Physiology and Biochemistry, 38(6):1785– 1794, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9675-4.

Liang:2012:COS

[377] Jing-Zhen Liang, Ying-Zhu Rao, and Ting-Bao Yang. Cathepsin L in the orange-spotted grouper, Epinephelus coioides: molecular cloning and gene expression after a Vibrio anguillarum challenge. Fish Physiology and Biochemistry, 38(6):1795–1806, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9676-3.

Saleh:2012:NDC

[378] M. C. Saleh and S. McConkey. NADH-dependent cytochrome b5 reductase and NADPH methemoglobin reductase activity in the erythrocytes of Oncorhynchus mykiss. Fish Physiology and Biochemistry, 38(6): 1807–1813, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9677-2.

Yao:2012:SCA

[379] Dongrui Yao, Saikun Pan, and Mingqian Zhou. Structural characterization and antitumor and mitogenic activity of a lectin from the gill of bighead carp (Aristichthys nobilis). Fish Physiology and Biochemistry, 38(6):1815–1824, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9678-1.

Feng:2012:MCC

[380] Ke Feng, Gui rong Zhang, and Hai chao Ping. Molecular characterization of cholecystokinin in grass carp (Ctenopharyngodon idellus): cloning, localization, developmental profile, and effect of fasting and refeeding on expression in the brain and intestine. Fish Physiology and Biochemistry, 38(6):1825–1834, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9679-0.

Pang:2012:IPO

[381] Qiuxiang Pang, Xuemei Liu, and Huanhuan Sun. Induction of phenoloxidase and other immunological activities in the humoral fluids of amphioxus Branchiostoma belcheri challenged with lipopolysaccharide (LPS). Fish Physiology and Biochemistry, 38(6):1835–1842, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9680-7.

Chen:2012:MCE

[382] Wenbo Chen, Haoran Lin, and Wensheng Li. Molecular characterization and expression pattern of insulin-like growth factor binding protein-3 (IGFBP-3) in common carp, Cyprinus carpio. Fish Physiology and Biochemistry, 38(6):1843–1854, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9681-6.

Roques:2012:EPB

[383] Jonathan A. C. Roques, Wout Abbink, and Gert Flik. Erratum to: Physiological and behavioral responses to an electrical stimulus in Mozambique tilapia (Oreochromis mossambicus). Fish Physiology and Biochemistry, 38(6):1855, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9636-y.

Abdel-Aziz:2012:EHH

[384] El-Saydah H. Abdel-Aziz, Suzan B. S. Abdu, and Huda F. Fouad. Erratum to: Haemopoiesis in the head kidney of tilapia, Oreochromis niloticus (Teleostei: Cichlidae): a morphological (optical and ultrastructural) study. Fish Physiology and Biochemistry, 38(6):1857, December 2012. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9637-x.

Joy:2013:F

[385] K. P. Joy, R. Moses Inbaraj, and R. Chaube. Foreword. Fish Physiology and Biochemistry, 39(1):1, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9654-9.

Lacerda:2013:GCT

[386] S. M. S. N. Lacerda, G. M. J. Costa, and L. R. França. Germ cell transplantation as a potential biotechnological approach to fish reproduction. Fish Physiology and Biochemistry, 39(1):3–11, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9606-4.

Kagawa:2013:MOM

[387] H. Kagawa, Y. Sakurai, and Y. Masuda. Mechanism of oocyte maturation and ovulation and its application to seed production in the Japanese eel. Fish Physiology and Biochemistry, 39(1):13–17, February 2013. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9607-3.

Cerda:2013:WHF

[388] J. Cerdà, C. Zapater, and R. N. Finn. Water homeostasis in the fish oocyte: new insights into the role and molecular regulation of a teleost-specific aquaporin. Fish Physiology and Biochemistry, 39(1): 19–27, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9608-2.

Hiramatsu:2013:MOL

[389] N. Hiramatsu, W. Luo, and A. Hara. Multiple ovarian lipoprotein receptors in teleosts. Fish Physiology and Biochemistry, 39(1):29–32, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9612-6.

Wu:2013:SSS

[390] Guan-Chung Wu and Ching-Fong Chang. The switch of secondary sex determination in protandrous black porgy, Acanthopagrus schlegeli. Fish Physiology and Biochemistry, 39(1):33–38, February 2013. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-012-9618-0.

Rawat:2013:VGF

[391] V. S. Rawat, K. V. Rani, and N. Sehgal. Vitellogenin genes in fish: differential expression on exposure to estradiol. Fish Physiology and Biochemistry, 39(1):39–46, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9626-0.

Nozu:2013:SOS

[392] Ryo Nozu, Ryo Horiguchi, and Masaru Nakamura. Survival of ovarian somatic cells during sex change in the protogynous wrasse, Halichoeres trimaculatus. Fish Physiology and Biochemistry, 39(1):47–51, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9632-2.

Desvignes:2013:NGF

[393] T. Desvignes, A. Fostier, and J. Bobe. The Nme gene family in fish. Fish Physiology and Biochemistry, 39(1):53–58, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9639-8.

Nocillado:2013:SAR

[394] J. N. Nocillado, A. S. Mechaly, and A. Elizur. In silico analysis of the regulatory region of the Yellowtail Kingfish and Zebrafish Kiss and Kiss receptor genes. Fish Physiology and Biochemistry, 39(1): 59–63, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9642-0.

Lal:2013:ENO

[395] B. Lal and N. Dubey. Existence of a nitric oxide synthase/nitric oxide system in fish testis and its role in modulation of androgenesis. Fish Physiology and Biochemistry, 39(1):65–69, February 2013. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-012-9648-7.

Blasco:2013:PKP

[396] Martín Blasco, Gustavo Manuel Somoza, and Denise Vizziano-Cantonnet. Presence of 11-ketotestosterone in pre-differentiated male gonads of Odontesthes bonariensis. Fish Physiology and Biochemistry, 39(1):71–74, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9651-z.

Uchida:2013:GHP

[397] Katsuhisa Uchida, Shunsuke Moriyama, and Masumi Nozaki. Glycoprotein hormone in the pituitary of hagfish and its evolutionary implications. Fish Physiology and Biochemistry, 39(1):75–83, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9657-6.

Saraiva:2013:EMR

[398] J. L. Saraiva, D. Gonçalves, and R. F. Oliveira. Ecological modulation of reproductive behaviour in the peacock blenny: a mini-review. Fish Physiology and Biochemistry, 39(1):85–89, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9658-5.

Berbejillo:2013:EED

[399] Julio Berbejillo, Anabel Martinez-Bengochea, and Denise Vizziano-Cantonnet. Expression of dmrt1 and sox9 during gonadal development in the Siberian sturgeon (Acipenser baerii). Fish Physiology and Biochemistry, 39(1):91–94, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9666-5.

Nagarajan:2013:IES

[400] Ganesan Nagarajan, Adimoolam Aruna, and Ching-Fong Chang. Increase in estrogen signaling in the early brain of orange-spotted grouper Epinephelus coioides: a mini-review. Fish Physiology and Biochemistry, 39(1):95–101, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9667-4.

Trombley:2013:LFP

[401] S. Trombley and M. Schmitz. Leptin in fish: possible role in sexual maturation in male Atlantic salmon. Fish Physiology and Biochemistry, 39 (1):103–106, February 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9731-0.

Chiu:2013:PCS

[402] Kuo-Hsun Chiu, Fu-Ming Hsieh, and Hin-Kiu Mok. Parvalbumin characteristics in the sonic muscle of a freshwater ornamental grunting toadfish (Allenbatrachus grunniens). Fish Physiology and Biochemistry, 39 (2):107–119, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9683-4.

Villalba-Villalba:2013:PCC

[403] Ana Gloria Villalba-Villalba, Juan Carlos Ramírez-Suárez, and Ciria Guadalupe Figueroa-Soto. Purification and characterization of chymotrypsin from viscera of vermiculated sailfin catfish, Pterygoplichthys disjunctivus, Weber, 1991. Fish Physiology and Biochemistry, 39(2):121–130, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9684-3.

Ma:2013:MCE

[404] Qian Ma, Yong-Quan Su, and Qi-Sheng Tang. Molecular cloning and expression analysis of major histocompatibility complex class IIB gene of the whitespotted bambooshark (Chiloscyllium plagiosum). Fish Physiology and Biochemistry, 39(2):131–142, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9685-2.

Costa:2013:MOH

[405] P. M. Costa, S. Caeiro, and M. H. Costa. Multi-organ histological observations on juvenile Senegalese soles exposed to low concentrations of waterborne cadmium. Fish Physiology and Biochemistry, 39 (2):143–158, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9686-1.

Enes:2013:EGG

[406] P. Enes, P. Pousão-Ferreira, and A. Oliva-Teles. Effect of guar gum on glucose and lipid metabolism in white sea bream Diplodus sargus. Fish Physiology and Biochemistry, 39(2):159–169, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9687-0.

Vetesnik:2013:EPT

[407] Lukáš Vetešník, Karel Halačka, and Andrea Šimková. The effect of ploidy and temporal changes in the biochemical profile of gibel carp (Carassius gibelio): a cyprinid fish species with dual reproductive strategies. Fish Physiology and Biochemistry, 39(2):171–180, April 2013. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-012-9688-z.

Wu:2013:EEC

[408] Zhi xin Wu, Su feng Pang, and Li jiao Pang. Effect of Coriolus versicolor polysaccharides on the hematological and biochemical parameters and protection against Aeromonas hydrophila in allogynogenetic crucian carp (Carassius auratus gibelio). Fish Physiology and Biochemistry, 39 (2):181–190, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9689-y.

Fazielawanie:2013:IPC

[409] N. M. R. Fazielawanie, S. S. Siraj, and M. Y. Ina-Salwany. Isolation and partial characterization of Asian sea bass (Lates calcarifer) vitellogenin. Fish Physiology and Biochemistry, 39(2):191–200, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9690-5.

Liu:2013:MCT

[410] Chun-Hung Liu, Ya-Huei Chen, and Ya-Li Shiu. Molecular characterization of two trypsinogens in the orange-spotted grouper, Epinephelus coioides, and their expression in tissues during early development. Fish Physiology and Biochemistry, 39(2):201–214, April 2013. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-012-9691-4.

Hachero-Cruzado:2013:SPQ

[411] Ismael Hachero-Cruzado, Asunción Forniés, and Gonzalo Martínez-Rodríguez. Sperm production and quality in brill Scophthalmus rhombus L.: relation to circulating sex steroid levels. Fish Physiology and Biochemistry, 39(2):215–220, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9692-3.

Kang:2013:IDB

[412] Duk-Young Kang and Hyo-Chan Kim. Influence of density and background color to stress response, appetite, growth, and blind-side hypermelanosis of flounder, Paralichthys olivaceus. Fish Physiology and Biochemistry, 39(2):221–232, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9693-2.

Hisar:2013:SDA

[413] Olcay Hisar, Adem Yavuz Sönmez, and Nejdet Gültepe. The sexually dimorphic adipose fin is an androgen target tissue in the brown trout (Salmo trutta fario). Fish Physiology and Biochemistry, 39(2): 233–241, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9694-1.

daCruz:2013:ABB

[414] André Luis da Cruz, Hugo Ribeiro da Silva, and Marisa Narciso Fernandes. Air-breathing behavior and physiological responses to hypoxia and air exposure in the air-breathing loricariid fish, Pterygoplichthys anisitsi. Fish Physiology and Biochemistry, 39(2):243–256, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9695-0.

Milinkovitch:2013:DUR

[415] Thomas Milinkovitch, Hélène Thomas-Guyon, and Nathalie Imbert. Dispersant use as a response to oil spills: toxicological effects on fish cardiac performance. Fish Physiology and Biochemistry, 39(2):257–262, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9696-z.

Noh:2013:CGE

[416] Gyeong Eon Noh, Han Kyu Lim, and Jong-Myoung Kim. Characterization of genes encoding prolactin and prolactin receptors in starry flounder Platichthys stellatus and their expression upon acclimation to freshwater. Fish Physiology and Biochemistry, 39(2):263–275, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9697-y.

Cheng:2013:STT

[417] Sha-Yen Cheng, Chih-Sung Chen, and Jiann-Chu Chen. Salinity and temperature tolerance of brown-marbled grouper Epinephelus fuscoguttatus. Fish Physiology and Biochemistry, 39(2):277–286, April 2013. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9698-x.

Mu:2013:CEAa

[418] Wei J. Mu, Hai S. Wen, and Bao X. Qi. Cloning and expression analysis of follicle-stimulating hormone and luteinizing hormone receptor during the reproductive cycle in Korean rockfish (Sebastes schlegeli). Fish Physiology and Biochemistry, 39(2):287–298, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9699-9.

Hirai:2013:CNE

[419] Mutsuko Hirai, Tadayoshi Muramatsu, and Hideaki Endo. Carbon nanotube enhanced label-free immunosensor for amperometric determination of oocyte maturation-inducing hormone in fish. Fish Physiology and Biochemistry, 39(2):299–308, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9700-7.

Pu:2013:MCC

[420] Lulu Pu, Kunhuang Han, and Yilei Wang. Molecular cloning, characterization, and gene expression of the androgen receptor in the large yellow croaker, Larimichthys crocea. Fish Physiology and Biochemistry, 39(2):309–324, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9701-6.

Martinez-Montano:2013:IAA

[421] Emmanuel Martínez-Montaño, Emyr Peña, and María Teresa Viana. Intestinal absorption of amino acids in the Pacific bluefin tuna (Thunnus orientalis): in vitro lysine–arginine interaction using the everted intestine system. Fish Physiology and Biochemistry, 39(2):325–334, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9702-5.

Benitez-Dorta:2013:TSF

[422] Vanessa Benítez-Dorta, María J. Caballero, and Daniel Montero. Total substitution of fish oil by vegetable oils in Senegalese sole (Solea senegalensis) diets: effects on fish performance, biochemical composition, and expression of some glucocorticoid receptor-related genes. Fish Physiology and Biochemistry, 39(2):335–349, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9703-4.

Amutha:2013:CAR

[423] C. Amutha and P. Subramanian. Cadmium alters the reproductive endocrine disruption and enhancement of growth in the early and adult stages of Oreochromis mossambicus. Fish Physiology and Biochemistry, 39(2):351–361, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9704-3.

Kwong:2013:EED

[424] Raymond W. M. Kwong, Charmain D. Hamilton, and Som Niyogi. Effects of elevated dietary iron on the gastrointestinal expression of Nramp genes and iron homeostasis in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 39(2):363–372, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9705-2.

Yamane:2013:CVD

[425] Kodai Yamane, Tomoki Yagai, and Akihiko Hara. Characterization of vitellogenin and its derived yolk proteins in cloudy catshark (Scyliorhinus torazame). Fish Physiology and Biochemistry, 39(2):373–390, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9706-1.

Kulac:2013:RAO

[426] Berna Kulac, Gülüzar Atli, and Mustafa Canli. Response of ATPases in the osmoregulatory tissues of freshwater fish Oreochromis niloticus exposed to copper in increased salinity. Fish Physiology and Biochemistry, 39(2):391–401, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9707-0.

Pham:2013:EDR

[427] Hung Quoc Pham and Augustine Arukwe. Effects of dopamine 2 receptor antagonist on sex steroid levels, oocyte maturation and spawning performances in waigieu seaperch (Psammoperca waigiensis). Fish Physiology and Biochemistry, 39(2):403–411, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9708-z.

Kumar:2013:EDC

[428] Neeraj Kumar, S. B. Jadhao, and R. S. Rana. Erratum to: Dietary choline, betaine and lecithin mitigates endosulfan-induced stress in Labeo rohita fingerlings. Fish Physiology and Biochemistry, 39(2): 413, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-012-9646-9.

Kumar:2013:EMD

[429] Neeraj Kumar, S. B. Jadhao, and R. S. Rana. Erratum to: Methyl donors potentiates growth, metabolic status and neurotransmitter enzyme in Labeo rohita fingerlings exposed to endosulfan and temperature. Fish Physiology and Biochemistry, 39(2):415, April 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9647-8.

Kim:2013:MCE

[430] Na Na Kim, Jehee Lee, and Cheol Young Choi. Molecular cloning and expression of caspase-3 in the protandrous cinnamon clownfish, Amphiprion melanopus, during sex change. Fish Physiology and Biochemistry, 39 (3):417–429, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9709-y.

Meena:2013:BGI

[431] D. K. Meena, Pronob Das, and S. C. Mukherjee. Beta-glucan: an ideal immunostimulant in aquaculture (a review). Fish Physiology and Biochemistry, 39(3):431–457, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9710-5.

Feng:2013:IEA

[432] Juan Feng, Youlu Su, and Yunxin Wang. Identification and expression analysis of a CC chemokine from cobia (Rachycentron canadum). Fish Physiology and Biochemistry, 39(3):459–469, June 2013. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-012-9711-4.

Hellstrom:2013:SPP

[433] Gustav Hellström, Tore Prestegaard, and K. Håkan Olsén. Sperm from pheromone primed brown trout (Salmo trutta L.) produce more larvae. Fish Physiology and Biochemistry, 39(3):471–478, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9712-3.

Oliveira:2013:EPC

[434] Catarina C. V. Oliveira, Rocio Aparício, and F. Javier Sánchez-Vazquez. Endocrine (plasma cortisol and glucose) and behavioral (locomotor and self-feeding activity) circadian rhythms in Senegalese sole (Solea senegalensis Kaup 1858) exposed to light/dark cycles or constant light. Fish Physiology and Biochemistry, 39(3):479–487, June 2013. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-012-9713-2.

Banaee:2013:BHC

[435] Mahdi Banaee, Antoni Sureda, and Kamal Ahmadi. Biochemical and histological changes in the liver tissue of rainbow trout (Oncorhynchus mykiss) exposed to sub-lethal concentrations of diazinon. Fish Physiology and Biochemistry, 39(3):489–501, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9714-1.

Zehra:2013:DLR

[436] Seemab Zehra and Mukhtar A. Khan. Dietary lysine requirement of fingerling Catla catla (Hamilton) based on growth, protein deposition, lysine retention efficiency, RNA/DNA ratio and carcass composition. Fish Physiology and Biochemistry, 39(3):503–512, June 2013. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-012-9715-0.

Luo:2013:EBS

[437] Yiping Luo, Wen Wang, and Qingda Huang. Effect of body size on organ-specific mitochondrial respiration rate of the largemouth bronze gudgeon. Fish Physiology and Biochemistry, 39(3):513–521, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9716-z.

Oliva:2013:HAS

[438] M. Oliva, J. J. Vicente-Martorell, and J. A. Perales. Histopathological alterations in Senegal sole, Solea senegalensis, from a polluted Huelva estuary (SW, Spain). Fish Physiology and Biochemistry, 39 (3):523–545, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9717-y.

Adebiyi:2013:PSS

[439] Fatimat Adenike Adebiyi, Siti Shapor Siraj, and Annie Christianus. Plasma sex steroid hormonal profile and gonad histology during the annual reproductive cycle of river catfish Hemibagrus nemurus (Valenciennes, 1840) in captivity. Fish Physiology and Biochemistry, 39 (3):547–557, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9718-x.

Feng:2013:EDH

[440] Lin Feng, Bo Zhao, and Xiaoqiu Zhou. Effects of dietary histidine on antioxidant capacity in juvenile Jian carp (Cyprinus carpio var. Jian). Fish Physiology and Biochemistry, 39(3):559–571, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9719-9.

Villeret:2013:PBA

[441] Mélanie Villeret, Sabrina Jolly, and Wilfried Sanchez. A potential biomarker of androgen exposure in European bullhead (Cottus sp.) kidney. Fish Physiology and Biochemistry, 39(3):573–580, June 2013. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9720-3.

Sanchez-Muros:2013:ECH

[442] María José Sánchez-Muros, Salvador Villacreces, and Fernando García-Barroso. Effects of chemical and handling exposure on fatty acids, oxidative stress and morphological welfare indicators in gilt-head sea bream (Sparus aurata). Fish Physiology and Biochemistry, 39(3):581– 591, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9721-2.

Chen:2013:EDV

[443] Yong-Jun Chen, Yong-Jian Liu, and Yun-Qiang Zhang. Effect of dietary vitamin E and selenium supplementation on growth, body composition, and antioxidant defense mechanism in juvenile largemouth bass (Micropterus salmoide) fed oxidized fish oil. Fish Physiology and Biochemistry, 39(3):593–604, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9722-1.

Li:2013:ECM

[444] Yun Li, Xiaochun Liu, and Haoran Lin. Effects of cysteamine on mRNA levels of growth hormone and its receptors and growth in orange-spotted grouper (Epinephelus coioides). Fish Physiology and Biochemistry, 39 (3):605–613, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9723-0.

Mosca:2013:HSP

[445] Francesco Mosca, Nicla Romano, and Pietro G. Tiscar. Heat shock protein 70 kDa (HSP70) increase in sea bass (Dicentrarchus labrax, L 1758) thymus after vaccination against Listonella anguillarum. Fish Physiology and Biochemistry, 39(3):615–626, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9724-z.

Ganesh:2013:NAS

[446] C. B. Ganesh and Ambarisha Chabbi. Naltrexone attenuates stressinduced suppression of LH secretion in the pituitary gland in the cichlid fish Oreochromis mossambicus: evidence for the opioidergic mediation of reproductive stress response. Fish Physiology and Biochemistry, 39 (3):627–636, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9725-y.

Hu:2013:PCG

[447] Wei Hu, Luo Zhi, and Cai-Xia Liu. Purification and characterization of glucose 6-phosphate dehydrogenase (G6PD) from grass carp (Ctenopharyngodon idella) and inhibition effects of several metal ions on G6PD activity in vitro. Fish Physiology and Biochemistry, 39 (3):637–647, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9726-x.

Sassi:2013:CEA

[448] Asma Sassi, Maria José Darias, and Enric Gisbert. Cadmium exposure affects the expression of genes involved in skeletogenesis and stress response in gilthead sea bream larvae. Fish Physiology and Biochemistry, 39(3):649–659, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9727-9.

Perez-Jimenez:2013:EDS

[449] Amalia Pérez-Jiménez, Helena Peres, and Aires Oliva-Teles. Effects of diet supplementation with white tea and methionine on lipid metabolism of gilthead sea bream juveniles (Sparus aurata). Fish Physiology and Biochemistry, 39(3):661–670, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9728-8.

Li:2013:CCE

[450] Wenge Li, Jiaren Zhang, and Haishen Wen. Cloning, characterization and expression of estrogen receptor beta in the male half-smooth tongue sole, Cynoglossus semilaevis. Fish Physiology and Biochemistry, 39(3):671–682, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9729-7.

Faria:2013:IPG

[451] M. T. Faria, R. F. Carvalho, and P. Bartolini. Isolation of the pituitary gonadotrophic α-subunit hormone of the giant Amazonian fish: pirarucu (Arapaima gigas). Fish Physiology and Biochemistry, 39(3): 683–693, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9730-1.

delPozo:2013:DRB

[452] A. del Pozo, L. M. Vera, and F. J. Sánchez-Vázquez. Daily rhythms of blood glucose differ in diurnal and nocturnal European sea bass (Dicentrarchus labrax L.) undergoing seasonal phase inversions. Fish Physiology and Biochemistry, 39(3):695–699, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9732-z.

Higaki:2013:ETO

[453] Shogo Higaki, Yoshie Koyama, and Tatsuyuki Takada. Establishment of testicular and ovarian cell lines from Honmoroko (Gnathopogon caerulescens). Fish Physiology and Biochemistry, 39(3):701–711, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9733-y. See erratum [509].

McConkey:2013:CND

[454] Sandra McConkey, Janet Saunders, and David J. Speare. Comparison of NADH-dependent cytochrome b5 reductase activity and in vitro methemoglobin induction by sodium nitrite in Oncorhynchus mykiss, Salmo salar, and Salvelinus fontinalis. Fish Physiology and Biochemistry, 39 (3):713–719, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9734-x.

Ahmad:2013:PMP

[455] Riaz Ahmad and Absar ul Hasnain. Peptide mapping of polymorphic myosin heavy chain isoforms in different muscle types of some freshwater teleosts. Fish Physiology and Biochemistry, 39(3):721–731, June 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9735-9.

Kumar:2013:EAH

[456] Saurav Kumar, R. P. Raman, and Abhay Kumar. Effect of azadirachtin on haematological and biochemical parameters of Argulus-infested goldfish Carassius auratus (Linn. 1758). Fish Physiology and Biochemistry, 39(4):733–747, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9736-8.

Madenjian:2013:RWE

[457] Charles P. Madenjian and Chunfang Wang. Reevaluation of a walleye (Sander vitreus) bioenergetics model. Fish Physiology and Biochemistry, 39(4):749–754, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9737-7.

Kondera:2013:CCR

[458] Elzbieta Kondera and Malgorzata Witeska. Cadmium and copper reduce hematopoietic potential in common carp (Cyprinus carpio L.) head kidney. Fish Physiology and Biochemistry, 39(4):755–764, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9738-6.

Fuentes:2013:ISS

[459] Eduardo N. Fuentes, Diego Safian, and Alfredo Molina. Isolation and selection of suitable reference genes for real-time PCR analyses in the skeletal muscle of the fine flounder in response to nutritional status: assessment and normalization of gene expression of growth-related genes. Fish Physiology and Biochemistry, 39(4):765–777, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9739-5.

Reza:2013:ERP

[460] A. H. M. M. Reza, S. F. Rakhi, and Z. Hossain. Enhancement of reproductive performances of gangetic leaffish, Nandus nandus through up regulation of serum ca^{2+} concentration, improved morphological alteration of liver and ovary with dietary polyunsaturated fatty acids. Fish Physiology and Biochemistry, 39(4):779–791, August 2013. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-012-9740-z.

Wang:2013:MCG

[461] Lu Wang, Na Shang, and Heping Dai. Molecular cloning of grass carp (Ctenopharyngodon idellus) T-bet and GATA-3, and their expression profiles with IFN-γ in response to grass carp reovirus (GCRV) infection. Fish Physiology and Biochemistry, 39(4):793–805, August 2013. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9741-y.

Ayala:2013:MDB

[462] Maria D. Ayala, Emilia Abellán, and Octavio M. López-Albors. Muscle development and body growth in larvae and early post-larvae of shi drum, Umbrina cirrosa L., reared under different larval photoperiod: muscle structural and ultrastructural study. Fish Physiology and Biochemistry, 39(4):807–827, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9742-x.

Yuan:2013:CCE

[463] Tian Yuan, Chuang Li, and Heng Xu. Cloning, characterization and expression of the LECT2 gene in grass carp. Fish Physiology and Biochemistry, 39(4):829–835, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9743-9.

Golombieski:2013:NPE

[464] Jaqueline Ineu Golombieski, Gessi Koakoski, and Bernardo Baldisserotto. Nitrogenous and phosphorus excretions in juvenile silver catfish (Rhamdia quelen) exposed to different water hardness, humic acid, and pH levels. Fish Physiology and Biochemistry, 39(4):837–849, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9744-8.

Palermo:2013:DEA

[465] Francesco A. Palermo, Paolo Cocci, and Gilberto Mosconi. Dietary Aloe vera components’ effects on cholesterol lowering and estrogenic responses in juvenile goldfish, Carassius auratus. Fish Physiology and Biochemistry, 39(4):851–861, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9745-7.

Gozdowska:2013:NIA

[466] Magdalena Gozdowska, Marek Ślebioda, and Ewa Kulczykowska. Neuropeptides isotocin and arginine vasotocin in urophysis of three fish species. Fish Physiology and Biochemistry, 39(4):863–869, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9746-6.

Bjorlykke:2013:SAS

[467] Gry Aletta Bjørlykke, Bjørn Olav Kvamme, and Erik Slinde. Slaughter of Atlantic salmon (Salmo salar L.) in the presence of carbon monoxide. Fish Physiology and Biochemistry, 39(4):871–879, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9747-5.

Xue:2013:CSE

[468] Yang Xue, Jinliang Zhao, and Wei Miao. Cloning and spatiotemporal expression of pepsinogen and gastric proton pump genes from mandarin fish (Siniperca chuatsi) during early ontogeny. Fish Physiology and Biochemistry, 39(4):881–893, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9748-4.

Murashita:2013:PCO

[469] Koji Murashita, Hirofumi Furuita, and Hideki Tanaka. Partial characterization and ontogenetic development of pancreatic digestive enzymes in Japanese eel Anguilla japonica larvae. Fish Physiology and Biochemistry, 39(4):895–905, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9749-3.

Hayashida:2013:EAS

[470] Kazufumi Hayashida, Hisaya Nii, and Hiroshi Ueda. Effects of anesthesia and surgery on U_{crit} performance and MO_{2} in chum salmon, Oncorhynchus keta. Fish Physiology and Biochemistry, 39(4):907–915, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9750-x.

Guerreiro:2013:DSR

[471] Pedro M. Guerreiro, Rita Costa, and Deborah M. Power. Dynamics of scale regeneration in seawater- and brackish water-acclimated sea bass, Dicentrarchus labrax. Fish Physiology and Biochemistry, 39(4): 917–930, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9751-9.

Xiao:2013:MCE

[472] Jun Xiao, Yongju Luo, and Xi Gan. Molecular cloning of vasa gene and the effects of LHRH-A on its expression in blue tilapia Oreochromis aureus. Fish Physiology and Biochemistry, 39(4):931–940, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9752-8.

Eroldogan:2013:FAM

[473] Tufan O. Eroldoğan, Asuman H. Yılmaz, and Pınar Mumoğullarında. Fatty acid metabolism in European sea bass (Dicentrarchus labrax): effects of n-6 PUFA and MUFA in fish oil replaced diets. Fish Physiology and Biochemistry, 39(4):941–955, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9753-7.

Paris-Palacios:2013:CAF

[474] Séverine Paris-Palacios, Laurence Delahaut, and Sylvie Biagianti-Risbourg. Catalasic activity in fish liver: improvement of the UV to visible analytic method. Fish Physiology and Biochemistry, 39(4): 957–966, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9754-6.

Suzer:2013:FCD

[475] Cüneyt Suzer, H. Okan Kamacı, and Şahin Saka. Functional changes in digestive enzyme activities of meagre (Argyrosomus regius; Asso, 1801) during early ontogeny. Fish Physiology and Biochemistry, 39(4): 967–977, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9755-5.

Volkoff:2013:EAI

[476] Hélène Volkoff. The effects of amphetamine injections on feeding behavior and the brain expression of orexin, CART, tyrosine hydroxylase (TH) and thyrotropin releasing hormone (TRH) in goldfish (Carassius auratus). Fish Physiology and Biochemistry, 39(4):979–991, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9756-4.

Reyes-Tomassini:2013:GIE

[477] José Reyes-Tomassini. GnRH isoforms expression in relation to the gonadal cycle and to dominance rank in the gilthead seabream, Sparus aurata. Fish Physiology and Biochemistry, 39(4):993–1005, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9757-3.

Maharana:2013:IMM

[478] Jitendra Maharana, Banikalyan Swain, and Mrinal Samanta. Identification of MDP (muramyl dipeptide)-binding key domains in NOD2 (nucleotide-binding and oligomerization domain-2) receptor of Labeo rohita. Fish Physiology and Biochemistry, 39(4):1007–1023, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9758-2.

Zhao:2013:SGF

[479] Feng Zhao, Yu Wang, and Jianyi Liu. Survival, growth, food conversion efficiency and plasma osmolality of juvenile Siganus guttatus (Bloch, 1787): experimental analyses of salinity effects. Fish Physiology and Biochemistry, 39(4):1025–1030, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9759-1.

Elseady:2013:AEB

[480] Y. Elseady and E. Zahran. Ameliorating effect of β-carotene on antioxidant response and hematological parameters of mercuric chloride toxicity in Nile tilapia (Oreochromis niloticus). Fish Physiology and Biochemistry, 39(4):1031–1041, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-012-9760-8.

Urbatzka:2013:TEP

[481] R. Urbatzka, S. Galante-Oliveira, and I. Cunha. Tissue expression of PPAR-alpha isoforms in Scophthalmus maximus and transcriptional response of target genes in the heart after exposure to WY-14643. Fish Physiology and Biochemistry, 39(4):1043–1055, August 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9761-7.

Shivashri:2013:HAC

[482] C. Shivashri, T. Rajarajeshwari, and P. Rajasekar. Hepatoprotective action of celery (Apium graveolens) leaves in acetaminophen-fed freshwater fish (Pangasius sutchi). Fish Physiology and Biochemistry, 39(5): 1057–1069, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9762-6.

Yang:2013:DTC

[483] Huijun Yang, Lixia Tian, and Yongjian Liu. Dietary taurine can improve the hypoxia-tolerance but not the growth performance in juvenile grass carp Ctenopharyngodon idellus. Fish Physiology and Biochemistry, 39(5): 1071–1078, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-012-9763-5.

Handeland:2013:PDS

[484] S. O. Handeland, A. K. Imsland, and M. Porter. Physiology during smoltification in Atlantic salmon: effect of melatonin implants. Fish Physiology and Biochemistry, 39(5):1079–1088, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-012-9765-3.

Williams:2013:UEP

[485] Laura Jay Williams, Matthew D. Campbell, and James A. Sulikowski. Using estradiol and progesterone concentrations to assess individual variability in the reproductive cyclicity of captive female little skates, Leucoraja erinacea, from the western Gulf of Maine. Fish Physiology and Biochemistry, 39(5):1089–1099, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9766-2.

Fridman:2013:SDA

[486] S. Fridman, K. J. Rana, and J. E. Bron. Structural differentiation of apical openings in active mitochondria-rich cells during early life stages of Nile tilapia (Oreochromis niloticus L.) as a response to osmotic challenge. Fish Physiology and Biochemistry, 39(5):1101–1114, October 2013. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-012-9767-1.

Torres:2013:EDE

[487] Leticia Torres, Carl E. Orazio, and Reynaldo Patiño. Effects of dietary exposure to brominated flame retardant BDE-47 on thyroid condition, gonadal development and growth of zebrafish. Fish Physiology and Biochemistry, 39(5):1115–1128, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-012-9768-0.

Geffroy:2013:NIR

[488] Benjamin Geffroy, Yann Guiguen, and Agnès Bardonnet. New insights regarding gonad development in European eel: evidence for a direct ovarian differentiation. Fish Physiology and Biochemistry, 39(5): 1129–1140, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9769-7.

Costas:2013:IEH

[489] Benjamín Costas, Cláudia Aragão, and Luís E. C. Conceição. Interactive effects of a high-quality protein diet and high stocking density on the stress response and some innate immune parameters of Senegalese sole Solea senegalensis. Fish Physiology and Biochemistry, 39(5): 1141–1151, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9770-1.

Liu:2013:IEH

[490] Shasha Liu, Kecheng Zhu, and Huanling Wang. Identification of HIF-1α promoter and expression regulation of HIF-1α gene by LPS and hypoxia in zebrafish. Fish Physiology and Biochemistry, 39(5):1153– 1163, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9771-0.

Edwards:2013:SRM

[491] Thea M. Edwards, Hilary D. Miller, and Louis J. Guillette, Jr. Seasonal reproduction of male Gambusia holbrooki (eastern mosquitofish) from two Florida lakes. Fish Physiology and Biochemistry, 39(5): 1165–1180, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9772-z.

Izvekova:2013:VAD

[492] Galina I. Izvekova, Mikhail M. Solovyev, and Evgeny I. Izvekov. Variations in the activity of digestive enzymes along the intestine of the burbot Lota lota expressed by different methods. Fish Physiology and Biochemistry, 39(5):1181–1193, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-013-9773-y.

Derayat:2013:GGD

[493] Amid Derayat, Árni Magnússon, and Björn Björnsson. Growth and gonadal development in diploid and triploid Atlantic cod (Gadus morhua). Fish Physiology and Biochemistry, 39(5):1195–1203, October 2013. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9775-9.

Bastos:2013:MAC

[494] F. F. Bastos, S. A. L. Tobar, and V. L. F. Cunha Bastos. Melatonin affects conjugation of 4-hydroxynonenal with glutathione in liver of pacu, a hypoxia-tolerant fish. Fish Physiology and Biochemistry, 39(5): 1205–1214, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9776-8.

Paredes:2013:IEP

[495] Marco Paredes, Katerina Gonzalez, and Enrique Montiel-Eulefi. Immunomodulatory effect of prolactin on Atlantic salmon (Salmo salar) macrophage function. Fish Physiology and Biochemistry, 39(5):1215– 1221, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9777-7.

Martins:2013:EDA

[496] Dulce Alves Martins, Filipa Rocha, and Gonzalo Martínez-Rodríguez. Effects of dietary arachidonic acid on cortisol production and gene expression in stress response in Senegalese sole (Solea senegalensis) post-larvae.

Fish Physiology and Biochemistry, 39(5):1223–1238, October 2013. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9778-6.

Sun:2013:MCE

[497] Airong Sun, Jian Li, and Qi Wang. Molecular cloning and expression analysis of cytochrome P450 3A gene in the turbot Scophthalmus maximus. Fish Physiology and Biochemistry, 39(5):1239–1251, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9779-5.

Sarameh:2013:PCM

[498] Sara Pourhosein Sarameh, Bahram Falahatkar, and Iraj Efatpanah. Physiological changes in male and female pikeperch Sander lucioperca (Linnaeus, 1758) subjected to different photoperiods and handling stress during the reproductive season. Fish Physiology and Biochemistry, 39(5): 1253–1266, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9780-z.

Mu:2013:LPA

[499] Yinnan Mu, Xiang Wan, and Xinhua Chen. Liver proteomic analysis of the large yellow croaker (Pseudosciaena crocea) following polyriboinosinic:polyribocytidylic acid induction. Fish Physiology and Biochemistry, 39(5):1267–1276, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9781-y.

Om:2013:AMT

[500] Ahmad Daud Om, Safiah Jasmani, and A. B. Abol-Munafi. Application MALDI TOF on protein identification of vitellogenin in giant grouper (Epinephelus lanceolatus). Fish Physiology and Biochemistry, 39(5): 1277–1286, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9782-x.

Cottone:2013:CRW

[501] Erika Cottone, Valentina Pomatto, and Maria Fosca Franzoni. Cannabinoid receptors are widely expressed in goldfish: molecular cloning of a CB2-like receptor and evaluation of CB1 and CB2 mRNA expression profiles in different organs. Fish Physiology and Biochemistry, 39(5): 1287–1296, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9783-9.

Ciji:2013:NIA

[502] A. Ciji, N. P. Sahu, and M. S. Akhtar. Nitrite-induced alterations in sex steroids and thyroid hormones of Labeo rohita juveniles: effects of dietary vitamin E and l-tryptophan. Fish Physiology and Biochemistry, 39(5): 1297–1307, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9784-8.

Bieczynski:2013:ABE

[503] Flavia Bieczynski, Virginia A. Bianchi, and Carlos M. Luquet. Accumulation and biochemical effects of microcystin-LR on the Patagonian pejerrey (Odontesthes hatcheri) fed with the toxic cyanobacteria Microcystis aeruginosa. Fish Physiology and Biochemistry, 39(5): 1309–1321, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9785-7.

Manju:2013:PED

[504] Maniyan Manju, Appiyathu Saraswathy Vijayasree, and Oommen Vilaverthottathil Oommen. Protective effect of dietary curcumin in Anabas testudineus (Bloch) with a special note on DNA fragmentation assay on hepatocytes and micronucleus assay on erythrocytes in vivo. Fish Physiology and Biochemistry, 39(5):1323–1330, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:

//link.springer.com/article/10.1007/s10695-013-9786-6.

Zacarias-Soto:2013:ODA

[505] Magali Zacarias-Soto, Benjamín Barón-Sevilla, and Juan P. Lazo. Ontogeny and distribution of alkaline and acid phosphatases in the digestive system of California halibut larvae (Paralichthys californicus). Fish Physiology and Biochemistry, 39(5):1331–1339, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9787-5.

Xia:2013:HES

[506] Xiaohua Xia, Ping Nan, and Zhongjie Chang. Homologue of Sox10 in Misgurnus anguillicaudatus: sequence, expression pattern during early embryogenesis. Fish Physiology and Biochemistry, 39(5):1341–1351, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-013-9788-4.

Shin:2013:PCR

[507] Jihye Shin, Mi Ae Kim, and Young Chang Sohn. Production and characterization of recombinant Manchurian trout thyrotropin. Fish Physiology and Biochemistry, 39(5):1353–1363, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9789-3.

deRodriganez:2013:VEE

[508] M. A. Sáenz de Rodrigáñez, J. Fuentes, and L. Ribeiro. In vitro evaluation of the effect of a high plant protein diet and nucleotide supplementation on intestinal integrity in meagre (Argyrosomus regius). Fish Physiology and Biochemistry, 39(5):1365–1370, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9790-x.

Higaki:2013:EET

[509] Shogo Higaki, Yoshie Koyama, and Tatsuyuki Takada. Erratum to: Establishment of testicular and ovarian cell lines from Honmoroko (Gnathopogon caerulescens). Fish Physiology and Biochemistry, 39(5):1371, October 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-012-9764-4. See [453].

Garina:2013:LTE

[510] D. V. Garina, A. K. Smirnov, and V. Kuz’mina. The long-term effect of serotonin on the thermoregulatory behavior in juvenile cyprinidae (Cyprinus carpio and Carassius auratus). Fish Physiology and Biochemistry, 39(6):1373–1376, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-013-9791-9.

Wang:2013:CMC

[511] T. T. Wang, N. Wang, and S. L. Chen. Cloning, molecular characterization and expression analysis of heat shock cognate 70 (Hsc70) cDNA from turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 39(6):1377–1386, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-013-9792-8.

Akhtar:2013:TTO

[512] M. S. Akhtar, A. K. Pal, and P. C. Mahanta. Thermal tolerance, oxygen consumption and haemato-biochemical variables of Tor putitora juveniles acclimated to five temperatures. Fish Physiology and Biochemistry, 39(6):1387–1398, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-013-9793-7.

Zhu:2013:ECF

[513] Dong-Mei Zhu, Kun Yang, and Wen Song. Establishment and characterization of a fin cell line from blunt snout bream, Megalobrama amblycephala. Fish Physiology and Biochemistry, 39(6):1399–1410, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-013-9794-6.

Ai-Jun:2013:CPC

[514] Ma Ai-Jun, Huang Zhi-hui, and Wang Xin-An. Changes in protein composition of epidermal mucus in turbot Scophthalmus maximus (L.) under high water temperature. Fish Physiology and Biochemistry, 39(6): 1411–1418, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9795-5.

Mu:2013:CEAb

[515] Wei J. Mu, Hai S. Wen, and Feng He. Cloning and expression analysis of Foxl2 during the reproductive cycle in Korean rockfish, Sebastes schlegeli. Fish Physiology and Biochemistry, 39(6):1419–1430, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9796-4.

Saravanan:2013:ACE

[516] Manoharan Saravanan, Mathan Ramesh, and Rakpong Petkam. Alteration in certain enzymological parameters of an Indian major carp, Cirrhinus mrigala exposed to short- and long-term exposure of clofibric acid and diclofenac. Fish Physiology and Biochemistry, 39(6):1431– 1440, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9797-3.

Hixson:2013:ERF

[517] Stefanie M. Hixson, Christopher C. Parrish, and Derek M. Anderson. Effect of replacement of fish oil with camelina (Camelina sativa) oil on growth, lipid class and fatty acid composition of farmed juvenile Atlantic cod (Gadus morhua). Fish Physiology and Biochemistry, 39(6): 1441–1456, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9798-2.

Lee:2013:DTT

[518] Kuan-Shern Lee, Kah-Hay Yuen, and Wing-Keong Ng. Deposition of tocopherol and tocotrienol in the tissues of red hybrid tilapia, Oreochromis sp., fed vitamin E-free diets supplemented with different plant oils. Fish Physiology and Biochemistry, 39(6):1457–1471, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9799-1.

Perini:2013:PSS

[519] Violeta da Rocha Perini, Alessandro Loureiro Paschoalini, and Elizete Rizzo. Profiles of sex steroids, fecundity and spawning of a migratory characiform fish from the Paraguay-paraná basin: a comparative study in a three-river system. Fish Physiology and Biochemistry, 39(6): 1473–1484, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9800-z.

Campos:2013:RTA

[520] Catarina Campos, M. Filipa Castanheira, and Luís E. C. Conceição. Rearing temperature affects Senegalese sole (Solea senegalensis) larvae protein metabolic capacity. Fish Physiology and Biochemistry, 39(6): 1485–1496, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9802-x.

Golovanova:2013:EAT

[521] I. L. Golovanova, V. K. Golovanov, and D. D. Pavlov. Effect of ambient temperature increase on intestinal mucosa amylolytic activity in freshwater fish. Fish Physiology and Biochemistry, 39(6):1497– 1504, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9803-9.

Hosseinifard:2013:IEC

[522] Seyyed Mehdi Hosseinifard, Asieh Ahmadpour, and Arezo Ebrahimpour. Immunomodulatory effect of cimetidine in common carp (Cyprinus carpio L.). Fish Physiology and Biochemistry, 39(6):1505–1511, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-013-9804-8.

Gupta:2013:SML

[523] S. K. Gupta, A. K. Pal, and A. K. Prusty. Supplementation of microbial levan in the diet of Cyprinus carpio fry (Linnaeus, 1758) exposed to sublethal toxicity of fipronil: effect on growth and metabolic responses. Fish Physiology and Biochemistry, 39(6):1513–1524, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9805-7.

Lin:2013:DCL

[524] Heng-Ju Lin, Shu-Hua Lee, and Jyh-Yih Chen. Development of cre–loxP technology in zebrafish to study the regulation of fish reproduction. Fish Physiology and Biochemistry, 39(6):1525–1539, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9806-6.

Zhang:2013:ILG

[525] Yuntong Zhang, Junling Zhang, and Xiaozhu Wang. Insulin-like growth factor binding protein-2 (IGFBP-2) in Japanese flounder, Paralichthys olivaceus: molecular cloning, expression patterns and hormonal regulation during metamorphosis. Fish Physiology and Biochemistry, 39(6): 1541–1554, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9807-5.

Latorre:2013:EPA

[526] Rocco Latorre, Maurizio Mazzoni, and Paolo Clavenzani. Enteroendocrine profile of α-transducin immunoreactive cells in the gastrointestinal tract of the European sea bass (Dicentrarchus labrax). Fish Physiology and Biochemistry, 39(6):1555–1565, December 2013. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-013-9808-4.

Robles:2013:EPP

[527] R. Robles, A. B. Lozano, and F. J. Moyano. Effect of partially protected butyrate used as feed additive on growth and intestinal metabolism in sea bream (Sparus aurata). Fish Physiology and Biochemistry, 39(6): 1567–1580, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9809-3.

Silveira:2013:FCJ

[528] Jenniffer Silveira, Carlos Peres Silva, and Débora Machado Fracalossi. Freshwater catfish jundiá (Rhamdia quelen) larvae are prepared to digest inert feed at the exogenous feeding onset: physiological and histological assessments. Fish Physiology and Biochemistry, 39(6):1581– 1590, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9810-x.

Rodrigues:2013:IWA

[529] E. Rodrigues, Jr., M. Feijó-Oliveira, and E. Rodrigues. Interaction of warm acclimation, low salinity, and trophic fluoride on plasmatic constituents of the Antarctic fish Notothenia rossii Richardson, 1844. Fish Physiology and Biochemistry, 39(6):1591–1601, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9811-9.

Olivares-Rubio:2013:POA

[530] Hugo F. Olivares-Rubio, M. Lysset Martínez-Torres, and Armando Vega-López. Pro-oxidant and antioxidant responses in the liver and kidney of wild Goodea gracilis and their relation with halomethanes bioactivation. Fish Physiology and Biochemistry, 39(6):1603–1617, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9812-8.

Hosono:2013:DCR

[531] Masahiro Hosono, Shigeki Sugawara, and Kazuo Nitta. Domain composition of rhamnose-binding lectin from shishamo smelt eggs and its carbohydrate-binding profiles. Fish Physiology and Biochemistry, 39(6): 1619–1630, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9814-6.

Ahmad:2013:DFO

[532] Wan A. R. Wan Ahmad, David A. J. Stone, and Kathryn A. Schuller. Dietary fish oil replacement with palm or poultry oil increases fillet oxidative stability and decreases liver glutathione peroxidase activity in barramundi (Lates calcarifer). Fish Physiology and Biochemistry, 39(6): 1631–1640, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9815-5.

Song:2013:ECC

[533] Yu-Feng Song, Zhi Luo, and Qi-Liang Chen. Effects of calcium and copper exposure on lipogenic metabolism, metal element compositions and histology in Synechogobius hasta. Fish Physiology and Biochemistry, 39(6):1641–1656, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-013-9816-4.

Ojopagogo:2013:PSR

[534] Yetunde Adedolapo Ojopagogo, Isaac Olusanjo Adewale, and Adeyinka Afolayan. Preliminary studies on the renaturation of denatured catfish (Clarias gariepinus) glutathione transferase. Fish Physiology and Biochemistry, 39(6):1657–1663, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9817-3.

Kaptaner:2013:AGC

[535] Burak Kaptaner and Ertuğrul Kankaya. Analysis of germ cell proliferation, apoptosis, and androgenesis in the Lake Van fish (Chalcalburnus tarichi) during testicular development. Fish Physiology and Biochemistry, 39(6):1665–1679, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-013-9818-2.

Reyes-Tomassini:2013:EGI

[536] Jose Reyes-Tomassini, Ten Tsao Wong, and Yonathan Zohar. Erratum to: GnRH isoforms expression in relation to the gonadal cycle and to dominance rank in the gilthead seabream, Sparus aurata. Fish Physiology and Biochemistry, 39(6):1681, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9801-y.

Izvekova:2013:EVA

[537] Galina I. Izvekova, Mikhail M. Solovyev, and Evgeny I. Izvekov. Erratum to: Variations in the activity of digestive enzymes along the intestine of the burbot Lota lota expressed by different methods. Fish Physiology and Biochemistry, 39(6):1683–1684, December 2013. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9813-7.

Talbott:2014:COH

[538] Mariah J. Talbott, Sarah A. Servid, and Molly A. H. Webb. Confirmation of ovarian homogeneity in post-vitellogenic cultured white sturgeon, Acipenser transmontanus. Fish Physiology and Biochemistry, 40 (1):1–7, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9774-x.

Kondera:2014:HAC

[539] Elżbieta Kondera, Katarzyna Ługowska, and Piotr Sarnowski. High affinity of cadmium and copper to head kidney of common carp (Cyprinus carpio L.). Fish Physiology and Biochemistry, 40(1):9–22, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9819-1.

Miyazaki:2014:RGC

[540] Taeko Miyazaki. Retinal ganglion cell topography in juvenile Pacific bluefin tuna Thunnus orientalis (Temminck and Schlegel). Fish Physiology and Biochemistry, 40(1):23–32, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9820-8.

Manuel:2014:SAC

[541] Remy Manuel, Jeroen Boerrigter, and Hans van de Vis. Stress in African catfish (Clarias gariepinus) following overland transportation. Fish Physiology and Biochemistry, 40(1):33–44, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9821-7.

Chen:2014:ICF

[542] Yong Chen, Narayan Prasad Pandit, and Jiale Li. Identification, characterization and feeding response of peptide YYb (PYYb) gene in grass carp (Ctenopharyngodon idellus). Fish Physiology and Biochemistry, 40 (1):45–55, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9822-6.

Alishahi:2014:HIR

[543] M. Alishahi, M. Karamifar, and M. Zarei. Hemato-immunological responses of Heros severus fed diets supplemented with different levels of Dunaliella salina. Fish Physiology and Biochemistry, 40(1): 57–65, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9823-5.

Arjamaa:2014:HIR

[544] Olli Arjamaa, Olli Vuolteenaho, and Mikko Nikinmaa. Hypoxia increases the release of salmon cardiac peptide (sCP) from the heart of rainbow trout (Oncorhynchus mykiss) under constant mechanical load in vitro. Fish Physiology and Biochemistry, 40(1):67–73, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9824-4.

Ovais:2014:PID

[545] M. Ovais, Fraz Ahmed, and Safia Sumoona. Prostanoids-induced dispersion in the melanophores of a carp Labeo rohita (Ham.). Fish Physiology and Biochemistry, 40(1):75–81, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9825-3.

Ghosh:2014:MCI

[546] Debasree Ghosh and Dipak Kumar Mandal. Mercuric chloride induced toxicity responses in the olfactory epithelium of Labeo rohita (Hamilton): a light and electron microscopy study. Fish Physiology and Biochemistry, 40(1):83–92, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9826-2.

Yang:2014:CMC

[547] Jixuan Yang, Qingsong Tan, and Lei Pan. Cloning and molecular characterization of cationic amino acid transporter y^{+} LAT1 in grass carp (Ctenopharyngodon idellus). Fish Physiology and Biochemistry, 40(1): 93–104, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9827-1.

Das:2014:VIO

[548] Puranjan Das, Kousik Pramanick, and B. R. Maiti. In vitro induction of oocyte maturation and steroidogenesis by gonadotropins, insulin, calcitonin and growth factor in an estuarine flat head grey mullet, Mugil cephalus L. Fish Physiology and Biochemistry, 40(1): 105–116, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9828-0.

Jourdehi:2014:CSD

[549] A. Yousefi Jourdehi, M. Sudagar, and M. A. Yazdani. Comparative study of dietary soy phytoestrogens genistein and equol effects on growth parameters and ovarian development in farmed female beluga sturgeon, Huso hus o. Fish Physiology and Biochemistry, 40(1): 117–128, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9829-z.

Gaulke:2014:ALO

[550] Greg L. Gaulke, Clark E. Dennis III, and Cory D. Suski. Acclimation to a low oxygen environment alters the hematology of largemouth bass (Micropterus salmoides). Fish Physiology and Biochemistry, 40(1): 129–140, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9830-6.

Menezes:2014:EEI

[551] Charlene Menezes, Jossiele Leitemperger, and Vania Lucia Loro. Evaluation of the effects induced by dietary diphenyl diselenide on common carp Cyprinus carpio. Fish Physiology and Biochemistry, 40(1): 141–149, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9831-5.

Witeska:2014:ECC

[552] Malgorzata Witeska, Piotr Sarnowski, and Ewelina Kowal. The effects of cadmium and copper on embryonic and larval development of ide Leuciscus idus L. Fish Physiology and Biochemistry, 40(1): 151–163, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9832-4.

Zupa:2014:OCR

[553] R. Zupa, M. Piccinno, and R. L. Sciorsci. Ovarian contractility in reared gilthead seabream (Sparus aurata L.) in different phases of the reproductive cycle. Fish Physiology and Biochemistry, 40(1):165–171, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-013-9833-3.

Ludvigsen:2014:EPA

[554] Stian Ludvigsen, Niels C. Stenklev, and Øyvind Aas-Hansen. Evoked potentials in the Atlantic cod following putatively innocuous and putatively noxious electrical stimulation: a minimally invasive approach. Fish Physiology and Biochemistry, 40(1):173–181, February 2014. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9834-2.

Leveelahti:2014:RRA

[555] Lotta Leveelahti, Kalle T. Rytkönen, and Mikko Nikinmaa. Revisiting redox-active antioxidant defenses in response to hypoxic challenge in both hypoxia-tolerant and hypoxia-sensitive fish species. Fish Physiology and Biochemistry, 40(1):183–191, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9835-1.

Huang:2014:GDF

[556] Chun Xiao Huang, Xin Lan Wei, and Huan Ling Wang. Growth differentiation factor 9 of Megalobrama amblycephala: molecular characterization and expression analysis during the development of early embryos and growing ovaries. Fish Physiology and Biochemistry, 40(1): 193–203, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9836-0.

Jeong:2014:SCA

[557] Su-Young Jeong, Jin-Hyoung Kim, and Kyung-Nam Han. Salinity changes in the anadromous river pufferfish, Takifugu obscurus, mediate gene regulation. Fish Physiology and Biochemistry, 40(1):205–219, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-013-9837-z.

Zhang:2014:CFH

[558] Lili Zhang, Chengfei Sun, and Yuanyuan Tian. Characterization of four heat-shock protein genes from Nile tilapia (Oreochromis niloticus) and demonstration of the inducible transcriptional activity of Hsp70 promoter. Fish Physiology and Biochemistry, 40(1):221–233, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9838-y.

Wang:2014:HED

[559] Meng-Yu Wang, Qiu-Hong Guo, and Yu-Feng Wang. HIRA is essential for the development of gibel carp. Fish Physiology and Biochemistry, 40 (1):235–244, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9839-x.

Mela:2014:MDT

[560] Maritana Mela, Francisco Filipak Neto, and Ciro Alberto de Oliveira Ribeiro. Mercury distribution in target organs and biochemical responses after subchronic and trophic exposure to neotropical fish Hoplias malabaricus. Fish Physiology and Biochemistry, 40(1):245–256, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9840-4.

Ditlecadet:2014:GSF

[561] Delphine Ditlecadet and William R. Driedzic. Glycerol synthesis in freeze-resistant rainbow smelt: towards the characterization of a key enzyme glycerol-3-phosphatase. Fish Physiology and Biochemistry, 40(1): 257–266, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9841-3.

Lavado:2014:ESA

[562] Ramon Lavado, Rosaura Aparicio-Fabre, and Daniel Schlenk. Effects of salinity acclimation on the expression and activity of Phase I enzymes (CYP450 and FMOs) in Coho salmon (Oncorhynchus kisutch). Fish Physiology and Biochemistry, 40(1):267–278, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9842-2.

Maltais:2014:PPC

[563] Domynick Maltais, Bernard-Antonin Dupont-Cyr, and Nathalie R. Le François. Purification and partial characterization of vitellogenin from spotted wolffish (Anarhichas minor) and development of an enzymelinked immunosorbent assay for the determination of gender and sexual maturity. Fish Physiology and Biochemistry, 40(1):279–294, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-013-9843-1.

Lobo:2014:DPS

[564] Carmen Lobo, Xabier Moreno-Ventas, and Inés García de La Banda. Dietary probiotic supplementation (Shewanella putrefaciens Pdp11) modulates gut microbiota and promotes growth and condition in Senegalese sole larviculture. Fish Physiology and Biochemistry, 40(1):295– 309, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9844-0.

Terova:2014:PPS

[565] Genciana Terova, Salvatore Pisanu, and Maria Filippa Addis. Proteomic profiling of sea bass muscle by two-dimensional gel electrophoresis and tandem mass spectrometry. Fish Physiology and Biochemistry, 40(1): 311–322, February 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9855-x.

Parodi:2014:AAE

[566] Thaylise V. Parodi, Mauro A. Cunha, and Bernardo Baldisserotto. Anesthetic activity of the essential oil of Aloysia triphylla and effectiveness in reducing stress during transport of albino and gray strains of silver catfish, Rhamdia quelen. Fish Physiology and Biochemistry, 40 (2):323–334, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9845-z.

Stewart:2014:SBF

[567] John Stewart and Julian M. Hughes. Swim bladder function and buoyancy control in pink snapper (Pagrus auratus) and mulloway (Argyrosomus japonicus). Fish Physiology and Biochemistry, 40(2):335–346, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-013-9846-y.

Adeyemi:2014:OSA

[568] Joseph A. Adeyemi. Oxidative stress and antioxidant enzymes activities in the African catfish, Clarias gariepinus, experimentally challenged with Escherichia coli and Vibrio fischeri. Fish Physiology and Biochemistry, 40(2):347–354, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9847-x.

Ozkan-Yilmaz:2014:EDS

[569] Ferbal Özkan-Yılmaz, Arzu Özlüer-Hunt, and Serap Yalın. Effects of dietary selenium of organic form against lead toxicity on the antioxidant system in Cyprinus carpio. Fish Physiology and Biochemistry, 40 (2):355–363, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9848-9.

Falahatkar:2014:LTE

[570] Bahram Falahatkar, Samaneh Poursaeid, and Iraj Efatpanah. Long-term effects of intraperitoneal injection of estradiol-17β on the growth and physiology of juvenile stellate sturgeon Acipenser stellatus. Fish Physiology and Biochemistry, 40(2):365–373, April 2014. CODEN FPBIEP.

ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9849-8.

Khemis:2014:CSP

[571] I. Ben Khemis, N. Hamza, and M. M’Hetli. Comparative study of pikeperch Sander lucioperca (Percidae; Linnaeus, 1758) eggs and larvae from wild females or from captive females fed chopped marine fish. Fish Physiology and Biochemistry, 40(2):375–384, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9850-2.

Takase:2014:DMT

[572] Mai Takase, Masataka Murata, and Hideaki Endo. Development of mediator-type biosensor to wirelessly monitor whole cholesterol concentration in fish. Fish Physiology and Biochemistry, 40(2):385–394, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-013-9851-1.

Williams:2014:MVP

[573] V. N. Williams, B. J. Reading, and C. V. Sullivan. Multiple vitellogenins and product yolk proteins in striped bass, Morone saxatilis: molecular characterization and processing during oocyte growth and maturation. Fish Physiology and Biochemistry, 40(2):395–415, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9852-0.

Vigliano:2014:ESS

[574] Fabricio A. Vigliano, Adolfo M. Araujo, and Lucrecia G. M. Dasso. Effects of sex and season in haematological parameters and cellular composition of spleen and head kidney of pejerrey (Odontesthes bonariensis). Fish Physiology and Biochemistry, 40(2):417–426, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9853-z.

Wade:2014:PMR

[575] Nicholas M. Wade, Sandrine Skiba-Cassy, and Brett D. Glencross. Postprandial molecular responses in the liver of the barramundi, Lates calcarifer. Fish Physiology and Biochemistry, 40(2):427–443, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9854-y.

deFaria:2014:GRO

[576] Marcos Tucunduva de Faria, Maria Fernanda Cury-Boaventura, and José Roberto Machado Cunha da Silva. Generation of reactive oxygen species by leukocytes of Prochilodus lineatus. Fish Physiology and Biochemistry, 40(2):445–455, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9856-9.

Biswas:2014:RCN

[577] Saikat P. Biswas, Arun G. Jadhao, and Nikhil V. Palande. Role of catecholamines and nitric oxide on pigment displacement of the chromatophores of freshwater snakehead teleost fish, Channa punctatus. Fish Physiology and Biochemistry, 40(2):457–467, April 2014. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-013-9857-8.

Lahnsteiner:2014:EKC

[578] Franz Lahnsteiner. The effect of K^{+}, Ca^{2+}, and Mg^{2+} on sperm motility in the perch, Perca fluviatilis. Fish Physiology and Biochemistry, 40 (2):469–480, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9858-7.

Kanani:2014:EGG

[579] Hosna Gholipour Kanani, Zahra Nobahar, and Hojatollah Jafarian. Effect of ginger- and garlic-supplemented diet on growth performance, some hematological parameters and immune responses in juvenile Huso huso. Fish Physiology and Biochemistry, 40(2):481–490, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9859-6.

Cai:2014:ETS

[580] Lu Cai, Guoyong Liu, and Yingping Huang. Effect of temperature on swimming performance of juvenile Schizothorax prenanti. Fish Physiology and Biochemistry, 40(2):491–498, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9860-0.

Feng:2014:ASS

[581] Lin Feng, Shu Zhao, and Xiaoqiu Zhou. Antioxidant status of serum, muscle, intestine and hepatopancreas for fish fed graded levels of biotin. Fish Physiology and Biochemistry, 40(2):499–510, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9861-z.

Murashita:2014:HGB

[582] Koji Murashita, Yasutoshi Yoshiura, and Takeshi Yamamoto. Homologue gene of bile acid transporters ntcp, asbt, and ost-alpha in rainbow trout Oncorhynchus mykiss: tissue expression, effect of fasting, and response to bile acid administration. Fish Physiology and Biochemistry, 40(2):511–525, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9862-y.

Iversen:2014:EAL

[583] Martin H. Iversen and Robert A. Eliassen. The effect of allostatic load on hypothalamic–pituitary–interrenal (HPI) axis before and after secondary vaccination in Atlantic salmon postsmolts (Salmo salar L.). Fish Physiology and Biochemistry, 40(2):527–538, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9863-x.

Korajoki:2014:SCS

[584] Hanna Korajoki and Matti Vornanen. Species- and chamber-specific responses of 12 kDa FK506-binding protein to temperature in fish heart. Fish Physiology and Biochemistry, 40(2):539–549, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9864-9.

Kieffer:2014:TRE

[585] James D. Kieffer, Faith M. Penny, and Vasoula Papadopoulos. Temperature has a reduced effect on routine metabolic rates of juvenile shortnose sturgeon (Acipenser brevirostrum). Fish Physiology and Biochemistry, 40(2):551–559, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9865-8.

Zhang:2014:IKF

[586] Wei-Ni Zhang, Ding-Ping Bai, and Yi-Fan Huang. Inactivation kinetics of formaldehyde on N-acetyl-β-d-glucosaminidase from Nile tilapia (Oreochromis niloticus). Fish Physiology and Biochemistry, 40(2):561– 569, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-013-9866-7.

Chehade:2014:MSL

[587] Chayrra Chehade, Mônica Cassel, and Fabiano Gonçalves Costa. Morphologic study of the liver of lambari (Astyanax altiparanae) with emphasis on the distribution of cytokeratin. Fish Physiology and Biochemistry, 40(2):571–576, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9867-6.

Francis:2014:LPD

[588] D. S. Francis, T. Thanuthong, and G. M. Turchini. n-3 LC-PUFA deposition efficiency and appetite-regulating hormones are modulated by the dietary lipid source during rainbow trout grow-out and finishing periods. Fish Physiology and Biochemistry, 40(2):577–593, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9868-5.

Zhu:2014:VEC

[589] Qing-Ling Zhu, Zhi Luo, and Qi-Liang Chen. In vitro exposure to copper influences lipid metabolism in hepatocytes from grass carp (Ctenopharyngodon idellus). Fish Physiology and Biochemistry, 40(2): 595–605, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9869-4.

Primavera-Tirol:2014:UAI

[590] Y. H. Primavera-Tirol, R. M. Coloso, and L. V. Laureta, Jr. Ultrastructure of the anterior intestinal epithelia of the orange-spotted grouper Epinephelus coioides larvae under different feeding regimes. Fish Physiology and Biochemistry, 40(2):607–624, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9870-y.

Yue:2014:CEA

[591] Pang Yue, Xiao Rong, and Qing Wei Li. Cloning and expression analysis of a novel high-mobility group box 2 homologue from Lampetra japonica. Fish Physiology and Biochemistry, 40(2):625–634, April 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9871-x.

Nigam:2014:CCS

[592] Ashwini Kumar Nigam, Usha Kumari, and Ajay Kumar Mittal. Characterization of carboxylesterase in skin mucus of Cirrhinus mrigala and its assessment as biomarker of organophosphate exposure. Fish Physiology and Biochemistry, 40(3):635–644, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9872-9.

Ribeiro:2014:LAE

[593] Andrea Martini Ribeiro, Wagner Ezequiel Risso, and Claudia B. R. Martinez. Lead accumulation and its effects on the branchial physiology of Prochilodus lineatus. Fish Physiology and Biochemistry, 40 (3):645–657, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9873-8.

Li:2014:EDL

[594] Xue-Yin Li, Ling Tang, and Xiao-Qiu Zhou. Effect of dietary lysine on growth, intestinal enzymes activities and antioxidant status of sub-adult grass carp (Ctenopharyngodon idella). Fish Physiology and Biochemistry, 40(3):659–671, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9874-7.

Li:2014:EDT

[595] Xue-Yin Li, Hui-Hua Huang, and Xiao-Qiu Zhou. The effects of dietary thiamin on oxidative damage and antioxidant defence of juvenile fish. Fish Physiology and Biochemistry, 40(3):673–687, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9875-6.

Cuenca-Soria:2014:PCD

[596] C. A. Cuenca-Soria, C. A. Álvarez-González, and E. Gisbert. Partial characterisation of digestive proteases of the Mayan cichlid Cichlasoma urophthalmus. Fish Physiology and Biochemistry, 40(3):689– 699, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-013-9876-5.

Toni:2014:FAE

[597] Cândida Toni, Alexssandro Geferson Becker, and Bernardo Baldisserotto. Fish anesthesia: effects of the essential oils of Hesperozygis ringens and Lippia alba on the biochemistry and physiology of silver catfish (Rhamdia quelen). Fish Physiology and Biochemistry, 40 (3):701–714, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9877-4.

Hedayati:2014:HGH

[598] Aliakbar Hedayati and Reza Tarkhani. Hematological and gill histopathological changes in iridescent shark, Pangasius hypophthalmus (Sauvage, 1878) exposed to sublethal diazinon and deltamethrin concentrations. Fish Physiology and Biochemistry, 40(3):715–720, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9878-3.

Cui:2014:EHS

[599] Yanting Cui, Bo Liu, and Yuanyuan Zhang. Effect of heat stress and recovery on viability, oxidative damage, and heat shock protein expression in hepatic cells of grass carp (Ctenopharyngodon idellus). Fish Physiology and Biochemistry, 40(3):721–729, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9879-2.

Cavaco:2014:TFO

[600] Sofia Cavaco, Matthew K. Williamson, and Dina C. Simes. Teleost fish osteocalcin 1 and 2 share the ability to bind the calcium mineral phase. Fish Physiology and Biochemistry, 40(3):731–738, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9880-9.

Palma:2014:EDO

[601] J. Palma, D. P. Bureau, and J. P. Andrade. The effect of diet on ontogenic development of the digestive tract in juvenile reared long snout seahorse Hippocampus guttulatus. Fish Physiology and Biochemistry, 40 (3):739–750, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9881-8.

Bermejo-Nogales:2014:TSG

[602] Azucena Bermejo-Nogales, Josep Alvar Calduch-Giner, and Jaume Pérez-Sánchez. Tissue-specific gene expression and functional regulation of uncoupling protein 2 (UCP2) by hypoxia and nutrient availability in gilthead sea bream (Sparus aurata): implications on the physiological significance of UCP1–3 variants. Fish Physiology and Biochemistry, 40(3):751–762, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9882-7.

Wu:2014:EPK

[603] Zhi xin Wu, Yan mei Yu, and Xiao xuan Chen. Effect of prebiotic konjac mannanoligosaccharide on growth performances, intestinal microflora, and digestive enzyme activities in yellow catfish, Pelteobagrus fulvidraco. Fish Physiology and Biochemistry, 40(3):763–771, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9883-6.

Martinez-Lagos:2014:CDE

[604] R. Martínez-Lagos, D. Tovar-Ramírez, and J. P. Lazo. Changes in digestive enzyme activities during larval development of leopard grouper (Mycteroperca rosacea). Fish Physiology and Biochemistry, 40(3):773– 785, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-013-9884-5.

El-Gawad:2014:EVC

[605] Eman A. Abd El-Gawad and Omnia M. Abdel Hamid. Effect of vitamin C dietary supplementation in reducing the alterations induced by fenitrothion in Oreochromis niloticus. Fish Physiology and Biochemistry, 40(3):787–796, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9885-4.

Chebaani:2014:IHI

[606] Nadjoua Chebaani, Francisco A. Guardiola, and Alberto Cuesta. Innate humoral immune parameters in Tilapia zillii under acute stress by low temperature and crowding. Fish Physiology and Biochemistry, 40 (3):797–804, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9886-3.

Zhang:2014:HEE

[607] Xiaozheng Zhang, Li Xiong, and Siyu Mao. Histopathological and estrogen effect of pentachlorophenol on the rare minnow (Gobiocypris rarus). Fish Physiology and Biochemistry, 40(3):805–816, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9887-2.

Deng:2014:ASP

[608] Wei Deng, Yuhua Zhao, and Rui Du. Anti-stress properties and two HSP70s mRNA expressions of blunt snout bream (Megalobrama amblycephala) fed with all-plant-based diet. Fish Physiology and Biochemistry, 40(3):817–825, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9888-1.

Guzman-Villanueva:2014:EDB

[609] Laura T. Guzmán-Villanueva, Felipe Ascencio-Valle, and Dariel Tovar-Ramírez. Effects of dietary β-1,3/1,6-glucan on the antioxidant and digestive enzyme activities of Pacific red snapper (Lutjanus peru) after exposure to lipopolysaccharides. Fish Physiology and Biochemistry, 40 (3):827–837, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9889-0.

Moguel-Hernandez:2014:DDE

[610] I. Moguel-Hernández, R. Peña, and I. Zavala-Leal. Development of digestive enzyme activity in spotted rose snapper, Lutjanus guttatus (Steindachner, 1869) larvae. Fish Physiology and Biochemistry, 40 (3):839–848, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9890-7.

Muriach:2014:CSB

[611] Borja Muriach, Manuel Carrillo, and José Miguel Cerdá-Reverter. Characterization of sea bass FSHβ 5′ flanking region: transcriptional control by 17β-estradiol. Fish Physiology and Biochemistry, 40(3):849– 864, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-013-9891-6.

Han:2014:BAC

[612] Junyan Han, Cheng Gao, and Dehong Tan. Betanin attenuates carbon tetrachloride (CCl_{4})-induced liver injury in common carp (Cyprinus carpio L.). Fish Physiology and Biochemistry, 40(3):865–874, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9892-5.

Galt:2014:HFD

[613] Nicholas J. Galt, Jacob Michael Froehlich, and Peggy R. Biga. Highfat diet reduces local myostatin-1 paralog expression and alters skeletal muscle lipid content in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 40(3):875–886, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9893-4.

Cofre:2014:PGE

[614] C. Cofre, R. Gonzalez, and R. Vidal. Phenotype gene expression differences between resistant and susceptible salmon families to IPNV. Fish Physiology and Biochemistry, 40(3):887–896, June 2014. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-013-9894-3.

deJesus:2014:PGM

[615] Lázaro Wender Oliveira de Jesus, Chayrra Chehade, and Maria Inês Borella. Pituitary gland morphogenesis and ontogeny of adenohypophyseal cells of Salminus brasiliensis (Teleostei, Characiformes). Fish Physiology and Biochemistry, 40(3):897–909, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9895-2.

Fatira:2014:DCP

[616] E. Fatira, N. Papandroulakis, and M. Pavlidis. Diel changes in plasma cortisol and effects of size and stress duration on the cortisol response in European sea bass (Dicentrarchus labrax). Fish Physiology and Biochemistry, 40(3):911–919, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-013-9896-1.

De:2014:PFS

[617] Bidhan C. De, D. K. Meena, and A. P. Sharma. Probiotics in fish and shellfish culture: immunomodulatory and ecophysiological responses. Fish Physiology and Biochemistry, 40(3):921–971, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-013-9897-0.

Zeng:2014:ECA

[618] Peng Zeng, Tianji Chen, and Jiang Shen. Effects of cold acclimation and storage temperature on crucian carp (Carassius auratus gibelio) in a waterless preservation. Fish Physiology and Biochemistry, 40 (3):973–982, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9898-z.

Chandrasekar:2014:ABW

[619] S. Chandrasekar, T. Nich, and S. Dasgupta. Acclimation of brackish water pearl spot (Etroplus suratensis) to various salinities: relative changes in abundance of branchial Na^{+}/K^{+}-ATPase and Na^{+}/K^{+}/2Cl^{−} co-transporter in relation to osmoregulatory parameters. Fish Physiology and Biochemistry, 40(3):983–996, June 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9899-y.

Albrizio:2014:MOR

[620] Maria Albrizio, Antonio C. Guaricci, and Giulio Aiudi. Mu opioid receptor in spermatozoa, eggs and larvae of gilthead sea bream (Sparus aurata) and its involvement in stress related to aquaculture. Fish Physiology and Biochemistry, 40(4):997–1009, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-013-9900-9.

Dabruzzi:2014:HEG

[621] Theresa F. Dabruzzi and Wayne A. Bennett. Hypoxia effects on gill surface area and blood oxygen-carrying capacity of the Atlantic stingray, Dasyatis sabina. Fish Physiology and Biochemistry, 40(4): 1011–1020, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9901-8.

Guerrero-Zarate:2014:PCD

[622] R. Guerrero-Zárate, C. A. Alvarez-González, and W. M. Contreras-Sánchez. Partial characterization of digestive proteases in tropical gar Atractosteus tropicus juveniles. Fish Physiology and Biochemistry, 40 (4):1021–1029, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-013-9902-7.

Liu:2014:HER

[623] Yifan Liu, Daoyuan Ma, and Jun Li. Histological and enzymatic responses of Japanese flounder (Paralichthys olivaceus) and its hybrids (P. olivaceus [female sign] × P. dentatus [male sign]) to chronic heat stress. Fish Physiology and Biochemistry, 40(4):1031–1041, August 2014. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9903-6.

Kiriake:2014:EPP

[624] Aya Kiriake, Mihoko Madokoro, and Kazuo Shiomi. Enzymatic properties and primary structures of hyaluronidases from two species of lionfish (Pterois antennata and Pterois volitans). Fish Physiology and Biochemistry, 40(4):1043–1053, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-013-9904-5.

Elia:2014:TOO

[625] Antonia Concetta Elia, Valentina Ciccotelli, and Maria Cesarina Abete. Transferability of oxytetracycline (OTC) from feed to carp muscle and evaluation of the antibiotic effects on antioxidant systems in liver and kidney. Fish Physiology and Biochemistry, 40(4):1055–1068, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-013-9905-4.

Makkapan:2014:EPR

[626] Walaiporn Makkapan, Goro Yoshizaki, and Wilaiwan Chotigeat. Expression profile of ribosomal protein L10a throughout gonadal development in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 40(4):1069–1081, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-013-9906-3.

Kanak:2014:EFS

[627] E. G. Kanak, Z. Dogan, and M. Canli. Effects of fish size on the response of antioxidant systems of Oreochromis niloticus following metal exposures. Fish Physiology and Biochemistry, 40(4):1083–1091, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-9907-x.

Jia:2014:BCQ

[628] Yudong Jia, Zhen Meng, and Jilin Lei. Biochemical composition and quality of turbot (Scophthalmus maximus) eggs throughout the reproductive season. Fish Physiology and Biochemistry, 40(4):1093–1104, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-9908-9.

Sadhu:2014:CSD

[629] Narasimhulu Sadhu, S. R. Krupesha Sharma, and K. K. Philipose. Chronic stress due to high stocking density in open sea cage farming induces variation in biochemical and immunological functions in Asian seabass (Lates calcarifer, Bloch). Fish Physiology and Biochemistry, 40 (4):1105–1113, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9909-8.

Grandi:2014:IUI

[630] Gilberto Grandi, Maria Gabriella Marchetti, and Milvia Chicca. Immunocytochemical and ultrastructural identification of adenohypophyseal cells in Ctenopharyngodon idella (Cypriniformes: Cyprinidae) during gonadal differentiation. Fish Physiology and Biochemistry, 40(4): 1115–1139, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9910-2.

Zheng:2014:MCP

[631] Yao Zheng, Hongwei Liang, and Zaizhao Wang. Molecular cloning of Pcc-dmrt1s and their specific expression patterns in Pengze crucian carp (Carassius auratus var. Pengze) affected by 17α-methyltestosterone. Fish Physiology and Biochemistry, 40(4):1141–1155, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-9911-1.

Ma:2014:ODD

[632] Zhenhua Ma, Huayang Guo, and Dianchang Zhang. Ontogenetic development of digestive functionality in golden pompano Trachinotus ovatus (Linnaeus 1758). Fish Physiology and Biochemistry, 40(4): 1157–1167, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9912-0.

Reyes-Becerril:2014:SCE

[633] Martha Reyes-Becerril, Felipe Ascencio, and María Ángeles Esteban. Single or combined effects of Lactobacillus sakei and inulin on growth, non-specific immunity and IgM expression in leopard grouper (Mycteroperca rosacea). Fish Physiology and Biochemistry, 40(4):1169–1180, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-9913-z.

Mei:2014:IGE

[634] Jie Mei, Wei Yan, and Yan He. Identification of a gonad-expression differential gene insulin-like growth factor-1 receptor (Igf1r) in the swamp eel (Monopterus albus). Fish Physiology and Biochemistry, 40(4): 1181–1190, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9914-y.

Piccinno:2014:VEI

[635] M. Piccinno, R. Zupa, and R. L. Sciorsci. In vitro effect of isotocin on ovarian tunica albuginea contractility of gilthead seabream (Sparus aurata L.) in different reproductive conditions. Fish Physiology and Biochemistry, 40(4):1191–1199, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9915-x.

Boerrigter:2014:EPT

[636] Jeroen G. J. Boerrigter, Hans W. van de Vis, and Gert Flik. Effects of Pro-tex on zebrafish (Danio rerio) larvae, adult common carp (Cyprinus carpio) and adult yellowtail kingfish (Seriola lalandi). Fish Physiology and Biochemistry, 40(4):1201–1212, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-9916-9.

Glencross:2014:IBD

[637] Brett D. Glencross, Douglas R. Tocher, and J. Gordon Bell. Interactions between dietary docosahexaenoic acid and other long-chain polyunsaturated fatty acids on performance and fatty acid retention in post-smolt Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 40 (4):1213–1227, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9917-8.

Shin:2014:SEL

[638] Hyun Suk Shin and Cheol Young Choi. The stimulatory effect of LED light spectra on genes related to photoreceptors and skin pigmentation in goldfish (Carassius auratus). Fish Physiology and Biochemistry, 40 (4):1229–1238, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9918-7.

Yuan:2014:GWI

[639] Jing Yuan, Wenjing Tao, and Deshou Wang. Genome-wide identification, phylogeny, and gonadal expression of fox genes in Nile tilapia, Oreochromis niloticus. Fish Physiology and Biochemistry, 40(4):1239– 1252, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-014-9919-6.

Orlich:2014:PDA

[640] José R. Quirós Orlich, Silvia Valverde Chavarría, and Juan B. Ulloa Rojas. The proteolytic digestive activity and growth during ontogeny of Parachromis dovii larvae (pisces: Cichlidae) using two feeding protocols.

Fish Physiology and Biochemistry, 40(4):1253–1261, August 2014. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-9920-0.

Ma:2014:CED

[641] Liman Ma, Wenji Wang, and Jie Qi. Characterization of the Dmrt1 gene in the black rockfish Sebastes schlegeli revealed a remarkable sex-dimorphic expression. Fish Physiology and Biochemistry, 40(4): 1263–1274, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9921-z.

Chang:2014:PST

[642] Zhi-Qiang Chang, Zhao-Xin Li, and Jian Li. Pharmacokinetics of sulfamonomethoxine in tongue sole (Cynoglossus semilaevis) after intravenous and oral administration. Fish Physiology and Biochemistry, 40 (4):1275–1279, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9922-y.

Lai:2014:CEP

[643] Xiao Jian Lai, Wan Shu Hong, and Shi Xi Chen. Cloning and expression of prostaglandin E_{2} receptor subtype 1 (ep_{1}) in Bostrichthys sinensis. Fish Physiology and Biochemistry, 40(4):1281–1288, August 2014. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-9923-x.

Tkachenko:2014:TSR

[644] Halyna Tkachenko, Natalia Kurhaluk, and Anastasiia Andriichuk. Tissue-specific responses of oxidative stress biomarkers and antioxidant defenses in rainbow trout Oncorhynchus mykiss during a vaccination against furunculosis. Fish Physiology and Biochemistry, 40(4): 1289–1300, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9924-9.

Ruan:2014:DQD

[645] Jiming Ruan, Kun Hu, and Xianle Yang. Distribution and quantitative detection of GABA_{A} receptor in Carassius auratus gibelio. Fish Physiology and Biochemistry, 40(4):1301–1311, August 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-9925-8.

Hao:2014:EDS

[646] Xiaofeng Hao, Qufei Ling, and Fashui Hong. Effects of dietary selenium on the pathological changes and oxidative stress in loach (Paramisgurnus dabryanus). Fish Physiology and Biochemistry, 40(5):1313–1323, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-9926-7.

Kumar:2014:SMG

[647] Prem Kumar, Shailesh Saurabh, and A. R. T. Arasu. Stress mitigating and growth enhancing effect of dietary tryptophan in rohu (Labeo rohita, Hamilton, 1822) fingerlings. Fish Physiology and Biochemistry, 40(5): 1325–1338, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9927-6.

Peres:2014:BCP

[648] Helena Peres, Sara Santos, and Aires Oliva-Teles. Blood chemistry profile as indicator of nutritional status in European seabass (Dicentrarchus labrax). Fish Physiology and Biochemistry, 40(5):1339–1347, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-9928-5.

Zheng:2014:DEA

[649] Jia-Lang Zheng, Zhi Luo, and Wei Hu. Differential effects of acute and chronic zinc exposure on lipid metabolism in three extrahepatic tissues of juvenile yellow catfish Pelteobagrus fulvidraco. Fish Physiology and Biochemistry, 40(5):1349–1359, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-9929-4.

Walker:2014:EUM

[650] Christina J. Walker, James Gelsleichter, and Charles A. Manire. Evaluation of the use of metallothionein as a biomarker for detecting physiological responses to mercury exposure in the bonnethead, Sphyrna tiburo. Fish Physiology and Biochemistry, 40(5):1361–1371, October 2014. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-9930-y.

Caruso:2014:CDE

[651] G. Caruso, M. G. Denaro, and G. Maricchiolo. Changes in digestive enzyme activities of red porgy Pagrus pagrus during a fasting– refeeding experiment. Fish Physiology and Biochemistry, 40(5):1373– 1382, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9931-x.

Wei:2014:ECO

[652] Jing Wei, WenChuang Qi, and Deshou Wang. Establishment and characterization of an ovarian cell line from Southern catfish (Silurus meridionalis). Fish Physiology and Biochemistry, 40(5):1383–1391, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-9932-9.

Alavi:2014:PSS

[653] Sayyed Mohammad Hadi Alavi, Pavla Postlerová-Maňásková, and Otomar Linhart. Protease in sturgeon sperm and the effects of protease inhibitors on sperm motility and velocity. Fish Physiology and Biochemistry, 40(5):1393–1398, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9933-8.

Sullivan:2014:DCD

[654] Melissa Sullivan, Alexandria C. Brown, and Ethan D. Clotfelter. Dietary carotenoids do not improve motility or antioxidant capacity in cichlid fish sperm. Fish Physiology and Biochemistry, 40(5):1399–1405, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-9934-7.

Wang:2014:ESP

[655] Tao Wang, Chaowei Zhou, and Zhiqiong Li. Schizothorax prenanti corticotropin-releasing hormone (CRH): molecular cloning, tissue expression, and the function of feeding regulation. Fish Physiology and Biochemistry, 40(5):1407–1415, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-9935-6.

Bondarenko:2014:RCM

[656] Olga Bondarenko, Borys Dzyuba, and Otomar Linhart. The role of Ca^{2+} and Na^{+} membrane transport in brook trout (Salvelinus fontinalis) spermatozoa motility. Fish Physiology and Biochemistry, 40(5): 1417–1421, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9936-5.

Liu:2014:EDA

[657] Xiang he Liu, Chao xia Ye, and An li Wang. Effects of dietary amylose/amylopectin ratio on growth performance, feed utilization, digestive enzymes, and postprandial metabolic responses in juvenile obscure puffer Takifugu obscurus. Fish Physiology and Biochemistry, 40(5): 1423–1436, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9937-4.

Hu:2014:ODR

[658] Kai Hu, Lin Feng, and Xiaoqiu Zhou. Oxidative damage repair by glutamine in fish enterocytes. Fish Physiology and Biochemistry, 40(5): 1437–1445, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9938-3.

Liu:2014:RAL

[659] Pin Liu, Chao Li, and Hong Ji. Regulation of adipocytes lipolysis by n-3 HUFA in grass carp (Ctenopharyngodon idellus) in vitro and in vivo. Fish Physiology and Biochemistry, 40(5):1447–1460, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-9939-2.

Shiozaki:2014:MCB

[660] Kazuhiro Shiozaki, Sena Ryuzono, and Taeko Miyagi. Molecular cloning and biochemical characterization of medaka (Oryzias latipes) lysosomal neu4 sialidase. Fish Physiology and Biochemistry, 40(5):1461– 1472, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9940-9.

Poochai:2014:CEA

[661] Watsida Poochai, Kiattawee Choowongkomon, and Nontawith Areechon. Characterization and expression analysis of the transferrin gene in Nile tilapia (Oreochromis niloticus) and its upregulation in response to Streptococcus agalactiae infection. Fish Physiology and Biochemistry, 40(5): 1473–1485, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9941-8.

Moretti:2014:HDI

[662] Débora B. Moretti, Wiolene M. Nordi, and Raul Machado-Neto. Histochemical distribution of intestinal enzymes of juvenile pacu (Piaractus mesopotamicus) fed lyophilized bovine colostrum. Fish Physiology and Biochemistry, 40(5):1487–1493, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-9942-7.

Ali:2014:SFEa

[663] T. El-Sayed Ali, S. H. Abdel-Aziz, and S. Zeid. Structural and functional effects of early exposure to 4-nonylphenol on gonadal development of Nile tilapia (Oreochromis niloticus): b-histological alterations in testes. Fish Physiology and Biochemistry, 40(5):1495–1507, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-9944-5.

Ali:2014:SFEb

[664] T. El-Sayed Ali, S. H. Abdel-Aziz, and S. Zeid. Structural and functional effects of early exposure to 4-nonylphenol on gonadal development of Nile tilapia (Oreochromis niloticus): a-histological alterations in ovaries. Fish Physiology and Biochemistry, 40(5):1509–1519, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-9943-6.

Luo:2014:ESI

[665] Hongwei Luo, Xinbin Duan, and Daqing Chen. Effects of surgically implanted dummy ultrasonic transmitters on physiological response of bighead carp Hypophthalmichthys nobilis. Fish Physiology and Biochemistry, 40(5):1521–1532, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9945-4.

Tang:2014:EMP

[666] Cheng-Hao Tang, Ming-Yih Leu, and Shu-Chuan Tsai. Exploration of the mechanisms of protein quality control and osmoregulation in gills of Chromis viridis in response to reduced salinity. Fish Physiology and Biochemistry, 40(5):1533–1546, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-9946-3.

Silva:2014:FMP

[667] Patricia I. M. Silva, Catarina I. M. Martins, and Øyvind Øverli. Feeding motivation as a personality trait in Nile tilapia (Oreochromis niloticus): role of serotonergic neurotransmission. Fish Physiology and Biochemistry, 40(5):1547–1557, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9947-2.

Hosono:2014:BPC

[668] Masahiro Hosono, Shigeki Sugawara, and Kazuo Nitta. Binding profiles and cytokine-inducing effects of fish rhamnose-binding lectins on Burkitt’s lymphoma raji cells. Fish Physiology and Biochemistry, 40(5): 1559–1572, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9948-1.

Padmini:2014:HSP

[669] Ekambaram Padmini and Jayachandran Tharani. Heat-shock protein 70 modulates apoptosis signal-regulating kinase 1 in stressed hepatocytes of Mugil cephalus. Fish Physiology and Biochemistry, 40(5): 1573–1585, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9949-0.

Hsu:2014:PTT

[670] Chih-Wei Hsu, Shu-Chuan Tsai, and Su Mei Wu. Profiles of thyrotropin, thyroid hormones, follicular cells and type I deiodinase gene expression during ontogenetic development of tilapia larvae and juveniles. Fish Physiology and Biochemistry, 40(5):1587–1599, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-9950-7.

Roosta:2014:EDV

[671] Zahra Roosta, Abdolmajid Hajimoradloo, and Seyed Hossein Hoseinifar. The effects of dietary vitamin C on mucosal immune responses and growth performance in Caspian roach (Rutilus rutilus caspicus) fry. Fish Physiology and Biochemistry, 40(5):1601–1607, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-9951-6.

Pacini:2014:MCC

[672] Nicole Pacini, Ambrosius Josef Martin Dörr, and Marino Prearo. Melamine–cyanurate complexes and oxidative stress markers in trout kidney following melamine and cyanuric acid long-term co-exposure and withdrawal. Fish Physiology and Biochemistry, 40(5):1609–1619, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-9952-5.

Hartviksen:2014:EEC

[673] Mali Hartviksen, Anne Marie Bakke, and Åshild Krogdahl. Evaluation of the effect of commercially available plant and animal protein sources in diets for Atlantic salmon (Salmo salar L.): digestive and metabolic investigations. Fish Physiology and Biochemistry, 40(5): 1621–1637, October 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9953-4.

Jia:2014:MCC

[674] Yudong Jia, Zhen Meng, and Jilin Lei. Molecular cloning, characterization, and expression analysis of luteinizing hormone receptor gene in turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 40(6): 1639–1650, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9954-3.

Kopp:2014:MBI

[675] Radovan Kopp, Miroslava Palíková, and Jan Mareš. Modulation of biochemical indices in common carp (Cyprinus carpio L.) under the influence of toxic cyanobacterial biomass in diet. Fish Physiology and Biochemistry, 40(6):1651–1658, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9955-2.

deAssis:2014:CCE

[676] Caio Rodrigo Dias de Assis, Amanda Guedes Linhares, and Luiz Bezerra Carvalho, Jr. Characterization of catalytic efficiency parameters of brain cholinesterases in tropical fish. Fish Physiology and Biochemistry, 40(6): 1659–1668, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9956-1.

Chen:2014:CMC

[677] Wenbo Chen, Weiguo Li, and Mengjie Li. Cloning, molecular characterization and expression analysis of insulin-like growth factor binding protein-2 (IGFBP-2) cDNA in goldfish, Carassius auratus. Fish Physiology and Biochemistry, 40(6):1669–1681, December 2014. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-014-9958-z.

Vargas-Chacoff:2014:SDE

[678] L. Vargas-Chacoff, E. Ortíz, and J. M. Mancera. Stocking density and Piscirickettsia salmonis infection effect on Patagonian blennie (Eleginops maclovinus, Cuvier 1830) skeletal muscle intermediate metabolism. Fish Physiology and Biochemistry, 40(6):1683–1691, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-9959-y.

Wang:2014:CEL

[679] Fei Wang, Weimin Chen, and Wensheng Li. Cloning, expression, and ligand-binding characterization of two neuropeptide Y receptor subtypes in orange-spotted grouper, Epinephelus coioides. Fish Physiology and Biochemistry, 40(6):1693–1707, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-9960-5.

Tang:2014:CEA

[680] Yongkai Tang, Hongxia Li, and Juhua Yu. Characterization and expression analysis of two distinct neuropeptide Ya paralogues in Jian carp (Cyprinus carpio var. Jian). Fish Physiology and Biochemistry, 40(6): 1709–1719, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9961-4.

Hu:2014:RGS

[681] Qing Hu, Wei Guo, and Dapeng Li. Reference gene selection for realtime RT-PCR normalization in rice field eel (Monopterus albus) during gonad development. Fish Physiology and Biochemistry, 40(6):1721– 1730, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9962-3.

Dzyuba:2014:ASS

[682] Viktoriya Dzyuba, Borys Dzyuba, and Marek Rodina. The antioxidant system of sterlet seminal fluid in testes and Wolffian ducts. Fish Physiology and Biochemistry, 40(6):1731–1739, December 2014. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-014-9963-2.

Murashita:2014:COD

[683] Koji Murashita, Hiroyuki Matsunari, and Takeshi Yamamoto. Characterization and ontogenetic development of digestive enzymes in Pacific bluefin tuna Thunnus orientalis larvae. Fish Physiology and Biochemistry, 40(6):1741–1755, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9964-1.

Oku:2014:CDE

[684] Hiromi Oku, Masaharu Tokuda, and Takeshi Yamamoto. Characterization of differentially expressed genes in liver in response to the rearing temperature of rainbow trout Oncorhynchus mykiss and their heritable differences. Fish Physiology and Biochemistry, 40(6):1757– 1769, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9965-0.

Shaliutina-Kolesova:2014:PCC

[685] A. Shaliutina-Kolešová, I. Gazo, and O. Linhart. Protection of common carp (Cyprinus carpio L.) spermatozoa motility under oxidative stress by antioxidants and seminal plasma. Fish Physiology and Biochemistry, 40(6):1771–1781, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9966-z.

Xu:2014:OMI

[686] Fan Xu, Yuan-Tu Ye, and Liang Song. Observation of the middle intestinal tight junction structure, cloning and studying tissue distribution of the four Claudin genes of the grass carp (Ctenopharyngodon idellus). Fish Physiology and Biochemistry, 40(6):1783–1792, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-9967-y.

Ferrando:2014:EUM

[687] Sara Ferrando, Lorenzo Gallus, and Maria Angela Masini. Effects of urea on the molecules involved in the olfactory signal transduction: a preliminary study on Danio rerio. Fish Physiology and Biochemistry, 40(6): 1793–1800, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9968-x.

Ni:2014:THG

[688] Meng Ni, Haishen Wen, and Houmeng Ding. Two HSPs gene from juvenile Amur sturgeon (Acipenser schrenckii): cloning, characterization and expression pattern to crowding and hypoxia stress. Fish Physiology and Biochemistry, 40(6):1801–1816, December 2014. CODEN FPBIEP.

ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-9969-9.

Rola:2014:RRI

[689] Regina Coimbra Rola, Luis Fernando Marins, and Juliana Zomer Sandrini. Responses to ROS inducer agents in zebrafish cell line: differences between copper and UV-b radiation. Fish Physiology and Biochemistry, 40(6):1817–1825, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9970-3.

Deng:2014:EDC

[690] Junming Deng, Xi Zhang, and Bin Kang. Effects of dietary cholesterol supplementation on growth and cholesterol metabolism of rainbow trout (Oncorhynchus mykiss) fed diets with cottonseed meal or rapeseed meal. Fish Physiology and Biochemistry, 40(6):1827–1838, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-9971-2.

Nguyen:2014:EAE

[691] Phuc Trong Hong Nguyen, Huong Thi Thanh Do, and David A. Hurwood. Experimental assessment of the effects of sublethal salinities on growth performance and stress in cultured tra catfish (Pangasianodon hypophthalmus). Fish Physiology and Biochemistry, 40(6):1839– 1848, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9972-1.

Sarkar:2014:BPI

[692] Shrabanti Sarkar, Debapriya Bhattacharya, and Panchanan Nath. Biological properties of Indian walking catfish (Clarias batrachus) (L.) gonadotropins in female reproduction. Fish Physiology and Biochemistry, 40(6):1849–1861, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9973-0.

Santana:2014:EMR

[693] Julio Cesar de Oliveira Santana and Irani Quagio-Grassiotto. Extracellular matrix remodeling of the testes through the male reproductive cycle in teleostei fish. Fish Physiology and Biochemistry, 40(6): 1863–1875, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9974-z.

Luo:2014:DEP

[694] Lei Luo, Aqin Chen, and Weiqun Lu. Dynamic expression pattern of corticotropin-releasing hormone, urotensin I and II genes under acute salinity and temperature challenge during early development of zebrafish. Fish Physiology and Biochemistry, 40(6):1877–1886, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-9975-y.

Li:2014:MGA

[695] Mao Li, Heather Christie, and John Leatherland. Modulation of GR activity does not affect the in vitro metabolism of cortisol by rainbow trout ovarian follicles. Fish Physiology and Biochemistry, 40(6): 1887–1897, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9976-x.

Gacic:2014:SSE

[696] Z. Gačić, A. Bajić, and I. Damjanović. Spectral sensitivity of the electroretinogram b-wave in dark-adapted Prussian carp (Carassius gibelio Bloch, 1782). Fish Physiology and Biochemistry, 40(6):1899–1906, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-9977-9.

Nakamura:2014:LPF

[697] Osamu Nakamura, Yuki Watabe, and Shigeyuki Tsutsui. Localization and possible function of nrF-AGP, an alpha-1-acid glycoproteinlike protein in viviparous fish Neoditrema ransonnetii (Perciformes, Embiotocidae). Fish Physiology and Biochemistry, 40(6):1907–1915, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-9978-8.

Rebl:2014:GET

[698] Alexander Rebl, Andreas Brietzke, and Hans-Martin Seyfert. GRP94 is encoded by two differentially expressed genes during development of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 40(6):1917–1926, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9979-7.

Song:2014:EIB

[699] Wen Ting Song, Zhi Jun Wang, and Hong Cai Liu. Effects of individual and binary mixtures of estrogens on male goldfish (Carassius auratus). Fish Physiology and Biochemistry, 40(6):1927–1935, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-9980-1.

Arockiaraj:2014:NME

[700] Jesu Arockiaraj, Rajesh Palanisamy, and Marimuthu Kasi. A novel murrel Channa striatus mitochondrial manganese superoxide dismutase: gene silencing, SOD activity, superoxide anion production and expression. Fish Physiology and Biochemistry, 40(6):1937–1955, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-9981-0.

Hosono:2014:EBP

[701] Masahiro Hosono, Shigeki Sugawara, and Kazuo Nitta. Erratum to: Binding profiles and cytokine-inducing effects of fish rhamnose-binding lectins on Burkitt’s lymphoma raji cells. Fish Physiology and Biochemistry, 40(6):1957, December 2014. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-014-9957-0.

Tian:2015:DLL

[702] Juan Tian, Fan Wu, and Hua Wen. Dietary lipid levels impact lipoprotein lipase, hormone-sensitive lipase, and fatty acid synthetase gene expression in three tissues of adult GIFT strain of Nile tilapia, Oreochromis niloticus. Fish Physiology and Biochemistry, 41(1):1–18, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-0001-1.

Pang:2015:ETR

[703] Xu Pang, Xing-Zhong Yuan, and Shi-Jian Fu. The effect of temperature on repeat swimming performance in juvenile qingbo (Spinibarbus sinensis). Fish Physiology and Biochemistry, 41(1):19–29, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-0002-0.

Pipil:2015:SVA

[704] S. Pipil, V. Kumar, and N. Sehgal. In silico and in vivo analysis of binding affinity of estrogens with estrogen receptor alpha in Channa punctatus (Bloch). Fish Physiology and Biochemistry, 41(1):31–40, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-0003-z.

Tirsgaard:2015:ETS

[705] Bjørn Tirsgaard, Jon Christian Svendsen, and John Fleng Steffensen. Effects of temperature on specific dynamic action in Atlantic cod Gadus morhua. Fish Physiology and Biochemistry, 41(1):41–50, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-0004-y.

Delgadin:2015:GIG

[706] Tomás Horacio Delgadin, Daniela Irina Pérez Sirkin, and Paula Gabriela Vissio. GH, IGF-i and GH receptors mRNA expression in response to growth impairment following a food deprivation period in individually housed cichlid fish Cichlasoma dimerus. Fish Physiology and Biochemistry, 41(1):51–60, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-0005-x.

Badruzzaman:2015:ENM

[707] Muhammad Badruzzaman, Satoshi Imamura, and Akihiro Takemura. Effects of neurotoxin 1-methyl-4-phenyl-1,2,3,6-tetrahydropyridine (MPTP) treatment on ovarian development of the sapphire devil, Chrysiptera cyanea. Fish Physiology and Biochemistry, 41(1):61–71, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-0006-9.

Dong:2015:CLA

[708] Gui-Fang Dong, Wen-Zuo Liu, and Yan ou Yang. Conjugated linoleic acid alters growth performance, tissue lipid deposition, and fatty acid composition of darkbarbel catfish (Pelteobagrus vachelli). Fish Physiology and Biochemistry, 41(1):73–89, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-0007-8.

Fang:2015:ADE

[709] Cheng Fang, Mingyang Ma, and Steven D. Mims. Alterations of digestive enzyme activities, intestinal morphology and microbiota in juvenile paddlefish, Polyodon spathula, fed dietary probiotics. Fish Physiology and Biochemistry, 41(1):91–105, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-0008-7.

Pipil:2015:CIV

[710] S. Pipil, V. S. Rawat, and N. Sehgal. Characterization of incomplete vitellogenin (VgC) in the Indian freshwater murrel, Channa punctatus (Bloch). Fish Physiology and Biochemistry, 41(1):107–117, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-0009-6.

Giannenas:2015:EDI

[711] Ilias Giannenas, Ioannis Karamaligas, and Evdokia Karagouni. Effect of dietary incorporation of a multi-strain probiotic on growth performance and health status in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 41(1):119–128, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-0010-0.

Toni:2015:SRS

[712] Cândida Toni, Juan Antonio Martos-Sitcha, and Bernardo Baldisserotto. Stress response in silver catfish (Rhamdia quelen) exposed to the essential oil of Hesperozygis ringens. Fish Physiology and Biochemistry, 41(1): 129–138, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-0011-z.

Xu:2015:ILG

[713] Yongjiang Xu, Kun Zang, and Xueying Shi. Insulin-like growth factors I and II in starry flounder (Platichthys stellatus): molecular cloning and differential expression during embryonic development. Fish Physiology and Biochemistry, 41(1):139–152, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-0012-y.

Zhou:2015:HEC

[714] Jishu Zhou, Chao Li, and Tianhe Zhu. Hepatoprotective effects of a Chinese herbal formulation, Yingchen decoction, on olaquindox-induced hepatopancreas injury in Jian carp (Cyprinus carpio var. Jian). Fish Physiology and Biochemistry, 41(1):153–163, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-0013-x.

Sonmez:2015:GPA

[715] Adem Yavuz Sönmez, Soner Bilen, and Gouranga Biswas. Growth performance and antioxidant enzyme activities in rainbow trout (Oncorhynchus mykiss) juveniles fed diets supplemented with sage, mint and thyme oils. Fish Physiology and Biochemistry, 41(1):165–175, February 2015. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-0014-9.

Kanjanaworakul:2015:CSE

[716] Poonmanee Kanjanaworakul, Prapansak Srisapoome, and Supawadee Poompuang. cDNA structure and the effect of fasting on myostatin expression in walking catfish (Clarias macrocephalus, Günther 1864). Fish Physiology and Biochemistry, 41(1):177–191, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-0015-8.

Viveiros:2015:MGM

[717] Ana T. M. Viveiros, Ariane F. Nascimento, and Jacky Cosson. Methyl glycol, methanol and DMSO effects on post-thaw motility, velocities, membrane integrity and mitochondrial function of Brycon orbignyanus and Prochilodus lineatus (Characiformes) sperm. Fish Physiology and Biochemistry, 41(1):193–201, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-0016-7.

Zhang:2015:EFI

[718] Chun-Nuan Zhang, Xiang-Fei Li, and Wen-Bin Liu. Effects of fructooligosaccharide on immune response, antioxidant capability and HSP70 and HSP90 expressions of blunt snout bream (Megalobrama amblycephala) under high ammonia stress. Fish Physiology and Biochemistry, 41(1):203–217, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-0017-6.

Papoutsoglou:2015:GSE

[719] Sofronios E. Papoutsoglou, Nafsika Karakatsouli, and N. Sakellaridis. Gilthead seabream (Sparus aurata) response to three music stimuli (Mozart — “Eine Kleine Nachtmusik,” Anonymous — “Romanza,” bach— “Violin Concerto No. 1”) and white noise under recirculating water conditions. Fish Physiology and Biochemistry, 41(1):219–232, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-0018-5.

Vanderplancke:2015:ECM

[720] Gwenaëlle Vanderplancke, Guy Claireaux, and José-Luis Zambonino-Infante. Exposure to chronic moderate hypoxia impacts physiological and developmental traits of European sea bass (Dicentrarchus labrax) larvae. Fish Physiology and Biochemistry, 41(1):233–242, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-0019-4.

Prabhu:2015:CEL

[721] P. Antony Jesu Prabhu, S. J. Kaushik, and I. Geurden. Comparison of endogenous loss and maintenance need for minerals in rainbow trout (Oncorhynchus mykiss) fed fishmeal or plant ingredient-based diets. Fish Physiology and Biochemistry, 41(1):243–253, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-0020-y.

Ferris:2015:EGR

[722] Jacqueline Ferris, Mao Li, and W. Allan King. Estrogen and glucocorticoid receptor agonists and antagonists in oocytes modulate the pattern of expression of genes that encode nuclear receptor proteins in very early stage rainbow trout (Oncorhynchus mykiss) embryos. Fish Physiology and Biochemistry, 41(1):255–265, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-0021-x.

Toledo-Solis:2015:CDE

[723] F. J. Toledo-Solís, A. Uscanga-Martínez, and C. A. Álvarez-González. Changes on digestive enzymes during initial ontogeny in the three-spot cichlid Cichlasoma trimaculatum. Fish Physiology and Biochemistry, 41 (1):267–279, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-0023-8.

Geda:2015:BAD

[724] F. Geda, A. Declercq, and G. P. J. Janssens. β-alanine does not act through branched-chain amino acid catabolism in carp, a species with low muscular carnosine storage. Fish Physiology and Biochemistry, 41 (1):281–287, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-0024-7.

Gerger:2015:AEB

[725] Courtney J. Gerger, Jith K. Thomas, and Lynn P. Weber. Acute effects of β-naphthoflavone on cardiorespiratory function and metabolism in adult zebrafish (Danio rerio). Fish Physiology and Biochemistry, 41 (1):289–298, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9982-z.

Venturini:2015:MET

[726] Francine P. Venturini, Fernanda D. Moraes, and Gilberto Moraes. Metabolic effects of trichlorfon (Masoten(R)) on the neotropical freshwater fish pacu (Piaractus mesopotamicus). Fish Physiology and Biochemistry, 41(1):299–309, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9983-y.

Seriani:2015:HCC

[727] Robson Seriani, Jakeline Galvão França, and Maria José Tavares Ranzani-Paiva. Hematological changes and cytogenotoxicity in the tilapia Oreochromis niloticus caused by sub-chronic exposures to mercury and selenium. Fish Physiology and Biochemistry, 41(1):311–322, February 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-014-9984-x.

Menezes:2015:CFC

[728] Charlene Menezes, Milene B. Fonseca, and Vania L. Loro. Commercial formulation containing 2,4-d affects biochemical parameters and morphological indices of silver catfish exposed for 90 days. Fish Physiology and Biochemistry, 41(2):323–330, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-9985-9.

Guo:2015:ELL

[729] Xiaoze Guo, Xu-Fang Liang, and Dan Shen. Effects of lipid-lowering pharmaceutical clofibrate on lipid and lipoprotein metabolism of grass carp (Ctenopharyngodon idellal val.) fed with the high non-protein energy diets. Fish Physiology and Biochemistry, 41(2):331–343, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-9986-8.

Ammar:2015:CLI

[730] I. Ben Ammar, F. Teletchea, and P. Fontaine. Continuous lighting inhibits the onset of reproductive cycle in pikeperch males and females. Fish Physiology and Biochemistry, 41(2):345–356, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-9987-7.

Low:2015:SEC

[731] Jennifer Low and Dennis M. Higgs. Sublethal effects of cadmium on auditory structure and function in fathead minnows (Pimephales promelas). Fish Physiology and Biochemistry, 41(2):357–369, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-014-9988-6.

Mehrim:2015:HAR

[732] Ahmed I. Mehrim, Fathy F. Khalil, and Montaha E. Hassan. Hydroyeast Aquaculture(R) as a reproductive enhancer agent for the adult Nile tilapia (Oreochromis niloticus Linnaeus, 1758). Fish Physiology and Biochemistry, 41(2):371–381, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9989-5.

Morin:2015:IEG

[733] Scott J. Morin, Wayne A. Decatur, and Stacia A. Sower. Identification and expression of GnRH2 and GnRH3 in the black sea bass (Centropristis striata), a hermaphroditic teleost. Fish Physiology and Biochemistry, 41 (2):383–395, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9990-z.

Altmann:2015:INS

[734] Simone Altmann, Alexander Rebl, and Tom Goldammer. Identification and de novo sequencing of housekeeping genes appropriate for gene expression analyses in farmed maraena whitefish (Coregonus maraena) during crowding stress. Fish Physiology and Biochemistry, 41(2):397– 412, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-014-9991-y.

Liu:2015:EAS

[735] Qinghua Liu, Xueying Wang, and Jun Li. Effect of the addition of six antioxidants on sperm motility, membrane integrity and mitochondrial function in red seabream (Pagrus major) sperm cryopreservation. Fish Physiology and Biochemistry, 41(2):413–422, April 2015. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-014-9993-9.

Li:2015:EDD

[736] Yang Li, Jian Gao, and Songqian Huang. Effects of different dietary phospholipid levels on growth performance, fatty acid composition, PPAR gene expressions and antioxidant responses of blunt snout bream Megalobrama amblycephala fingerlings. Fish Physiology and Biochemistry, 41 (2):423–436, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9994-8.

Xu:2015:SAE

[737] Mengxia Xu, Chunxiao Huang, and Huanling Wang. Sequence analysis and expression regulation of rbp4 by 9- cis -RA in Megalobrama amblycephala. Fish Physiology and Biochemistry, 41(2):437–447, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-9995-7.

Chi:2015:MCF

[738] Jing-Ruei Chi, Long-Si Liao, and Shao-Yang Hu. Molecular cloning and functional characterization of the hepcidin gene from the convict cichlid (Amatitlania nigrofasciata) and its expression pattern in response to lipopolysaccharide challenge. Fish Physiology and Biochemistry, 41 (2):449–461, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-014-9996-6.

Gressler:2015:HMB

[739] Luciane Tourem Gressler, Fernando Jonas Sutili, and Bernardo Baldisserotto. Hematological, morphological, biochemical and hydromineral responses in Rhamdia quelen sedated with propofol. Fish Physiology and Biochemistry, 41(2):463–472, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9997-5.

Zhou:2015:MCC

[740] Ze jun Zhou, Reng Qiu, and Jian Zhang. Molecular characterization of the cathepsin B of turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 41(2):473–483, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-014-9998-4.

Ghaedi:2015:OSR

[741] Gholamreza Ghaedi, Bahram Falahatkar, and Gholamreza Nikbakht Broujeni. The onset of stress response in rainbow trout Oncorhynchus mykiss embryos subjected to density and handling. Fish Physiology and Biochemistry, 41(2):485–493, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-014-9999-3.

Cheng:2015:EAA

[742] Chang-Hong Cheng, Fang-Fang Yang, and An-Li Wang. Effect of acute ammonia exposure on expression of GH/IGF axis genes GHR1, GHR2 and IGF-1 in pufferfish (Takifugu obscurus). Fish Physiology and Biochemistry, 41(2):495–507, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0025-1.

Rawat:2015:MCS

[743] Arpana Rawat, Radha Chaube, and Keerikkattil P. Joy. Molecular cloning, sequencing and phylogeny of vasotocin receptor genes in the air-breathing catfish Heteropneustes fossilis with sex dimorphic and seasonal variations in tissue expression. Fish Physiology and Biochemistry, 41(2):509–532, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0026-0.

Manor:2015:DGF

[744] Meghan L. Manor, Beth M. Cleveland, and Tim Leeds. Differences in growth, fillet quality, and fatty acid metabolism-related gene expression between juvenile male and female rainbow trout. Fish Physiology and Biochemistry, 41(2):533–547, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0027-z.

Teles:2015:ODT

[745] Andressa Teles, Wanessa Melo Costa, and Vinicius Ronzani Cerqueira. Ontogeny of the digestive tract of Centropomus parallelus larvae. Fish Physiology and Biochemistry, 41(2):549–559, April 2015. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0028-y.

Duan:2015:GCS

[746] Jundan Duan, Gongqing Feng, and Haobin Zhao. Germ cell-specific expression of dead end (dnd) in rare minnow (Gobiocypris rarus). Fish Physiology and Biochemistry, 41(2):561–571, April 2015. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0029-x.

Li:2015:MCE

[747] Jiuxuan Li, Haibin Zhang, and Zhaobin Song. Molecular cloning and expression of two heat-shock protein genes (HSC70/HSP70) from Prenant’s schizothoracin (Schizothorax prenanti). Fish Physiology and Biochemistry, 41(2):573–585, April 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-015-0030-4.

Chi:2015:MCC

[748] Mei L. Chi, Meng Ni, and Hai S. Wen. Molecular cloning and characterization of gonadotropin subunits (GTHα, FSHβ and LHβ) and their regulation by hCG and GnRHa in Japanese sea bass (Lateolabrax japonicas) in vivo. Fish Physiology and Biochemistry, 41(3):587–601, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-014-9992-x.

Mazzoni:2015:ATA

[749] Maurizio Mazzoni, Alessio Bonaldo, and Paolo Clavenzani. α-transducin and α-gustducin immunoreactive cells in the stomach of common sole (Solea solea) fed with mussel meal. Fish Physiology and Biochemistry, 41(3):603–612, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0031-3.

Zhao:2015:IDB

[750] Fei Zhao, Bai Wang, and Shaoguo Ru. Induction of DNA base damage and strand breaks in peripheral erythrocytes and the underlying mechanism in goldfish (Carassius auratus) exposed to monocrotophos. Fish Physiology and Biochemistry, 41(3):613–624, June 2015. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0032-2.

Topal:2015:NEN

[751] Ahmet Topal, Muhammed Atamanalp, and Bahar Yılmaz. Neurotoxic effects of nickel chloride in the rainbow trout brain: Assessment of c-Fos activity, antioxidant responses, acetylcholinesterase activity, and histopathological changes. Fish Physiology and Biochemistry, 41 (3):625–634, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0033-1.

Hu:2015:EDG

[752] Kai Hu, Jing-Xiu Zhang, and Xiao-Qiu Zhou. Effect of dietary glutamine on growth performance, non-specific immunity, expression of cytokine genes, phosphorylation of target of rapamycin (TOR), and antioxidative system in spleen and head kidney of Jian carp (Cyprinus carpio var. Jian). Fish Physiology and Biochemistry, 41(3):635–649, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0034-0.

Hashish:2015:NAS

[753] Emad A. Hashish, Shimaa A. Elgaml, and Ryad Khalil. Nephroprotective and antioxidant significance of selenium and α-tocopherol on lead acetate-induced toxicity of Nile tilapia (Oreochromis niloticus). Fish Physiology and Biochemistry, 41(3):651–660, June 2015. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0035-z.

Li:2015:ECF

[754] Xia Li, Chen Ma, and Ai-Jun Pei. Establishment and characterization of fin cell lines from diploid, triploid, and tetraploid oriental weatherfish (Misgurnus anguillicaudatus). Fish Physiology and Biochemistry, 41(3):661–672, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0036-y.

Sun:2015:ECG

[755] Ai Sun, Song-Lin Chen, and Zhen-Xia Sha. Establishment and characterization of a gonad cell line from half-smooth tongue sole Cynoglossus semilaevis pseudomale. Fish Physiology and Biochemistry, 41(3): 673–683, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0037-x.

Rajan:2015:SLA

[756] Koilmani Emmanuvel Rajan, Subramanian Thangaleela, and Chellam Balasundaram. Spatial learning associated with stimulus response in goldfish Carassius auratus: relationship to activation of CREB signalling. Fish Physiology and Biochemistry, 41(3):685–694, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0038-9.

Sun:2015:EPI

[757] Bo guang Sun and Yong hua Hu. Evaluation of potential internal references for quantitative real-time RT-PCR normalization of gene expression in red drum (Sciaenops ocellatus). Fish Physiology and Biochemistry, 41(3):695–704, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0039-8.

Betancor:2015:RSA

[758] M. B. Betancor, P. F. Almaida-Pagán, and D. R. Tocher. Roles of selenoprotein antioxidant protection in zebrafish, Danio rerio, subjected to dietary oxidative stress. Fish Physiology and Biochemistry, 41(3):705–720, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0040-2.

Nofouzi:2015:IEL

[759] Katayoon Nofouzi, Najmeh Sheikhzadeh, and Javad Ashrafi-Helan. Influence of extremely low frequency electromagnetic fields on growth performance, innate immune response, biochemical parameters and disease resistance in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 41(3):721–731, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0041-1.

Zheng:2015:EDO

[760] Qiaoran Zheng, Yinglong Wu, and Huailiang Xu. Effect of dietary oxidized konjac glucomannan on Schizothorax prenanti growth performance, body composition, intestinal morphology and intestinal microflora. Fish Physiology and Biochemistry, 41(3):733–743, June 2015. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0042-0.

Fernandez:2015:ZVK

[761] Ignacio Fernández, Parameswaran Vijayakumar, and Vincent Laizé. Zebrafish vitamin K epoxide reductases: expression in vivo, along extracellular matrix mineralization and under phylloquinone and warfarin in vitro exposure. Fish Physiology and Biochemistry, 41(3):745–759, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-015-0043-z.

Kumar:2015:MBO

[762] Ravi Kumar and Keerikkattil P. Joy. Melanins as biomarkers of ovarian follicular atresia in the catfish Heteropneustes fossilis: biochemical and histochemical characterization, seasonal variation and hormone effects. Fish Physiology and Biochemistry, 41(3):761–772, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0044-y.

Tian:2015:EPS

[763] Juan Tian, Gen He, and Chengdong Liu. Effects of postprandial starvation on mRNA expression of endocrine-, amino acid and peptide transporter-, and metabolic enzyme-related genes in zebrafish (Danio rerio). Fish Physiology and Biochemistry, 41(3):773–787, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0045-x.

Mohanty:2015:PPW

[764] Bimal Prasanna Mohanty, Tandrima Mitra, and Sasmita Mohanty. Proteomic profiling of white muscle from freshwater catfish Rita rita. Fish Physiology and Biochemistry, 41(3):789–802, June 2015. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0046-9.

Li:2015:CPO

[765] Ru-Qin Li, Yu-Wei Ren, and Zhi-Xin Wu. Comparative pharmacokinetics of oxytetracycline in blunt-snout bream (Megalobrama amblycephala) with single and multiple-dose oral administration. Fish Physiology and Biochemistry, 41(3):803–809, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0047-8.

Raymond:2015:TPF

[766] James A. Raymond. Two potential fish glycerol-3-phosphate phosphatases. Fish Physiology and Biochemistry, 41(3):811–818, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0048-7.

Nie:2015:EDG

[767] Qin Nie, Huijun Miao, and Kangsen Mai. Effects of dietary glucose and dextrin on activity and gene expression of glucokinase and fructose-1,6-bisphosphatase in liver of turbot Scophthalmus maximus. Fish Physiology and Biochemistry, 41(3):819–832, June 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0049-6.

Mulhollem:2015:RLB

[768] Joshua J. Mulhollem, Cory D. Suski, and David H. Wahl. Response of largemouth bass (Micropterus salmoides) from different thermal environments to increased water temperature. Fish Physiology and Biochemistry, 41(4):833–842, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0050-0.

Yorio:2015:GHL

[769] M. P. Di Yorio, T. H. Delgadin, and P. G. Vissio. Growth hormone, luteinizing hormone, and follicle-stimulating hormone regulation by neuropeptide Y in both sexes of the cichlid fish, Cichlasoma dimerus. Fish Physiology and Biochemistry, 41(4):843–852, August 2015. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0051-z.

Zhao:2015:PDM

[770] J. L. Zhao, Y. F. Si, and S. L. Chen. Polymorphisms and DNA methylation level in the CpG site of the GHR1 gene associated with mRNA expression, growth traits and hormone level of half-smooth tongue sole (Cynoglossus semilaevis). Fish Physiology and Biochemistry, 41(4): 853–865, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0052-y.

Liu:2015:ODA

[771] Pin Liu, Hong Ji, and Ping Yu. Ontogenetic development of adipose tissue in grass carp (Ctenopharyngodon idellus). Fish Physiology and Biochemistry, 41(4):867–878, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0053-x.

Valdes:2015:AOS

[772] Joaquín Valdés, Jesús Olivares, and Oliver Schmachtenberg. Analysis of olfactory sensitivity in rainbow trout (Oncorhynchus mykiss) reveals their ability to detect lactic acid, pyruvic acid and four B vitamins. Fish Physiology and Biochemistry, 41(4):879–885, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0054-9.

Newton:2015:DEA

[773] Kyle C. Newton, James Wraith, and Kathryn A. Dickson. Digestive enzyme activities are higher in the shortfin mako shark, Isurus oxyrinchus, than in ectothermic sharks as a result of visceral endothermy. Fish Physiology and Biochemistry, 41(4):887–898, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0055-8.

Lu:2015:TDO

[774] Aijun Lü, Xiucai Hu, and Jingfeng Sun. Tissue distribution of olive flounder VDAC2 and its expression in fish cell lines. Fish Physiology and Biochemistry, 41(4):899–907, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0056-7.

Shahi:2015:SRP

[775] Neetu Shahi, Sumanta Kumar Mallik, and Atul Kumar Singh. Spermatogenesis and related plasma androgen and progestin level in wild male golden mahseer, Tor putitora (Hamilton, 1822), during the spawning season. Fish Physiology and Biochemistry, 41(4):909–920, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0057-6.

Lopez:2015:EFR

[776] Lus M. López, Maricela Flores-Ibarra, and Conal D. True. Effect of fishmeal replacement by soy protein concentrate with taurine supplementation on growth performance, hematological and biochemical status, and liver histology of totoaba juveniles (Totoaba macdonaldi). Fish Physiology and Biochemistry, 41(4):921–936, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0058-5.

Ma:2015:CPP

[777] X. Y. Ma, J. Qiang, and P. Xu. Changes in the physiological parameters, fatty acid metabolism, and SCD activity and expression in juvenile GIFT tilapia (Oreochromis niloticus) reared at three different temperatures. Fish Physiology and Biochemistry, 41(4):937–950, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0059-4.

Wang:2015:FME

[778] Jing jing Wang and Li Sun. Ferritin M of Paralichthys olivaceus possesses antimicrobial and antioxidative properties. Fish Physiology and Biochemistry, 41(4):951–959, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0060-y.

Wang:2015:EIZ

[779] Xueqian Wang, Xin Wang, and Dong Liu. Egfl6 is involved in zebrafish notochord development. Fish Physiology and Biochemistry, 41 (4):961–969, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0061-x.

Patterson:2015:PMD

[780] Joshua T. Patterson and Christopher C. Green. Physiological management of dietary deficiency in n-3 fatty acids by spawning Gulf killifish (Fundulus grandis). Fish Physiology and Biochemistry, 41(4): 971–979, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0062-9.

Lu:2015:EGI

[781] Rong-Hua Lu, Yi Zhou, and Yu-Hua Zhao. Effects of glucose, insulin and triiodothyroxine on leptin and leptin receptor expression and the effects of leptin on activities of enzymes related to glucose metabolism in grass carp (Ctenopharyngodon idella) hepatocytes. Fish Physiology and Biochemistry, 41(4):981–989, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0063-8.

Kaneko:2015:CIL

[782] Nobuto Kaneko, Natsumi Taniyama, and Munetaka Shimizu. Circulating insulin-like growth factor I in juvenile chum salmon: relationship with growth rate and changes during downstream and coastal migration in northeastern Hokkaido, Japan. Fish Physiology and Biochemistry, 41 (4):991–1003, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0064-7.

Simcic:2015:EIT

[783] Tatjana Simčič, Dušan Jesenšek, and Anton Brancelj. Effects of increased temperature on metabolic activity and oxidative stress in the first life stages of marble trout (Salmo marmoratus). Fish Physiology and Biochemistry, 41(4):1005–1014, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0065-6.

Gonzalez:2015:PRA

[784] Carlos Aguilera González, Julio Cruz, and Roberto Mendoza Alfaro. Physiological response of alligator gar juveniles (Atractosteus spatula) exposed to sub-lethal doses of pollutants. Fish Physiology and Biochemistry, 41(4):1015–1027, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0066-5.

Hannesson:2015:SGP

[785] Kirsten O. Hannesson, Elisabeth Ytteborg, and Mona E. Pedersen. Sulphated glycosaminoglycans and proteoglycans in the developing vertebral column of juvenile Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 41(4):1029–1051, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0067-4.

Cheng:2015:ICF

[786] Chang-Hong Cheng, Fang-Fang Yang, and Li-Wei Liu. Identification, characterization and functional analysis of anti-apoptotic protein BCL-2-like gene from pufferfish, Takifugu obscurus, responding to bacterial challenge. Fish Physiology and Biochemistry, 41(4):1053–1064, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0068-3.

Hernandez:2015:PII

[787] A. Hernández, B. García García, and M. D. Hernández. Preliminary insights into the incorporation of rosemary extract (Rosmarinus officinalis L.) in fish feed: influence on performance and physiology of gilthead seabream (Sparus aurata). Fish Physiology and Biochemistry, 41 (4):1065–1074, August 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0069-2.

Frias-Quintana:2015:DDT

[788] C. A. Frías-Quintana, G. Márquez-Couturier, and E. Gisbert. Development of digestive tract and enzyme activities during the early ontogeny of the tropical gar Atractosteus tropicus. Fish Physiology and Biochemistry, 41(5):1075–1091, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0070-9.

Su:2015:IES

[789] Yanfang Su, Yuanshuai Fu, and Lina Gao. Identification and expression of SRF targeted by miR-133a during early development of Paralichthys olivaceus. Fish Physiology and Biochemistry, 41(5):1093–1104, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-015-0071-8.

Ikegami:2015:ITD

[790] Taro Ikegami, Akihiro Takemura, and Mitsuhiro Furuse. Increase in telencephalic dopamine and cerebellar norepinephrine contents by hydrostatic pressure in goldfish: the possible involvement in hydrostatic pressure-related locomotion. Fish Physiology and Biochemistry, 41(5): 1105–1115, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0072-7.

Galaviz:2015:DSD

[791] Mario A. Galaviz, Lus M. López, and Enric Gisbert. Digestive system development and study of acid and alkaline protease digestive capacities using biochemical and molecular approaches in totoaba (Totoaba macdonaldi) larvae. Fish Physiology and Biochemistry, 41(5): 1117–1130, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0073-6.

Dalmolin:2015:FIA

[792] Camila Dalmolin, Daniela Volcan Almeida, and Luis Fernando Marins. Food intake and appetite control in a GH-transgenic zebrafish. Fish Physiology and Biochemistry, 41(5):1131–1141, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0074-5.

Jiang:2015:EGG

[793] Jun Jiang, Dan Shi, and Ye Zhao. Effects of glutamate on growth, antioxidant capacity, and antioxidant-related signaling molecule expression in primary cultures of fish enterocytes. Fish Physiology and Biochemistry, 41(5):1143–1153, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0076-3.

Brzuzan:2015:CMW

[794] P. Brzuzan, C. Kramer, and M. Woźny. c-myc in whitefish (Coregonus lavaretus): structure, expression, and insights into possible posttranscriptional regulatory mechanism. Fish Physiology and Biochemistry, 41 (5):1155–1171, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0077-2.

Chaube:2015:CVL

[795] Radha Chaube, Rahul Kumar Singh, and Keerikattil P. Joy. Changes in vasotocin levels in relation to ovarian development in the catfish Heteropneustes fossilis exposed to altered photoperiod and temperature. Fish Physiology and Biochemistry, 41(5):1173–1186, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0078-1.

Betancor:2015:EDF

[796] M. B. Betancor, P. F. Almaida-Pagán, and D. R. Tocher. Effects of dietary fatty acids on mitochondrial phospholipid compositions, oxidative status and mitochondrial gene expression of zebrafish at different ages. Fish Physiology and Biochemistry, 41(5):1187–1204, October 2015. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0079-0.

Rodrigues:2015:MRA

[797] Edson Rodrigues, Jr., Mariana Feijó-Oliveira, and Helena Passeri Lavrado. Metabolic responses of the Antarctic fishes Notothenia rossii and Notothenia coriiceps to sewage pollution. Fish Physiology and Biochemistry, 41(5):1205–1220, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0080-7.

Fu:2015:KGS

[798] Yuanshuai Fu, Junling Zhang, and Liang Jia. A key gene of the small RNA pathway in the flounder, Paralichthys olivaceus: identification and functional characterization of dicer. Fish Physiology and Biochemistry, 41(5):1221–1231, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0081-6.

Novelli:2015:DSE

[799] B. Novelli, J. A. Socorro, and L. Molina Domínguez. Development of seahorse (Hippocampus reidi, Ginsburg 1933): histological and histochemical study. Fish Physiology and Biochemistry, 41(5):1233–1251, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-015-0082-5.

Hoffmayer:2015:CSM

[800] Eric R. Hoffmayer, Jill M. Hendon, and Matthew D. Campbell. A comparison of single and multiple stressor protocols to assess acute stress in a coastal shark species, Rhizoprionodon terraenovae. Fish Physiology and Biochemistry, 41(5):1253–1260, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0083-4.

Pijanowski:2015:AHP

[801] L. Pijanowski, P. Jurecka, and M. Chadzinska. Activity of the hypothalamus–pituitary–interrenal axis (HPI axis) and immune response in carp lines with different susceptibility to disease. Fish Physiology and Biochemistry, 41(5):1261–1278, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0084-3.

Liu:2015:MCH

[802] Qiuping Liu, Shuting Huang, and Yan Liu. Molecular characterization of heat–shock protein 90 gene and its expression in Gobiocypris rarus juveniles exposed to pentachlorophenol. Fish Physiology and Biochemistry, 41(5):1279–1291, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0085-2.

Xia:2015:BMS

[803] Ji-Gang Xia, Li-Juan Nie, and Shi-Jian Fu. Behavior, metabolism and swimming physiology in juvenile Spinibarbus sinensis exposed to PFOS under different temperatures. Fish Physiology and Biochemistry, 41(5): 1293–1304, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0086-1.

Li:2015:CCI

[804] HuaTao Li, Lin Feng, and XiaoQiu Zhou. Ca^{2+} and caspases are involved in hydroxyl radical-induced apoptosis in erythrocytes of Jian carp (Cyprinus carpio var. Jian). Fish Physiology and Biochemistry, 41(5): 1305–1319, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0087-0.

Gabriel:2015:DEA

[805] Ndakalimwe Naftal Gabriel, Jun Qiang, and Kai Liu. Dietary Aloe vera improves plasma lipid profile, antioxidant, and hepatoprotective enzyme activities in GIFT-tilapia (Oreochromis niloticus) after Streptococcus iniae challenge. Fish Physiology and Biochemistry, 41(5): 1321–1332, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0088-z.

Guerreiro:2015:ESC

[806] Inês Guerreiro, Paula Enes, and Aires Oliva-Teles. Effects of short-chain fructooligosaccharides (scFOS) and rearing temperature on growth performance and hepatic intermediary metabolism in gilthead sea bream (Sparus aurata) juveniles. Fish Physiology and Biochemistry, 41(5): 1333–1344, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0089-y.

Luo:2015:EED

[807] Sheng-Wei Luo, Wei-Na Wang, and Liang-Biao Chen. Effects of a Dissostichus mawsoni-CaM recombinant proteins feed additive on the juvenile orange-spotted grouper (Epinephelus coioides) under the acute low temperature challenge. Fish Physiology and Biochemistry, 41(5): 1345–1358, October 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0090-5.

Kuzmina:2015:RPI

[808] V. V. Kuz’mina, E. G. Skvortsova, and K. E. Kovalenko. Role of peptidases of the intestinal microflora and prey in temperature adaptations of the digestive system in planktivorous and benthivorous fish. Fish Physiology and Biochemistry, 41(6):1359–1368, December 2015. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0091-4.

Vargas-Chacoff:2015:EMG

[809] L. Vargas-Chacoff, E. Saavedra, and C. Bertrán. Effects on the metabolism, growth, digestive capacity and osmoregulation of juvenile of Sub-Antarctic Notothenioid fish Eleginops maclovinus acclimated at different salinities. Fish Physiology and Biochemistry, 41(6):1369– 1381, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0092-3.

Seibt:2015:IEA

[810] Kelly Juliana Seibt, Renata da Luz Oliveira, and Carla Denise Bonan. Investigation into effects of antipsychotics on ectonucleotidase and adenosine deaminase in zebrafish brain. Fish Physiology and Biochemistry, 41(6):1383–1392, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0093-2.

Hu:2015:OKC

[811] Wei Hu, Zhi Luo, and Jia-Lang Zheng. Ontogeny and kinetics of carnitine palmitoyltransferase I in hepatopancreas and skeletal muscle of grass carp (Ctenopharyngodon idella). Fish Physiology and Biochemistry, 41(6): 1393–1401, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0094-1.

Kohn:2015:PAE

[812] Yair Y. Kohn, Jane E. Symonds, and P. Mark Lokman. Proteomic analysis of early-stage embryos: implications for egg quality in hapuku (Polyprion oxygeneios). Fish Physiology and Biochemistry, 41(6): 1403–1417, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0095-0.

Wang:2015:MIE

[813] Dan-Dan Wang, Gui-Rong Zhang, and Kun-Ci Chen. Molecular identification and expression of the Foxl2 gene during gonadal sex differentiation in northern snakehead Channa argus. Fish Physiology and Biochemistry, 41(6):1419–1433, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0096-z.

Moreira:2015:IGG

[814] Renata Guimarães Moreira, Renato Massaaki Honji, and Alexandre Wagner Silva Hilsdorf. The involvement of gonadotropins and gonadal steroids in the ovulatory dysfunction of the potamodromous Salminus hilarii (Teleostei: Characidae) in captivity. Fish Physiology and Biochemistry, 41(6):1435–1447, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0097-y.

Yang:2015:MCE

[815] Qibin Yang, Panlong Zheng, and Jian G. Qin. Molecular cloning and expression analysis of the retinoid X receptor (RXR) gene in golden pompano Trachinotus ovatus fed Artemia nauplii with different enrichments. Fish Physiology and Biochemistry, 41(6):1449–1461, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0098-x.

Sun:2015:IME

[816] Bo guang Sun and Yong hua Hu. Identification, mRNA expression profiling and activity characterization of cathepsin L from red drum (Sciaenops ocellatus). Fish Physiology and Biochemistry, 41(6):1463– 1473, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0099-9.

Mohammadian:2015:EDS

[817] Takavar Mohammadian, Pedram Malekpouri, and Mohammad Anwar Zainodini. Effects of different spawning agents on serum levels of reproductive steroid hormones and cortisol level in adult female Barbus sharpeyi (Gunther, 1874). Fish Physiology and Biochemistry, 41(6): 1475–1489, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0100-7.

Yao:2015:CAL

[818] Zongli Yao, Qifang Lai, and Hui Wang. Carbonic anhydrase 2-like and Na^{+}-K^{+}-ATPase α gene expression in medaka (Oryzias latipes) under carbonate alkalinity stress. Fish Physiology and Biochemistry, 41(6): 1491–1500, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0101-6.

Wolkers:2015:SCI

[819] Carla Patrícia Bejo Wolkers, Mônica Serra, and Elisabeth Criscuolo Urbinati. Social challenge increases cortisol and hypothalamic monoamine levels in matrinxã (Brycon amazonicus). Fish Physiology and Biochemistry, 41(6):1501–1508, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0102-5.

Soares:2015:CAN

[820] Florbela Soares, Laura Ribeiro, and Pedro Pousão-Ferreira. Comparative analysis on natural spawning of F1 meagre, Argyrosomus regius, with wild broodstock spawns in Portugal. Fish Physiology and Biochemistry, 41(6):1509–1514, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0103-4.

Chen:2015:EOG

[821] Dong Chen, Jiang Liu, and Lihong Zhang. Expression and ontogeny of growth hormone (Gh) in the protogynous hermaphroditic ricefield eel (Monopterus albus). Fish Physiology and Biochemistry, 41(6):1515– 1525, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0104-3.

Jensen:2015:ETD

[822] Linda B. Jensen, Thomas Wahli, and Carolina Tafalla. Effect of temperature and diet on wound healing in Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 41(6):1527–1543, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0105-2.

Shang:2015:SSC

[823] Mei Shang, Baofeng Su, and Rex A. Dunham. Spermatogonial stem cells specific marker identification in channel catfish, Ictalurus punctatus and blue catfish, I. furcatus. Fish Physiology and Biochemistry, 41(6): 1545–1556, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0106-1.

Lovatto:2015:NEP

[824] Naglezi de Menezes Lovatto, Fernanda Rodrigues Goulart, and Leila Picolli da Silva. Nutritional evaluation of phosphorylated pumpkin seed (Cucurbita moschata) protein concentrate in silver catfish Rhamdia quelen (Quoy and Gaimard, 1824). Fish Physiology and Biochemistry, 41(6): 1557–1567, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0107-0.

Lee:2015:IES

[825] Dongwook Lee, Jun Hyung Ryu, and Seung Pyo Gong. Identification of embryonic stem cell activities in an embryonic cell line derived from marine medaka (Oryzias dancena). Fish Physiology and Biochemistry, 41(6):1569–1576, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-015-0108-z.

Zhang:2015:CDE

[826] Yun-Long Zhang, Qiao-Wan Wu, and Qi-Xue Fan. Changes in digestive enzyme activities during larval development of Chinese loach Paramisgurnus dabryanus (Dabry de Thiersant, 1872). Fish Physiology and Biochemistry, 41(6):1577–1585, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0109-y.

Abdelmalek:2015:EXO

[827] Baha Eddine Abdelmalek, Dorra Driss, and Ali Bougatef. Effect of xylan oligosaccharides generated from corncobs on food acceptability, growth performance, haematology and immunological parameters of Dicentrarchus labrax fingerlings. Fish Physiology and Biochemistry, 41(6): 1587–1596, December 2015. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0110-5.

Guest:2016:DMC

[828] Taylor W. Guest, Reginald B. Blaylock, and Andrew N. Evans. Development of a modified cortisol extraction procedure for intermediately sized fish not amenable to whole-body or plasma extraction methods. Fish Physiology and Biochemistry, 42(1):1–6, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0111-4.

Wood:2016:DPE

[829] Richard K. Wood, Emma Crowley, and Christopher J. Martyniuk. Developmental profiles and expression of the DNA methyltransferase genes in the fathead minnow (Pimephales promelas) following exposure to di-2-ethylhexyl phthalate. Fish Physiology and Biochemistry, 42(1): 7–18, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0112-3.

Wu:2016:DLF

[830] Haiyun Wu, Hitoshi Ohnuki, and Hideaki Endo. Development of a labelfree immunosensor system for detecting plasma cortisol levels in fish. Fish Physiology and Biochemistry, 42(1):19–27, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0113-2.

Vo:2016:DPC

[831] Nguyen T. K. Vo and Niels C. Bols. Demonstration of primary cilia and acetylated α-tubulin in fish endothelial, epithelial and fibroblast cell lines. Fish Physiology and Biochemistry, 42(1):29–38, February 2016.

CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0114-1.

Wu:2016:YFE

[832] Hongwei Wu, Fangjun Lin, and Zhiqiong Li. Ya-fish (Schizothorax prenanti) spexin: identification, tissue distribution and mRNA expression responses to periprandial and fasting. Fish Physiology and Biochemistry, 42(1):39–49, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0115-0.

Chen:2016:CEE

[833] Pei Chen, Qin Tang, and Chunfang Wang. Characterizing and evaluating the expression of the type IIb sodium-dependent phosphate cotransporter (slc34a2) gene and its potential influence on phosphorus utilization efficiency in yellow catfish (Pelteobagrus fulvidraco). Fish Physiology and Biochemistry, 42(1):51–64, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0116-z.

Shabanzadeh:2016:GPI

[834] Sadigheh Shabanzadeh, Maryam Shapoori, and Amir Ali Shahbazfar. Growth performance, intestinal histology, and biochemical parameters of rainbow trout (Oncorhynchus mykiss) in response to dietary inclusion of heat-killed Gordonia bronchialis. Fish Physiology and Biochemistry, 42(1):65–71, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0117-y.

Becker:2016:PST

[835] Alexssandro G. Becker, Thaylise V. Parodi, and Bernardo Baldisserotto. Pre-sedation and transport of Rhamdia quelen in water containing essential oil of Lippia alba: metabolic and physiological responses. Fish Physiology and Biochemistry, 42(1):73–81, February 2016. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0118-x.

Zhao:2016:ICT

[836] Yini Zhao, Qi Sun, and Xianle Yang. Isolation, characterization, and tissue-specific expression of GABA A receptor α 1 subunit gene of Carassius auratus gibelio after avermectin treatment. Fish Physiology and Biochemistry, 42(1):83–92, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0119-9.

daSilva:2016:EOS

[837] Janaína Camacho da Silva, Antonio Sergio Varela Junior, and Carine Dahl Corcini. The effects of osmolality on sperm quality in Jenynsia multidentata (Cyprinodontiformes: Anablepidae). Fish Physiology and Biochemistry, 42(1):93–102, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0120-3.

Forwood:2016:PTY

[838] James M. Forwood, Erin J. Bubner, and Marty R. Deveney. Praziquantel treatment for yellowtail kingfish (Seriola lalandi): dose and duration safety study. Fish Physiology and Biochemistry, 42(1):103– 109, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0121-2.

Liu:2016:MCB

[839] Xiao-Hong Liu, Zhi-Jian Wang, and Yao-Guang Zhang. Molecular characterization of beclin 1 in rare minnow (Gobiocypris rarus) and its expression after waterborne cadmium exposure. Fish Physiology and Biochemistry, 42(1):111–123, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0122-1.

Purohit:2016:EHG

[840] Gopal Krishna Purohit, Arabinda Mahanty, and Sasmita Mohanty. Evaluation of housekeeping genes as references for quantitative real-time PCR analysis of gene expression in the murrel Channa striatus under high-temperature stress. Fish Physiology and Biochemistry, 42(1): 125–135, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0123-0.

Li:2016:ACE

[841] Keshuai Li, Rolf Erik Olsen, and Yngvar Olsen. Atlantic cod (Gadus morhua) larvae can biosynthesis phospholipid de novo from 2-oleoyl-glycerol and glycerol precursors. Fish Physiology and Biochemistry, 42 (1):137–147, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0125-y.

Crovetto:2016:SBC

[842] Cecilia Alejandra Crovetto and Osvaldo León Córdoba. Structural and biochemical characterization and evolutionary relationships of the fatty acid-binding protein 10 (fabp10) of hake (Merluccius hubbsi). Fish Physiology and Biochemistry, 42(1):149–165, February 2016. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0126-x.

Niu:2016:CGS

[843] Jingjing Niu, Conghui Liu, and Jie Qi. Characterization and genomic structure of Dnah9 , and its roles in nodal signaling pathways in the Japanese flounder (Paralichthys olivaceus). Fish Physiology and Biochemistry, 42(1):167–178, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0127-9.

Jia:2016:MCQ

[844] Yudong Jia, Ai Sun, and Jilin Lei. Molecular characterization and quantification of the follicle-stimulating hormone receptor in turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 42(1): 179–191, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0128-8.

Xu:2016:MSS

[845] Gefeng Xu, Tianqing Huang, and Zhenbo Mu. Morphology, sex steroid level and gene expression analysis in gonadal sex reversal of triploid female (XXX) rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 42(1):193–202, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0129-7.

Castro:2016:EFO

[846] Carolina Castro, Ana Couto, and Aires Oliva-Teles. Effects of fish oil replacement by vegetable oil blend on digestive enzymes and tissue histomorphology of European sea bass (Dicentrarchus labrax) juveniles. Fish Physiology and Biochemistry, 42(1):203–217, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0130-1.

Kammerer:2016:PES

[847] Brittany D. Kammerer, Tien-Chieh Hung, and Swee J. Teh. Physiological effects of salinity on Delta Smelt, Hypomesus transpacificus. Fish Physiology and Biochemistry, 42(1):219–232, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0131-0.

Luckenbach:2016:GSD

[848] J. Adam Luckenbach and William T. Fairgrieve. Gonadal sex differentiation and effects of dietary methyltestosterone treatment in sablefish (Anoplopoma fimbria). Fish Physiology and Biochemistry, 42(1): 233–248, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0132-z.

Hu:2016:DED

[849] Jun-Ru Hu, Yan-Hua Huang, and Jun-Ming Cao. Deficient and excess dietary selenium levels affect growth performance, blood cells apoptosis and liver HSP70 expression in juvenile yellow catfish Pelteobagrus fulvidraco. Fish Physiology and Biochemistry, 42(1):249–261, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-015-0133-y.

Gonzalez:2016:SRE

[850] Margarita P. González, Luis Vargas-Chacoff, and Sandra L. Marín. Stress response of Salmo salar (Linnaeus 1758) when heavily infested by Caligus rogercresseyi (Boxshall & Bravo 2000) copepodids. Fish Physiology and Biochemistry, 42(1):263–274, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0134-x.

Das:2016:EPO

[851] Sumana Das, Suravi Majumder, and Dilip Mukherjee. Effects of phenol on ovarian P450arom gene expression and aromatase activity in vivo and antioxidant metabolism in common carp Cyprinus carpio. Fish Physiology and Biochemistry, 42(1):275–286, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0135-9.

Li:2016:ACP

[852] Zhi-Hua Li, Li-Qiao Zhong, and Wei-Na Mu. Alteration of cytochrome P450 1 regulation and HSP 70 level in brain of juvenile common carp (Cyprinus carpio) after chronic exposure to tributyltin. Fish Physiology and Biochemistry, 42(1):287–294, February 2016. CODEN FPBIEP.

ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0136-8.

Li:2016:EBC

[853] Ge Li, Hang Xie, and Yiping Luo. Effects of body chemical components on the allometric scaling of the resting metabolic rate in four species of cyprinids. Fish Physiology and Biochemistry, 42(1): 295–301, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0137-7.

Jing:2016:ESK

[854] Hongli Jing, Longying Gao, and Shaoqiang Wu. Establishment from the snout and kidney of goldfish, Carassius auratus, of two new cell lines and their susceptibility to infectious pancreatic necrosis virus. Fish Physiology and Biochemistry, 42(1):303–311, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0138-6.

Chen:2016:ICU

[855] Liyong Chen, Fenfang Wu, and Bo Feng. Identification and characterization of a ubiquitin-conjugating enzyme UBE2A gene from lamprey. Fish Physiology and Biochemistry, 42(1):313–320, February 2016. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0139-5.

Pes:2016:EDE

[856] Tanise S. Pês, Etiane M. H. Saccol, and Maria A. Pavanato. Effect of diets enriched with rutin on blood parameters, oxidative biomarkers and pituitary hormone expression in silver catfish (Rhamdia quelen). Fish Physiology and Biochemistry, 42(1):321–333, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0140-z.

Fang:2016:PSA

[857] Xingxing Fang, Jianguo Zhou, and Xiuhong Liu. Pharmacokinetics of sarafloxacin in allogynogenetic silver crucian carp, Carassius auratus gibelio. Fish Physiology and Biochemistry, 42(1):335–341, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0141-y.

Coutinho:2016:MRD

[858] Filipe Coutinho, Helena Peres, and Paula Enes. Metabolic responses to dietary protein/carbohydrate ratios in zebra sea bream (Diplodus cervinus, Lowe, 1838) juveniles. Fish Physiology and Biochemistry, 42(1): 343–352, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0142-x.

Hamed:2016:GRE

[859] R. R. Hamed, N. S. M. Saleh, and S. S. Abdel-Ghany. Glutathione and its related enzymes in the gonad of Nile tilapia (Oreochromis niloticus). Fish Physiology and Biochemistry, 42(1):353–364, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0143-9.

Mohammed-Geba:2016:ILG

[860] Khaled Mohammed-Geba, J. A. Martos-Sitcha, and G. Martínez-Rodríguez. Insulin-like growth factor 1 (IGF-1) regulates prolactin, growth hormone, and IGF-1 receptor expression in the pituitary gland of the gilthead sea bream Sparus aurata. Fish Physiology and Biochemistry, 42(1):365–377, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0144-8.

Liu:2016:GPR

[861] Dongwu Liu, Kangsen Mai, and Qinghui Ai. GSK-3β participates in the regulation of hepatic lipid deposition in large yellow croaker (Larmichthys crocea). Fish Physiology and Biochemistry, 42(1):379– 388, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0145-7.

Madureira:2016:ACO

[862] Tânia Vieira Madureira, L. Filipe C. Castro, and Eduardo Rocha. Acylcoenzyme a oxidases 1 and 3 in brown trout (Salmo trutta f. fario): Can peroxisomal fatty acid β-oxidation be regulated by estrogen signaling? Fish Physiology and Biochemistry, 42(1):389–401, February 2016. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0146-6.

Anonymous:2016:A

[863] Anonymous. Acknowledgments. Fish Physiology and Biochemistry, 42 (1):403–406, February 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0198-2.

Si:2016:GPD

[864] Yufeng Si, Feng He, and Songlin Chen. Genetic polymorphisms and DNA methylation in exon 1 CpG-rich regions of PACAP gene and its effect on mRNA expression and growth traits in half smooth tongue sole (Cynoglossus semilaevis). Fish Physiology and Biochemistry, 42 (2):407–421, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0147-5.

Wang:2016:RRG

[865] Xinyan Wang, Yafei Guo, and Hong Zhou. Regulatory roles of grass carp EpCAM in cell morphology, proliferation and migration. Fish Physiology and Biochemistry, 42(2):423–430, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0148-4.

Suely:2016:THE

[866] A. Suely, H. Zabed, and P. Ganesan. Toxicological and hematological effect of Terminalia arjuna bark extract on a freshwater catfish, Heteropneustes fossilis. Fish Physiology and Biochemistry, 42 (2):431–444, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0149-3.

Murussi:2016:EDG

[867] Camila R. Murussi, Maiara D. Costa, and Vania L. Loro. Exposure to different glyphosate formulations on the oxidative and histological status of Rhamdia quelen. Fish Physiology and Biochemistry, 42 (2):445–455, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0150-x.

Najdegerami:2016:EBG

[868] Ebrahim H. Najdegerami, Farideh Bakhshi, and Forouzan Bagherzadeh Lakani. Effects of biofloc on growth performance, digestive enzyme activities and liver histology of common carp (Cyprinus carpio L.) fingerlings in zero-water exchange system. Fish Physiology and Biochemistry, 42 (2):457–465, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0151-9.

Liu:2016:SAE

[869] Conghui Liu, Wei Liu, and Xubo Wang. Sequences analyses and expression profiles in tissues and embryos of Japanese flounder (Paralichthys olivaceus) PRDM1. Fish Physiology and Biochemistry, 42 (2):467–482, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0152-8.

Kuzu:2016:PGP

[870] Muslum Kuzu, Abdulselam Aslan, and Naim Uzun. Purification of glucose-6-phosphate dehydrogenase and glutathione reductase enzymes from the gill tissue of Lake Van fish and analyzing the effects of some chalcone derivatives on enzyme activities. Fish Physiology and Biochemistry, 42(2):483–491, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0153-7.

Kuzu:2016:EPG

[871] Muslum Kuzu, Abdulselam Aslan, and Naim Uzun. Erratum to: Purification of glucose-6-phosphate dehydrogenase and glutathione reductase enzymes from the gill tissue of Lake Van fish and analyzing the effects of some chalcone derivatives on enzyme activities. Fish Physiology and Biochemistry, 42(2):493, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0209-3.

Hamza:2016:EVP

[872] Ahlem Hamza, Kais Fdhila, and Ahmed Sleheddine Masmoudi. Virgibacillus proomii and Bacillus mojavensis as probiotics in sea bass (Dicentrarchus labrax) larvae: effects on growth performance and digestive enzyme activities. Fish Physiology and Biochemistry, 42(2): 495–507, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0154-6.

Pena:2016:IES

[873] B. Peña, A. Isla, and J. Figueroa. Immunostimulatory effect of salmon prolactin on expression of toll-like receptors in Oncorhynchus mykiss infected with Piscirickettsia salmonis. Fish Physiology and Biochemistry, 42(2):509–516, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0155-5.

Frias-Quintana:2016:UCM

[874] C. A. Frías-Quintana, J. Domínguez-Lorenzo, and R. Martínez-García. Using cornstarch in microparticulate diets for larvicultured tropical gar (Atractosteus tropicus). Fish Physiology and Biochemistry, 42(2): 517–528, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0156-4.

Shan:2016:MHC

[875] Xiujuan Shan, Hanfeng Quan, and Shuozeng Dou. Morphological and histological changes in digestive tract development during starvation in the miiuy croaker. Fish Physiology and Biochemistry, 42 (2):529–546, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0157-3.

Yuan:2016:IET

[876] Junqing Yuan, Jiajun Jiang, and Quanqi Zhang. Insights into Trx1, TRP14, and Prx1 homologs of Paralichthys olivaceus: molecular profiles and transcriptional responses to immune stimulations. Fish Physiology and Biochemistry, 42(2):547–561, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0158-2.

Dzyuba:2016:ASS

[877] Viktoriya Dzyuba, Jacky Cosson, and Otomar Linhart. The antioxidant system of seminal fluid during in vitro storage of sterlet Acipenser ruthenus sperm. Fish Physiology and Biochemistry, 42(2):563–568, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-015-0159-1.

Vargas:2016:CSP

[878] Rafael Vargas and Isabel Cristina Vásquez. Cardiac and somatic parameters in zebrafish: tools for the evaluation of cardiovascular function. Fish Physiology and Biochemistry, 42(2):569–577, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0160-8.

Pal:2016:GMR

[879] Palash Kumar Pal, Kazi Nurul Hasan, and Saumen Kumar Maitra. Gut melatonin response to microbial infection in carp Catla catla. Fish Physiology and Biochemistry, 42(2):579–592, April 2016. CODEN FPBIEP.

ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0161-7.

Trabelsi:2016:EHT

[880] Awatef Trabelsi, Andrzej Jaworski, and Pascal Fontaine. The effect of hatching time on the bioenergetics of northern pike (Esox lucius) larvae from a single egg batch during the endogenous feeding period. Fish Physiology and Biochemistry, 42(2):593–606, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0162-6.

Schnitzler:2016:TGT

[881] Joseph G. Schnitzler, Peter H. M. Klaren, and Krishna Das. The thyroid gland and thyroid hormones in sheepshead minnow (Cyprinodon variegatus) during early development and metamorphosis. Fish Physiology and Biochemistry, 42(2):607–616, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0163-5.

El-Barbary:2016:DAE

[882] Manal I. El-Barbary. Detoxification and antioxidant effects of garlic and curcumin in Oreochromis niloticus injected with aflatoxin B_{1} with reference to gene expression of glutathione peroxidase (GPx) by RT-PCR. Fish Physiology and Biochemistry, 42(2):617–629, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0164-4.

Celi:2016:VNP

[883] Monica Celi, Francesco Filiciotto, and Giuseppa Buscaino. Vessel noise pollution as a human threat to fish: assessment of the stress response in gilthead sea bream (Sparus aurata, Linnaeus 1758). Fish Physiology and Biochemistry, 42(2):631–641, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0165-3.

Guernic:2016:SEA

[884] Antoine Le Guernic, Wilfried Sanchez, and Béatrice Gagnaire. In situ experiments to assess effects of constraints linked to caging on ecotoxicity biomarkers of the three-spined stickleback (Gasterosteus aculeatus L.). Fish Physiology and Biochemistry, 42(2):643–657, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0166-2.

Chen:2016:CRC

[885] Yadong Chen, Shuhong Zhou, and Yang Liu. Chemokine receptor CXCR3 in turbot (Scophthalmus maximus): cloning, characterization and its responses to lipopolysaccharide. Fish Physiology and Biochemistry, 42(2):659–671, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-015-0167-1.

Singh:2016:OSI

[886] Samar Pal Singh, JaiGopal Sharma, and Rina Chakrabarti. Oxygen stress: impact on innate immune system, antioxidant defence system and expression of HIF-1α and ATPase 6 genes in Catla catla. Fish Physiology and Biochemistry, 42(2):673–688, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0168-0.

Xu:2016:FRA

[887] Chao Xu, Xiang-Fei Li, and Wen-Bin Liu. Feeding rates affect growth, intestinal digestive and absorptive capabilities and endocrine functions of juvenile blunt snout bream Megalobrama amblycephala. Fish Physiology and Biochemistry, 42(2):689–700, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0169-z.

Wang:2016:EHS

[888] Yanni Wang, Zhe Liu, and Li Jiang. Effects of heat stress on respiratory burst, oxidative damage and serpinh1 (hsp47 ) mRNA expression in rainbow trout Oncorhynchus mykiss. Fish Physiology and Biochemistry, 42(2):701–710, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0170-6.

Lushchak:2016:CIO

[889] Volodymyr I. Lushchak. Contaminant-induced oxidative stress in fish: a mechanistic approach. Fish Physiology and Biochemistry, 42(2): 711–747, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0171-5.

Zhao:2016:AAG

[890] Feng Zhao, Beibei Wu, and Ping Zhuang. Adaptive alterations on gill Na^{+}, K^{+}-ATPase activity and mitochondrion-rich cells of juvenile Acipenser sinensis acclimated to brackish water. Fish Physiology and Biochemistry, 42(2):749–756, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0172-4.

Kareem:2016:ESD

[891] Zana H. Kareem, Yasser M. Abdelhadi, and Nicholas Romano. Effects of some dietary crude plant extracts on the growth and gonadal maturity of Nile tilapia (Oreochromis niloticus) and their resistance to Streptococcus agalactiae infection. Fish Physiology and Biochemistry, 42 (2):757–769, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0173-3.

Veron:2016:LTF

[892] Vincent Véron, Stéphane Panserat, and Françoise Médale. Long-term feeding a plant-based diet devoid of marine ingredients strongly affects certain key metabolic enzymes in the rainbow trout liver. Fish Physiology and Biochemistry, 42(2):771–785, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0174-2.

Rajaram:2016:IBS

[893] Shailja Rajaram, Hiral Murawala, and Suresh Balakrishnan. Inhibition of BMP signaling reduces MMP-2 and MMP-9 expression and obstructs wound healing in regenerating fin of teleost fish Poecilia latipinna. Fish Physiology and Biochemistry, 42(2):787–794, April 2016. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0175-1.

Caruso:2016:ETO

[894] Catherine C. Caruso, Timothy S. Breton, and David L. Berlinsky. The effects of temperature on ovarian aromatase (cyp19a1a) expression and sex differentiation in summer flounder (Paralichthys dentatus). Fish Physiology and Biochemistry, 42(2):795–805, April 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0176-0.

Metochis:2016:EID

[895] C. Metochis, V. O. Crampton, and K. D. Thompson. The effects of increasing dietary levels of amino acid-supplemented soy protein concentrate and constant dietary supplementation of phosphorus on growth, composition and immune responses of juvenile Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 42(3):807–829, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-015-0177-z.

Gonzalez:2016:MSA

[896] R. Gonzalez and S. Unniappan. Mass spectrometry-assisted confirmation of the inability of dipeptidyl peptidase-4 to cleave goldfish peptide YY(1–36) and the lack of anorexigenic effects of peptide YY(3–36) in goldfish (Carassius auratus). Fish Physiology and Biochemistry, 42 (3):831–844, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0178-y.

Wang:2016:ECA

[897] Qiuju Wang, Xue Ju, and Dongming Zhang. Effects of l-carnitine against H_{2}O_{2}-induced oxidative stress in grass carp ovary cells (Ctenopharyngodon idellus). Fish Physiology and Biochemistry, 42(3):845–857, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-015-0179-x.

Rodriguez-Estrada:2016:ESA

[898] Jesús Rodríguez-Estrada, Alma Socorro Sobrino-Figueroa, and Fernando Martínez-Jerónimo. Effect of sublethal α-cypermethrin exposure on main macromolecules concentration, energy content, and malondialdehyde concentration in free-feeding Danio rerio larvae. Fish Physiology and Biochemistry, 42(3):859–868, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0180-4.

Saavedra:2016:AAE

[899] L. M. Saavedra, R. A. Quiñones, and E. J. Niklitschek. Aerobic and anaerobic enzymatic activity of orange roughy (Hoplostethus atlanticus) and alfonsino (Beryx splendens) from the juan fernandez seamounts area. Fish Physiology and Biochemistry, 42(3):869–882, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0181-3.

Wang:2016:MCM

[900] Tao Wang, Dengyue Yuan, and Zhiqiong Li. Molecular characterization of melanin-concentrating hormone (MCH) in Schizothorax prenanti: cloning, tissue distribution and role in food intake regulation. Fish Physiology and Biochemistry, 42(3):883–893, June 2016. CODEN FPBIEP.

ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0182-2.

Vikingstad:2016:ETF

[901] Erik Vikingstad, Eva Andersson, and Geir Lasse Taranger. Effects of temperature on the final stages of sexual maturation in Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 42(3): 895–907, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0183-1.

Penghan:2016:ESF

[902] Liu-Yi Penghan, Xu Pang, and Shi-Jian Fu. The effects of starvation on fast-start escape and constant acceleration swimming performance in rose bitterling (Rhodeus ocellatus) at two acclimation temperatures. Fish Physiology and Biochemistry, 42(3):909–918, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-015-0184-0.

Ekambaram:2016:DRP

[903] Padmini Ekambaram, Parimala Parasuraman, and Tharani Jayachandran. Differential regulation of pro- and antiapoptotic proteins in fish adipocytes during hypoxic conditions. Fish Physiology and Biochemistry, 42(3):919–934, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0185-z.

Jia:2016:IPA

[904] Yihe Jia, Shaowu Yin, and Xinhua Su. iTRAQ proteomic analysis of salinity acclimation proteins in the gill of tropical marbled eel (Anguilla marmorata). Fish Physiology and Biochemistry, 42(3):935–946, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-015-0186-y.

Bai:2016:DMP

[905] Jin Bai, Weida Gong, and Shi Xi Chen. Dynamic methylation pattern of cyp19a1a core promoter during zebrafish ovarian folliculogenesis. Fish Physiology and Biochemistry, 42(3):947–954, June 2016. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0187-x.

Elabd:2016:EAM

[906] Hiam Elabd, Han-Ping Wang, and Amany Abbass. Astragalus membranaceus (AM) enhances growth performance and antioxidant stress profiles in bluegill sunfish (Lepomis macrochirus). Fish Physiology and Biochemistry, 42(3):955–966, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-015-0188-9.

Akhavan:2016:CVL

[907] Sobhan Ranay Akhavan, Amir Parviz Salati, and Seyed Amir Hossein Jalali. Changes of vitellogenin and lipase in captive sterlet sturgeon Acipenser ruthenus females during previtellogenesis to early atresia. Fish Physiology and Biochemistry, 42(3):967–978, June 2016. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-015-0189-8.

Chen:2016:NCL

[908] Qi-Liang Chen, Zhi Luo, and Kun Wu. De novo characterization of the liver transcriptome of javelin goby Synechogobius hasta and analysis of its transcriptomic profile following waterborne copper exposure. Fish Physiology and Biochemistry, 42(3):979–994, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0190-2.

Lacy:2016:ILR

[909] E. R. Lacy, E. Reale, and L. Luciano. Immunohistochemical localization of renin-containing cells in two elasmobranch species. Fish Physiology and Biochemistry, 42(3):995–1004, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0191-1.

Leitemperger:2016:EBB

[910] Jossiele Leitemperger, Charlene Menezes, and Vania Lucia Loro. Early biochemical biomarkers for zinc in silver catfish (Rhamdia quelen) after acute exposure. Fish Physiology and Biochemistry, 42(3):1005– 1014, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-015-0192-0.

Jang:2016:DLP

[911] Jun-Chul Jang, Mi-Jin Choi, and Jong-Myoung Kim. Dim-light photoreceptor of chub mackerel Scomber japonicus and the photoresponse upon illumination with LEDs of different wavelengths. Fish Physiology and Biochemistry, 42(3):1015–1025, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0193-z.

Ayala:2016:EET

[912] María D. Ayala, Juan M. Martínez, and Rosa Cal. Effect of the early temperature on the growth of larvae and postlarvae turbot, Scophthalmus maximus l.: muscle structural and ultrastructural study. Fish Physiology and Biochemistry, 42(3):1027–1042, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-015-0194-y.

Chen:2016:ILG

[913] Wenbo Chen, Zhen Zhang, and Fangfang Yan. Insulin-like growth factorbinding protein-1 (IGFBP-1) in goldfish, Carassius auratus: molecular cloning, tissue expression, and mRNA expression responses to periprandial changes and cadmium exposure. Fish Physiology and Biochemistry, 42(3):1043–1052, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-015-0195-x.

Cheng:2016:EEM

[914] Nana Cheng, Maomao Guo, and Haobin Zhao. Expression of mep50 in adult and embryos of medaka fish (Oryzias latipes). Fish Physiology and Biochemistry, 42(3):1053–1061, June 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0196-4.

Najafabad:2016:EDC

[915] Masume Kamali Najafabad, Mohammad Reza Imanpoor, and Alireza Alishahi. Effect of dietary chitosan on growth performance, hematological parameters, intestinal histology and stress resistance of Caspian kutum (Rutilus frisii kutum Kamenskii, 1901) fingerlings. Fish Physiology and Biochemistry, 42(4):1063–1071, August 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0197-3.

Song:2016:ICE

[916] Huayu Song, Mengxun Wang, and Quanqi Zhang. Identification and characterization of kiss2 and kissr2 homologs in Paralichthys olivaceus. Fish Physiology and Biochemistry, 42(4):1073–1092, August 2016. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0199-1.

Hu:2016:EWZ

[917] Wei Hu, Kang-Sen Mai, and Ya-Xiong Pan. Effect of waterborne zinc exposure on lipid deposition and metabolism in hepatopancreas and muscle of grass carp Ctenopharyngodon idella. Fish Physiology and Biochemistry, 42(4):1093–1105, August 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0200-z.

Jaya-Ram:2016:EOI

[918] Annette Jaya-Ram, Alexander Chong Shu-Chien, and Meng-Kiat Kuah. Echium oil increased the expression of a Δ 4 Fads2 fatty acyl desaturase and the deposition of n−3 long-chain polyunsaturated fatty acid in comparison with linseed oil in striped snakehead (Channa striata) muscle. Fish Physiology and Biochemistry, 42(4):1107–1122, August 2016. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0201-y.

Wang:2016:THG

[919] Peng-Fei Wang, Shuang Zeng, and Gui-Feng Li. Two HSP90 genes in mandarin fish Siniperca chuatsi: identification, characterization and their specific expression profiles during embryogenesis and under stresses. Fish Physiology and Biochemistry, 42(4):1123–1136, August 2016. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0202-x.

Garina:2016:DSM

[920] D. V. Garina, V. A. Nepomnyashchikh, and A. A. Mekhtiev. Does serotonin-modulating anticonsolidation protein (SMAP) influence the choice of turning direction in carps, Cyprinus carpio, in a T-maze? Fish Physiology and Biochemistry, 42(4):1137–1141, August 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0203-9.

Taufek:2016:EDC

[921] Norhidayah Mohd Taufek, Firdaus Aspani, and Zazali Alias. The effect of dietary cricket meal (Gryllus bimaculatus) on growth performance, antioxidant enzyme activities, and haematological response of African catfish (Clarias gariepinus). Fish Physiology and Biochemistry, 42(4): 1143–1155, August 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0204-8.

Golshan:2016:TGE

[922] Mahdi Golshan, Hamid R. Habibi, and Sayyed Mohammad Hadi Alavi. Transcripts of genes encoding reproductive neuroendocrine hormones and androgen receptor in the brain and testis of goldfish exposed to vinclozolin, flutamide, testosterone, and their combinations. Fish Physiology and Biochemistry, 42(4):1157–1165, August 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0205-7.

Weng:2016:MCS

[923] Shenda Weng, Feng You, and Yuxia Zou. Molecular cloning and sexually dimorphic expression of wnt4 in olive flounder (Paralichthys olivaceus). Fish Physiology and Biochemistry, 42(4):1167–1176, August 2016. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0206-6.

Mohammed-Geba:2016:MEC

[924] Khaled Mohammed-Geba, Manuel Yúfera, and Juan Miguel Mancera. Molecular endocrine changes of Gh/Igf1 axis in gilthead sea bream (Sparus aurata L.) exposed to different environmental salinities during larvae to post-larvae stages. Fish Physiology and Biochemistry, 42(4): 1177–1186, August 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0207-5.

Borey:2016:PKG

[925] Marion Borey, Stephane Panserat, and Christine Burel. Postprandial kinetics of gene expression of proteins involved in the digestive process in rainbow trout (O. mykiss) and impact of diet composition. Fish Physiology and Biochemistry, 42(4):1187–1202, August 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0208-4.

Pang:2016:ESR

[926] Xu Pang, Shi-Jian Fu, and Yao-Guang Zhang. The effects of starvation and re-feeding on growth and swimming performance of juvenile black carp (Mylopharyngodon piceus). Fish Physiology and Biochemistry, 42 (4):1203–1212, August 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0210-x.

Armelin:2016:GDE

[927] Vinicius Araújo Armelin, Victor Hugo da Silva Braga, and Ana Lúcia Kalinin. Gill denervation eliminates the barostatic reflex in a neotropical teleost, the tambaqui (Colossoma macropomum). Fish Physiology and Biochemistry, 42(4):1213–1224, August 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0211-9.

Abdel-Khalek:2016:TEC

[928] Amr A. Abdel-Khalek, Shereen R. Badran, and Mohamed-Assem S. Marie. Toxicity evaluation of copper oxide bulk and nanoparticles in Nile tilapia, Oreochromis niloticus, using hematological, bioaccumulation and histological biomarkers. Fish Physiology and Biochemistry, 42 (4):1225–1236, August 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0212-8.

Furne:2016:EDM

[929] M. Furné, M. García-Gallego, and A. Sanz. Effect of dietary macronutrient proportion on intermediate metabolism and oxidative status in sturgeon (Acipenser naccarii) and trout (Oncorhynchus mykiss): comparative study. Fish Physiology and Biochemistry, 42(4):1237–1248, August 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-016-0213-7.

Basak:2016:SRM

[930] Reetuparna Basak, Alivia Roy, and Umesh Rai. Seasonality of reproduction in male spotted murrel Channa punctatus: correlation of environmental variables and plasma sex steroids with histological changes in testis. Fish Physiology and Biochemistry, 42(5):1249–1258, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0214-6.

Santos:2016:DEA

[931] Juliana Ferreira Santos, Karollina Lopes Siqueira Soares, and Ranilson Souza Bezerra. Digestive enzyme activity in the intestine of Nile tilapia (Oreochromis niloticus L.) under pond and cage farming systems. Fish Physiology and Biochemistry, 42(5):1259–1274, October 2016. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0215-5.

Liu:2016:FCJ

[932] Wei Liu, Huayu Song, and Jie Qi. Functional characterization of the Japanese flounder (Paralichthys olivaceus) Sox2 gene promoter. Fish Physiology and Biochemistry, 42(5):1275–1285, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0216-4.

Kumar:2016:MHH

[933] Ravi Kumar, K. P. Joy, and S. M. Singh. Morpho-histology of head kidney of female catfish Heteropneustes fossilis: seasonal variations in melano-macrophage centers, melanin contents and effects of lipopolysaccharide and dexamethasone on melanins. Fish Physiology and Biochemistry, 42(5):1287–1306, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0218-2.

Baptista:2016:AHR

[934] R. B. Baptista, N. Souza-Castro, and V. M. F. Almeida-Val. Acute hypoxia up-regulates HIF-1α and VEGF mRNA levels in Amazon hypoxiatolerant Oscar (Astronotus ocellatus). Fish Physiology and Biochemistry, 42(5):1307–1318, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0219-1.

Moguel-Hernandez:2016:OCW

[935] I. Moguel-Hernández, R. Peña, and E. Gisbert. Ontogeny changes and weaning effects in gene expression patterns of digestive enzymes and regulatory digestive factors in spotted rose snapper (Lutjanus guttatus) larvae. Fish Physiology and Biochemistry, 42(5):1319–1334, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0220-8.

Jung:2016:EMG

[936] Seo Jin Jung, Na Na Kim, and Cheol Young Choi. Effects of melatonin and green-wavelength LED light on the physiological stress and immunity of goldfish, Carassius auratus, exposed to high water temperature. Fish Physiology and Biochemistry, 42(5):1335–1346, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0221-7.

Kida:2016:EAW

[937] Bianca Mayumi Silva Kida, Raisa Pereira Abdalla, and Renata Guimarães Moreira. Effects of acidic water, aluminum, and manganese on testicular steroidogenesis in Astyanax altiparanae. Fish Physiology and Biochemistry, 42(5):1347–1356, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0222-6.

Menezes:2016:EDD

[938] Charlene Menezes, Jossiele Leitemperger, and Vania Lucia Loro. Effect of diphenyl diselenide diet supplementation on oxidative stress biomarkers in two species of freshwater fish exposed to the insecticide fipronil. Fish Physiology and Biochemistry, 42(5):1357–1368, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0223-5.

Wang:2016:CCE

[939] Hao Wang, Libo He, and Yaping Wang. Cloning and characterization of Bax1 and Bax2 genes of Ctenopharyngodon idellus and evaluation of transcript expression in response to grass carp reovirus infection. Fish Physiology and Biochemistry, 42(5):1369–1382, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0225-3.

Solovyev:2016:ITS

[940] Mikhail Solovyev and Enric Gisbert. Influence of time, storage temperature and freeze/thaw cycles on the activity of digestive enzymes from gilthead sea bream (Sparus aurata). Fish Physiology and Biochemistry, 42 (5):1383–1394, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0226-2.

Zeng:2016:EBG

[941] Lin Zeng, Yong-Hong Wang, and Rong Cai. Effects of β-glucan on ROS production and energy metabolism in yellow croaker (Pseudosciaena crocea) under acute hypoxic stress. Fish Physiology and Biochemistry, 42(5): 1395–1405, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0227-1.

Nurilmala:2016:MCS

[942] Mala Nurilmala and Yoshihiro Ochiai. Molecular characterization of southern bluefin tuna myoglobin (Thunnus maccoyii). Fish Physiology and Biochemistry, 42(5):1407–1416, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0228-0.

Yousefi:2016:SBN

[943] Morteza Yousefi, Mehdi Paktinat, and Seyyed Morteza Hoseini. Serum biochemical and non-specific immune responses of rainbow trout (Oncorhynchus mykiss) to dietary nucleotide and chronic stress. Fish Physiology and Biochemistry, 42(5):1417–1425, October 2016. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-016-0229-z.

Doan:2016:EEM

[944] Hien Van Doan, Sompong Doolgindachbaporn, and Amnuaysilpa Suksri. Effects of eryngii mushroom (Pleurotus eryngii) and Lactobacillus plantarum on growth performance, immunity and disease resistance of Pangasius catfish (Pangasius bocourti, Sauvage 1880). Fish Physiology and Biochemistry, 42(5):1427–1440, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0230-6.

Castanheira:2016:CSC

[945] Maria Filipa Castanheira, Sonia Martínez Páramo, and Luís E. C. Conceição. Are coping styles consistent in the teleost fish Sparus aurata through sexual maturation and sex reversal? Fish Physiology and Biochemistry, 42(5):1441–1452, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0231-5.

Wang:2016:CCA

[946] Jian-Wei Wang, Zhen-Dong Cao, and Shi-Jian Fu. A comparison of constant acceleration swimming speeds when acceleration rates are different with critical swimming speeds in Chinese bream under two oxygen tensions. Fish Physiology and Biochemistry, 42(5):1453–1461, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0232-4.

Sanden:2016:MLM

[947] Monica Sanden, Nina S. Liland, and Nini H. Sissener. Minor lipid metabolic perturbations in the liver of Atlantic salmon (Salmo salar L.) caused by suboptimal dietary content of nutrients from fish oil. Fish Physiology and Biochemistry, 42(5):1463–1480, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0233-3.

Hu:2016:SFC

[948] Wenxian Hu, Jie Zhang, and Bin Kang. Structure and function of corneal surface of mudskipper fishes. Fish Physiology and Biochemistry, 42(5): 1481–1489, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0234-2.

Akhavan:2016:ECV

[949] Sobhan Ranay Akhavan, Amir Parviz Salati, and Seyed Amir Hossein Jalali. Erratum to: Changes of vitellogenin and lipase in captive sterlet sturgeon Acipenser ruthenus females during previtellogenesis to early atresia. Fish Physiology and Biochemistry, 42(5):1491, October 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0217-3.

Canada:2016:DIA

[950] Paula Canada, Sofia Engrola, and Luís E. C. Conceição. Dietary indispensable amino acids profile affects protein utilization and growth of Senegalese sole larvae. Fish Physiology and Biochemistry, 42(6): 1493–1508, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0235-1.

Babaei:2016:EDC

[951] Sedigheh Babaei, Abdolmohammad Abedian Kenari, and Isidoro Metón. Effect of diet composition on growth performance, hepatic metabolism and antioxidant activities in Siberian sturgeon (Acipenser baerii, Brandt, 1869) submitted to starvation and refeeding. Fish Physiology and Biochemistry, 42(6):1509–1520, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0236-0.

Henderson:2016:ASN

[952] C. J. Henderson, T. F. Stevens, and S. Y. Lee. Assessing the suitability of a non-lethal biopsy punch for sampling fish muscle tissue. Fish Physiology and Biochemistry, 42(6):1521–1526, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0237-z.

Fu:2016:CPA

[953] Xiaoqin Fu, Zhujin Ding, and Hong Liu. Characterization, promoter analysis and expression of the interleukin-6 gene in blunt snout bream, Megalobrama amblycephala. Fish Physiology and Biochemistry, 42(6): 1527–1540, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0238-y.

Vutukuru:2016:EDB

[954] S. S. Vutukuru, Jayasree Ganugapati, and Ravindra Babu Potti. Endocrine disruption by bisphenol a, polychlorinated biphenyls and polybrominated diphenyl ether, in zebra fish (Danio rerio) model: an in silico approach. Fish Physiology and Biochemistry, 42(6):1541–1555, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-016-0239-x.

Yuan:2016:FDP

[955] Xiaochen Yuan, Xu-Fang Liang, and Qingchao Wang. Fat deposition pattern and mechanism in response to dietary lipid levels in grass carp, Ctenopharyngodon idellus. Fish Physiology and Biochemistry, 42(6): 1557–1569, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0240-4.

Hossain:2016:SSR

[956] Md. Afzal Hossain, Shefali Aktar, and Jian G. Qin. Salinity stress response in estuarine fishes from the Murray Estuary and Coorong, South Australia. Fish Physiology and Biochemistry, 42(6):1571–1580, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-016-0241-3.

Zhang:2016:TAG

[957] Wei Zhang, Yuezhong Liu, and Yan He. Transcriptome analysis of the gonads of olive flounder (Paralichthys olivaceus). Fish Physiology and Biochemistry, 42(6):1581–1594, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0242-2.

Zeng:2016:MCE

[958] Lin Zeng, Bin Liu, and Wan-Shu Hong. Molecular characterization and expression analysis of AMPK α subunit isoform genes from Scophthalmus maximus responding to salinity stress. Fish Physiology and Biochemistry, 42(6):1595–1607, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0243-1.

Li:2016:SER

[959] Yixue Li, Zhi He, and Lihong Zhang. Scp3 expression in relation to the ovarian differentiation in the protogynous hermaphroditic ricefield eel Monopterus albus. Fish Physiology and Biochemistry, 42(6):1609– 1619, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0244-0.

Boonanuntanasarn:2016:CEV

[960] Surintorn Boonanuntanasarn, Paiboon Bunlipatanon, and Goro Yoshizaki. Characterization of a vasa homolog in the brown-marbled grouper (Epinephelus fuscoguttatus) and its expression in gonad and germ cells during larval development. Fish Physiology and Biochemistry, 42(6): 1621–1636, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0245-z.

Zhou:2016:EEG

[961] Chaowei Zhou, Jinfa Zheng, and Benhe Zeng. Evidence that ghrelin may be associated with the food intake of gibel carp (Carassius auratus gibelio). Fish Physiology and Biochemistry, 42(6):1637– 1646, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0246-y.

Blondeau-Bidet:2016:MCE

[962] Eva Blondeau-Bidet, Maryline Bossus, and Catherine Lorin-Nebel. Molecular characterization and expression of Na^{+} /K^{+}-ATPase α 1 isoforms in the European sea bass Dicentrarchus labrax osmoregulatory tissues following salinity transfer. Fish Physiology and Biochemistry, 42(6): 1647–1664, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0247-x.

Han:2016:LRT

[963] Dongdong Han, Huijun Miao, and Kangsen Mai. Leptin and its receptor in turbot Scophthalmus maximus: cloning, characterization and expression response to ratios of dietary carbohydrate–lipid. Fish Physiology and Biochemistry, 42(6):1665–1679, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0248-9.

Sharma:2016:HCD

[964] Prakash Sharma, M. S. Akhtar, and Debajit Sarma. Histomorphological changes in digestive tract of golden mahseer (Tor putitora) during different developmental stages. Fish Physiology and Biochemistry, 42(6): 1681–1698, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0249-8.

Santos:2016:ELA

[965] R. A. Santos, S. Caldas, and J. M. Monserrat. Effects of lipoic acid on growth and biochemical responses of common carp fed with carbohydrate diets. Fish Physiology and Biochemistry, 42(6):1699–1707, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-016-0250-2.

Wu:2016:AOF

[966] Su Mei Wu, Li-Hsin Shu, and Jia-Hao Liu. Anti-oxidative functions of mt2 and smtb mRNA expression in the gills and brain of zebrafish (Danio rerio) upon cadmium exposure. Fish Physiology and Biochemistry, 42(6): 1709–1720, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0251-1.

Costa:2016:CSS

[967] Daniele Dietrich Moura Costa, Dandie Antunes Bozza, and Ciro Alberto de Oliveira Ribeiro. Characterization, specificity and sensibility of produced anti- Rhamdia quelen vitellogenin in Brazilian fish species. Fish Physiology and Biochemistry, 42(6):1721–1732, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0252-0.

Safari:2016:MAD

[968] Roghieh Safari, Seyed Hossein Hoseinifar, and Morteza Kavandi. Modulation of antioxidant defense and immune response in zebra fish (Danio rerio) using dietary sodium propionate. Fish Physiology and Biochemistry, 42(6):1733–1739, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0253-z.

Shirangi:2016:SEO

[969] Seyedeh Ainaz Shirangi, Mohammad Reza Kalbassi, and Jehan-Hervé Lignot. Salinity effects on osmoregulation and gill morphology in juvenile Persian sturgeon (Acipenser persicus). Fish Physiology and Biochemistry, 42(6):1741–1754, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0254-y.

Dzyuba:2016:CPA

[970] Viktoriya Dzyuba, Mariola Słowińska, and Borys Dzyuba. Characterization of proteolytic and anti-proteolytic activity involvement in sterlet spermatozoon maturation. Fish Physiology and Biochemistry, 42(6): 1755–1766, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0255-x.

Valverde-Chavarria:2016:VDP

[971] Silvia Valverde-Chavarría, Carlos A. Álvarez-González, and Juan B. Ulloa-Rojas. In vitro digestibility and proteases inhibitory effect of several feedstuffs for Parachromis dovii juveniles and P. dovii hybrid larvae. Fish Physiology and Biochemistry, 42(6):1767–1776, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0256-9.

Mesa-Rodriguez:2016:BDS

[972] A. Mesa-Rodríguez, C. M. Hernández-Cruz, and J. Roo. Bone development of the skull, pectoral and pelvic fins in Seriola rivoliana (Valenciennes, 1833) larvae. Fish Physiology and Biochemistry, 42(6):1777– 1789, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0257-8.

Dey:2016:SEP

[973] Sangeeta Dey, Manabendra Dutta Choudhury, and Suchismita Das. Sublethal effects of pulp and paper mill effluent on two commonly cultured carps: a SEM- and EDS-based hematological biomarker analysis. Fish Physiology and Biochemistry, 42(6):1791–1805, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0258-7.

Zhu:2016:EST

[974] Zhiming Zhu, Bolan Song, and Zhongneng Xu. Effect of sustained training on glycolysis and fatty acids oxidation in swimming muscles and liver in juvenile tinfoil barb Barbonymus schwanenfeldii (Bleeker, 1854). Fish Physiology and Biochemistry, 42(6):1807–1817, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0259-6.

Tkachenko:2016:EOS

[975] Halyna Tkachenko and Joanna Grudniewska. Evaluation of oxidative stress markers in the heart and liver of rainbow trout (Oncorhynchus mykiss Walbaum) exposed to the formalin. Fish Physiology and Biochemistry, 42(6):1819–1832, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0260-0.

Safarian:2016:MCL

[976] Mina Safarian, Mohammad Reza Tabandeh, and Ebrahim Rajabzadeh Ghotrami. Molecular characteristics of lysozyme G in Euryglossa orientalis; cDNA cloning, phylogenic analysis, physicochemical properties and tissue gene expression. Fish Physiology and Biochemistry, 42(6): 1833–1844, December 2016. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0261-z.

Ren:2017:TAZ

[977] Tao Ren, Gui-Hong Fu, and Xian-Le Yang. Toxicity and accumulation of zinc pyrithione in the liver and kidneys of Carassius auratus gibelio: association with P-glycoprotein expression. Fish Physiology and Biochemistry, 43(1):1–9, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0262-y.

Li:2017:SEN

[978] Meijie Li, Xungang Tan, and Feng You. The stimulatory effect of neuropeptide Y on growth hormone expression, food intake, and growth in olive flounder (Paralichthys olivaceus). Fish Physiology and Biochemistry, 43(1):11–18, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0263-x.

Vuong:2017:EPC

[979] Tram T. Vuong, Sissel B. Rønning, and Mona E. Pedersen. The enzyme profiles in the connective tissue attaching pin bones to the surrounding tissue is specific in farmed salmon (Salmo salar) and cod (Gadus morhua L.). Fish Physiology and Biochemistry, 43(1):19–25, February 2017. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0264-9.

Kowalska:2017:DRI

[980] Agata Kowalska, Andrzej K. Siwicki, and Radosław K. Kowalski. Dietary resveratrol improves immunity but reduces reproduction of broodstock medaka Oryzias latipes (Temminck & Schlegel). Fish Physiology and Biochemistry, 43(1):27–37, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0265-8.

Schmitz:2017:OIS

[981] Mélodie Schmitz, Sébastien Baekelandt, and Patrick Kestemont. Osmoregulatory and immunological status of the pond-raised striped catfish (Pangasianodon hypophthalmus s.) as affected by seasonal runoff and salinity changes in the Mekong Delta, Vietnam. Fish Physiology and Biochemistry, 43(1):39–49, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0266-7.

Nobrega:2017:CGC

[982] Rafael Henrique Nóbrega, Lázaro Wender Oliveira de Jesus, and Maria Inês Borella. Characterization of gonadotropic cells during continuous and seasonal spermatogenesis of two freshwater fish species: a histochemical and immunohistochemical study. Fish Physiology and Biochemistry, 43(1):51–63, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0267-6.

Lu:2017:BAO

[983] Kang-Le Lu, Li-Na Wang, and Wei-Na Xu. Berberine attenuates oxidative stress and hepatocytes apoptosis via protecting mitochondria in blunt snout bream Megalobrama amblycephala fed high-fat diets. Fish Physiology and Biochemistry, 43(1):65–76, February 2017. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-016-0268-5.

Hangzo:2017:ASU

[984] Hnunlalliani Hangzo, Bodhisattwa Banerjee, and Nirmalendu Saha. Ammonia stress under high environmental ammonia induces Hsp70 and Hsp90 in the mud eel, Monopterus cuchia. Fish Physiology and Biochemistry, 43(1):77–88, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0269-4.

Mahanty:2017:EHE

[985] Arabinda Mahanty, Gopal Krishna Purohit, and Bimal Prasanna Mohanty. hsp90 and hsp47 appear to play an important role in minnow Puntius sophore for surviving in the hot spring run-off aquatic ecosystem. Fish Physiology and Biochemistry, 43(1):89–102, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0270-y.

Liu:2017:MCG

[986] Hongyu Liu, Xiaohui Dong, and Beiping Tan. Molecular cloning of glucose transporter 1 in grouper Epinephelus coioides and effects of an acute hyperglycemia stress on its expression and glucose tolerance. Fish Physiology and Biochemistry, 43(1):103–114, February 2017. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-016-0271-x.

Dong:2017:ICS

[987] Haiyan Dong, Wenbo Chen, and Jinyun Ye. Identification, characterization of selenoprotein W and its mRNA expression patterns in response to somatostatin 14, cysteamine hydrochloride, 17β-estradiol and a binary mixture of 17β-estradiol and cysteamine hydrochloride in topmouth culter (Erythroculter ilishaeformis). Fish Physiology and Biochemistry, 43 (1):115–126, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0272-9.

Liu:2017:NAD

[988] Yu Liu, Yanfang Zhang, and Xingguo Liang. Nucleic acids digestion by enzymes in the stomach of snakehead (Channa argus) and banded grouper (Epinephelus awoara). Fish Physiology and Biochemistry, 43 (1):127–136, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0273-8.

Hano:2017:ETS

[989] Takeshi Hano, Katsutoshi Ito, and Kazuhiko Mochida. Effect of taurine supplementation on hepatic metabolism and alleviation of cadmium toxicity and bioaccumulation in a marine teleost, red sea bream, Pagrus major. Fish Physiology and Biochemistry, 43(1):137–152, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0274-7.

Zheng:2017:DEL

[990] Jia-Lang Zheng, Lin Zeng, and Chang-Wen Wu. Different effects of low- and high-dose waterborne zinc on Zn accumulation, ROS levels, oxidative damage and antioxidant responses in the liver of large yellow croaker Pseudosciaena crocea. Fish Physiology and Biochemistry, 43(1): 153–163, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0275-6.

Wosnick:2017:TDU

[991] Natascha Wosnick, Hugo Bornatowski, and Carolina Arruda Freire. Talking to the dead: using Post-mortem data in the assessment of stress in tiger sharks (Galeocerdo cuvier) (Péron and Lesueur, 1822). Fish Physiology and Biochemistry, 43(1):165–178, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0276-5.

Dawood:2017:PRB

[992] Mahmoud A. O. Dawood, Shunsuke Koshio, and Adissin Olivier. Physiological response, blood chemistry profile and mucus secretion of red sea bream (Pagrus major) fed diets supplemented with Lactobacillus rhamnosus under low salinity stress. Fish Physiology and Biochemistry, 43 (1):179–192, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0277-4.

Barreto:2017:MAA

[993] Rodrigo Egydio Barreto. Mianserin affects alarm reaction to conspecific chemical alarm cues in Nile tilapia. Fish Physiology and Biochemistry, 43 (1):193–201, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0279-2.

Mohammed-Geba:2017:MPP

[994] Khaled Mohammed-Geba, Antonio Astola González, and Juan Miguel Mancera. Molecular performance of Prl and Gh/Igf1 axis in the Mediterranean meager, Argyrosomus regius, acclimated to different rearing salinities. Fish Physiology and Biochemistry, 43(1):203–216, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0280-9.

Wattanakul:2017:FCE

[995] Wattana Wattanakul, Uraiwan Wattanakul, and Chutchawan Muenpo. Fish condensate as effective replacer of fish meal protein in diet for striped snakehead, Channa striata (Bloch). Fish Physiology and Biochemistry, 43(1):217–228, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0281-8.

Chen:2017:CIT

[996] Pei Chen, Yanqing Huang, and Chunfang Wang. Characterization of the isoforms of type IIb sodium-dependent phosphate cotransporter (slc34a2) in yellow catfish, Pelteobagrus fulvidraco, and their vitamin D_{3}-regulated expression under low-phosphate conditions. Fish Physiology and Biochemistry, 43(1):229–244, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0282-7.

Xiao:2017:DSS

[997] Peizhen Xiao, Hong Ji, and Zhenyu Du. Dietary silymarin supplementation promotes growth performance and improves lipid metabolism and health status in grass carp (Ctenopharyngodon idellus) fed diets with elevated lipid levels. Fish Physiology and Biochemistry, 43(1): 245–263, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0283-6.

Zhao:2017:GAS

[998] Chunyan Zhao, Shihong Xu, and Jun Li. Gonadogenesis analysis and sex differentiation in cultured turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 43(1):265–278, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0284-5.

Cai:2017:EFD

[999] Lu Cai, David Johnson, and Yingping Huang. Effects of feeding, digestion and fasting on the respiration and swimming capability of juvenile sterlet sturgeon (Acipenser ruthenus, Linnaeus 1758). Fish Physiology and Biochemistry, 43(1):279–286, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0285-4.

Lv:2017:GEM

[1000] Feng Lv, Chenwen Zhu, and Dong Liu. Generation of a mef2aa:egfp transgenic zebrafish line that expresses EGFP in muscle cells. Fish Physiology and Biochemistry, 43(1):287–294, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0286-3.

Li:2017:EEB

[1001] Ge Li, Hang Xie, and Yiping Luo. Erratum to: Effects of body chemical components on the allometric scaling of the resting metabolic rate in four species of cyprinids. Fish Physiology and Biochemistry, 43 (1):295, February 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0278-3.

Li:2017:MAL

[1002] Jiao Li, Liwei Liu, and Jia Wang. Modulation of appetite, lipid and glucose metabolism of juvenile grass carp (Ctenopharyngodon idellus) by different dietary protein levels. Fish Physiology and Biochemistry, 43(2):297–307, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0287-2.

Dong:2017:AAS

[1003] Yuan Dong, Pengfei Cui, and Shicui Zhang. Aging asymmetry: systematic survey of changes in age-related biomarkers in the annual fish Nothobranchius guentheri. Fish Physiology and Biochemistry, 43(2): 309–319, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0288-1.

Yu:2017:CER

[1004] Jie Yu, Yuanshuai Fu, and Zhiyi Shi. Coordinated expression and regulation of deiodinases and thyroid hormone receptors during metamorphosis in the Japanese flounder (Paralichthys olivaceus). Fish Physiology and Biochemistry, 43(2):321–336, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0289-0.

Preston:2017:ETF

[1005] Andrew C. Preston, John F. Taylor, and Hervé Migaud. Effects of temperature on feed intake and plasma chemistry after exhaustive exercise in triploid brown trout (Salmo trutta l). Fish Physiology and Biochemistry, 43(2):337–350, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0290-7.

Niedzwiecka:2017:NPM

[1006] Natalia Niedźwiecka, Jadwiga Gronczewska, and Edward F. Skorkowski. NAD-preferring malic enzyme: localization, regulation and its potential role in herring (Clupea harengus) sperm cells. Fish Physiology and Biochemistry, 43(2):351–360, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0291-6.

Zheng:2017:EDS

[1007] Qingmei Zheng, Chunyan Han, and Ming Zhong. Effects of dietary supplementation with green tea waste on growth, digestive enzyme and lipid metabolism of juvenile hybrid tilapia, Oreochromis niloticus × O. aureus. Fish Physiology and Biochemistry, 43(2):361–371, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0292-5.

Kuzmina:2017:PAS

[1008] V. V. Kuz’mina, G. V. Zolotareva, and V. A. Sheptitskiy. Proteolytic activity in some freshwater animals and associated microflora in a wide pH range. Fish Physiology and Biochemistry, 43(2):373–383, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0293-4.

Jiao:2017:CEP

[1009] Shuang Jiao, Zhihao Wu, and Feng You. Characterization of pax3a and pax3b genes in artificially induced polyploid and gynogenetic olive flounder (Paralichthys olivaceus) during embryogenesis. Fish Physiology and Biochemistry, 43(2):385–395, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0294-3.

Gao:2017:DCD

[1010] Xiao-Qiang Gao, Zhi-Feng Liu, and Lei Hong. Developmental changes in digestive enzyme activity in American shad, Alosa sapidissima, during early ontogeny. Fish Physiology and Biochemistry, 43(2):397–409, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-016-0295-2.

Yuan:2017:DEB

[1011] Shuang-Shuang Yuan, Huan-Zhi Xu, and Jia-Lang Zheng. Different effects of blue and red light-emitting diodes on antioxidant responses in the liver and ovary of zebrafish Danio rerio. Fish Physiology and Biochemistry, 43(2):411–419, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0296-1.

Huyben:2017:EDY

[1012] David Huyben, Aleksandar Vidakovic, and Anders Kiessling. Effects of dietary yeast inclusion and acute stress on post-prandial whole blood profiles of dorsal aorta-cannulated rainbow trout. Fish Physiology and Biochemistry, 43(2):421–434, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0297-0.

Ma:2017:CEP

[1013] Qian Ma, Wenrong Feng, and Shufang Liu. Cloning, expression profiling and promoter functional analysis of bone morphogenetic protein 6 and 7 in tongue sole (Cynoglossus semilaevis). Fish Physiology and Biochemistry, 43(2):435–454, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0298-z.

Albendin:2017:CVS

[1014] G. Albendín, J. M. Arellano, and M. I. Arufe. Characterization and in vitro sensitivity of cholinesterases of gilthead seabream (Sparus aurata) to organophosphate pesticides. Fish Physiology and Biochemistry, 43 (2):455–464, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0299-y.

He:2017:PCH

[1015] Yan He, Jie Fang, and Jie Mei. Potential contributions of heat shock proteins and related genes in sexual differentiation in yellow catfish (Pelteobagrus fulvidraco). Fish Physiology and Biochemistry, 43(2): 465–475, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0303-6.

Mahapatra:2017:PDE

[1016] S. Mahapatra, Sk. Kabita, and P. Nath. Purification and development of ELISAs for two forms of vitellogenin in Indian walking catfish, Clarias batrachus (L.). Fish Physiology and Biochemistry, 43 (2):477–491, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0304-5.

Betancor:2017:LMR

[1017] Mónica B. Betancor, Aurelio Ortega, and Gabriel Mourente. Lipid metabolism-related gene expression pattern of Atlantic bluefin tuna (Thunnus thynnus L.) larvae fed on live prey. Fish Physiology and Biochemistry, 43(2):493–516, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0305-4.

Liu:2017:ZPD

[1018] Yuan-Yuan Liu, Xiu-Rong Su, and Shicui Zhang. Zebrafish phosvitinderived peptide pt5 inhibits melanogenesis via cAMP pathway. Fish Physiology and Biochemistry, 43(2):517–525, April 2017. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-016-0306-3.

Xu:2017:EIG

[1019] Yongjiang Xu, Bin Wang, and Kun Zang. Evidences for involvement of growth hormone and insulin-like growth factor in ovarian development of starry flounder (Platichthys stellatus). Fish Physiology and Biochemistry, 43(2):527–537, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0307-2.

Sun:2017:TIH

[1020] Jian Sun, Zhou Yang, and LiQiao Chen. Two isoforms of hormonesensitive lipase b are generated by alternative exons usage and transcriptional regulation by insulin in grass carp (Ctenopharyngodon idella). Fish Physiology and Biochemistry, 43(2):539–547, April 2017. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-016-0308-1.

Ronning:2017:REM

[1021] Sissel B. Rønning, Tone-Kari Østbye, and Mona E. Pedersen. The role of extracellular matrix components in pin bone attachments during storage — a comparison between farmed Atlantic salmon (Salmo salar) and cod (Gadus morhua L.). Fish Physiology and Biochemistry, 43(2):549–562, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0309-0.

Perera:2017:ESM

[1022] Erick Perera and Manuel Yúfera. Effects of soybean meal on digestive enzymes activity, expression of inflammation-related genes, and chromatin modifications in marine fish (Sparus aurata L.) larvae. Fish Physiology and Biochemistry, 43(2):563–578, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0310-7.

Babaheydari:2017:MPR

[1023] Samad Bahrami Babaheydari, Saeed Keyvanshokooh, and Seyed Ali Johari. Modifications in the proteome of rainbow trout (Oncorhynchus mykiss) embryo and fry as an effect of triploidy induction. Fish Physiology and Biochemistry, 43(2):579–589, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0312-5.

Herrera:2017:EAA

[1024] Marcelino Herrera, María Antonia Herves, and Velmurugu Puvanendran. Effects of amino acid supplementations on metabolic and physiological parameters in Atlantic cod (Gadus morhua) under stress. Fish Physiology and Biochemistry, 43(2):591–602, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0314-3.

Schrama:2017:EDF

[1025] Denise Schrama, Nadège Richard, and Pedro M. L. Rodrigues. Enhanced dietary formulation to mitigate winter thermal stress in gilthead sea bream (Sparus aurata): a 2d-DIGE plasma proteome study. Fish Physiology and Biochemistry, 43(2):603–617, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0315-2.

Zhuo:2017:IIM

[1026] Mei-Qin Zhuo, Ya-Xiong Pan, and Zhi Luo. IRS1 and IRS2: molecular characterization, tissue expression and transcriptional regulation by insulin in yellow catfish Pelteobagrus fulvidraco. Fish Physiology and Biochemistry, 43(2):619–630, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-016-0316-1.

Zhang:2017:CAU

[1027] Yun-Long Zhang, Hai-Long Zhang, and Qi-Xue Fan. Changes of ammonia, urea contents and transaminase activity in the body during aerial exposure and ammonia loading in Chinese loach Paramisgurnus dabryanus. Fish Physiology and Biochemistry, 43(2):631–640, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0317-0.

Guan:2017:CDH

[1028] Wen-Zhi Guan, Dan-Dan Guo, and Shu-Ming Zou. Characterization of duplicated heme oxygenase-1 genes and their responses to hypoxic stress in blunt snout bream (Megalobrama amblycephala). Fish Physiology and Biochemistry, 43(2):641–651, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0318-z.

Zheng:2017:OSE

[1029] Jia-Lang Zheng, Shuang-Shuang Yuan, and Chang-Wen Wu. Organspecific effects of low-dose zinc pre-exposure on high-dose zinc induced mitochondrial dysfunction in large yellow croaker Pseudosciaena crocea. Fish Physiology and Biochemistry, 43(2):653–661, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0319-y.

Shen:2017:ILM

[1030] Yung-Chi Shen, Todd Hsu, and Chia-Wei Liu. Identification of lowmolecular-weight vitellogenin 1 (vg1)-like proteins as nucleotide excision repair (NER) factors in developing zebrafish (Danio rerio) using a transcription-based DNA repair assay. Fish Physiology and Biochemistry, 43(2):663–676, April 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0321-4.

Fawole:2017:HBN

[1031] Femi John Fawole, N. P. Sahu, and Nalini Poojary. Haematobiochemical, non-specific immunity, antioxidant capacity and histopathological changes in Labeo rohita fingerlings fed rubber protein isolate. Fish Physiology and Biochemistry, 43(3):677–690, June 2017. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-016-0322-3.

Ferreira:2017:ECL

[1032] Pollyanna de Moraes França Ferreira, Maria Tatiana Soares Martins, and Jener Alexandre Sampaio Zuanon. Curcuma longa as additive in the diet for Astyanax aff. bimaculatus. Fish Physiology and Biochemistry, 43(3):691–702, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0325-0.

Tian:2017:RCM

[1033] Jing-Jing Tian, Cai-Xia Lei, and Ai Jin. Role of cyclooxygenase-mediated metabolites in lipid metabolism and expression of some immune-related genes in juvenile grass carp (Ctenopharyngodon idellus) fed arachidonic acid. Fish Physiology and Biochemistry, 43(3):703–717, June 2017. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0326-z.

Zhuo:2017:APM

[1034] Mei-Qin Zhuo, Ya-Xiong Pan, and Zhi Luo. AKTs/PKBs: molecular characterization, tissue expression and transcriptional responses to insulin and/or wortmannin in yellow catfish Pelteobagrus fulvidraco. Fish Physiology and Biochemistry, 43(3):719–730, June 2017. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-016-0327-y.

Song:2017:CEK

[1035] Huayu Song, Mengxun Wang, and Quanqi Zhang. Characterization of kiss2 and kissr2 genes and the regulation of kisspeptin on the HPG axis in Cynoglossus semilaevis. Fish Physiology and Biochemistry, 43 (3):731–753, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0328-x.

Wolkers:2017:AFT

[1036] Carla Patricia Bejo Wolkers, Mônica Serra, and Elisabeth Criscuolo Urbinati. Acute fluoxetine treatment increases aggressiveness in juvenile matrinxã (Brycon amazonicus). Fish Physiology and Biochemistry, 43(3):755–759, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0329-9.

Liu:2017:EDS

[1037] Zhi-Feng Liu, Xiao-Qiang Gao, and Lei Hong. Effects of different salinities on growth performance, survival, digestive enzyme activity, immune response, and muscle fatty acid composition in juvenile American shad (Alosa sapidissima). Fish Physiology and Biochemistry, 43 (3):761–773, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0330-3.

Ilham:2017:GEG

[1038] I. Ilham and Ravi Fotedar. Growth, enzymatic glutathione peroxidase activity and biochemical status of juvenile barramundi (Lates calcarifer) fed dietary fermented soybean meal and organic selenium. Fish Physiology and Biochemistry, 43(3):775–790, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0331-2.

Ge:2017:DEM

[1039] Yinghua Ge, Jiewen Zhang, and Lili Liu. Differential expression and miRNA regulation of the GSTP1 gene in the regenerating liver of Chiloscyllium plagiosum. Fish Physiology and Biochemistry, 43(3): 791–802, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0332-1.

Song:2017:EFT

[1040] Yi Song, Cheng Zhao, and Ya-Xiong Tao. Effects of fasting, temperature, and photoperiod on preproghrelin mRNA expression in Chinese perch. Fish Physiology and Biochemistry, 43(3):803–812, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-016-0335-y.

Lei:2017:SGK

[1041] Caixia Lei, Jingjing Tian, and Hong Ji. Stimulation of glycerol kinase in grass carp preadipocytes by EPA. Fish Physiology and Biochemistry, 43(3):813–822, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0336-x.

Reyes-Tomassini:2017:SEA

[1042] José J. Reyes-Tomassini, Ten-Tsao Wong, and Yonathan Zohar. Seasonal expression of arginine vasotocin mRNA and its correlations to gonadal steroidogenic enzymes and sexually dimorphic coloration during sex reversal in the gilthead seabream (Sparus aurata). Fish Physiology and Biochemistry, 43(3):823–832, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0338-3.

Novelli:2017:DSS

[1043] B. Novelli, F. Otero-Ferrer, and L. Molina Domínguez. Development of short-snouted seahorse (Hippocampus hippocampus, l. 1758): osteological and morphological aspects. Fish Physiology and Biochemistry, 43(3):833–848, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0339-2.

Kwasek:2017:EHI

[1044] Karolina Kwasek, Simona Rimoldi, and Genciana Terova. The expression of hypoxia-inducible factor-1α gene is not affected by low-oxygen conditions in yellow perch (Perca flavescens) juveniles. Fish Physiology and Biochemistry, 43(3):849–862, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0340-9.

Huang:2017:IMC

[1045] Cui-Hong Huang, Nan Chen, and Huan-Ling Wang. Involvement of the miR-462/731 cluster in hypoxia response in Megalobrama amblycephala. Fish Physiology and Biochemistry, 43(3):863–873, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0341-8.

Tian:2017:MCG

[1046] Juan Tian, Wei Liu, and Hua Wen. Molecular cloning and gene/protein expression of FAT/CD36 from grass carp (Ctenopharyngodon idella) and the regulation of its expression by dietary energy. Fish Physiology and Biochemistry, 43(3):875–888, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0342-7.

Lin:2017:PLI

[1047] Tingting Lin, Xin Liu, and Dong Zhang. Plasma levels of immune factors and sex steroids in the male seahorse Hippocampus erectus during a breeding cycle. Fish Physiology and Biochemistry, 43(3):889– 899, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-017-0343-6.

Wang:2017:ESV

[1048] Liansheng Wang, Hong Xu, and Qiyou Xu. Effects of the supplementation of vitamin D_{3} on the growth and vitamin D metabolites in juvenile Siberian sturgeon (Acipenser baerii). Fish Physiology and Biochemistry, 43(3):901–909, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0344-5.

Hano:2017:EET

[1049] Takeshi Hano, Katsutoshi Ito, and Kazuhiko Mochida. Erratum to: Effect of taurine supplementation on hepatic metabolism and alleviation of cadmium toxicity and bioaccumulation in a marine teleost, red sea bream, Pagrus major. Fish Physiology and Biochemistry, 43 (3):911–912, June 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0323-2.

Zhang:2017:EDV

[1050] Yin Zhang, Yang Li, and Jian Gao. Effects of dietary vitamin E supplementation on growth performance, fatty acid composition, lipid peroxidation and peroxisome proliferator-activated receptors (PPAR) expressions in juvenile blunt snout bream Megalobrama amblycephala. Fish Physiology and Biochemistry, 43(4):913–922, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0224-4.

Sesay:2017:EDF

[1051] Daniella Fatmata Sesay, Habte-Michael Habte-Tsion, and Liangkun Pan. The effect of dietary folic acid on biochemical parameters and gene expression of three heat shock proteins (HSPs) of blunt snout bream (Megalobrama amblycephala) fingerling under acute high temperature stress. Fish Physiology and Biochemistry, 43(4):923–940, August 2017. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-016-0311-6.

Deng:2017:ERF

[1052] Junming Deng, Kun Wang, and Haifeng Mi. Effects of replacing fish meal with rubber seed meal on growth, nutrient utilization, and cholesterol metabolism of tilapia (Oreochromis niloticus × O. aureus). Fish Physiology and Biochemistry, 43(4):941–954, August 2017. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-016-0313-4.

Zeng:2017:ASS

[1053] Lin Zeng, Chun-Xiang Ai, and Chang-Wen Wu. Abrupt salinity stress induces oxidative stress via the Nrf2-Keap1 signaling pathway in large yellow croaker Pseudosciaena crocea. Fish Physiology and Biochemistry, 43(4):955–964, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0334-z.

Cheng:2017:MIR

[1054] Chang-Hong Cheng, Sheng-Wei Luo, and Zhi-Xun Guo. Molecular and immune response characterizations of a novel Bax inhibitor-1 gene in pufferfish, Takifugu obscurus. Fish Physiology and Biochemistry, 43 (4):965–975, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-016-0337-9.

Jing:2017:ECH

[1055] Hongli Jing, Xiangmei Lin, and Shaoqiang Wu. Establishment and characterization of a heart-derived cell line from goldfish (Carassius auratus). Fish Physiology and Biochemistry, 43(4):977–986, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0345-4.

Teng:2017:MCE

[1056] Tao Teng, Bingwen Xi, and Liangkun Pan. Molecular cloning and expression analysis of Megalobrama amblycephala transferrin gene and effects of exposure to iron and infection with Aeromonas hydrophila. Fish Physiology and Biochemistry, 43(4):987–997, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0346-3.

Rozynski:2017:ESI

[1057] Maciej Rożyński, Andrzej Kapusta, and Zdzisław Zakeś. The effects of surgically implanted dummy tags on the survival, growth performance, and physiology of pikeperch (Sander lucioperca). Fish Physiology and Biochemistry, 43(4):999–1010, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0347-2.

Xie:2017:EAC

[1058] Shiwei Xie, Lixia Tian, and Yongjian Liu. Effect of N-acetyl cysteine and glycine supplementation on growth performance, glutathione synthesis, and antioxidative ability of grass carp, Ctenopharyngodon idella. Fish Physiology and Biochemistry, 43(4):1011–1020, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0348-1.

Zhao:2017:PRE

[1059] Juan Zhao, Pei Wu, and Lin Feng. Preventive and reparative effects of isoleucine against copper-induced oxidative damage in primary fish enterocytes. Fish Physiology and Biochemistry, 43(4):1021–1032, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-017-0349-0.

Lopes:2017:UCG

[1060] F. P. Lopes, B. F. Pereira, and F. H. Caetano. Ultramorphological changes in gill rakers of Astyanax altiparanae (Characidae) kept in contaminated environments. Fish Physiology and Biochemistry, 43(4): 1033–1041, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0350-7.

Castanho:2017:ELF

[1061] S. Castanho, G. Califano, and L. Ribeiro. The effect of live feeds bathed with the red seaweed Asparagopsis armata on the survival, growth and physiology status of Sparus aurata larvae. Fish Physiology and Biochemistry, 43(4):1043–1054, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0351-6.

Zhang:2017:ICM

[1062] Runfeng Zhang, Ruiwen Li, and Yaqiu Lin. Identification and characterization of microRNAs in the muscle of Schizothorax prenanti. Fish Physiology and Biochemistry, 43(4):1055–1064, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0352-5.

Arnemo:2017:EDE

[1063] Marianne Arnemo, Arturas Kavaliauskis, and Tor Gjøen. Effects of dietary n-3 fatty acids on toll-like receptor activation in primary leucocytes from Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 43(4):1065–1080, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0353-4.

Chen:2017:MRC

[1064] Xiu-Mei Chen, Gui-Liang Guo, and Dong-Ming Zhang. Modulatory role of L-carnitine against microcystin-LR-induced immunotoxicity and oxidative stress in common carp. Fish Physiology and Biochemistry, 43(4): 1081–1093, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0354-3.

Sayed:2017:EDO

[1065] Alaa El-Din H. Sayed and Rania F. K. Ismail. Endocrine disruption, oxidative stress, and testicular damage induced by 4-nonylphenol in Clarias gariepinus: the protective role of Cydonia oblonga. Fish Physiology and Biochemistry, 43(4):1095–1104, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0355-2.

Hidalgo:2017:TSD

[1066] M. C. Hidalgo, C. E. Trenzado, and A. Sanz. Tissue-specific daily variation in the oxidative status of sturgeon (Acipenser naccarii) and rainbow trout (Oncorhynchus mykiss): a comparative study. Fish Physiology and Biochemistry, 43(4):1105–1115, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0356-1.

Churova:2017:AME

[1067] Maria V. Churova, Olga V. Meshcheryakova, and Nina N. Nemova. Activity of metabolic enzymes and muscle-specific gene expression in parr and smolts Atlantic salmon Salmo salar l. of different age groups. Fish Physiology and Biochemistry, 43(4):1117–1130, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0357-0.

Mahanty:2017:DSC

[1068] Arabinda Mahanty, Sasmita Mohanty, and Bimal P. Mohanty. Dietary supplementation of curcumin augments heat stress tolerance through upregulation of nrf-2 -mediated antioxidative enzymes and hsps in Puntius sophore. Fish Physiology and Biochemistry, 43(4):1131–1141, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-017-0358-z.

Mirghaed:2017:HIM

[1069] Ali Taheri Mirghaed, Melika Ghelichpour, and Kourosh Amini. Hemolysis interference in measuring fish plasma biochemical indicators. Fish Physiology and Biochemistry, 43(4):1143–1151, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0359-y.

Mekhtiev:2017:DIP

[1070] Arif A. Mekhtiev, Turan N. Allahverdiyeva, and Sevda K. Movsumzadeh. DNA integrity-protecting and survival-promoting activity of serotonergic system in sturgeon juveniles and sazans. Fish Physiology and Biochemistry, 43(4):1153–1160, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0360-5.

Dioguardi:2017:VDA

[1071] M. Dioguardi, F. A. Guardiola, and M. Cammarata. Vitamin D_{3} affects innate immune status of European sea bass (Dicentrarchus labrax L.). Fish Physiology and Biochemistry, 43(4):1161–1174, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0362-3.

Yang:2017:SMH

[1072] Lizhu Yang, Jing Fang, and Zhuangzhi Yang. Study on the morphology, histology and enzymatic activity of the digestive tract of Gymnocypris eckloni Herzenstein. Fish Physiology and Biochemistry, 43(4): 1175–1185, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0363-2.

Kantserova:2017:PDS

[1073] Nadezda P. Kantserova, Liudmila A. Lysenko, and Nina N. Nemova. Protein degradation systems in the skeletal muscles of parr and smolt Atlantic salmon Salmo salar l. and brown trout Salmo trutta L. Fish Physiology and Biochemistry, 43(4):1187–1194, August 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0364-1.

Ebrahimi:2017:CED

[1074] Mahdi Ebrahimi, Nor Hafizah Daeman, and Nicholas Romano. Comparing the effects of different dietary organic acids on the growth, intestinal short-chain fatty acids, and liver histopathology of red hybrid tilapia (Oreochromis sp.) and potential use of these as preservatives. Fish Physiology and Biochemistry, 43(4):1195–1207, August 2017. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0365-0.

Liu:2017:PEE

[1075] Yingjuan Liu, Chunyun Zhang, and Guojun Yin. Protective effect of Ganoderma lucidum polysaccharide against carbon tetrachloride-induced hepatic damage in precision-cut carp liver slices. Fish Physiology and Biochemistry, 43(5):1209–1221, October 2017. CODEN FPBIEP.

ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-016-0333-0.

Kong:2017:SDI

[1076] Wei-Guang Kong, Si-Si Li, and Zhi-Xin Wu. A study of the damage of the intestinal mucosa barrier structure and function of Ctenopharyngodon idella with Aeromonas hydrophila. Fish Physiology and Biochemistry, 43 (5):1223–1235, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0366-z.

Silva:2017:CEC

[1077] R. C. Silva, S. Liebel, and F. Filipak Neto. Cylindrospermopsin effects on cell viability and redox milieu of neotropical fish Hoplias malabaricus hepatocytes. Fish Physiology and Biochemistry, 43(5):1237– 1244, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0367-y.

Zhang:2017:ETR

[1078] Yueyang Zhang and James D. Kieffer. The effect of temperature on the resting and post-exercise metabolic rates and aerobic metabolic scope in shortnose sturgeon Acipenser brevirostrum. Fish Physiology and Biochemistry, 43(5):1245–1252, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0368-x.

Vijayalaxmi:2017:ILE

[1079] Vijayalaxmi and C. B. Ganesh. Influence of leucine-enkephalin on pituitary-ovary axis of the cichlid fish Oreochromis mossambicus. Fish Physiology and Biochemistry, 43(5):1253–1264, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0369-9.

Wang:2017:IED

[1080] Liping Wang, Yuzhe Han, and Ning Bao. Interactive effects of dietary leucine and isoleucine on growth, blood parameters, and amino acid profile of Japanese flounder Paralichthys olivaceus. Fish Physiology and Biochemistry, 43(5):1265–1278, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0370-3.

Yuan:2017:ETF

[1081] Xi Yuan, Yi hong Zhou, and Zhi ying Tu. Effects of temperature and fatigue on the metabolism and swimming capacity of juvenile Chinese sturgeon (Acipenser sinensis). Fish Physiology and Biochemistry, 43(5): 1279–1287, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0371-2.

Wang:2017:CLD

[1082] Na Wang, Ruoqing Wang, and Songlin Chen. Characterization of a lowdensity lipoprotein receptor, lrp13, in Chinese tongue sole (Cynoglossus semilaevis) and medaka (Oryzias latipes). Fish Physiology and Biochemistry, 43(5):1289–1298, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0372-1.

Wang:2017:CMP

[1083] Di Wang, Yong-Qiang Zhao, and Jun-Quan Zhu. Characterization of mitochondrial prohibitin from Boleophthalmus pectinirostris and evaluation of its possible role in spermatogenesis. Fish Physiology and Biochemistry, 43(5):1299–1313, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0373-0.

Hasanpour:2017:EDG

[1084] Soleiman Hasanpour, Amir Parviz Salati, and Hamid Mohammadi Azarm. Effects of dietary green tea (Camellia sinensis L.) supplementation on growth performance, lipid metabolism, and antioxidant status in a sturgeon hybrid of sterlet (Huso huso [male sign] × Acipenser ruthenus [female sign]) fed oxidized fish oil. Fish Physiology and Biochemistry, 43 (5):1315–1323, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0374-z.

Alimba:2017:DAA

[1085] Chibuisi G. Alimba, Raphael D. Ajiboye, and Olakunle S. Fagbenro. Dietary ascorbic acid reduced micronucleus and nuclear abnormalities in Clarias gariepinus (Burchell 1822) exposed to hospital effluent. Fish Physiology and Biochemistry, 43(5):1325–1335, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0375-y.

Li:2017:MCF

[1086] Xiang-Fei Li, Chao Xu, and Wen-Bin Liu. Molecular characterization of fructose-1,6-bisphosphatase 1b in blunt snout bream Megalobrama amblycephala and the transcriptional response to glucose loading after the adaptation to high-carbohydrate diets. Fish Physiology and Biochemistry, 43(5):1337–1349, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0376-x.

Zhang:2017:CTK

[1087] Dan-Dan Zhang, Xin-Ming Gao, and Jun-Quan Zhu. The C-terminal kinesin motor KIFC1 may participate in nuclear reshaping and flagellum formation during spermiogenesis of Larimichthys crocea. Fish Physiology and Biochemistry, 43(5):1351–1371, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0377-9.

Naderi:2017:ECH

[1088] Mahdi Naderi, Saeed Keyvanshokooh, and Alireza Ghaedi. Effects of chronic high stocking density on liver proteome of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 43(5):1373– 1385, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0378-8.

Zhang:2017:CTA

[1089] Yongqin Zhang, Jinhui Liu, and Yamei Xiao. Comparative transcriptome analysis of molecular mechanism underlying gray-to-red body color formation in red crucian carp (Carassius auratus, red var.). Fish Physiology and Biochemistry, 43(5):1387–1398, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0379-7.

Pereira:2017:ECP

[1090] Raquel Tatiane Pereira, Thaiza Rodrigues de Freitas, and Priscila Vieira Rosa. Endocrine cells producing peptide hormones in the intestine of Nile tilapia: distribution and effects of feeding and fasting on the cell density. Fish Physiology and Biochemistry, 43(5):1399–1412, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0380-1.

Yu:2017:RFO

[1091] Deng-Hang Yu, Jia-Zhi Chang, and Jun Liu. Replacement of fish oil with soybean oil in diets for juvenile Chinese sucker (Myxocyprinus asiaticus): effects on liver lipid peroxidation and biochemical composition. Fish Physiology and Biochemistry, 43(5):1413–1420, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0381-0.

Kim:2017:GPO

[1092] Jun-Hwan Kim, Hee-Ju Park, and Ju-Chan Kang. Growth performance, oxidative stress, and non-specific immune responses in juvenile sablefish, Anoplopoma fimbria, by changes of water temperature and salinity. Fish Physiology and Biochemistry, 43(5):1421–1431, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0382-z.

Jiang:2017:EBG

[1093] Danli Jiang, Yubo Wu, and Yan Wang. Effect of blood glucose level on acute stress response of grass carp Ctenopharyngodon idella. Fish Physiology and Biochemistry, 43(5):1433–1442, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0383-y.

Jiang:2017:ICE

[1094] Yonghua Jiang, Kunhuang Han, and Ziping Zhang. Isolation, characterization, and expression of proto-oncogene cmyc in large yellow croaker Larimichthys crocea. Fish Physiology and Biochemistry, 43(5): 1443–1461, October 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0384-x.

Wu:2017:SFA

[1095] Minglin Wu, Jiaqi Wang, and Xiaowu Chen. Sequence and functional analysis of intestinal alkaline phosphatase from Lateolabrax maculatus. Fish Physiology and Biochemistry, 43(6):1463–1476, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0385-9.

Wang:2017:TSC

[1096] Ruoqing Wang, Nianwei Zhang, and Na Wang. Two skin cell lines from wild-type and albino Japanese flounder (Paralichthys olivaceus): establishment, characterization, virus susceptibility, efficient transfection, and application to albinism study. Fish Physiology and Biochemistry, 43(6): 1477–1486, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0386-8.

Xiao:2017:SIA

[1097] Peizhen Xiao, Zhou Yang, and Jun Xie. Silymarin inhibits adipogenesis in the adipocytes in grass carp Ctenopharyngodon idellus in vitro and in vivo. Fish Physiology and Biochemistry, 43(6):1487– 1500, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0387-7.

Fuchs:2017:IIP

[1098] V. I. Fuchs, J. Schmidt, and D. Steinhagen. Influence of immunostimulant polysaccharides, nucleic acids, and Bacillus strains on the innate immune and acute stress response in turbots (Scophthalmus maximus) fed soy bean- and wheat-based diets. Fish Physiology and Biochemistry, 43(6):1501–1515, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0388-6.

Stell:2017:AIP

[1099] Sonja K. Stell and Peter Moller. Androgen-induced pseudo-hermaphroditic phenotypes in female Brevimyrus niger Günther 1866 (Teleostei, Mormyridae). Fish Physiology and Biochemistry, 43(6):1517–1529, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-017-0389-5.

Nie:2017:MBL

[1100] Li-Juan Nie and Shi-Jian Fu. Metabolic, behavioral, and locomotive effects of feeding in five cyprinids with different habitat preferences. Fish Physiology and Biochemistry, 43(6):1531–1542, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0390-z.

Kondo:2017:ESV

[1101] Fumitaka Kondo, Takashi Ohta, and Takeshi Miura. Effect of the squid viscera hydrolysate on growth performance and digestion in the red sea bream Pagrus major. Fish Physiology and Biochemistry, 43(6):1543– 1555, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0391-y.

Wang:2017:DES

[1102] Wei Wang, Hua Zhu, and CuiJuan Niu. Dimorphic expression of sexrelated genes in different gonadal development stages of sterlet, Acipenser ruthenus, a primitive fish species. Fish Physiology and Biochemistry, 43(6):1557–1569, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-017-0392-x.

Li:2017:MCC

[1103] Huan Li, Yujin Li, and Zhijian Wang. Molecular characterization of the CD79a and CD79b and its role against Aeromonas hydrophila infection in Chinese sucker (Myxocyprinus asiaticus). Fish Physiology and Biochemistry, 43(6):1571–1585, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0394-8.

Li:2017:LSA

[1104] Siping Li, Feng He, and Lingcai Meng. Low salinity affects cellularity, DNA methylation, and mRNA expression of igf1 in the liver of half smooth tongue sole (Cynoglossus semilaevis). Fish Physiology and Biochemistry, 43(6):1587–1602, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0395-7.

Ohkubo:2017:DMA

[1105] Nobuyuki Ohkubo, Yuji Tomaru, and Kazuhiko Mochida. Development of a method to assess the ichthyotoxicity of the harmful marine microalgae Karenia spp. using gill cell cultures from red sea bream (Pagrus major). Fish Physiology and Biochemistry, 43(6):1603–1612, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-017-0396-6.

Teles:2017:HSG

[1106] Andressa Teles, Joan Salas-Leiva, and Dariel Tovar-Ramírez. Histological study of the gastrointestinal tract in longfin yellowtail (Seriola rivoliana) larvae. Fish Physiology and Biochemistry, 43(6):1613– 1628, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0397-5.

Hossain:2017:DSF

[1107] Md. Sakhawat Hossain and Shunsuke Koshio. Dietary substitution of fishmeal by alternative protein with guanosine monophosphate supplementation influences growth, digestibility, blood chemistry profile, immunity, and stress resistance of red sea bream, Pagrus major. Fish Physiology and Biochemistry, 43(6):1629–1644, December 2017. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0398-4.

Soltanian:2017:IRC

[1108] Siyavash Soltanian and Mohammad Saeed Fereidouni. Immunotoxic responses of chronic exposure to cypermethrin in common carp. Fish Physiology and Biochemistry, 43(6):1645–1655, December 2017. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-017-0399-3.

Xiao:2017:SET

[1109] Weiwei Xiao, Weidan Jiang, and Xiaoqiu Zhou. Supplementation of enzyme-treated soy protein saves dietary protein and promotes digestive and absorptive ability referring to TOR signaling in juvenile fish. Fish Physiology and Biochemistry, 43(6):1657–1675, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0400-1.

Bera:2017:DCH

[1110] Aritra Bera, Paramita Banerjee Sawant, and Asim Kumar Pal. Diel cyclic hypoxia alters plasma lipid dynamics and impairs reproduction in goldfish (Carassius auratus). Fish Physiology and Biochemistry, 43(6): 1677–1688, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0401-0.

Khan:2017:OOI

[1111] Kifayat Ullah Khan, Amina Zuberi, and Huda Sarwar. An overview of the ongoing insights in selenium research and its role in fish nutrition and fish health. Fish Physiology and Biochemistry, 43(6):1689– 1705, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0402-z.

Ren:2017:ESD

[1112] Yuanyuan Ren, Haishen Wen, and Meng Ni. Effects of stocking density on lipid deposition and expression of lipid-related genes in Amur sturgeon (Acipenser schrenckii). Fish Physiology and Biochemistry, 43(6): 1707–1720, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0403-y.

Hamed:2017:MEL

[1113] Heba S. Hamed and Alaa G. M. Osman. Modulatory effect of lycopene against carbofuran toxicity in African catfish, Clarias gariepinus. Fish Physiology and Biochemistry, 43(6):1721–1731, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0404-x.

Adjoumani:2017:EDB

[1114] Jean-Jacques Yao Adjoumani, Kaizhou Wang, and Dingdong Zhang. Effect of dietary betaine on growth performance, antioxidant capacity and lipid metabolism in blunt snout bream fed a high-fat diet. Fish Physiology and Biochemistry, 43(6):1733–1745, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0405-9.

Maltez:2017:AES

[1115] Lucas Campos Maltez, Giovanna Rodrigues Stringhetta, and Luciano Garcia. Ammonia exposure and subsequent recovery trigger oxidative stress responses in juveniles of Brazilian flounder Paralichthys orbignyanus. Fish Physiology and Biochemistry, 43(6):1747–1759, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-017-0406-8.

Vargas:2017:EOH

[1116] Rafael Vargas and Isabel Cristina Vásquez. Effects of overfeeding and high-fat diet on cardiosomatic parameters and cardiac structures in young and adult zebrafish. Fish Physiology and Biochemistry, 43(6): 1761–1773, December 2017. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0407-7.

Bruning:2018:ILI

[1117] Anika Brüning, Franz Hölker, and Werner Kloas. Influence of light intensity and spectral composition of artificial light at night on melatonin rhythm and mRNA expression of gonadotropins in roach Rutilus rutilus. Fish Physiology and Biochemistry, 44(1):1–12, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0408-6.

Huang:2018:FPG

[1118] Songqian Huang, Xiaojuan Cao, and Mona Nasr. Fertility and ploidy of gametes of allodiploid and allotriploid loaches produced by diploid Misgurnus anguillicaudatus females and Paramisgurnus dabryanus males. Fish Physiology and Biochemistry, 44(1):13–20, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0409-5.

Souza:2018:CLC

[1119] Carine de Freitas Souza, Matheus Dellaméa Baldissera, and Bernardo Baldisserotto. Citral and linalool chemotypes of Lippia alba essential oil as anesthetics for fish: a detailed physiological analysis of side effects during anesthetic recovery in silver catfish (Rhamdia quelen). Fish Physiology and Biochemistry, 44(1):21–34, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0410-z.

Feng:2018:PIT

[1120] Chengcheng Feng, Shihong Xu, and Jun Li. Progestin is important for testicular development of male turbot (Scophthalmus maximus) during the annual reproductive cycle through functionally distinct progestin receptors. Fish Physiology and Biochemistry, 44(1):35–48, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0411-y.

Roberts:2018:EEE

[1121] Jordan C. Roberts and Douglas A. Syme. Effects of epinephrine exposure on contractile performance of compact and spongy myocardium from rainbow trout (Oncorhynchus mykiss) during hypoxia. Fish Physiology and Biochemistry, 44(1):49–62, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0412-x.

Zhao:2018:RRG

[1122] Zhigang Zhao, Shuanglin Dong, and Qiyou Xu. Respiratory response of grass carp Ctenopharyngodon idellus to dissolved oxygen changes at three acclimation temperatures. Fish Physiology and Biochemistry, 44 (1):63–71, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0413-9.

Klanian:2018:PPG

[1123] Mariel Gullian Klanian, Omar Zapata Pérez, and Miguel Angel Vela-Magaña. Phenotypic plasticity in gene expression and physiological response in red drum Sciaenops ocellatus exposed to a long-term freshwater environment. Fish Physiology and Biochemistry, 44(1):73–85, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-017-0414-8.

Liu:2018:ECK

[1124] Xiao-Feng Liu, Ya-Hong Wu, and Song-Lin Chen. Establishment and characterization of a kidney cell line from kelp grouper Epinephelus moara. Fish Physiology and Biochemistry, 44(1):87–93, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0415-7.

Kumari:2018:GSC

[1125] Kavita Kumari, Gireesh-Babu Pathakota, and Gopal Krishna. Gene structure and comparative and phylogenetic analyses of Catla catla CYP1A full-length cDNA and its responsiveness to benzo(a)pyrene and copper sulphate at early developmental stages. Fish Physiology and Biochemistry, 44(1):95–108, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0416-6.

Bulfon:2018:MPE

[1126] Chiara Bulfon, Marco Galeotti, and Donatella Volpatti. Medicinal plant extracts modulate respiratory burst and proliferation activity of rainbow trout (Oncorhynchus mykiss) leukocytes. Fish Physiology and Biochemistry, 44(1):109–117, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0417-5.

Kim:2018:EGH

[1127] Jin-Hyoung Kim, Satid Chatchaiphan, and Robert H. Devlin. Effect of growth hormone overexpression on gastric evacuation rate in Coho salmon. Fish Physiology and Biochemistry, 44(1):119–135, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-017-0418-4.

Birnie-Gauvin:2018:AMF

[1128] Kim Birnie-Gauvin, Martin H. Larsen, and Steven J. Cooke. Nacetylcysteine manipulation fails to elicit an increase in glutathione in a teleost model. Fish Physiology and Biochemistry, 44(1):137–142, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-017-0419-3.

Nepal:2018:CNV

[1129] Sunil Nepal, Vikas Kumar, and Klaus Becker. Comparative nutritional value of Jatropha curcas protein isolate and soy protein isolate in common carp. Fish Physiology and Biochemistry, 44(1):143–162, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0420-x.

Zhang:2018:RSI

[1130] Dingdong Zhang, Yanan Yan, and Wenbin Liu. Resveratrol supplementation improves lipid and glucose metabolism in high-fat diet-fed blunt snout bream. Fish Physiology and Biochemistry, 44(1):163–173, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-017-0421-9.

He:2018:INS

[1131] Yu-Hui He, Ling Li, and Yan-Peng Zhang. Inhibitory neurotransmitter serotonin and excitatory neurotransmitter dopamine both decrease food intake in Chinese perch (Siniperca chuatsi). Fish Physiology and Biochemistry, 44(1):175–183, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0422-8.

Huang:2018:EFC

[1132] Tianqing Huang, Huizhi Sun, and Ying Han. Effect of follicle cell autophagy on gonadal development of triploid female rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 44(1):185–196, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-017-0423-7.

Jin:2018:DDA

[1133] Ai Jin, Cai xia Lei, and Hong Ji. Dietary docosahexaenoic acid decreased lipid accumulation via inducing adipocytes apoptosis of grass carp, Ctenopharygodon idella. Fish Physiology and Biochemistry, 44(1):

197–207, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0424-6.

Cheng:2018:EDA

[1134] Chang-Hong Cheng, Zhi-Xun Guo, and An-Li Wang. Effect of dietary astaxanthin on the growth performance, non-specific immunity, and antioxidant capacity of pufferfish (Takifugu obscurus) under high temperature stress. Fish Physiology and Biochemistry, 44(1):209–218, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0425-5.

Castillo:2018:GRA

[1135] A. Castillo, A. Alvarez, and G. Gaxiola. Glycemic response after glucose oral administration of wild juvenile red grouper Epinephelus morio fed two different diets. Fish Physiology and Biochemistry, 44(1): 219–226, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0426-4.

Oguz:2018:DOT

[1136] Ahmet R. Oğuz. Development of osmoregulatory tissues in the Lake Van fish (Alburnus tarichi) during larval development. Fish Physiology and Biochemistry, 44(1):227–233, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0427-3.

Khoshroo:2018:SIR

[1137] Mahsa Mohammadi-Zadeh Khoshroo, Mehdi Shamsaie Mehrjan, and Seyed Pezhman Hosseini Shekarabi. Some immunological responses of common carp (Cyprinus carpio) fingerling to acute extremely lowfrequency electromagnetic fields (50 Hz). Fish Physiology and Biochemistry, 44(1):235–243, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0429-1.

Fujimoto:2018:COI

[1138] Rodrigo Yudi Fujimoto, Débora Martins Pereira, and Luis André Luz Barbas. Clove oil induces anaesthesia and blunts muscle contraction power in three Amazon fish species. Fish Physiology and Biochemistry, 44 (1):245–256, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0430-8.

Cortes:2018:EAH

[1139] Raul Cortés, Mariana Teles, and José Miguel Cerdá-Reverter. Effects of acute handling stress on short-term central expression of orexigenic/ anorexigenic genes in zebrafish. Fish Physiology and Biochemistry, 44 (1):257–272, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0431-7.

Matias:2018:PSH

[1140] Ana C. Matias, Laura Ribeiro, and Pedro Pousão-Ferreira. Preliminary studies on haematological and plasmatic parameters in gilthead sea bream (Sparus aurata) held under day/night temperature variations. Fish Physiology and Biochemistry, 44(1):273–282, February 2018. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0432-6.

Torrecillas:2018:SAA

[1141] S. Torrecillas, M. B. Betancor, and D. Montero. Supplementation of arachidonic acid rich oil in European sea bass juveniles (Dicentrarchus labrax) diets: effects on growth performance, tissue fatty acid profile and lipid metabolism. Fish Physiology and Biochemistry, 44(1): 283–300, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0433-5.

Ammar:2018:DCP

[1142] I. Ben Ammar, S. Milla, and P. Fontaine. Does constant photoperiod inhibit the onset of the reproductive cycle in northern pike (Esox lucius) males? Fish Physiology and Biochemistry, 44(1):301–310, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0435-3.

Yu:2018:EAT

[1143] Xiaoming Yu, Lei Chen, and Guosheng Zhang. Effects of acute temperature and salinity changes, body length and starvation on the critical swimming speed of juvenile tiger puffer, Takifugu rubripes. Fish Physiology and Biochemistry, 44(1):311–318, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0436-2.

Jiang:2018:LMS

[1144] Jun Jiang, Shangxiao Xu, and Xiaoqiu Zhou. Lysine and methionine supplementation ameliorates high inclusion of soybean meal inducing intestinal oxidative injury and digestive and antioxidant capacity decrease of yellow catfish. Fish Physiology and Biochemistry, 44(1): 319–328, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0437-1.

Khan:2018:STF

[1145] J. R. Khan, C. C. Lazado, and P. V. Skov. Short-term feed and light deprivation reduces voluntary activity but improves swimming performance in rainbow trout Oncorhynchus mykiss. Fish Physiology and Biochemistry, 44(1):329–341, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0438-0.

Milinkovitch:2018:ITA

[1146] Thomas Milinkovitch, Christel Lefrançois, and Hélène Thomas-Guyon. Influence of temperature on age-related lipid peroxidation in a shortlived vertebrate (Nothobranchius furzeri). Fish Physiology and Biochemistry, 44(1):343–347, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0439-z.

Pujante:2018:EDS

[1147] I. M. Pujante, F. J. Moyano, and G. Martínez-Rodríguez. Effect of different salinities on gene expression and activity of digestive enzymes in the thick-lipped grey mullet (Chelon labrosus). Fish Physiology and Biochemistry, 44(1):349–373, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0440-6.

Jia:2018:IES

[1148] Yongfang Jia, Fang Wang, and Zhongjie Chang. Identification of suh gene and evidence for involvement of notch signaling pathway on gonadal differentiation of Yellow River carp (Cyprinus carpio). Fish Physiology and Biochemistry, 44(1):375–386, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0441-5.

Huang:2018:EFA

[1149] Hongli Huang, Yu Zhang, and Weiliang Shen. Effects of fasting on the activities and mRNA expression levels of lipoprotein lipase (LPL), hormone-sensitive lipase (HSL) and fatty acid synthetase (FAS) in spotted seabass Lateolabrax maculatus. Fish Physiology and Biochemistry, 44 (1):387–400, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0442-4.

Lin:2018:EDP

[1150] Shi-Mei Lin, Fa-Jian Li, and Sompong Doolgindachbaporn. Effect of dietary phospholipid levels on growth, lipid metabolism, and antioxidative status of juvenile hybrid snakehead (Channa argus×Channa maculata). Fish Physiology and Biochemistry, 44(1):401–410, February 2018. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0443-3.

Zhang:2018:OEC

[1151] Xin Zhang, Yundi Gao, and Zhiqiong Li. One evidence of cocaine- and amphetamine-regulated transcript (CART) has the bidirectional effects on appetite in Siberian sturgeon (Acipenser baerii). Fish Physiology and Biochemistry, 44(1):411–422, February 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0444-2.

Cao:2018:PMO

[1152] Quanquan Cao, Jie Gu, and Shaowu Yin. Physiological mechanism of osmoregulatory adaptation in anguillid eels. Fish Physiology and Biochemistry, 44(2):423–433, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0464-6.

Wu:2018:BAZ

[1153] Tianli Wu, Yunying Cheng, and Deshou Wang. Bioinformatic analyses of zona pellucida genes in vertebrates and their expression in Nile tilapia. Fish Physiology and Biochemistry, 44(2):435–449, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0434-4.

Fu:2018:CEE

[1154] Yuanshuai Fu, Lina Gao, and Wenjuan Li. Characterization and expression of lin-28a involved in lin28/let-7 signal pathway during early development of P. olivaceus. Fish Physiology and Biochemistry, 44 (2):451–463, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0445-1.

Zago:2018:EAT

[1155] Daniane C. Zago, Alessandro C. Santos, and Mauro A. Cunha. Aloysia triphylla in the zebrafish food: effects on physiology, behavior, and growth performance. Fish Physiology and Biochemistry, 44(2):465–474, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-017-0446-0.

Phulia:2018:SSM

[1156] Vikas Phulia, Parimal Sardar, and Subodh Gupta. Substitution of soybean meal with fermented Jatropha kernel meal: effect on growth performance, body composition, and metabolic enzyme activity of Labeo rohita. Fish Physiology and Biochemistry, 44(2):475–487, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0447-z.

Tsai:2018:FLS

[1157] Jeng-Wei Tsai, Hon-Jung Liew, and Cheng-Hao Tang. A field and laboratory study of the responses of cytoprotection and osmoregulation to salinity stress in mosquitofish (Gambusia affinis). Fish Physiology and Biochemistry, 44(2):489–502, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0448-y.

Higaki:2018:CMF

[1158] Shogo Higaki, Takaaki Todo, and Tatsuyuki Takada. Cryopreservation of male and female gonial cells by vitrification in the critically endangered cyprinid honmoroko Gnathopogon caerulescens. Fish Physiology and Biochemistry, 44(2):503–513, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0449-x.

Ammar:2018:CET

[1159] Asmma Y. Ammar, Abeer F. El Nahas, and Asmaa M. Hassan. Characterization of Type IV antifreeze gene in

Nile tilapia (Oreochromis niloticus) and influence of cold and hot weather on its expression and some immune-related genes. Fish Physiology and Biochemistry, 44(2):515–525, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0450-4.

Chatzifotis:2018:ELT

[1160] Stavros Chatzifotis, Sofia Clavero, and Efthimia Antonopoulou. Effects of long-term feed deprivation on body weight loss, muscle composition, plasma metabolites, and intermediate metabolism of meagre (Argyrosomus regius) under different water temperatures. Fish Physiology and Biochemistry, 44(2):527–542, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0451-3.

Na-Phatthalung:2018:IEE

[1161] Pinanong Na-Phatthalung, Mariana Teles, and Camino Fierro-Castro. Immunomodulatory effects of Rhodomyrtus tomentosa leaf extract and its derivative compound, rhodomyrtone, on head kidney macrophages of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 44(2):543–555, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0452-2.

Qi:2018:MEM

[1162] Delin Qi, Yan Chao, and Rongrong Wu. Molecular evolution of myoglobin in the Tibetan Plateau endemic schizothoracine fish (Cyprinidae, Teleostei) and tissue-specific expression changes under hypoxia. Fish Physiology and Biochemistry, 44(2):557–571, April 2018. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-017-0453-1.

Xu:2018:EAK

[1163] Qiyou Xu and Delbert M. Gatlin III. Effects of alpha-ketoglutarate (AKG) on growth performance and non-specific immunity of juvenile red drum fed diets with low or adequate phosphorus levels. Fish Physiology and Biochemistry, 44(2):573–582, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-017-0454-0.

Menzikov:2018:BPS

[1164] Sergey Menzikov. Biochemical properties of the sensitivity to GABA_{A} ergic ligands, Cl^{−} /HCO_{3}^{−}-ATPase isolated from fish (Cyprinus carpio) olfactory mucosa and brain. Fish Physiology and Biochemistry, 44 (2):583–597, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0455-z.

Schneebauer:2018:TLM

[1165] Gabriel Schneebauer, David Mauracher, and Bernd Pelster. Transcript levels of members of the SLC2 and SLC5 families of glucose transport proteins in eel swimbladder tissue: the influence of silvering and the influence of a nematode infection. Fish Physiology and Biochemistry, 44(2):599–613, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0456-y.

Ruiz-Jarabo:2018:GET

[1166] Ignacio Ruiz-Jarabo, J. A. Martos-Sitcha, and F. J. Arjona. Gene expression of thyrotropin- and corticotrophin-releasing hormones is regulated by environmental salinity in the euryhaline teleost Sparus aurata. Fish Physiology and Biochemistry, 44(2):615–628, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0457-x.

Srivastava:2018:AAC

[1167] Ayan Srivastava, Usha Kumari, and Ajay Kumar Mittal. Alterations in the activity of certain enzymes in the gills of a carp Labeo rohita exposed to an azo dye, eriochrome black t: a biochemical investigation. Fish Physiology and Biochemistry, 44(2):629–637, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-017-0458-9.

Abdel-Tawwab:2018:SED

[1168] Mohsen Abdel-Tawwab and Mohamed N. Monier. Stimulatory effect of dietary taurine on growth performance, digestive enzymes activity, antioxidant capacity, and tolerance of common carp, Cyprinus carpio L., fry to salinity stress. Fish Physiology and Biochemistry, 44 (2):639–649, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-017-0459-8.

Wu:2018:DGE

[1169] Fengrui Wu, Limin Wu, and Deshou Wang. Duplication and gene expression patterns of β-catenin in Nile tilapia. Fish Physiology and Biochemistry, 44(2):651–659, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-017-0460-2.

Vizcaino:2018:GPB

[1170] A. J. Vizcaíno, A. Rodiles, and F. Javier Alarcón. Growth performance, body composition, and digestive functionality of Senegalese sole (Solea senegalensis Kaup, 1858) juveniles fed diets including microalgae freeze-dried biomass. Fish Physiology and Biochemistry, 44 (2):661–677, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0462-8.

Kinoshita:2018:PAF

[1171] Shigeharu Kinoshita, Saltuk Bugrahan Ceyhun, and Shugo Watabe. Promoter analysis of the fish gene of slow/cardiac-type myosin heavy chain implicated in specification of muscle fiber types. Fish Physiology and Biochemistry, 44(2):679–691, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0463-7.

You:2018:PYC

[1172] Wen-Jing You, Xiao-Ying Tan, and Dan-Dan Li. PPARβ in yellow catfish Pelteobagrus fulvidraco: molecular characterization, tissue expression and transcriptional regulation by dietary Cu and Zn. Fish Physiology and Biochemistry, 44(2):693–702, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0465-5.

Arge:2018:ETA

[1173] Regin Arge, Jens-Erik Dessen, and Kjell-Arne Rørvik. Effects of tetradecylthioacetic acid (TTA) treatment on lipid metabolism in salmon hearts — in vitro and in vivo studies. Fish Physiology and Biochemistry, 44 (2):703–716, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0466-4.

Shang:2018:TGL

[1174] Mei Shang, Baofeng Su, and Rex A. Dunham. Testicular germ line cell identification, isolation, and transplantation in two North American catfish species. Fish Physiology and Biochemistry, 44(2):717–733, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-018-0467-3.

Cheng:2018:GPP

[1175] Chang-Hong Cheng, Zhi-Xun Guo, and An-Li Wang. Growth performance and protective effect of vitamin E on oxidative stress pufferfish (Takifugu obscurus) following by ammonia stress. Fish Physiology and Biochemistry, 44(2):735–745, April 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0468-2.

Du:2018:ERA

[1176] Jinliang Du, Rui Jia, and Guojun Yin. Effects of rhizoma alismatis extract on biochemical indices and adipose gene expression in oleic acidinduced hepatocyte injury in Jian carp (Cyprinus carpio var. Jian). Fish Physiology and Biochemistry, 44(3):747–768, June 2018. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-017-0428-2.

Zhao:2018:AFK

[1177] Yong-Qiang Zhao, Dan-Li Mu, and Jun-Quan Zhu. Analysis of the function of KIF3A and KIF3B in the spermatogenesis in Boleophthalmus pectinirostris. Fish Physiology and Biochemistry, 44(3):769–788, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-017-0461-1.

Liu:2018:DAS

[1178] Zhi-Hao Liu, Qi-Liang Chen, and Ying-Wen Li. Diethylstilbestrol arrested spermatogenesis and somatic growth in the juveniles of yellow catfish (Pelteobagrus fulvidraco), a fish with sexual dimorphic growth. Fish Physiology and Biochemistry, 44(3):789–803, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0469-1.

Wang:2018:HIR

[1179] Shuqi Wang, Junliang Chen, and Yuanyou Li. Hnf4α is involved in the regulation of vertebrate LC-PUFA biosynthesis: insights into the regulatory role of Hnf4α on expression of liver fatty acyl desaturases in the marine teleost Siganus canaliculatus. Fish Physiology and Biochemistry, 44(3):805–815, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0470-8.

Bouchekioua:2018:ETM

[1180] Selma Bouchekioua, Sung-Pyo Hur, and Akihiro Takemura. Effects of temperature and melatonin on day–night expression patterns of arginine vasotocin and isotocin mRNA in the diencephalon of a temperate wrasse Halichoeres tenuispinis. Fish Physiology and Biochemistry, 44 (3):817–828, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0471-7.

Wang:2018:SHS

[1181] Shengpeng Wang, Bin Wang, and Songlin Chen. Spexin in the halfsmooth tongue sole (Cynoglossus semilaevis): molecular cloning, expression profiles, and physiological effects. Fish Physiology and Biochemistry, 44(3):829–839, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0472-6.

Husak:2018:AEC

[1182] Viktor V. Husak, Nadia M. Mosiichuk, and Volodymyr I. Lushchak. Acute exposure to copper induces variable intensity of oxidative stress in goldfish tissues. Fish Physiology and Biochemistry, 44(3):841–852, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-018-0473-5.

Yaghoubi:2018:GHE

[1183] Morteza Yaghoubi, Mansour Torfi Mozanzadeh, and Jasem G. Marammazi. Gastrointestinal and hepatic enzyme activities in juvenile silveryblack porgy (Sparidentex hasta) fed essential amino acid-deficient diets. Fish Physiology and Biochemistry, 44(3):853–868, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0475-3.

Liang:2018:MCI

[1184] Hualiang Liang, Ahmed Mokrani, and Mingchun Ren. Molecular characterization and identification of facilitative glucose transporter 2 (GLUT2) and its expression and of the related glycometabolism enzymes in response to different starch levels in blunt snout bream (Megalobrama amblycephala). Fish Physiology and Biochemistry, 44 (3):869–883, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0477-1.

Oliveira:2018:MMP

[1185] Catarina C. V. Oliveira, Filipe Figueiredo, and Maria Teresa Dinis. Meagre’s melatonin profiles under captivity: circadian rhythmicity and light sensitiveness. Fish Physiology and Biochemistry, 44(3):885–893, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-018-0478-0.

Candelma:2018:CCM

[1186] Michela Candelma, Luisa Dalla Valle, and Oliana Carnevali. Cloning, characterization, and molecular expression of gonadotropin receptors in European hake (Merluccius merluccius), a multiple-spawning species. Fish Physiology and Biochemistry, 44(3):895–910, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0479-z.

Castro:2018:NHD

[1187] Carolina Castro, Amalia Peréz-Jiménez, and Paula Enes. Nutritional history does not modulate hepatic oxidative status of European sea bass (Dicentrarchus labrax) submitted to handling stress. Fish Physiology and Biochemistry, 44(3):911–918, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0480-6.

Mirghaed:2018:RTE

[1188] Ali Taheri Mirghaed, Mahyar Yasari, and Seyyed Morteza Hoseini. Rainbow trout (Oncorhynchus mykiss) anesthesia with myrcene: efficacy and physiological responses in comparison with eugenol. Fish Physiology and Biochemistry, 44(3):919–926, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0481-5.

Rozynski:2018:IIG

[1189] Maciej Rożyński, Krystyna Demska-Zakeś, and Zdzisław Zakeś. Impact of inducing general anesthesia with propiscin (etomidate) on the physiology and health of European perch (Perca fluviatilis L.). Fish Physiology and Biochemistry, 44(3):927–937, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0482-4.

Pinedo-Gil:2018:EDI

[1190] Julia Pinedo-Gil, Ana Belén Martín-Diana, and Ana Tomás-Vidal. Effects of dietary inclusions of red beet and betaine on the acute stress response and muscle lipid peroxidation in rainbow trout. Fish Physiology and Biochemistry, 44(3):939–948, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0483-3.

Bernal:2018:TEB

[1191] Diego Bernal, Joseph P. Reid, and Jeffrey B. Graham. Temperature effects on the blood oxygen affinity in sharks. Fish Physiology and Biochemistry, 44(3):949–967, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0484-2.

Abd-Elkareem:2018:HRN

[1192] Mahmoud Abd-Elkareem, Nasser S. Abou Khalil, and Alaa H. Sayed. Hepatotoxic responses of 4-nonylphenol on African catfish (Clarias gariepinus): antixoidant and histochemical biomarkers. Fish Physiology and Biochemistry, 44(3):969–981, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0485-1.

Liu:2018:CEA

[1193] Xiao-Hong Liu, Bi-Wen Xie, and Yao-Guang Zhang. Characterization and expression analyses of somatolactin-α and -β genes in rare minnows (Gobiocypris rarus) following waterborne cadmium exposure. Fish Physiology and Biochemistry, 44(3):983–995, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0487-z.

Rehberger:2018:ITH

[1194] Kristina Rehberger, Lisa Baumann, and Thomas Braunbeck. Intrafollicular thyroid hormone staining in whole-mount zebrafish (Danio rerio) embryos for the detection of thyroid hormone synthesis disruption. Fish Physiology and Biochemistry, 44(3):997–1010, June 2018. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-018-0488-y.

Kondera:2018:EGB

[1195] E. Kondera, B. Teodorczuk, and M. Witeska. Effect of glyphosatebased herbicide on hematological and hemopoietic parameters in common carp (Cyprinus carpio l). Fish Physiology and Biochemistry, 44 (3):1011–1018, June 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0489-x.

Shi:2018:CPM

[1196] Xiao chen Shi, Jian Sun, and Ji qin Huang. Cytochrome P450 2AA molecular clone, expression pattern, and different regulation by fish oil and lard oil in diets of grass carp (Ctenopharyngodon idella). Fish Physiology and Biochemistry, 44(4):1019–1026, August 2018. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-018-0486-0.

Choi:2018:ERV

[1197] Ji Yong Choi and Cheol Young Choi. Effects of recombinant vertebrate ancient long opsin on reproduction in goldfish, Carassius auratus: profiling green-wavelength light. Fish Physiology and Biochemistry, 44(4): 1027–1036, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0490-4.

Ranjan:2018:CEF

[1198] Amit Ranjan, Narottam Prasad Sahu, and Kamal Kant Jain. Comparative evaluation of fermented and non-fermented de-oiled rice bran with or without exogenous enzymes supplementation in the diet of Labeo rohita (Hamilton, 1822). Fish Physiology and Biochemistry, 44 (4):1037–1049, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0492-2.

Baldissera:2018:CCP

[1199] Matheus D. Baldissera, Carine F. Souza, and Bernardo Baldisserotto. Changes in the cerebral phosphotransfer network impair energetic homeostasis in an aflatoxin B_{1}-contaminated diet. Fish Physiology and Biochemistry, 44(4):1051–1059, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0493-1.

Pal:2018:RGM

[1200] Palash Kumar Pal and Saumen Kumar Maitra. Response of gastrointestinal melatonin, antioxidants, and digestive enzymes to altered feeding conditions in carp (Catla catla). Fish Physiology and Biochemistry, 44 (4):1061–1073, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0494-0.

Saekhow:2018:MWV

[1201] Suktianchai Saekhow, Karun Thongprajukaew, and Harit Sae-khoo. Minimal water volume for intensively producing male Siamese fighting fish (Betta splendens Regan, 1910). Fish Physiology and Biochemistry, 44 (4):1075–1085, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0495-z.

Saffari:2018:EDO

[1202] Sadegh Saffari, Saeed Keyvanshokooh, and Mansour Torfi Mozanzadeh. Effects of dietary organic, inorganic, and nanoparticulate selenium sources on growth, hemato-immunological, and serum biochemical parameters of common carp (Cyprinus carpio). Fish Physiology and Biochemistry, 44(4):1087–1097, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0496-y.

Ghiasi:2018:IEP

[1203] Maryam Ghiasi, Mohammad Binaii, and Atefeh Amerizadeh. Inclusion of Pediococcus acidilactici as probiotic candidate in diets for beluga (Huso huso) modifies biochemical parameters and improves immune functions. Fish Physiology and Biochemistry, 44(4):1099–1107, August 2018. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0497-x.

Lugowska:2018:ERG

[1204] Katarzyna Lugowska. The effects of roundup on gametes and early development of common carp (Cyprinus carpio l). Fish Physiology and Biochemistry, 44(4):1109–1117, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0498-9.

Kuzu:2018:IPS

[1205] Müslüm Kuzu, Veysel Çomaklı, and Ömer İrfan Küfrevioğlu. Inhibitory properties of some heavy metals on carbonic anhydrase I and II isozymes activities purified from Van Lake fish (Chalcalburnus tarichi) gill. Fish Physiology and Biochemistry, 44(4):1119–1125, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0499-8.

Wu:2018:ICE

[1206] Chenglong Wu, Jun’e Gao, and Jinyun Ye. Identification, characterization, and expression analysis of adiponectin receptors in black carp Mylopharyngodon piceus in response to dietary carbohydrate. Fish Physiology and Biochemistry, 44(4):1127–1141, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0501-5.

Dai:2018:MCA

[1207] Yong-Jun Dai, Wen-Bin Liu, and Guang-Zhen Jiang. Molecular cloning of adipose triglyceride lipase (ATGL) gene from blunt snout bream and its expression after LPS-induced TNF-α factor. Fish Physiology and Biochemistry, 44(4):1143–1157, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0502-4.

Cai:2018:LSB

[1208] Wan-Cun Cai, Wen-Bin Liu, and Xiang-Fei Li. Lysine supplement benefits the growth performance, protein synthesis, and muscle development of Megalobrama amblycephala fed diets with fish meal replaced by rice protein concentrate. Fish Physiology and Biochemistry, 44(4): 1159–1174, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0503-3.

Sun:2018:ERC

[1209] Yaling Sun, Yingwen Li, and Qiliang Chen. Environmentally relevant concentrations of mercury exposure alter thyroid hormone levels and gene expression in the hypothalamic–pituitary–thyroid axis of zebrafish larvae. Fish Physiology and Biochemistry, 44(4):1175–1183, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0504-2.

Li:2018:EAH

[1210] Songlin Li, Chunyan Sang, and Xuxiong Huang. Effects of acute hyperglycemia stress on plasma glucose, glycogen content, and expressions of glycogen synthase and phosphorylase in hybrid grouper (Epinephelus fuscoguttatus [female sign] × E. lanceolatus [male sign]). Fish Physiology and Biochemistry, 44(4):1185–1196, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0508-y.

Pumputis:2018:RRT

[1211] Patrick G. Pumputis, Vivian R. Dayeh, and Niels C. Bols. Responses of rainbow trout intestinal epithelial cells to different kinds of nutritional deprivation. Fish Physiology and Biochemistry, 44(4):1197–1214, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-018-0511-3.

Oclon:2018:PBA

[1212] Ewa Ocłoń, Gili Solomon, and Arieh Gertler. Preparation of biologically active monomeric recombinant zebrafish (Danio rerio) and rainbow trout (Oncorhynchus mykiss) recombinant growth hormones. Fish Physiology and Biochemistry, 44(4):1215–1222, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0512-2.

Mansour:2018:EDI

[1213] Abdallah Tageldein Mansour, Liang Miao, and M. Ángeles Esteban. Effects of dietary inclusion of Moringa oleifera leaves on growth and some systemic and mucosal immune parameters of seabream. Fish Physiology and Biochemistry, 44(4):1223–1240, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0515-z.

AbouShabana:2018:EBH

[1214] N. M. AbouShabana, R. AbdelKader, and A. M. Abdel-Galil. Enhancement of broodstock health and maternal immunity in gilthead seabream (Sparus aurata L.) using ExcelMOS(R). Fish Physiology and Biochemistry, 44(4):1241–1251, August 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0517-x.

Souza:2018:EDM

[1215] Carine de Freitas Souza, Cristine Rampelotto, and Bernardo Baldisserotto. Effects of dietary microencapsulated Cymbopogon flexuosus essential oil on reproductive-related parameters in male Rhamdia quelen. Fish Physiology and Biochemistry, 44(4):1253–1264, August 2018. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0518-9.

Li:2018:ECP

[1216] Hongxia Li, Xiao Meng, and Jiting Wang. Effects of chromium picolinate supplementation on growth, body composition, and biochemical parameters in Nile tilapia Oreochromis niloticus. Fish Physiology and Biochemistry, 44(5):1265–1274, October 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0514-0.

Yan:2018:IGI

[1217] Hongwei Yan, Xufang Shen, and Yusheng Jiang. Identification of genes involved in gonadal sex differentiation and the dimorphic expression pattern in Takifugu rubripes gonad at the early stage of sex differentiation. Fish Physiology and Biochemistry, 44(5):1275–1290, October 2018. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0519-8.

Satora:2018:PEH

[1218] Leszek Satora, Jennifer Mytych, and Anna Bilska-Kos. The presence and expression of the HIF-1α in the respiratory intestine of the bronze Corydoras Corydoras aeneus (Callichthyidae Teleostei). Fish Physiology and Biochemistry, 44(5):1291–1297, October 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0520-2.

Arsyad:2018:EOL

[1219] Muh Ali Arsyad, Takashi Akazawa, and Masahiro Ogawa. Effects of olive leaf powder supplemented to fish feed on muscle protein of red sea bream. Fish Physiology and Biochemistry, 44(5):1299–1308, October 2018. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0521-1.

Shahjahan:2018:IWT

[1220] Md. Shahjahan, Md. Helal Uddin, and Md. Mahfuzul Haque. Increased water temperature altered hemato-biochemical parameters and structure of peripheral erythrocytes in striped catfish Pangasianodon hypophthalmus. Fish Physiology and Biochemistry, 44(5):1309–1318, October 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0522-0.

Gonzalez-Felix:2018:RDE

[1221] Mayra L. González-Félix, Delbert M. Gatlin III, and Michael Hume. Red drum Sciaenops ocellatus growth and expression of bile salt-dependent lipase in response to increasing dietary lipid supplementation. Fish Physiology and Biochemistry, 44(5):1319–1331, October 2018. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-018-0523-z.

Liu:2018:TAD

[1222] Wei Liu, Xiaoxu Liu, and Lihua Jiang. Transcriptome analysis demonstrates that long noncoding RNA is involved in the hypoxic response in Larimichthys crocea. Fish Physiology and Biochemistry, 44(5):1333– 1347, October 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0525-x.

Maltez:2018:OSA

[1223] Lucas Campos Maltez, Luis André Luz Barbas, and Luciano Garcia. Oxidative stress and antioxidant responses in juvenile Brazilian flounder Paralichthys orbignyanus exposed to sublethal levels of nitrite. Fish Physiology and Biochemistry, 44(5):1349–1362, October 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0526-9.

Jafari:2018:GPP

[1224] Naghmeh Jafari, Bahram Falahatkar, and Mir Masoud Sajjadi. Growth performance and plasma metabolites in juvenile Siberian sturgeon Acipenser baerii (Brandt, 1869) subjected to various feeding strategies at different sizes. Fish Physiology and Biochemistry, 44(5):1363– 1374, October 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0527-8.

Terova:2018:NDT

[1225] Genciana Terova, Simona Rimoldi, and Giovanni Bernardini. Nanodelivery of trace minerals for marine fish larvae: influence on skeletal ossification, and the expression of genes involved in intestinal transport of minerals, osteoblast differentiation, and oxidative stress response. Fish Physiology and Biochemistry, 44(5):1375–1391, October 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0528-7.

Maugars:2018:EAT

[1226] Gersende Maugars, Marie-Chanteuse Manirafasha, and Jehan-Hervé Lignot. The effects of acute transfer to freshwater on ion transporters of the pharyngeal cavity in European seabass (Dicentrarchus labrax). Fish Physiology and Biochemistry, 44(5):1393–1408, October 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0529-6.

Alak:2018:NED

[1227] Gonca Alak, Arzu Ucar, and Hasan Türkez. Neuroprotective effects of dietary borax in the brain tissue of rainbow trout (Oncorhynchus mykiss) exposed to copper-induced toxicity. Fish Physiology and Biochemistry, 44 (5):1409–1420, October 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0530-0.

Neves:2018:EER

[1228] J. Das Neves, I. E. J. Barnhoorn, and G. M. Wagenaar. The effects of environmentally relevant concentrations of aldrin and methoxychlor on the testes and sperm of male Clarias gariepinus (Burchell, 1822) after short-term exposure. Fish Physiology and Biochemistry, 44(6): 1421–1434, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0474-4.

Cejko:2018:OSP

[1229] Beata Irena Cejko, Ákos Horváth, and Radosław Kajetan Kowalski. Optimisation of sodium and potassium concentrations and pH in the artificial seminal plasma of common carp Cyprinus carpio L. Fish Physiology and Biochemistry, 44(6):1435–1442, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0491-3.

Diogo:2018:EUM

[1230] Patrícia Diogo, Gil Martins, and Elsa Cabrita. Electric ultrafreezer (−150^{o}c) as an alternative for zebrafish sperm cryopreservation and storage. Fish Physiology and Biochemistry, 44(6):1443–1455, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0500-6.

Gallego:2018:SOA

[1231] V. Gallego, J. G. Herranz-Jusdado, and Juan F. Asturiano. Subjective and objective assessment of fish sperm motility: when the technique and technicians matter. Fish Physiology and Biochemistry, 44(6): 1457–1467, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0505-1.

deSiqueira-Silva:2018:BAF

[1232] Diógenes Henrique de Siqueira-Silva, Taiju Saito, and George Shigueki Yasui. Biotechnology applied to fish reproduction: tools for conservation. Fish Physiology and Biochemistry, 44(6):1469–1485, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0506-0.

Lujic:2018:IGC

[1233] Jelena Lujić, Zoran Marinović, and Ákos Horváth. Interspecific germ cell transplantation: a new light in the conservation of valuable Balkan trout genetic resources? Fish Physiology and Biochemistry, 44(6):1487– 1498, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0510-4.

Kasa:2018:DSV

[1234] Eszter Kása, Jelena Lujić, and Ákos Horváth. Development of sperm vitrification protocols for two endangered salmonid species: the Adriatic grayling, Thymallus thymallus, and the marble trout, Salmo marmoratus. Fish Physiology and Biochemistry, 44(6):1499–1507, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0516-y.

Yilmaz:2018:MVZ

[1235] Ozlem Yilmaz, Amélie Patinote, and Julien Bobe. Multiple vitellogenins in zebrafish (Danio rerio): quantitative inventory of genes, transcripts and proteins, and relation to egg quality. Fish Physiology and Biochemistry, 44(6):1509–1525, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0524-y.

Xin:2018:PRA

[1236] Miaomiao Xin, Jan Sterba, and Otomar Linhart. Protective role of antifreeze proteins on sterlet (Acipenser ruthenus) sperm during cryopreservation. Fish Physiology and Biochemistry, 44(6):1527–1533, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-018-0538-5.

Psenicka:2018:RAN

[1237] Martin Pšenička. Rapid de-adhesion of northern pike Esox lucius eggs using sodium hypochlorite. Fish Physiology and Biochemistry, 44(6): 1535–1539, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0544-7.

Iegorova:2018:FEV

[1238] Viktoriia Iegorova, Milos Havelka, and Taiju Saito. First evidence of viable progeny from three interspecific parents in sturgeon. Fish Physiology and Biochemistry, 44(6):1541–1550, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0553-6.

Momin:2018:SQA

[1239] Momin Momin and Devrim Memiş. Sperm quality analysis of normal season (NG) and out-season by photoperiod manipulation (PG) of male rainbow trout broodstock (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 44(6):1551–1560, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0564-3.

Carnevali:2018:EDC

[1240] Oliana Carnevali, Stefania Santangeli, and Francesca Maradonna. Endocrine-disrupting chemicals in aquatic environment: what are the risks for fish gametes? Fish Physiology and Biochemistry, 44(6):1561– 1576, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0507-z.

Kollar:2018:VTT

[1241] Tímea Kollár, Eszter Kása, and Ákos Horváth. In vitro toxicology test system based on common carp (Cyprinus carpio) sperm analysis. Fish Physiology and Biochemistry, 44(6):1577–1589, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0541-x.

Bondarenko:2018:CSM

[1242] Volodymyr Bondarenko, Miroslav Blecha, and Tomas Policar. Changes of sperm morphology, volume, density, and motility parameters in northern pike during the spawning period. Fish Physiology and Biochemistry, 44(6):1591–1597, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0546-5.

Maouche:2018:NIE

[1243] Ahmed Maouche, Edouard Curran, and Jean-Jacques Lareyre. New insights into the evolution, hormonal regulation, and spatiotemporal expression profiles of genes involved in the Gfra1/Gdnf and Kit/Kitlg regulatory pathways in rainbow trout testis. Fish Physiology and Biochemistry, 44(6):1599–1616, December 2018. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0547-4.

Senarat:2019:SGS

[1244] Sinlapachai Senarat, Jes Kettratad, and Sasipong Tipdomrongpong. The sbGnRH–GTH system in the female short mackerel, Rastrelliger brachysoma (Bleeker, 1851), during breeding season: implications for low gamete production in captive broodstock. Fish Physiology and Biochemistry, 45(1):1–18, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0509-x.

Senarat:2019:CSG

[1245] Sinlapachai Senarat, Jes Kettratad, and Sasipong Tipdomrongpong. Correction to: The sbGnRH–GTH system in the female short mackerel, Rastrelliger brachysoma (Bleeker, 1851), during breeding season: implications for low gamete production in captive broodstock. Fish Physiology and Biochemistry, 45(1):19, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0549-2.

Liu:2019:EDM

[1246] Wei Liu, Xing Lu, and Hua Wen. Effects of dietary manipulation on compensatory growth of juvenile genetically improved farmed tilapia (Oreochromis niloticus). Fish Physiology and Biochemistry, 45(1):21– 32, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-018-0531-z.

Faheem:2019:MBK

[1247] Mehwish Faheem, Nusrat Jahan, and Khalid Parvez Lone. Modulation of brain kisspeptin expression after bisphenol-A exposure in a teleost fish, Catla catla. Fish Physiology and Biochemistry, 45(1): 33–42, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0532-y.

Li:2019:EEE

[1248] HuaTao Li, SiYi Tang, and SiShun Zhou. The effects of ethoxyquin and Angelica sinensis extracts on lipid oxidation in fish feeds and growth, digestive and absorptive capacities and antioxidant status in juvenile red carp (Cyprinus carpio var. xingguonensis): a comparative study. Fish Physiology and Biochemistry, 45(1):43–61, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0533-x.

Souza:2019:OSM

[1249] Carine F. Souza, Matheus D. Baldissera, and Aleksandro S. da Silva. Oxidative stress mediated the inhibition of cerebral creatine kinase activity in silver catfish fed with aflatoxin B_{1}-contaminated diet. Fish Physiology and Biochemistry, 45(1):63–70, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0534-9.

Hamed:2019:AAE

[1250] Heba S. Hamed and Yasser S. El-Sayed. Antioxidant activities of Moringa oleifera leaf extract against pendimethalin-induced oxidative stress and genotoxicity in Nile tilapia, Oreochromis niloticus (L.). Fish Physiology and Biochemistry, 45(1):71–82, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0535-8.

Zhou:2019:EBG

[1251] Wenhao Zhou, Samad Rahimnejad, and Wenbin Liu. Effects of berberine on growth, liver histology, and expression of lipid-related genes in blunt snout bream (Megalobrama amblycephala) fed high-fat diets. Fish Physiology and Biochemistry, 45(1):83–91, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0536-7.

Wang:2019:IET

[1252] Jun Wang, Xinying Hou, and Zhou Yang. Interactive effects of temperature and salinity on the survival, oxidative stress, and Na^{+} /K^{+}-ATPase activity of newly hatched obscure puffer (Takifugu obscurus) larvae. Fish Physiology and Biochemistry, 45(1):93–103, February 2019. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-018-0537-6.

Rashid:2019:ESP

[1253] Ibrahim Rashid, Md Shakhawate Hossain, and S. M. Rafiquzzaman. Evaluation of seminal plasma composition and spermatozoa quality parameters of silver barb, Barbonymus gonionotus Bleeker, 1850. Fish Physiology and Biochemistry, 45(1):105–114, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0539-4.

Zhang:2019:MLM

[1254] Muzi Zhang, Chengdong Hou, and Rixin Wang. Modulation of lipid metabolism in juvenile yellow catfish (Pelteobagrus fulvidraco) as affected by feeding frequency and environmental ammonia. Fish Physiology and Biochemistry, 45(1):115–122, February 2019. CODEN FPBIEP.

ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0540-y.

Kumar:2019:ECC

[1255] Raj Kumar, Charan Ravi, and Thangaraj Raja Swaminathan. Establishment and characterization of a caudal fin-derived cell line, AOF, from the Oscar, Astronotus ocellatus. Fish Physiology and Biochemistry, 45 (1):123–131, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0542-9.

Chen:2019:IMF

[1256] Xiaowu Chen, Yawei Shen, and Jinliang Zhao. Irf3 from mandarin fish thymus initiates interferon transcription. Fish Physiology and Biochemistry, 45(1):133–144, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0543-8.

Parolini:2019:ASD

[1257] Marco Parolini, Rocco Iacobuzio, and Nicola Saino. Age- and sexdependent variation in the activity of antioxidant enzymes in the brown trout (Salmo trutta). Fish Physiology and Biochemistry, 45(1):145– 154, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0545-6.

Souza:2019:IHA

[1258] Carine de Freitas Souza, Sharine Descovi, and Gonzalo Martinez-Rodríguez. Involvement of HPI-axis in anesthesia with Lippia alba essential oil citral and linalool chemotypes: gene expression in the secondary responses in silver catfish. Fish Physiology and Biochemistry, 45 (1):155–166, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0548-3.

Prado:2019:EPI

[1259] Paula Suzanna Prado, Ana Paula Barbosa Pinheiro, and Elizete Rizzo. Expression patterns and immunolocalisation of IGF-i and IGF-II in male and female gonads of the neotropical characid fish Astyanax fasciatus. Fish Physiology and Biochemistry, 45(1):167–176, February 2019. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0550-9.

Du:2019:CST

[1260] Jinxing Du, Xiaowen Chen, and Chenghui Wang. Comparative skin transcriptome of two Oujiang color common carp (Cyprinus carpio var. color) varieties. Fish Physiology and Biochemistry, 45(1):177–185, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-018-0551-8.

Yuan:2019:EPR

[1261] Xiang-Yang Yuan, Guang-Zhen Jiang, and Wen-Bin Liu. Effects of partial replacement of fish meal by yeast hydrolysate on antioxidant capability, intestinal morphology, and inflammation-related gene expression of juvenile Jian carp (Cyprinus carpio var. Jian). Fish Physiology and Biochemistry, 45(1):187–197, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0552-7.

Ma:2019:IEC

[1262] Rong-Rong Ma, Jing Sun, and Kun Hu. Identification of Carassius auratus gibelio liver cell proteins interacting with the GABA_{A} receptor γ 2 subunit using a yeast two-hybrid system. Fish Physiology and Biochemistry, 45(1):199–208, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0554-5.

Nofouzi:2019:BEK

[1263] Katayoon Nofouzi, Najmeh Sheikhzadeh, and Amir Ali Shahbazfar. Beneficial effects of killed Tsukamurella inchonensis on rainbow trout (Oncorhynchus mykiss) growth, intestinal histology, immunological, and biochemical parameters. Fish Physiology and Biochemistry, 45(1):209–217, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-018-0555-4.

Dawood:2019:DSS

[1264] Mahmoud A. O. Dawood, Shunsuke Koshio, and Mohamed S. Hassaan. Dietary supplementation of selenium nanoparticles modulated systemic and mucosal immune status and stress resistance of red sea bream (Pagrus major). Fish Physiology and Biochemistry, 45(1):219– 230, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0556-3.

Liu:2019:AHI

[1265] Jun Liu, Juan juan Pang, and Guang xian Liu. The accumulation, histopathology, and intestinal microorganism effects of waterborne cadmium on Carassius auratus gibelio. Fish Physiology and Biochemistry, 45 (1):231–243, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0557-2.

Jia:2019:MIE

[1266] Yongyi Jia, Jianbo Zheng, and Liqiao Chen. Molecular identification of dmrt1 and its promoter CpG methylation in correlation with gene expression during gonad development in Culter alburnus. Fish Physiology and Biochemistry, 45(1):245–252, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0558-1.

Bermejo-Poza:2019:PMR

[1267] Rubén Bermejo-Poza, Montserrat Fernández-Muela, and Morris Villarroel. Physio-metabolic response of rainbow trout during prolonged food deprivation before slaughter. Fish Physiology and Biochemistry, 45(1): 253–265, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0559-0.

Xu:2019:PTR

[1268] Wenjie Xu, Junyan Jin, and Shouqi Xie. Physiological and transcriptomic responses to fishmeal-based diet and rapeseed meal-based diet in two strains of gibel carp (Carassius gibelio). Fish Physiology and Biochemistry, 45(1):267–286, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0560-7.

Zhou:2019:GPL

[1269] Ji Shu Zhou, Pan Guo, and Yi An Chen. Growth performance, lipid metabolism, and health status of grass carp (Ctenopharyngodon idella) fed three different forms of sodium butyrate. Fish Physiology and Biochemistry, 45(1):287–298, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0561-6.

Yu:2019:PVM

[1270] Jie Yu, Yuanshuai Fu, and Zhiyi Shi. Proteomic variation in metamorphosing Paralichthys olivaceus induced by exogenous thyroid hormone.

Fish Physiology and Biochemistry, 45(1):299–309, February 2019. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0562-5.

Liao:2019:MCT

[1271] Shengchen Liao, Jingjing Dong, and Jun Xie. Molecular cloning, tissue distribution, and pharmacological characterization of blunt snout bream (Megalobrama amblycephala) melanocortin-5 receptor. Fish Physiology and Biochemistry, 45(1):311–321, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0563-4.

Li:2019:IEM

[1272] Xinru Li, Tao Wang, and Wenxu Zhu. The improved energy metabolism and blood oxygen-carrying capacity for pufferfish, Takifugu fasciatus, against acute hypoxia under the regulation of oxygen sensors. Fish Physiology and Biochemistry, 45(1):323–340, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0565-2.

Venturini:2019:MBA

[1273] F. P. Venturini, F. D. de Moraes, and G. Moraes. A multi-biomarker approach to lambda-cyhalothrin effects on the freshwater teleost matrinxa Brycon amazonicus: single-pulse exposure and recovery. Fish Physiology and Biochemistry, 45(1):341–353, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0566-1.

Bai:2019:MFG

[1274] Yan Bai, Chengdong Liu, and Hongying Wang. Molecular, functional, and gene expression analysis of zebrafish ror1 receptor. Fish Physiology and Biochemistry, 45(1):355–363, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0567-0.

Jiang:2019:HPA

[1275] Yu Jiang, Bin Chen, and Guo-Xing Zhu. Hepcidin protects against iron overload-induced inhibition of bone formation in zebrafish. Fish Physiology and Biochemistry, 45(1):365–374, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0568-z.

Wang:2019:MCE

[1276] Bin Wang, Yongjiang Xu, and Bao Shi. Molecular characterization and expression profiles of insulin-like growth factors in yellowtail kingfish (Seriola lalandi) during embryonic development. Fish Physiology and Biochemistry, 45(1):375–390, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0570-5.

Song:2019:EBL

[1277] Jin Ah Song and Cheol Young Choi. Effects of blue light spectra on retinal stress and damage in goldfish (Carassius auratus). Fish Physiology and Biochemistry, 45(1):391–400, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0571-4.

Jiang:2019:GPD

[1278] Guang-Zhen Jiang, Hua-Juan Shi, and Xiang-Fei Li. Glucose-6-phosphate dehydrogenase in blunt snout bream Megalobrama amblycephala: molecular characterization, tissue distribution, and the responsiveness to dietary carbohydrate levels. Fish Physiology and Biochemistry, 45(1):401–415, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0572-3.

Barbieri:2019:EMC

[1279] Edison Barbieri, Alessandra Maria Tegon Ferrarini, and Oswaldo Luiz Alves. Effects of multiwalled carbon nanotubes and carbofuran on metabolism in Astyanax ribeirae, a native species. Fish Physiology and Biochemistry, 45(1):417–426, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0573-2.

Chen:2019:MCE

[1280] Wenbo Chen, Fangfang Yan, and Haiyan Dong. Molecular cloning, expression analysis, and the immune-related role of a thymosin β in the goldfish, Carassius auratus. Fish Physiology and Biochemistry, 45(1): 427–437, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0574-1.

Sony:2019:EDF

[1281] Nadia Mahjabin Sony, Manabu Ishikawa, and Saichiro Yokoyama. The effect of dietary fucoidan on growth, immune functions, blood characteristics and oxidative stress resistance of juvenile red sea bream, Pagrus major. Fish Physiology and Biochemistry, 45(1):439–454, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0575-0.

Mattioli:2019:PMR

[1282] Cristiano Campos Mattioli, Rodrigo Takata, and Ronald Kennedy Luz. Physiological and metabolic responses of juvenile Lophiosilurus alexandri catfish to air exposure. Fish Physiology and Biochemistry, 45(1): 455–467, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0576-z.

Arantzamendi:2019:LDC

[1283] Leire Arantzamendi, Francisco Roo, and Marisol Izquierdo. Lipid digestion capacity in gilthead seabream (Sparus aurata) from first feeding to commercial size. Fish Physiology and Biochemistry, 45(1): 469–484, February 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0577-y.

Allais:2019:NWT

[1284] Laetitia Allais, Chao Zhao, and Zhenhua Ma. Nutrition and water temperature regulate the expression of heat-shock proteins in golden pompano larvae (Trachinotus ovata, limmaeus 1758). Fish Physiology and Biochemistry, 45(2):485–497, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0578-x.

Jimenez:2019:HDC

[1285] Ana Gabriela Jimenez, Evan Braun, and Kailey Tobin. How does chronic temperature exposure affect hypoxia tolerance in sheepshead minnows’ (Cyprinodon variegatus variegatus) ability to tolerate oxidative stress? Fish Physiology and Biochemistry, 45(2):499–510, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0583-0.

Boguslawska-Was:2019:RER

[1286] Elżbieta Bogusławska-Was, Alicja Dłubała, and Maria Laskowska. The role of Rhodotorula mucilaginosa in selected biological process of wild fish. Fish Physiology and Biochemistry, 45(2):511–521, April 2019. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0591-0.

Khieokhajonkhet:2019:MCH

[1287] Anurak Khieokhajonkhet, Niran Aeksiri, and Gen Kaneko. Molecular characterization and homology modeling of liver X receptor in Asian seabass, Lates calcarifer: predicted functions in reproduction and lipid metabolism. Fish Physiology and Biochemistry, 45(2):523– 538, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-019-00617-6.

Chen:2019:AAG

[1288] Jiaojiao Chen, Dianfu Zhang, and Pengcheng Hu. Arginine affects growth and integrity of grass carp enterocytes by regulating TOR signaling pathway and tight junction proteins. Fish Physiology and Biochemistry, 45 (2):539–549, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00613-w.

Baldissera:2019:CMB

[1289] Matheus D. Baldissera, Carine F. Souza, and Bernardo Baldisserotto. Caffeine modulates brain purinergic signaling in Nile tilapia (Oreochromis niloticus) under hypoxia conditions: improvement of immune and inflammatory responses. Fish Physiology and Biochemistry, 45(2):551– 560, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-018-0592-z.

Kanno:2019:TTF

[1290] Gaku Kanno, Sappasith Klomklao, and Hideki Kishimura. A thermostable trypsin from freshwater fish Japanese dace (Tribolodon hakonensis): a comparison of the primary structures among fish trypsins. Fish Physiology and Biochemistry, 45(2):561–571, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0600-3.

Domingues:2019:ETD

[1291] P. Domingues, J. Hernández-Urcera, and P. Gallardo. Effect of triploidy on digestive enzyme activity of early stages of turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 45(2):573–582, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00610-z.

Nie:2019:NMT

[1292] Miaomiao Nie, Xungang Tan, and Feng You. Network of microRNAtranscriptional factor-mRNA in cold response of turbot Scophthalmus maximus. Fish Physiology and Biochemistry, 45(2):583–597, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00611-y.

Hekmatpour:2019:CSB

[1293] Fatemeh Hekmatpour, Preeta Kochanian, and Seyed-Mohammad Mousavi. Changes in serum biochemical parameters and digestive enzyme activity of juvenile sobaity sea bream (Sparidentex hasta) in response to partial replacement of dietary fish meal with poultry by-product meal. Fish Physiology and Biochemistry, 45(2):599–611, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00619-4.

Hossain:2019:FRS

[1294] Md. Sakhawat Hossain, Shunsuke Koshio, and Md. Jakiul Islam. Fishmeal replacement by soya protein concentrate with inosine monophosphate supplementation influences growth, digestibility, immunity, blood health, and stress resistance of red sea bream, Pagrus major. Fish Physiology and Biochemistry, 45(2):613–629, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0581-2.

Tian:2019:LAG

[1295] Jing jing Tian, Hong Ji, and Wang bao Gong. Lipid accumulation in grass carp (Ctenopharyngodon idellus) fed faba beans (Vicia faba L.). Fish Physiology and Biochemistry, 45(2):631–642, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0589-7.

Zhong:2019:DSE

[1296] Xiao-Qun Zhong, Ming-Yang Liu, and Xiang-Fei Li. Dietary supplementation of Streptococcus faecalis benefits the feed utilization, antioxidant capability, innate immunity, and disease resistance of blunt snout bream (Megalobrama amblycephala). Fish Physiology and Biochemistry, 45(2):643–656, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0595-9.

Feng:2019:MCE

[1297] Xiu Feng, Xiaomu Yu, and Jingou Tong. Molecular characterization and expression regulation of the factor-inhibiting HIF-1 (FIH-1) gene under hypoxic stress in bighead carp (Aristichthys nobilis). Fish Physiology and Biochemistry, 45(2):657–665, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0597-7.

Kumar:2019:ODT

[1298] Aditya Kumar, P. K. Pradhan, and E. Gisbert. Ontogeny of the digestive tract in stinging catfish, Heteropneustes fossilis (Bloch) larvae. Fish Physiology and Biochemistry, 45(2):667–679, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00618-5.

Castro:2019:VOC

[1299] Carolina Castro, Ana Couto, and Aires Oliva-Teles. Vegetable oil and carbohydrate-rich diets marginally affected intestine histomorphology, digestive enzymes activities, and gut microbiota of gilthead sea bream juveniles. Fish Physiology and Biochemistry, 45(2):681–695, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0579-9.

Betancor:2019:PFU

[1300] Mónica B. Betancor, Aurelio Ortega, and Gabriel Mourente. Performance, feed utilization, and hepatic metabolic response of weaned juvenile Atlantic bluefin tuna (Thunnus thynnus L.): effects of dietary lipid level and source. Fish Physiology and Biochemistry, 45 (2):697–718, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0587-9.

Liang:2019:DLA

[1301] Hualiang Liang, Ahmed Mokrani, and Ajun Sun. Dietary leucine affects glucose metabolism and lipogenesis involved in TOR/PI3K/Akt signaling pathway for juvenile blunt snout bream Megalobrama amblycephala. Fish Physiology and Biochemistry, 45(2):719–732, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0594-x.

Teles:2019:CDE

[1302] Andressa Teles, Joan Salas-Leiva, and Dariel Tovar-Ramírez. Changes in digestive enzyme activities during early ontogeny of Seriola rivoliana.

Fish Physiology and Biochemistry, 45(2):733–742, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0598-6.

Xu:2019:IPD

[1303] Zhan-Ning Xu, Guo-Dong Zheng, and Shu-Ming Zou. Identification of proteins differentially expressed in the gills of grass carp (Ctenopharyngodon idella) after hypoxic stress by two-dimensional gel electrophoresis analysis. Fish Physiology and Biochemistry, 45(2):743–752, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0599-5.

Amano:2019:LTF

[1304] Masafumi Amano, Noriko Amiya, and Yoshitaka Sakakura. Localization of three forms of gonadotropin-releasing hormone in the brain and pituitary of the self-fertilizing fish, Kryptolebias marmoratus. Fish Physiology and Biochemistry, 45(2):753–771, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0601-2.

Aliper:2019:UFS

[1305] Alexey T. Aliper, Alisa A. Zaichikova, and Elena M. Maximova. Updated functional segregation of retinal ganglion cell projections in the tectum of a cyprinid fish — further elaboration based on microelectrode recordings. Fish Physiology and Biochemistry, 45(2):773–792, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0603-0.

Kohshahi:2019:DCC

[1306] Ayyoub Jamali Kohshahi, Iman Sourinejad, and Seyed Ali Johari. Dietary cosupplementation with curcumin and different selenium sources (nanoparticulate, organic, and inorganic selenium): influence on growth performance, body composition, immune responses, and glutathione peroxidase activity of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 45(2):793–804, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0585-y.

Xu:2019:CEM

[1307] Jing Xu, Fuyuan Hou, and Guangzhong Yang. Characterization and expression of melanin-concentrating hormone (MCH) in common carp (Cyprinus carpio) during fasting and reproductive cycle. Fish Physiology and Biochemistry, 45(2):805–817, April 2019. CODEN FPBIEP.

ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0586-x.

Nakatsuji:2019:LTS

[1308] Nobuyoshi Nakatsuji, Kohsuke Adachi, and Katsuji Morioka. Longterm stability of RNA isolated from muscle of red seabream (Pagrus major) during ice storage. Fish Physiology and Biochemistry, 45(2): 819–828, April 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0588-8.

Gao:2019:MVG

[1309] Xin-Ming Gao, Yang Zhou, and Jun-Quan Zhu. Multiple vitellogenin genes (vtgs) in large yellow croaker (Larimichthys crocea): molecular characterization and expression pattern analysis during ovarian development. Fish Physiology and Biochemistry, 45(3):829–848, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0569-y.

Martins:2019:MFA

[1310] Elói Martins, Pedro R. Almeida, and Maria João Lança. Muscle fatty acid profiles of sea lamprey (Petromyzon marinus L.) indicate the use of fast metabolized energy during ontogenesis. Fish Physiology and Biochemistry, 45(3):849–862, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0580-3.

Chao:2019:MCE

[1311] Yan Chao, Mingzhe Xia, and Delin Qi. Molecular characterization and expression changes of cytoglobin genes in response to hypoxia in a Tibetan schizothoracine fish, Schizopygopsis pylzovi. Fish Physiology and Biochemistry, 45(3):863–872, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0582-1.

Lee:2019:BAI

[1312] Jin Wuk Lee, Jae-Woo Lee, and Kyunghwa Park. n-butyl acrylateinduced antioxidant system alteration through two generations in Oryzias latipes. Fish Physiology and Biochemistry, 45(3):873–883, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-018-0584-z.

Rawat:2019:SLV

[1313] Arpana Rawat, Radha Chaube, and Keerrikkattil P. Joy. In situ localization of vasotocin receptor gene transcripts in the brain-pituitary-gonadal axis of the catfish Heteropneustes fossilis: a morpho-functional study. Fish Physiology and Biochemistry, 45(3):885–905, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-018-0590-1.

Kumari:2019:HAS

[1314] Suman Kumari, Mahua G. Choudhury, and Nirmalendu Saha. Hyperammonia stress causes induction of inducible nitric oxide synthase gene and more production of nitric oxide in air-breathing magur catfish, Clarias magur (Hamilton). Fish Physiology and Biochemistry, 45 (3):907–920, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0593-y.

Zhou:2019:TFC

[1315] Chaowei Zhou, Luo Lei, and Zonglin Zheng. Three forms of cocaine- and amphetamine-regulated transcript may be involved in food intake regulation in gibel carp (Carassius auratus gibelio). Fish Physiology and Biochemistry, 45(3):921–933, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-018-0596-8.

Li:2019:EAE

[1316] Li Li, Qian Xiao, and Zhongjie Chang. Expression analysis of And4 during fin regeneration in Misgurnus anguillicaudatus provides insights into its function. Fish Physiology and Biochemistry, 45(3):935–942, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-018-0602-1.

Mania:2019:EEH

[1317] Manuela Mania, Giuseppe Bruschetta, and Salvatore Campo. Evidence for embryonic haemoglobins from Sparus aurata under normal and hypoxic conditions. Fish Physiology and Biochemistry, 45 (3):943–954, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0605-y.

Jia:2019:IEG

[1318] Yudong Jia, Qiqi Jing, and Bin Huang. Involvement and expression of growth hormone/insulin-like growth factor member mRNAs in the ovarian development of turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 45(3):955–964, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-018-0604-z.

Wang:2019:ICE

[1319] Dan Wang, Quanquan Cao, and Tao Wang. Individual and combined effects of salinity and lipopolysaccharides on the immune response of juvenile Takifugu fasciatus. Fish Physiology and Biochemistry, 45 (3):965–976, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-018-0607-9.

Teimouri:2019:EES

[1320] Mahdi Teimouri, Sakineh Yeganeh, and Soleiman Mahjoub. The effect of Spirulina platensis meal on antioxidant gene expression, total antioxidant capacity, and lipid peroxidation of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 45(3):977–986, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-0608-3.

Yang:2019:ESR

[1321] Song Yang, Kuo He, and Quan Gong. Effect of starvation and refeeding on oxidative stress and antioxidant defenses in Yangtze sturgeon (Acipenser dabryanus). Fish Physiology and Biochemistry, 45(3):987– 995, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-0609-2.

Abdel-Tawwab:2019:FRH

[1322] Mohsen Abdel-Tawwab, Mohamed N. Monier, and Caterina Faggio. Fish response to hypoxia stress: growth, physiological, and immunological biomarkers. Fish Physiology and Biochemistry, 45(3):997–1013, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00614-9.

Li:2019:NRP

[1323] Ruixin Li, Hongyu Liu, and Liqiao Chen. Nutritional regulation of pyruvate kinase and phosphoenolpyruvate carboxykinase at the enzymatic and molecular levels in cobia Rachycentron canadum. Fish Physiology and Biochemistry, 45(3):1015–1028, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00612-x.

Rahimi:2019:HNM

[1324] Ruhollah Rahimi, Hamid Farahmand, and Amirreza Abed-Elmdoust. ^{1} H NMR metabolic profiling of the cryopreserved spermatozoa of the wild endangered Persian sturgeon (Acipenser persicus) with the use of betacyclodextrin as an external cryoprotectant. Fish Physiology and Biochemistry, 45(3):1029–1040, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-019-00615-8.

Baldissera:2019:CSD

[1325] Matheus D. Baldissera, Carine F. Souza, and Bernardo Baldisserotto. A caffeine-supplemented diet modulates oxidative stress markers and prevents oxidative damage in the livers of Nile tilapia (Oreochromis niloticus) exposed to hypoxia. Fish Physiology and Biochemistry, 45 (3):1041–1049, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00616-7.

Ota:2019:HGA

[1326] Erika do Carmo Ota, Claucia Aparecida Honorato, and Claudia Andrea Lima Cardoso. Hepatic and gastroprotective activity of Serjania marginata leaf aqueous extract in Nile tilapia (Oreochromis niloticus). Fish Physiology and Biochemistry, 45(3):1051–1065, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00622-9.

Valente:2019:DME

[1327] Luísa Maria Pinheiro Valente, Marco Custódio, and Viswanath Kiron. Defatted microalgae (Nannochloropsis sp.) from biorefinery as a potential feed protein source to replace fishmeal in European sea bass diets. Fish Physiology and Biochemistry, 45(3):1067–1081, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00621-w.

Muttray:2019:ERS

[1328] Annette F. Muttray, Edward M. Donaldson, and Robert H. Devlin. Effects of recombinant salmon type II growth hormone and bovine growth hormone on growth of Coho salmon (Oncorhynchus kisutch). Fish Physiology and Biochemistry, 45(3):1083–1090, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00620-x.

Jin:2019:AED

[1329] Ai Jin, Xiao chen Shi, and Hong Ji. Ameliorative effect of docosahexaenoic acid on hepatocyte apoptosis and inflammation induced by oleic acid in grass carp, Ctenopharyngodon idella. Fish Physiology and Biochemistry, 45(3):1091–1099, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-019-00623-8.

Lan:2019:CEA

[1330] Tian Lan, Yu-Long Chen, and Ze-Xia Gao. Comparative expression analysis of let-7 microRNAs during ovary development in Megalobrama amblycephala. Fish Physiology and Biochemistry, 45(3):1101– 1115, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-019-00624-7.

Zhang:2019:IFG

[1331] Dian-Guang Zhang, Jie Cheng, and Zhi Luo. Identification of five genes in endoplasmic reticulum (ER) stress–apoptosis pathways in yellow catfish Pelteobagrus fulvidraco and their transcriptional responses to dietary lipid levels. Fish Physiology and Biochemistry, 45(3):1117– 1127, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-019-00627-4.

Gao:2019:ECF

[1332] Ya Gao, Huihui Zhou, and Gen He. Establishment and characterization of a fibroblast-like cell line from the muscle of turbot (Scophthalmus maximus L.). Fish Physiology and Biochemistry, 45(3):1129– 1139, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-019-00628-3.

Guo:2019:IDE

[1333] Hong-Hong Guo, Yuan Sun, and Shu-Ming Zou. Identification of duplicated Cited3 genes and their responses to hypoxic stress in blunt snout bream (Megalobrama amblycephala). Fish Physiology and Biochemistry, 45(3):1141–1152, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00625-6.

Khan:2019:SAE

[1334] Nawaz Alam Khan, JaiGopal Sharma, and Rina Chakrabarti. The study of ameliorative effect of dietary supplementation of vitamin C, vitamin E, and tryptophan on Labeo rohita (Cyprinidae) fry exposed to intense light. Fish Physiology and Biochemistry, 45(3):1153–1165, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00626-5.

Penny:2019:LCS

[1335] F. M. Penny and J. D. Kieffer. Lack of change in swimming capacity (u_{crit}) following acute salinity exposure in juvenile shortnose sturgeon (Acipenser brevirostrum). Fish Physiology and Biochemistry, 45 (3):1167–1175, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00629-2.

Santos:2019:TLT

[1336] Thamiris Pinheiro Santos, Priscila Rafaela Leão Soares, and Pabyton Gonçalves Cadena. Thyroxine, levothyroxine, and thyroxine complexed into cyclodextrin changed animal behavior, oxygen consumption, and photopic electroretinogram of Colossoma macropomum. Fish Physiology and Biochemistry, 45(3):1177–1187, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00630-9.

Huang:2019:TTF

[1337] Ming-Chih Huang, Cheng-Linn Lee, and Shugo Watabe. Thermostability of tropomyosins from the fast skeletal muscles of tropical fish species. Fish Physiology and Biochemistry, 45(3):1189–1202, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00632-7.

Nurdalila:2019:MPE

[1338] A’wani Aziz Nurdalila, Yosmetha Mayalvanan, and Syarul Nataqain Baharum. Metabolite profiling of Epinephelus fuscoguttatus infected with vibriosis reveals Omega 9 as potential metabolite biomarker. Fish Physiology and Biochemistry, 45(3):1203–1215, June 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00633-6.

El-Leithy:2019:OSN

[1339] Ahmed A. A. El-Leithy, Shaaban A. Hemeda, and Zeinab A. Helmy. Optimum salinity for Nile tilapia (Oreochromis niloticus) growth and mRNA transcripts of ion-regulation, inflammatory, stress- and immunerelated genes. Fish Physiology and Biochemistry, 45(4):1217–1232, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00640-7.

Wang:2019:RSM

[1340] Na Wang, Renkai Wang, and Songlin Chen. RNA-seq and microRNAseq analysis of Japanese flounder (Paralichthys olivaceus) larvae treated by thyroid hormones. Fish Physiology and Biochemistry, 45(4):1233– 1244, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-019-00654-1.

Bernard:2019:TSM

[1341] Bernoît Bernard, Syaghalirwa N. M. Mandiki, and Patrick Kestemont. A temperature shift on the migratory route similarly impairs hypoosmoregulatory capacities in two strains of Atlantic salmon (Salmo salar L.) smolts. Fish Physiology and Biochemistry, 45(4):1245–1260, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00666-x.

Catteau:2019:ICF

[1342] Audrey Catteau, Antoine Le Guernic, and Jean-Marc Porcher. Impact of confinement and food access restriction on the three-spined stickleback (Gasterosteus aculeatus, L.) during caging: a multi-biomarker approach. Fish Physiology and Biochemistry, 45(4):1261–1276, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00670-1.

Safari:2019:DAF

[1343] Omid Safari, Mehrdad Sarkheil, and Marina Paolucci. Dietary administration of ferula (Ferula asafoetida) powder as a feed additive in diet of koi carp, Cyprinus carpio koi: effects on hemato-immunological parameters, mucosal antibacterial activity, digestive enzymes, and growth performance. Fish Physiology and Biochemistry, 45(4):1277–1288, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00674-x.

Souza:2019:AMZ

[1344] Jaqueline P. Souza, Adrislaine S. Mansano, and Valtencir Zucolotto. Antioxidant metabolism of zebrafish after sub-lethal exposure to graphene oxide and recovery. Fish Physiology and Biochemistry, 45(4):1289–1297, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00678-7.

Hechter:2019:RBS

[1345] Drake T. Hechter and Caleb T. Hasler. Repeatability of burst swimming performance in medaka (Oryzias latipes). Fish Physiology and Biochemistry, 45(4):1299–1307, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-019-00679-6.

Rapatsa:2019:EAH

[1346] Mmaditshaba M. Rapatsa and Ngonidzashe A. G. Moyo. Enzyme activity and histological analysis of Clarias gariepinus fed on Imbrasia belina meal used for partial replacement of fishmeal. Fish Physiology and Biochemistry, 45(4):1309–1320, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00652-3.

Shaalan:2019:ECM

[1347] Walaa M. Shaalan, Nassr Allah Abd El-Hameid, and Mohamed Salem. Expressions and characterization of MuRFs, Atrogin-1, F-box25 genes in tilapia, Oreochromis niloticus, in response to starvation. Fish Physiology and Biochemistry, 45(4):1321–1330, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00667-w.

Xu:2019:EMP

[1348] Mengmeng Xu, Tingting Wang, and Huiwen Sun. An evaluation of mixed plant protein in the diet of Yellow River carp (Cyprinus carpio): growth, body composition, biochemical parameters, and growth hormone/insulin-like growth factor 1. Fish Physiology and Biochemistry, 45 (4):1331–1342, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00641-6.

Wang:2019:AIW

[1349] Yunhao Wang, Guofeng Han, and Vishwajit S. Chowdhury. An acute increase in water temperature can increase free amino acid concentrations in the blood, brain, liver, and muscle in goldfish (Carassius auratus). Fish Physiology and Biochemistry, 45(4):1343–1354, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00642-5.

Nigam:2019:CCM

[1350] Ashwini Kumar Nigam, Neeraj Verma, and Ajay Kumar Mittal. Characterisation of cholinesterases in mucous secretions and their localisation in epidermis of Labeo rohita and Cirrhinus mrigala. Fish Physiology and Biochemistry, 45(4):1355–1366, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00663-0.

Sun:2019:HDF

[1351] Jingfeng Sun, Yize Wang, and Kezhi Xing. Histochemical distribution of four types of enzymes and mucous cells in the intestine of koi carp (Cyprinus carpio var. koi). Fish Physiology and Biochemistry, 45(4): 1367–1376, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00673-y.

Batalhao:2019:EMA

[1352] Isabela Gertrudes Batalhão, Daína Lima, and Eduardo Alves de Almeida. Effects of methylphenidate on the aggressive behavior, serotonin and dopamine levels, and dopamine-related gene transcription in brain of male Nile tilapia (Oreochromis niloticus). Fish Physiology and Biochemistry, 45(4):1377–1391, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-019-00645-2.

Li:2019:EMS

[1353] Zequn Li, Ning Bao, and Jingyun Ding. The effect of a multi-strain probiotic on growth performance, non-specific immune response, and intestinal health of juvenile turbot, Scophthalmus maximus L. Fish Physiology and Biochemistry, 45(4):1393–1407, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00635-4.

Mahanty:2019:HSI

[1354] Arabinda Mahanty, Gopal Krishna Purohit, and Bimal Prasanna Mohanty. Heat stress–induced alterations in the expression of genes associated with gonadal integrity of the teleost Puntius sophore. Fish Physiology and Biochemistry, 45(4):1409–1417, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00643-4.

Banavreh:2019:IPA

[1355] Akbar Banavreh, Mehdi Soltani, and Mehdi Shamsaie. Immunophysiological and antioxidant responses of Siberian sturgeon (Acipenser baerii) fed with different levels of olive pomace. Fish Physiology and Biochemistry, 45(4):1419–1429, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-019-00649-y.

Pan:2019:CMC

[1356] Wen-Qian Pan, Jun-Peng Wang, and Xiao-Qin Li. Cloning, molecular characterization, and tissue differential expression of connective tissue growth factor (ctgf ) of grass carp. Fish Physiology and Biochemistry, 45 (4):1431–1443, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00653-2.

Kandalski:2019:ELT

[1357] Priscila Krebsbach Kandalski, Tania Zaleski, and Lucélia Donatti. Effect of long-term thermal challenge on the Antarctic notothenioid Notothenia rossii. Fish Physiology and Biochemistry, 45(4):1445–1461, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00660-3.

Hemalatha:2019:DDP

[1358] Devan Hemalatha, Bojan Nataraj, and Mathan Ramesh. DNA damage and physiological responses in an Indian major carp Labeo rohita exposed to an antimicrobial agent triclosan. Fish Physiology and Biochemistry, 45 (4):1463–1484, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00661-2.

Young:2019:IAH

[1359] Tim Young, Seumas P. Walker, and Jane Symonds. Impact of acute handling stress, anaesthesia, and euthanasia on fish plasma biochemistry: implications for veterinary screening and metabolomic sampling. Fish Physiology and Biochemistry, 45(4):1485–1494, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00669-8.

Falcao:2019:BBE

[1360] Bruno Falcão, Márcia Marques, and Bruno Nunes. Behavioral and biochemical effects of the antifouler and antidandruff zinc pyrithione on the freshwater fish Gambusia holbrooki. Fish Physiology and Biochemistry, 45(4):1495–1512, August 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-019-00634-5.

Shi:2019:LOC

[1361] Chao-Ming Shi, Hang Zhao, and Shi-Mei Lin. Linseed oil can decrease liver fat deposition and improve antioxidant ability of juvenile largemouth bass, Micropterus salmoides. Fish Physiology and Biochemistry, 45(5):1513–1521, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-019-00636-3.

Kisten:2019:PRJ

[1362] Y. Kisten, N. A. Strydom, and S. Paul. Physiological responses of a juvenile marine estuarine-dependent fish (family Sparidae) to changing salinity. Fish Physiology and Biochemistry, 45(5):1523–1531, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00637-2.

Liu:2019:GSE

[1363] Qi Liu, Hongwei Yan, and Ying Liu. Growth and survival of Takifugu rubripes larvae cultured under different light conditions. Fish Physiology and Biochemistry, 45(5):1533–1549, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00639-0.

Borissov:2019:ESP

[1364] Radoslav S. Borissov, Sølvi Espeland, and Martin H. Iversen. Evaluation of the i-STAT (portable clinical analyser) for measuring haematological parameters in Atlantic cod (Gadus morhua) at different CO_{2} and temperature conditions. Fish Physiology and Biochemistry, 45(5): 1551–1562, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00638-1.

Meng:2019:RGP

[1365] Xiao-Lin Meng, Zhen-Xiang Zhu, and Guo-Xing Nie. Regulation of growth performance and lipid metabolism in juvenile grass carp (Ctenopharyngodon idella) with honeysuckle (Lonicera japonica) extract. Fish Physiology and Biochemistry, 45(5):1563–1573, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00644-3.

Garza-Rodriguez:2019:OLB

[1366] María Lourdes Garza-Rodríguez, Rafael González-Álvarez, and Iram Pablo Rodriguez Sanchez. Olfactomedin-like 2 a and B (OLFML2A and OLFML2B) profile expression in the retina of spotted gar (Lepisosteus oculatus) and bioinformatics mining. Fish Physiology and Biochemistry, 45(5):1575–1587, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-019-00647-0.

Zhang:2019:SGD

[1367] Ya-Lin Zhang, Xu-Dong Duan, and Xiao-Qiu Zhou. Soybean glycinin decreased growth performance, impaired intestinal health, and amino acid absorption capacity of juvenile grass carp (Ctenopharyngodon idella). Fish Physiology and Biochemistry, 45(5):1589–1602, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00648-z.

Wang:2019:CAG

[1368] Qi Wang, Hua Mu, and Kangsen Mai. Comparative analysis of glucose metabolism responses of large yellow croaker Larimichthys crocea fed diet with fish oil and palm oil. Fish Physiology and Biochemistry, 45(5): 1603–1614, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00646-1.

Gronczewska:2019:BFS

[1369] J. Gronczewska, N. Niedźwiecka, and E. F. Skorkowski. Bioenergetics of fish spermatozoa with focus on some herring (Clupea harengus) enzymes. Fish Physiology and Biochemistry, 45(5):1615–1625, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00650-5.

Zhao:2019:DTA

[1370] Ye Zhao, Xiao yun Wu, and Jun Jiang. Dietary tryptophan affects growth performance, digestive and absorptive enzyme activities, intestinal antioxidant capacity, and appetite and GH–IGF axis-related gene expression of hybrid catfish (Pelteobagrus vachelli [female sign] × Leiocassis longirostris [male sign]). Fish Physiology and Biochemistry, 45(5): 1627–1647, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00651-4.

Xu:2019:CEA

[1371] Yingyan Xu, Qingsong Tan, and Junpeng Yao. Characterization and expression analysis of FGF6 (fibroblast growth factor 6) genes of grass carp (Ctenopharyngodon idellus) reveal their regulation on muscle growth. Fish Physiology and Biochemistry, 45(5):1649–1662, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00655-0.

Liu:2019:SRG

[1372] Jie Liu, Erteng Jia, and Dingdong Zhang. Selection of reference genes for miRNA quantitative PCR and its application in miR-34a/Sirtuin-1 mediated energy metabolism in Megalobrama amblycephala. Fish Physiology and Biochemistry, 45(5):1663–1681, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00658-x.

Nissen:2019:IBS

[1373] Andrew C. Nissen, Brooke J. Vetter, and Allen F. Mensinger. Impacts of broadband sound on silver (Hypophthalmichthys molitrix) and bighead (H. nobilis) carp hearing thresholds determined using auditory evoked potential audiometry. Fish Physiology and Biochemistry, 45(5): 1683–1695, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00657-y.

Gao:2019:MCC

[1374] Yunhong Gao, Qiqi Jing, and Yudong Jia. Molecular cloning, characterization, and mRNA expression of gonadotropins during larval development in turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 45(5):1697–1707, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-019-00656-z.

Lv:2019:EAF

[1375] Liyuan Lv, Xu-Fang Liang, and Shan He. Effect of agmatine on food intake in mandarin fish (Siniperca chuatsi). Fish Physiology and Biochemistry, 45(5):1709–1716, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-019-00659-w.

Andreeva:2019:IAM

[1376] Alla M. Andreeva, Alexey S. Vasiliev, and Alexandra Filippova. Involvement of apolipoprotein a in maintaining tissue fluid balance in goldfish Carassius auratus. Fish Physiology and Biochemistry, 45(5): 1717–1730, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00662-1.

Zhang:2019:MCE

[1377] Xinyi Zhang, Zhiyong Zhang, and Shixia Xu. Molecular cloning and expression pattern of IGFBP-2a in black porgy (Acanthopagrus schlegelii) and evolutionary analysis of IGFBP-2s in the species of Perciformes. Fish Physiology and Biochemistry, 45(5):1731–1745, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00665-y.

Qin:2019:EDP

[1378] Qin Qin, Xiu-Fei Cao, and Wen-Bin Liu. Effects of dietary protein level on growth performance, digestive enzyme activity, and gene expressions of the TOR signaling pathway in fingerling Pelteobagrus fulvidraco. Fish Physiology and Biochemistry, 45(5):1747–1757, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00664-z.

Jang:2019:SSP

[1379] Jun-Chul Jang, Gyeong Eon Noh, and Jong-Myoung Kim. Spectral sensitivity and photoresponse in the rock bream Oplegnathus fasciatus and their relationships with the absorption maximum of the photoreceptor. Fish Physiology and Biochemistry, 45(5):1759–1769, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00672-z.

Xi:2019:CIC

[1380] Yuan Xi, Xu Meng, and Tu Zhi-ying. Color-induced changes in oxygen consumption and swimming performance of juvenile bighead carp (Aristichthys nobilis). Fish Physiology and Biochemistry, 45(5):1771– 1777, October 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00671-0.

Rong:2019:EDH

[1381] Hua Rong, Yunlong Zhang, and Xiaobo Wen. Effects of dietary hydroxyproline on collagen metabolism, proline 4-hydroxylase activity, and expression of related gene in swim bladder of juvenile Nibea diacanthus. Fish Physiology and Biochemistry, 45(6):1779–1790, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00676-9.

Dadras:2019:VAE

[1382] Hadiseh Dadras, Viktoriya Dzyuba, and Borys Dzyuba. In vitro antioxidant enzyme activity and sperm motility at different temperatures in sterlet Acipenser ruthenus and rainbow trout Oncorhynchus mykiss. Fish Physiology and Biochemistry, 45(6):1791–1800, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00675-w.

deMattos:2019:CEH

[1383] Bruno Olivetti de Mattos, José Fernando López-Olmeda, and Rodrigo Fortes-Silva. Coping with exposure to hypoxia: modifications in stress parameters in gilthead seabream (Sparus aurata) fed spirulina (Arthrospira platensis) and brewer’s yeast (Saccharomyces cerevisiae). Fish Physiology and Biochemistry, 45(6):1801–1812, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00677-8.

Lawrence:2019:CMM

[1384] Michael J. Lawrence, Erika J. Eliason, and Steven J. Cooke. Cortisol modulates metabolism and energy mobilization in wild-caught pumpkinseed (Lepomis gibbosus). Fish Physiology and Biochemistry, 45(6): 1813–1828, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00680-z.

Xu:2019:NCH

[1385] Yang Xu, Yang Tian, and Qingwei Li. A novel CDK-2 homolog identified in lamprey, Lampetra japonica, with roles in apoptosis. Fish Physiology and Biochemistry, 45(6):1829–1843, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00683-w. See correction

Liang:2019:GSC

[1386] Yinyin Liang, Huayang Guo, and Dianchang Zhang. Genomic structure and characterization of growth hormone receptors from golden pompano Trachinotus ovatus and their expression regulation by feed types. Fish Physiology and Biochemistry, 45(6):1845–1865, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00682-x.

Kaneko:2019:UEG

[1387] Nobuto Kaneko, Meredith L. Journey, and Munetaka Shimizu. Utilization of an endocrine growth index, insulin-like growth factor binding protein (IGFBP)-1b, for postsmolt Coho salmon in the Strait of Georgia, British Columbia, Canada. Fish Physiology and Biochemistry, 45(6): 1867–1878, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00681-y.

Liu:2019:GPM

[1388] Bo Liu, Hua-Yang Guo, and Dian-Chang Zhang. Growth, physiological, and molecular responses of golden pompano Trachinotus ovatus (Linnaeus, 1758) reared at different salinities. Fish Physiology and Biochemistry, 45(6):1879–1893, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-019-00684-9.

Mahmoud:2019:SNI

[1389] Usama M. Mahmoud, Imam A. A. Mekkawy, and Alaa El-Din H. Sayed. Silver nanoparticle–induced nephrotoxicity in Clarias gariepinus: physio-histological biomarkers. Fish Physiology and Biochemistry, 45(6): 1895–1905, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-019-00686-7.

Abdel-Razek:2019:EDP

[1390] Nashwa Abdel-Razek, Somayah M. Awad, and Mohsen Abdel-Tawwab. Effect of dietary purslane (Portulaca oleracea L.) leaves powder on growth, immunostimulation, and protection of Nile tilapia, Oreochromis niloticus against Aeromonas hydrophila infection. Fish Physiology and Biochemistry, 45(6):1907–1917, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00685-8.

Fan:2019:PKG

[1391] J. J. Fan, X. H. Tang, and P. Jiang. Pyruvate kinase genes in grass carp Ctenopharyngodon idella: molecular characterization, expression patterns, and effects of dietary carbohydrate levels. Fish Physiology and Biochemistry, 45(6):1919–1931, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00688-5.

Andreyeva:2019:HDM

[1392] A. Y. Andreyeva, A. A. Soldatov, and S. Gambaryan. Hemoglobin deoxygenation and methemoglobinemia prevent regulatory volume decrease in crucian carp (Carassius carassius) red blood cells. Fish Physiology and Biochemistry, 45(6):1933–1940, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-019-00689-4.

Ma:2019:LPA

[1393] Rongrong Ma, Wenhong Fang, and Kun Hu. Liver proteome analysis of grass carp (Ctenopharyngodon idellus) following treatment with enrofloxacin. Fish Physiology and Biochemistry, 45(6):1941–1952, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1007/
s10695-019-00690-x.

Wang:2019:GIS

[1394] Yan Wang, Xiaochen Zhan, and Linjie Wang. GSK3β inhibition suppresses the hepatic lipid accumulation in Schizothorax prenanti. Fish Physiology and Biochemistry, 45(6):1953–1961, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/s10695-019-00691-w.

He:2019:CTA

[1395] Fei-Xiang He, Dong-Neng Jiang, and Chun-Hua Zhu. Comparative transcriptome analysis of male and female gonads reveals sex-biased genes in spotted scat (Scatophagus argus). Fish Physiology and Biochemistry, 45(6):1963–1980, December 2019. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-019-00693-8.

Xu:2021:CNC

[1396] Yang Xu, Yang Tian, and Qingwei Li. Correction to: A novel CDK-2 homolog identified in lamprey, Lampetra japonica, with roles in apoptosis.

Fish Physiology and Biochemistry, 47(1):189, February 2021. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:

//link.springer.com/article/10.1007/s10695-020-00890-w. See

[1385].

Huang:2025:CDE

[1397] Weiren Huang, Yong Zhang, Xiaoping Jia, Xilan Ma, Shuisheng Li, Yun Liu, Pei Zhu, Danqi Lu, Huihong Zhao, Wenna Luo, Shibai Yi, Xiaochun Liu, and Haoran Lin. Correction to: Distinct expression of three estrogen receptors in response to bisphenol A and nonylphenol in male Nile tilapias (Oreochromis niloticus). Fish Physiology and Biochemistry, 51(1):??, February 2025. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/
10.1007/s10695-024-01432-4. See [26].